A Selected Bibliography of Publications by, and about, Paul Adrien Maurice Dirac

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 March 2025 Version 1.212

Title word cross-reference

(1 + 1) [Das93]. (2 + 1) [JM99]. −1/2 [CT67a]. −1/2, 1/2, 3/2, 5/2 [Mac98]. −\frac{1}{2}, \frac{1}{2} [CT67b]. 1 [CG07]. 1 + 1 [Fei02]. 1/2 [CT67a, FA01, Ogu96, dB51, dB52a]. 2.7^{∘} [Noe73]. $250.00 [War96]. $26.00 [Ryc17]. $29.95 [RS10, Dys10]. 3 [RRS06]. 3 + 1 [Nak00]. 3 + 2 [Dir63b]. 3/2 [CT67a]. $44.50 [Dre90]. $49.50 [Sch85]. $7.50 [Suc67]. _{6} [DN96]. α [BM98, LT04]. α, β, γ, δ [Som36]. c [IC04, MBS02, Mag15]. D [DSP03]. δ [DEŠ89, IP56, IP57, Lut07, Sch72]. e [IC04, MBS02]. e^{2} [CP68]. ϵ [GGS04]. \frac{3}{2} [CT67b]. G [AWW^{+}81, Dir75f, Dir78b, Dir80g, Dir82f]. G_{F} [FA10]. Gr_{2}(C^{m+2}) [Mil98]. H^{s}(R^{3}) [EV97]. h̄ [IC04]. m_{p}/m_{e} [BBSF10, FK07]. F_{p}(η) = 1/Γ(p + 1) ∫_{0}^{∞}[ϵ^{p}/(1 + e^{ϵ−η})]dϵ [GFG01]. SU (2) [yRSBSV99]. H_{2} [SBK^{+}08]. N [Ata89, KY01, MHN98, MB90, Mor85, Won90]. ∇^{2}n/n [MHN98]. ∇n/n [MHN98]. π [JT10]. q [CPV10, JT10]. R [Szm98, Szm99]. R^{n} [Bet83]. r^{t} [BBR85]. s > 1 [EV97]. SL(2, C) [Hos82, Hos83]. SO(4) [LH86]. SO(4, 1) [BC69]. sp(4, R) [Lee95]. SU(N) [LH86]. SU_{2} [Mal75c, Yan78b, Yan78a].

-algorithm [GGS04]. -dimensional [Das93, KY01, MB90, Nak00, Won90]. -exponential [JT10]. -exponentials [CPV10]. -function [Lut07]. -functions [IP56, IP57]. -invariant [Hos82, Hos83]. -matrix [Szm99, Szm98]. -shell [DEŠ89]. -space [JM99]. -Uniform [RRS06].

/van [Sch85]. /Vigier [Sch85].

0-521-31911-0 [Ble86, Sch85]. 0-521-34013-6 [Hen89]. 0-521-34017-9 [Hen89]. 0-521-36231-8 [War96]. 0-521-38089-8 [Hen91].

1 [Dir43a]. 1911 [Meh75]. 1933 [CCJ^{+}34]. 1933/Chadwick [Ano51]. 1935/1936 [Ano51]. 1945 [Pau45]. 1948 [Hov96, Dar96]. 1959 [LT62]. 1960s [Mla98]. 1968 [Sal69]. 1971 [DPI71]. 1972 [Dys74, IPM73, Meh73]. 1973 [Kur73]. 1975 [DHS78]. 1978 [KPS^{+}78]. 1979 [Ruf82]. 1984 [Cas85, DP86]. 1988 [SDEW90]. 1993 [BCEP94]. 1994 [dG97]. 1999 [F^{+}00]. 1D [BJ03].

2 [UM86]. 20.X.1984 [Smo84]. 2002 [BB03]. 2009 [Pet11]. 2013 [Ano13]. 20th [Gun15, Meh73, SS03, Wal09, Koz90, Med90]. 21st [RCC^{+}06]. 22 [CCJ^{+}34]. 25 [Ble86].

3 [Ryc17]. 30.00 [Hen89]. 35 [Hen91]. 39 [Szm99].

42 [Pei42].

50 [Ben34]. 50th [DS83].

7.80 [Deh36]. 745 [Goa95, Goa97]. 779 [Mac98].

8.VIII.1902 [Smo84].

’91 [Wat91]. 978 [Ryc17]. 978-0-300-20998-3 [Ryc17].

AB4 [Pei42]. AB4/42 [Pei42]. Abb [Deh36]. Abdus [Bel73, Ell91]. abelian [KM84]. abermals [Haa34]. Abnahme [Tak31]. Abschlusses [Pau25]. absolute [Dir53b]. Absorption [Dir27d, MWF07, Sei36, SCPA04, DH32a]. abstract [Jär64, Tha88]. abstrakten [Jär64]. academia [Buc80]. Academic [Suc67]. Academy [Dir80c]. accelerating [UK82]. acceptance [Dir71a]. Acclaim [Ano30]. According [Inf53a, MRB75]. account [Alp73]. Accuracy [Neu33]. Accurate [BCJT16, Rob85, SM91, SFVW95, YG77, MN97, NM01]. achievement [Hoy92]. Achinstein [L.88]. Across [Bus10, DP03, II08]. action [Dir74a, Tol98]. actions [Lan01]. activity [Meh87b]. adaptation [SM35]. Address [Dir80a, Pei87a, Wig87a, Wig96a, Wig96b, Haw98, Sno73]. ADI [RSK99]. adiabatic [Dir25a, Dir25b]. adjoint [Zor80]. adjointness [KY01, LR79]. admitting [KM84]. Adrien [Ano52, Ano83a, Ano83b, DP86, Dir74d, Hal85, Hen89, Kra06a, KW87, Pei85, Ram89, Sch88, Stu88, Wig88]. Advance [Ano34b]. Advanced [Dys07a]. adventure [MB11b]. Aesthetic [McA90]. aether [Dir54a, Dir51b, Dir52a, Dir52b, Dir54d, CO03]. Aethers [Dir53a, Dir53a]. af [BR64, RRK^{+}64]. After [Dir60b, Kae48]. against [Noe73]. Age [Stu18, Meh72, Meh87b]. Aharonov [Tam03, Vil95]. Aid [CMO00]. Airplane [GD10]. Akademische [Deh36]. al [Sch85]. Alapveto [Dir78a]. Albert [Ano05, HE82, Kle05, PS79]. alcohol [Mag15]. algebra [Bor85, CDK90, Cra85, Dir26b, DW09, Ham84, IKRT89b, Lee95, MB91, Mar88, yRSBSV99, Nas84, Ten68, Vrb94]. Algebraic [BN72, PSA94, QGW01, RR93, Wat91, AT94]. Algebraical [Ano00]. algebras [ML93]. Algorithm [IKRT89a, GGS04, PSA94, TM85, Goa95, Goa97, Mac98]. állandôk [Dir78a]. Allen [Tiw12]. allows [GGS04]. Almost [Tau61b]. alpha [BMS02b, KM04, MBS02, Bar02]. alsidighed [Dir64g]. Alternative [Aug95, Bar81, Dra23]. Alwyn [Ble86, Tor87]. Amateur [Har01]. Ambition [Jon08]. amplitudes [Eeg80, NM84]. Anachronism [Sta12]. analogie [dB32b]. analogies [FG22]. analogy [Dir45b, dB32b]. Analyses [Dar91, Dar92]. Analysis [LU31, Tau63b, BNB94, CK90, CK91, Lig87, LW75, MB90, MB91, Mar88, Pon05, de 89]. Analytic [Tu91, Fer86, KY93]. Analytical [BDM81a, Kra81a, Lan05b, Nak97]. Anderson [Ano51]. anisotropic [Por95]. Annahmen [Dir77a]. annihilation [Dir30e]. anniversary [DS83]. announcement [Lub75]. anomalous [BB85, Beh85, Dat04, Ogu96, Rou84]. Anpassung [SM35]. Ansatz [CH02]. Anschluss [Jär64]. anti [Dit15]. anti-Dirac [Dit15]. antilinear [HW90]. Antimatter [Clo10, Dir78h, Jac98]. Anton [Fis10a, Fis12a, Fis10a, Fis12a]. apart [Far09a]. Application [Dav98, DFPP43, Dir45a, LU31, Tiw12, Dir26f, LR79, LRK80, OP95, SMI83, Wig01, WY85, Dav98, PFD43]. Applications [CO03, ERW75, FLS76, RL68, SS11, Tha88]. Applied [F^{+}00, LR59, Kim80a, Kim80b, MD88]. Appointed [Ano34b]. appreciation [Har01]. Approach [Bro05, Fey48, FBD05, CH02, Dir78f, DSP03, IT84, Jun97, Leh91, Nak97, OP95, RdO07, RdO16, SBZ90, Bri18]. approaches [Dat04]. approximants [Rod80]. Approximate [Dir30a, Dir43a, IS10a, PFD43, SM35]. Approximating [CK90, CK91]. Approximation [QGW01, Inf53b, SS11]. Approximations [CT67b, SBB^{+}62, CT67a, Fer86]. Äquivalenzverbot [JW28]. Arbeit [Hei34b]. Arbitrary [Dir43a, IKRT89b, Kat00, PFD43]. Archimedes [Pic08]. Aristole [vW02]. Aristoteles [vW02]. arrival [Dir82b]. Arthur [Deh36]. Asim [Ble86, Tor87]. Aspects [SW72, Meh75, Pai98, Bel73]. assessment [SS03]. Associated [GZM05]. association [HC87]. Assumptions [Dir77a]. astounding [Haw11]. astronomers [Car01]. Astronomia [RCC^{+}06]. astronomical [SV90]. astronomy [FF91]. Astrophysics [Tau63c, IPM73]. Asymptotic [BG91, Bou08, Din58, IM95, Nov74, Wal76, CK78, GFG01]. asymptotically [Ogu96]. asymptotics [Doa91, NP96, SM03]. atmosphere [Dir25d]. Atom [Ano30, Dir26e, Dir67c, EKI94, Far09b, Har28, Pau25, Rod19, Bun52b, Dat04, Dir30c, MR10, Sto94, RS10, Dys10, Cas10, Bla11]. atomaren [Sei36]. Atomic [BBD^{+}46, Dir84b, FR13, HSD34a, Ish90, McW73, Pei97, Rob85, Sei36, SH59, WBR76, Bel91, CCJ^{+}34, DSK72, Dir36a, Dir55f, LBNZ98]. atomique [Dir48e]. atomiques [CCJ^{+}34]. átomo [Bun52b]. Atoms [CR12, Dat95, Csa81, Dir25c, DH32a, DFW99, Tiw12]. Atomtheorie [Ben34, HSD34a, Wil36]. attached [Pei42]. Attraction [KK10]. Aufl [Deh36, Haa34]. August [DP86, F^{+}00, Zic88, DHS78]. Ausgestaltung [Lan30]. auspices [CCJ^{+}34]. Australia [DHS78]. autographs [PKKL09]. autoionization [ACM10]. Automata [Tau63b]. automatic [Nil03b]. automorphism [Cra93]. autres [Kle05, Ano05]. Award [Ano33, LW34]. Axisymmetric [JP73].

b [Sch85]. background [CGKS94, CV02, Fei02, FSY00, Han99, Por95, SS98, YS03]. bag [Fei02]. Baltimore [L.88]. bang [Das96, YS03]. banquet [Sno73]. Barbara [GG76]. Bargmann [MS95]. Bargmann- [MS95]. Barut [Ble86, Tor87, Sch85]. Based [EKI94, AT07, FA10, Haa34]. Basic [Dys10, RS10, Dir69a, Jor52, Jor55, MS87]. Basis [Dat95, QGW01, Dir29a, Dir38b]. Be [Ano30, Far03, Dir62a, Dir69b, Dir70a]. Beautiful [Far03, Wri16]. Beauty [Dir54b, Dir79a, HK93, McC04, Smo93]. became [Cri18]. Beck [Nis35a, Nis35b]. Bedeutung [Jor59a, Smo93, Som36]. before [Hen83]. Began [Kae48]. behavior [Son99]. behaviour [CK78]. behind [Pic08]. Behram [L.88, Ram89, Hen89, Sch88, Stu88]. Belfer [Suc67]. Belief [Dir76a]. believe [Dir80h]. belleza [Bas08]. Bemerkungen [Hei34b, vN28, Hei34a]. Bergmann [GNH78, Kru94, TM85]. Berichtigung [Hei34b]. Besprechung [Hei30a]. Bessel [Car92]. Bethe [CH02]. Between [Stu18, Bal85, Bok04, Dir25c, Dir39c, Dir45b, dB32b]. Beyond [Rod19]. Bibliography [RZ80]. Bicentennial [Pet11]. Big [Hür22, YS03, Das96]. Binary [VKA21]. biographical [Dal87b]. biographies [Ano65]. Biography [Dre90, Hen91, Deb13, Kra90, Sch91b, Bro91, Dar92]. biology [BP79b]. Birth [BH83a, BH83b, RCC^{+}06, Sch94b, ACM10, BH82]. birthday [BvdM83, Cha73]. bis [Fis10a, Fis12a, vW02]. bispinor [Cra85]. black [FSY00]. Blackett [Dir84a]. Body [Bus10, CV90, GL92, II08, LC98, SCSB93, Sem93]. Bohm [AW17, Col12, Fab22, Tam03, Vil95]. Bohr [Lak96, Bal85, BR64, Bok08, Dir24b, Dir28a, Dir64g, Dir67d, Kra17, MR82d, MR82e, RRK^{+}64, Roz67]. Bohrs [Dir64g]. Boltzmann [Faz15]. Bondi [Dir52a]. Bonn [Wat91]. Book [And88, Bel73, Ble86, Bri18, Bro91, Cas10, Dar96, Dav36, Dra23, Dre90, Dys74, Dys10, Ell91, Eps35, Ett31, Goe16, Hen89, Hen91, Hov96, Kib98, Kil76, L.88, LJ31, Pau30, Pol58, Ram89, RS10, Ryc17, Sch69, Sch85, Sch88, Sch91b, Sti13, Stu88, Suc67, Tem35, Tor87, War96, Bla11, Jär64]. Books [Dys10, L.88, K.L88, RS10]. boomerons [Leo80]. Boosting [CKPT04]. Born [Dir60d, Gre05]. Boronlike [VIK98]. Bose [BBR03, Din58, Faz15, Gau93, Kap40, Kos86]. bosons [Bel85, RM96]. both [Cra88]. Bound [BB85, RU81, Cor70, IS10a, ZQ90]. bound-state [Cor70]. Boundary [GZM05, JP73, FGM^{+}96, FSS98, RS87]. Bounds [SV90, SV95, MS95]. Bra [Rob66, D’A02, Pru73]. bracket [FK70, Lan72, MS68, Rug88, Wot91]. brackets [FLS76, Sei99]. brain [Lip87]. Branch [Bun82]. bras [Jau72]. Brave [HW06]. Brecht [Sch99]. Breit [EKI94, Ish90]. brief [Pol87b]. British [Wal09]. Broglie [AW17, Ble86, Bou01, Col12, Deh36, Fab22, Haa34, Sch85, Tor87, Bou01, BVV84, Dav98, Haa34, dBDW^{+}84]. Broken [Pei92, Dir73c, Dir73d]. Brownian [AW95]. Brueckner [MHN98]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34, Far01]. Buch [Jär64]. Buchbesprechung [Deh36, Hei30b, Wil36]. Bücher [Ben34]. bucket [BP95]. bulk [vN96]. Bundle [Bri18, Oev88, RdO07, RdO16]. bundles [Gre77, UB83]. Bunge [Hin77]. Bunin [Ano33].

C.W [Hen89]. Cabrera [Lub82]. Calculated [Dat95]. Calculation [IKRT89a, SBB^{+}62, Vul03, NM84, Pic89, Sag91]. Calculations [DFW99, Eeg80, Ish90, QGW01, VIK98, Dir55c, EG84, IKRT89b, IKRT91, YG77]. calculus [Eeg80]. California [Ano34a]. called [Gol05]. Calogero [MS95]. Calogero-type [MS95]. Cambridge [Ble86, Dre90, Hen89, Hen91, Pei92, Sch85, War96, Ano30, Ano24c, Dir37c, EP72, Hen84, KN03, Pei87a]. Can [Dir69b, Dir70a]. canonical [Bel85, BC69, Olv86, PSA94]. Carolina [BCEP94]. Cartan [DMH85]. Cartesian [SC91]. Case [Bou08, Pas12, Pic81, Kra81a, RSK99]. case-study [Kra81a]. cases [All03]. cat [Lam87]. Catania [RCC^{+}06]. Cauchy [Cha72]. Celebration [Hen89]. Centenary [BCEP94, Hen89, GC97, SS03, Wil05]. Centennial [RCC^{+}06, BB03, HE82]. Center [KPS^{+}78, KP83]. Central [Har28, SM91, SCSB93]. Centre [Meh73, Ruf82, Sal69]. centrifugal [IU86, YI90]. centrifuge [Dir42c]. Century [Gun15, RCC^{+}06, Bra09, CM86, Meh73, SS03, Wal09, Wei77, Koz90, Med90]. Certain [Dir79c, Wal76, Gre05]. certificates [Ano10]. Chadwick [Ano51, Dir75e]. challenge [CT12]. challenges [F^{+}00]. champ [Dir39a]. Change [Dir79e, Gam67a, Tel48, Wes80, Wil58]. Changing [Gun15]. chapter [Kra15a]. Characterization [YYH^{+}01]. characters [Coq85]. Charge [Dys67, Gam67d, Per67]. charged [AW17, Bor85, BG90, FA01, LMV91, RU81]. che [Fer24]. Chebyshev [CT67a, CT67b]. Chemistry [Bun82, DN96, Len14]. Chemists [Sim02]. chiral [Fin00, Wot91]. chromodynamics [F^{+}00]. chronology [Rec97]. Chronon [GF02]. Chunk [BMP14]. CI [DN96]. circle [All03, BR64]. Circulation [Bun52a]. Claim [Sim02]. Clarifying [GG11]. Class [BG91, DK89, EF77, FM75, IV85, IK91, KL90, Wil91]. Classical [BG90, Dir38a, LBNZ98, Roh60, Bok04, Dir33a, Dir45b, Dir49a, Dir51c, Dir51d, Dir52d, Dir54c, FK70, GS79, MS68, RU81, Sno73, IU86]. Classically [GPS86]. classification [KLMW88]. classmate [Cri18]. Clifford [Bri18, CDK90, ML93, PP71, RdO07, RdO16, UB83, Vrb94]. Closed [BBR85, Mor85, Bok04]. Closed-Form [BBR85]. closer [BBSF10]. Cloud [Inf53a]. Cluster [EKI94]. clusters [EG84]. co [RRK^{+}64]. co-workers [RRK^{+}64]. codes [DSK72]. colleagues [BR64, Roz67]. Collected [Dar96, Hov96, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, War96, Dal95, Wig96d, Dir02, Dir03b, Dir04, Dir05b]. Collection [CMO00, Bus10, Koz90, CMO00]. College [Dir37c, Jef87]. collisions [Dir27e]. Colloquium [ACM10, Bat07, UM86]. combined [HD77]. combs [Ric03]. Commemoration [RCC^{+}06, Buc80]. Comment [BF12, Kha86, MWF07]. commentary [DP97]. Commission [BBD^{+}46]. Committe [Ano67]. commutation [Bel85]. commuting [KLMW88]. compact [BT90]. Comparison [QGW01, Dat04]. Complete [GFG01, Goa95, RZ80, KLMW88, KL92]. completion [MR00, Pau25]. Complex [Dir37a, Gre77, Sla29, Eeg80, Hes75, Mar88, Nas86, Pau25, RA11]. comprehensive [Her18]. Compton [Dir25d, Dir26f, Dir27a]. Computation [Fuk14, Goa95, MS38, Wat91, Doa91, Gau93, TM85]. computational [Len14]. computations [AT07]. Computer [F^{+}00, IKRT89b]. Computers [Tau63b, Faz15, Len14]. Computing [Ano18, Bel91, F^{+}00, BDM81b, BDM84]. concept [Dir78e, Kra81a]. Conception [Dir79c, Dys74, Dir73a, Meh73]. Conceptual [Dun12]. Concluding [Wig96b]. conclusions [Jor66]. condition [Dir24b, Her18, YI90]. conditions [Dir25c, Dir62a, FGM^{+}96, FSS98]. Conduction [DuM28]. cone [Fin00]. conféerence [Dir48e]. Conference [BCEP94, Dor82, IPM73, Kur73, Pet11, RCC^{+}06, DPI71, Far01]. conferences [Meh75]. confined [Das93]. Confirmation [DuM28, Jor62b]. Conformal [MSB92, Son99, Dir36c, IU86]. congruences [KM84]. conjecture [CT83, CPV10]. conseil [CCJ^{+}34]. Consequences [Dir78b, HE06, Jor62a, Jor71, Dir75d, Jär64, HE35]. Conservation [Bou85, GS79, GLR14, Dir36a, LCH10]. Conserved [Ker83, CV02, SES82]. Considerations [Coh70, Fer24]. Considerazioni [Fer24]. Consistency [GH75]. Consistent [Ish90, EG84, MHN98, YL09]. Consistently [BSW88]. Constancy [Irv83, Per67]. Constant [Ano12, BS67, Bek86, BF12, Dir72f, Dir75b, Dir79e, FK07, JRF06, KM12, Mar84, MWF07, Noe73, Rob83, Sch91a, SCPA04, Tel72, VKA21, WB78, WFC^{+}99, Wil93a, WW86, Alp73, AG03, BBSF10, Car01, CKYY07, Cho81, Dir78c, FA10, Hal78, Han99, Inf61, Inf63, Lev80, MG87, Nor86, Ogu96, Peg77, Sch99, ZdSS91]. Constants [Gam68, Kra19, SV95, Tel48, Tel72, Wil58, Alp73, Bar02, BMS02a, Car01, CGK72, Dic59, Dir37b, Dir73b, Dir78a, DOV02, Dys72, Dys78a, Dys78b, DFW99, ES10, GT85a, GT85b, GT86, IC04, Jor37, Kha86, KPW86, LMSV08, Nil03a, Nor86, Pet85, Pet88, SV90, Uza03]. Constrained [Tay87a, BNB94, GB90, Sei99]. Constraining [JRF06, SBK^{+}08]. constraint [CY91]. Constraints [sC78, Lee21, Pic81, VKA21, GNH78, Han99, IC04, Kru94, LMSV08, Oev88, Olv86, TM85, YS03]. construction [Nas97]. Constructive [Sta12]. Contact [Löw64, DEŠ92]. contain [Fer24]. container [Dir42a, Dir42b]. containing [DW09]. Contemporary [Sal69, Sal69]. contengono [Fer24]. Context [Sch91a, Pet11]. Continues [Ble86, Sch85, Tor87, BVV84, dBDW^{+}84]. Continuous [Tau62]. continuum [Bel91]. Contribution [Ano30, Meh87a]. contributions [ES10]. Control [GZM05]. Controversy [Pic81]. convection [CG07]. Convergence [Rod80, Tam03]. Converging [Din58]. conversations [Bri95a, BP79b]. Coordinate [Neu33]. coordinates [Dir59d, SC91, SC92a, Vil90b]. Copley [Ano24a]. Coral [IPM73, KPS^{+}78, KP83, Per80, DPI71, Kur73, Kur87b]. Corepresentations [HW90]. Cornelius [BCEP94]. Correction [Ehb23, Hei34b]. Corrections [Kir97, Csa81, GF02]. corrector [Doa91]. Correlation [EKI94, MHN98]. Correspondence [Ano67]. Corrigendum [CT67a]. coset [BILP03, OP95]. Cosmic [BS67, WBR76, Han99, YS03, Mil97]. Cosmo [Kra82]. Cosmo-Physics [Kra82]. Cosmological [CO03, Dir74b, IRV^{+}08, Sch91a, BM98, BMS02a, DSW08, Dir37b, Dir78c, Dir78f, Hal78, Jor59b, Por95, ZdSS91]. Cosmologies [Kra06b, GT85a, GT85b, GT86]. Cosmology [Alp73, BBBM90, Dic61, Dir61a, Dir78g, Dir05b, Gam67c, Gam67b, Goe16, Kra82, Rei72a, Rei72b, Dir38b, Dir72b, Jor52, Jor55, Kra16, Nor65, Nor90, PD62, ZdSS91]. Cosmonumerology [Kra91b, Kra91a]. cosmos [HW06]. cotangent [Oev88]. Coulomb [EKI94, Har28, WY85, BK76, BM70, BBR85, BJ03, CH02, DEŠ92, DSP03, Hos83, Hos87, Hyl84, IS10a, IS10b, IM95, Kim80b, yRSBSV99, RU81, ROB64, SWO76, Vil95, WY84, Won90, de 89]. Coulomb-like [IS10a, IS10b]. counts [Lub82]. Coupled [BSW88, EKI94, BB85, Cha72, LC98]. Coupling [Löw64, DST92, GT85a, GT85b, GT86]. couplings [DSW08, YZ08, YL09]. course [Zic88]. covariance [Bel85, UB83]. Covariant [UL31, CV90, CY91, LC98]. Creation [HND^{+}62, CM86]. Creativity [Sch94c]. creators [Str11]. criteria [DSK72]. Critique [Ano05]. cuántica [Dir67b, Dir68c, Dir69e]. cuántico [Dir54a]. cubic [Pen99]. cultural [HE82]. Curies [Kae48]. Current [Pas12, Bor85, Jor59b, LW78]. Currents [Ker83, SES82]. curved [BO92, Ran87, Rüd84, Vul03]. curvilinear [SC92a, Vil90b]. cylindrical [BR97, Vil90b]. Cylindrically [BSW88]. Czech [Dir59b].

d [Jor66, CG07]. d’Albert [Buc80]. Dalitz [Dar96, Hov96, War96]. Danish [BR64, Dir64g, RRK^{+}64]. dans [dB32a, dB52b]. darkest [HW06]. darling [Dir87a]. Data [SV95, CG07, Cor70, Han99, SV90]. dataset [Lee21]. De-mystifying [Lut07]. dead [Dir65a]. Debate [Pas12, Bok08, HSZ76]. Debye [BCK97]. December [BB03, BCEP94]. declining [Kra15a]. Decoding [RS10]. decrease [Tak31]. Deduction [BI34, BI34]. déduit [Bor49]. déecembre [Dir48e]. Déecouverte [Dir48e]. Deep [Kol67]. Deep-Sea [Kol67]. deepest [Bar02]. déeveloppements [Dir48e]. défient [CT12]. Defining [Mla98]. Definite [Dir79b]. Deformations [FLS76]. degeneracies [LMV91]. Degeneracy [BM70, JM99, MC70]. Degenerate [DuM28, Can98]. degli [Fer24]. d’Einstein [PS32a, PS32b]. del [Bun52b, Maj37, Wei77]. delayed [Rod80]. Delivered [Wig96a, DHS78, Wig87a]. dell [Maj37]. della [Wei77]. Delta [JP73, Lam69, Lam70, Nov74, Wal76, ABR06, BMP14, CPV10, Deb13, GGS04, JT10, LW75]. Denied [Ano54]. Dennis [Bin00]. Dennison [Ano67]. Dense [Ric03]. densités [dB32a]. densities [Cra85, dB32a]. Density [Csa92, Dir43a, Rij66, Csa81, Dir30d, Dir42d, Kos86, MHN98, PFD43, SS11, Wig01]. Department [Hal78]. dependence [Rob85]. dependent [AME14, BK10, IS10b, RSK99]. Derivation [Eva04, Wot91, Riv94]. derivatives [Cor70, GZDA01, MKK05]. derive [Cra88]. derived [Bor49]. description [Car94]. design [DSK72, Tau63b]. detect [GGS04]. Detecting [IRV^{+}08]. Determinants [FGM^{+}96, Lan01]. Deuterium [CR12]. Development [Aur71, Dir73a, Jor49, Meh73, Deb13, Dir71a, Dir73b, Dir74c, Dir78a, Dir81b, Meh75, Meh87a]. Developments [Ano62, Dir46, Dir79d, Dir73g, Per80]. devices [MZF^{+}12]. diagonal [SCSB93, Wig01]. diagonalization [Rou84]. Diagrams [Wüt11b]. Dialogue [Kra91b, Kra91a, Bel99]. Diatomic [FK07, BBSF10, YG77]. dice [MB11b]. Dicke [Dir61a]. Did [Far10]. Difference [LR59, QGW01]. Differential [Coq85, GZM05, Olv86]. differentiation [Nil03b]. Difficulties [Ano34a, Dir47a, MR82d, MR82e]. diffraction [Ric03]. diffusion [CG07]. dilations [All03]. dilaton [CKYY07]. dimension [IKRT89b]. Dimensional [Sch31a, Sch31b, AKV83, BV85, Das93, FA01, IT88, IU86, KY01, Kam86, Kis98, Lég95, MB90, Nak00, NP96, NdlC08, RSK99, Ten68, Won90]. dimensions [Car01, DSP03, Fei02, Kha86, KPW86, Zas89]. diminishing [Jor62b]. dinner [Jef87]. Dipartimento [RCC^{+}06]. Dipole [Roz20, Pos84]. Dirac [And88, Ano51, Ano13, BB03, Ben34, Bla11, Ble86, Bri18, Bus10, Cas85, Cas10, CT67a, CMO00, DP86, Dal87b, Dar96, Deh36, Dys10, Ehb23, Ell91, Fis10b, Fis12b, Goe16, Haa34, Hei30b, Hen89, Hen91, Hov96, Jär64, Kib98, KBN83, Lak96, Lan87a, Meh91, Neu33, PR86, Pau30, Pei92, RS10, Sch85, Sch91b, Smo93, Smo84, Stu88, Suc67, Szm99, Tiw12, Tor87, Tu91, War96, Wig96a, Wil36, dB51, dG97, AD87, AW95, All03, All04, AME14, AKV83, AW17, Ama68, AC72, AT07, Ano30, Ano33, Ano34a, Ano34b, Ano52, Ano54, Ano67, Ano75, Ano83a, Ano83b, Ano85, Ano94, Ano02, Ano08, Ano10, Anoxx, Ant69a, Ant69b, Ara93, ABR06, AT94, Ata89, Ati98, AT99, Aug95, BSW88, BO92, BOS96, BK10, BILP03]. Dirac [Bal85, BC78, BDM81a, BDM81b, BDM84, BCJT16, BNB94, Bar81, BK76, BvdM83, BVV84, BB85, Bas08, Bat00, Bat07, BSW98, BN72, Beh85, Bel85, Bel91, Bel73, Bel82, Bel94, BMP14, BCK97, BM70, BBR85, Bet83, BBR03, BtGW95, BG91, Bok04, Bok08, BR97, Bor85, BI34, Bor52, Bou85, Bou01, Bou08, BC69, Bri95b, Bro77, BR87, Bro06, BJ03, BP79a, Bue05, BT90, Bun52a, Bun52b, BV85, CT83, Can98, Car92, Car94, CG07, CO03, Caw69, Cha72, Cha73, CGK72, sC78, CPV10, CH02, CDK90, CGKS94, CH16, Cla94, CT67b, CK90, CK91, Coh70, Col12, Coq85, Cor70, Cor07, CV02, CV90, CY91, Cra85, Cra88, Cra93, Cri18, Cri76, Csa81, Csa92, D’A02, DJ95, Dal87a, Dal95, Dar90, DK89, Das93, Das96, Dat95, Dat04]. Dirac [Dav98, DSK72, Dav36, DN96, DL82, Deb13, Dic61, DMH85, Din58, Dir61a, Dir74d, Dir80a, DRFS80, DS83, DEŠ89, DEŠ92, Dit15, DST92, Doa91, DSP03, DW09, Dra23, DuM28, DRF17, DP03, DR93, Dys74, Dys87, Dys07b, EP72, Eeg80, Ehb22, EM33b, EM33a, EF77, EKI94, Eli87, ERW75, EG84, Eps35, EGT21, EV97, Esp12, Ett31, Eva04, Fab22, FA01, FG87, FGM^{+}96, FSS98, Far09a, Far09b, Far10, Far16, Faz15, Fei02, FM75, Fer49, Fer86, FW87, Fin00, FSY00, Fio06, FLS76, FK70, Fre87, Fre95, Fuk14, FR86, FG22, FZ91, GG02, GLMR94, GL92, GF02, GB90, Gam33, Gar84, GFG01, Gau93, GJEF10, Gin09, GB13, GS84, GS79, GD10, Goa95, God98, GG11, Gol05, GG51, GZDA01, Goo55, GZM05, GNH78, Got11, Gru00, GH94]. Dirac [GGS04, GH75, Haa34, Hal87, Hal78, Hal85, HD77, Ham84, HW90, HC87, Hau72, Haw98, Hei30a, Hei34b, Hei34a, HE35, HDW75, HE06, Hen83, Her18, HJL00, Hes73, Hes75, HL38, Hil92, Hin77, HJKS91, Hor11, Hos82, Hos83, Hos87, Hos89, HM99b, HM99a, HK93, Hov96, Hoy92, Hyl84, IT84, IT88, IJ02, IS10a, IS10b, IKRT89a, IKRT89b, IKRT91, IM95, Inf34a, Inf53a, Inf53b, IP56, IP57, Ish90, II08, IU86, IV85, IV87, IK91, Jac25, JM99, JT10, Jin91, JP73, Jor59a, Jor59b, Jor62b, Jor62a, Jor66, Jor71, Jun97, KY01, KMW86, KM92, Kam86, KM84, KLMW88, Kap40, Kat00, KRR90, KR92, Kem87, Ker90, Ker83, Kil76, Kim80a, Kim80b, Kis98, Kla90, KK10, Köp06, Kos86, KL90, KL92, Koz90, Kra81b, Kra82, Kra90]. Dirac [Kra91b, Kra91a, Kra06a, Kra13, Kra16, Kri87, Kru94, KY93, Kur87b, Kur87a, KW87, L.88, K.L88, LW34, Lam69, Lam70, Lan30, Lan05a, Lan05b, LR79, LRK80, Lan01, LW78, Lan87b, LU31, LR59, Lee95, LH86, Lég95, Leh91, LJ31, Leo80, LCH10, LW75, Lin10, LC98, Löw64, Lub82, Lut07, MD88, Mac98, Mal75c, MHN98, MG87, Mar65, MB90, MB91, Mar88, MS95, yRSBSV99, MSB92, Mat87, MR10, McA90, McC04, MS38, MS87, Med89, Med90, Meh72, Meh87a, Meh87b, Meh87c, MR51, ML93, MRB75, Mil98, Mil97, Mit10, MN97, MKK05, Moh07, Mon95, MB11a, Mor85, Mot87, Moy81a, Moy81b, Moy81c, MS68, Mur71, Nak97, Nak00, NM84, Nas84, Nas86, Nas97, NK93, NM01, NP96, NdlC08, Neu00, Neu33, Nil82, Nil03b, Nis32]. Dirac [Nis35a, Nis35b, Nov74, Oak00, Oev88, OS81, Ogu96, Olv86, OP95, Paĭ90, Pai87, Pai98, Pai00b, PT86, Pas12, PS32a, PS32b, Pau43, PD62, Pei34, Pei35, Pei42, Pei72, Pei85, Pei87a, Pei87b, Pen99, PP71, Pic89, Plo15, Pod97, Pol58, Pol87b, Pol87c, Pol87a, Pon05, Por95, Pos84, PSA94, QGW01, Rad96, RU81, Ran87, RSK99, RA11, RK96, RM95, RM96, Reu99, ROB64, Rho50, Ric03, Rij66, Riv94, Rob66, RRS06, Rod80, RdO07, RdO16, RR93, RL68, Ros83, Rou84, Roz20, Rüd84, RTV99, Rug88, Rum86, RS87, RZ80, Sag91, Sal87a, Sal87b, SM91, SFVW95, SS03, San79, Sch69, Sch38, Sch31a, Sch31b, Sch41, Sch52, Sch94c, Sch94b, Sei36, Sei99, Sel94, SCSB93, Sem93, SBZ90, Sha87, SMI83, SV89b, SV89a]. Dirac [SC91, SC92b, SC92a, SV93, SS98, SS11, SG11, Sim02, Sky71, Smo93, Smo87, Sno73, SM35, Som36, Son99, SM03, Sri83, Sri89, Sta12, Sta67, Sta87, SES81, SES82, SOS84, Str71, SWO76, Szm98, Szm01, Tak79, Tal04, Tam03, TVvN34, Tau49, Tay87b, Tel73, Tem35, Ten68, Tha88, Tha92, Tha06, Tjo75, Tol98, TM85, TCC97, Tra92, TKS19, UL31, UK82, UB83, Van72, VIK98, Vil90a, VP90, Vil90b, Vil95, VC02, Vrb94, Vul03, Wal09, Wal76, WP96, WBD70, Web87, WBSB14, Wie29, Wig72a, Wig01, Wig88, Wig87b, Wig96b, Wig96c, Wil03, Wil98, Wil91, WY84, WY85, Won90, Wot91, Wri16, Wüt11a, zXZ06, Yan78b, YYH^{+}01, YG77, YI90, YZ08, YL09, Zas89, Zec96, ZQ90, Zic00, Zor80, dB32b, dB32a, dB33, dB34]. Dirac [dB51, dB52b, dB52a, dBDW^{+}84, dG80, de 89, de 91, vEdG85, vSJLEdJG86, vN28, vN40, Dar91, Ram89, Sch88, Sti13, Stu88, Bro91, Dar92]. Dirac- [Gar84]. Dirac-félé [Neu33]. Dirac-Gleichung [EM33b, SM35]. Dirac-Like [MRB75, WP96]. Dirac-spinor [Eeg80]. Dirac-Theorie [Som36]. Dirac-Type [RRS06, MD88]. Dirac. [Dre90]. Dirac/Gamow [Kra91a]. Diracgleichung [Sch38]. Diracs [CKPT04]. diracsche [Pei35]. Diracschen [EM33a, Hei34b, Hei34a, HE35, Jor59a, Jor59b, Jor66, Neu33, vN28]. DiracSolver [TKS19]. Dirak [Koz90, Med89, Med90]. Diraka [Bel82]. directed [RR93]. Directions [DHS78]. Disciplines [Len14]. Discontinuities [Szm01]. Discover [Ser12]. discoveries [Bra09]. Discovery [CR12, BBC^{+}94, Fal97, GC97, Got11, Kae48, Kox97, MR82a, Rob97]. Discrete [Dir72a, All03, BK10]. discretizations [CG07]. Discussion [Dir28a, Dir34a, HND^{+}62, SBB^{+}62, vN40]. discussions [CCJ^{+}34]. disk [FSS98]. dispensability [Bue05]. Dispersion [Sei36, Dir27c]. Dispose [Ano34a]. Dissipative [All04]. Dissociation [Dir24a]. Distant [MWF07, SCPA04]. distinctions [GG11]. Distributed [QGW01]. distribution [CK90, CK91, Dir34a, Fuk14]. distributions [Bet83, Cra88, RS87, Zor80]. divergent [Bel91]. DNB [Ano24d]. Do [Wil58, Alp73, Ros83, zXZ06]. documents [PKKL09]. Does [BS67, Dir36a, Dir75b, Dir79e, Dir81a, Gam67a, Kap40, Wes80, MB11b]. d’onde [dB52b]. d’onde-pilote [dB52b]. d’ondes [BI34]. Doppler [Dir24b]. double [dB52b]. double-solution [dB52b]. d’ouvrages [Dar91, Dar92]. Dr. [Ano30, Dir40]. dreams [Haw11]. Drehelektrons [vN28]. driven [AG03]. drogi [Inf33]. d’une [Dav98]. during [DHS78]. dynamic [DSK72]. dynamical [Dir48c, Dir49a, PSA94]. Dynamics [Dir24c, Dir27b, Dir33a, Dir49b, Dir50b, Dir51a, Dir55e, Dir58b, Dir77b, Son99, Tay87a]. dyon [ZQ90]. Dyson [Sch94a].

E. [Bel82]. Early [Dir82c, LMSV08, Dir73g, Meh87a, Pai72]. earth [FG22, Jor66, Jor71, Kra15a, Kra15b]. École [Bou01]. ed [Hen89, War96]. Eddington [Bek86]. Edge [YYH^{+}01]. edition [Jär64]. editors [Tor87]. eds [Ble86, Hen89, Sch85]. Effect [FK08, Bat00, Bat07, Dat04, Dir25d, Dir27a, Kox97]. effects [DFW99, Gar84]. efficiency [Dir42a]. Ehrenhaft [Dir77c]. eigenfunction [Szm01]. eigenfunctions [Cla94]. Eigenparameter [BK10]. eigenvalue [All04]. Eigenvalues [Cla94, AT07, YI90]. eightieth [BvdM83, KMPD85]. einatomigen [Fer26]. einem [Dir60b]. Einführung [Deh36, Haa34]. Einheit [Sch38]. Einige [Jär64, vN28]. Einstein [Ano05, Bal85, Bri18, Buc80, HD77, HE82, SH99, BOS96, BBR03, Bok08, Buc80, Cra88, DMH85, Din58, Dir79a, Dir79d, Dir79c, Dir79b, Dir79f, Dir79g, Dir80e, Dir80d, Dir80c, Dir80h, Dir81b, Dir79m, Faz15, Gau93, GMW80, Gru00, HW96, Jin91, Kap40, Kle05, Kos86, KBN83, MR82d, MR82e, PR86, PS32a, PS32b, PS79, RA11, RM95, Rob15, RdO07, RdO16, RTV99, Tol98, Vil90a, Wie29, Wil93b, Wil91, Zic00]. elastic [Gar84]. Elected [DP86]. Electric [Inf53a, Per67, BM70, DH32a]. Electrically [SV93]. Electricity [Gam67c, Gam67b, Bun52a, Sto94]. electrodynamic [Jär64]. Electrodynamics [Ano75, HDW75, Sch52, Sch94b, Sch58, Sch03, BI34, Bro77, DFP32, Dir35b, Dir43b, Dir46, Dir47a, Dir50c, Dir52c, Dir55a, Dir57b, Dir60d, Dir65a, Dir68b, Inf53b]. électrodynamique [BI34, Dir52c]. Electromagnetic [MWF07, SCPA04, BN72, Dir31b, Dir39a, ZQ90, dB32b]. électromagnétique [Dir39a, dB32b]. Electron [Ano30, Ano34a, Csa92, Dir77h, Dir79h, Dir79i, Dir79j, HL38, Neu33, Roh60, SBK^{+}08, vN28, Bat07, BN72, Bor52, CH02, CKYY07, CDK90, Dir28f, Dir28b, Dir28c, Dir29b, Dir30d, Dir35a, Dir39a, Dir59a, Dir62c, Dir77g, Dir77i, Dir79k, DFW99, Fal97, GC97, Gol05, Hin77, Kim80a, Kim80b, Kox97, Lan30, Lan05a, Maj37, Mil97, Moy81a, Moy81b, Moy81c, Pai72, Pei34, Pen97, Rec97, Rob97, Som34, Sto94, dB32b, dB33, dB34, dB52b, Ano34a, Dir39a, dB32b, dB33, dB52b, Gar84]. electrona [Dir59f]. Electronic [Inf53a]. Electrons [Dir30i, Dir34d, Dir58a, DuM28, Dir25c, Dir25d, Dir28d, Dir28e, Dir30e, Dir30j, Dir34a, Dir38a, Dir48b, Dir51c, Dir52d, Dir54c, Dir59b, Dir59f, Dir65b, Dir91b, Gam33, KD33, Pau25, Seb19, dB52a, Kra81b, dB52a, dB51, dB51]. Electrophoresis [CKPT04]. electroweak [KM04]. elektron [Neu33]. Elektronengruppen [Pau25]. Elektroni [Dir57a]. Elektrons [Lan30, Dir28f, Dir28d, Neu33, Dir28e]. Elektrony [Dir59b]. element [Rig02, RS87]. Elementary [And88, BH83a, Dir86, FW87, Gam67d, BH82, Fer86, Kat00, KP83, PB68]. elementary-particle [BH82]. elementi [Fer24]. Elements [Dir32a, Fer24, Hos87, WY84, ZQ90]. elettrone [Maj37]. Elimination [Dir26a]. elliptic [ABR06]. elméletében [Neu33]. embedded [EG84]. EMC2006 [RCC^{+}06]. Emeritus [Ano13]. Emission [Dir27d]. Empirical [Jor62b]. Empiricism [Kra91b, Kra91a]. enclosed [Dir42b]. encode [Bar02]. end [Gre05]. Energies [EKI94]. Energy [Ano30, BBD^{+}46, Csa92, Dir59c, Dir71f, Dir72c, Dir73h, GLR14, IPM73, MWF07, Roh60, SCPA04, Ano08, BtGW95, Csa81, DPI71, DMH85, Dir36a, Dir42a, Dir62b, Dir70a, Dir71d, Dir71e, Kra17, KPW75, KPS^{+}78, NP96, Per80, SWO76, Wig72b, KMPD85]. energy-momentum [DMH85]. Enhanced [BBSF10, FK07]. enquiry [Alp73]. Enrico [Mil07]. Entgegennahme [Wil36, HSD34a]. entre [Bal85, dB32b]. episodes [Bra09, Smo87]. epistemology [Wri16]. Epoch [Dir90c, Dic59, Dir88a, Dir88b, Dir90d]. Epokhe [Dir90c]. Equation [Aug95, BCJT16, Dir71f, Dir72c, Dir77h, Dir79h, Dir79i, Dir79j, Eva04, GB13, GG51, GH94, Inf34a, KK10, Löw64, MRB75, Plo15, QGW01, RK96, Sta67, Tiw12, Wüt11a, vN40, AD87, AME14, AKV83, AW17, Ata89, Ati98, BO92, BK76, Beh85, BCK97, BM70, BtGW95, BI34, BC69, BJ03, CG07, sC78, CGKS94, Coh70, CY91, Cri76, Dat04, Dav98, Dir35a, Dir59a, Dir71e, Dir77i, Dir79k, DSP03, EF77, EGT21, EV97, Esp12, FA01, FM75, FSY00, FZ91, GLMR94, GL92, GF02, GS84, Ham84, Hil92, Hos82, Hos83, Hos87, Hos89, HM99b, HM99a, IJ02, IS10a, IS10b, IV85, IV87, IK91, Jac25, Jin91, Jun97, KMW86, KM92, KRR90, KR92, Kis98, Köp06, KL90, KL92, Lan05b, Lin10, MB90, MB91, Mar88, MS95, yRSBSV99]. equation [MR51, Mil97, Nak00, RSK99, Riv94, RL68, Rüd84, San79, Sch38, Sel94, SCSB93, Sem93, SMI83, SV89b, SV89a, SC91, SC92b, SC92a, Sky71, SM35, Son99, Sri89, SES81, TVvN34, Tha92, Tha06, TCC97, TKS19, UK82, Vil90a, Vil90b, Vil95, Vrb94, Vul03, Wig72a, Wil03, WY84, Won90, Zas89, Zec96, BI34, Dav98]. Equations [Bri18, Dir25e, Dir67a, EM33b, Far03, Ker83, LU31, LR59, Sch31a, Sch31b, Sch41, Tau49, UL31, Wie29, BOS96, BK10, BC78, BR97, Caw69, Cha72, CV90, DK89, Das93, Das96, DMH85, Dir36b, Dir36c, Dir69b, Dir70a, Dir71d, Dir77d, EM33a, Fre95, GS79, HD77, Hau72, Jin91, KM84, KBN83, LH86, Lég95, LC98, MD88, Mal75a, Mal75b, Mal75c, MR10, MR82c, Nak97, Pai87, PT86, PS32a, PS32b, PP71, Por95, Rad96, RA11, RdO07, RdO16, SM91, SFVW95, SBZ90, SG11, SES82, SOS84, Szm98, Szm99, Tak79, Tu91, VP90, VC02, Wil91, de 89, PS32a, PS32b]. equilibrium [BSW98, Col12, Dir25c]. Equivalence [Dir64a, Bro77, JW28, Nas86]. equivalent [AW17]. Era [Dir77f, Dir87e]. Erde [Jor66]. Erforschung [Fis10a, Fis12a]. Ergodic [Tau61b]. Erice [Zic88]. Erratum [Gam67c, GT85a, Mal75a, Szm99]. error [ABR06, CG07]. Erwin [Dir61b, MR87]. Essays [Ble86, Sch85, Tor87, BVV84, SH99, dBDW^{+}84]. Essential [KY01, Ogu96, LR79, LRK80, Rig02]. established [Ano85]. Estimate [Dir42a]. estimates [ABR06, BtGW95, CG07]. était [Ano05, Kle05]. etc [Jor59a]. etc. [Jor59a]. éter [Dir54a]. Ettore [RCC^{+}06, ACM10, Esp08, ES10, EAW15, MRB75]. Euclidean [FA01, Lég95, Leh91, Ric03, dG80]. Eugene [Ble86, Hen89, L.88, Ram89, Sch85, Sch88, Stu88, Tor87, BVV84, Wig96d, dBDW^{+}84]. Eulogy [Lan87a]. Europe [Rod19]. European [Dor82]. Evaluation [McA90, WBD70, BBR03, GZDA01, MN97, MKK05, NK93]. Evaluations [Moy81a]. Evans [Eva04]. even [SMI83]. Evening [BDH70]. Evidence [CT83, Noe73]. evidential [Fra04]. Evolution [Dir63a, Dir71b]. Evolutionary [Dir72b]. Exact [CGKS94, FA01, HM99b, HM99a, IV87, Jin91, Por95, ROB64, Rou84, RTV99, Sta67, Vil90a, Vil95, WP96, Won90, BC78, DK89, FM75, KBN83, PR86, SV89a, SV93, VP90, Vil90b, VC02, Wil91]. examination [BBSF10]. example [Coq85]. examples [AT94]. Excellence [Dir80e, Dir79f]. exceptional [Nas86]. Exchange [Csa92, Dir30c]. Exchange-Energy [Csa92]. Exciting [Dir77f, Dir87e]. Exile [Ano33]. exist [Ros83]. existence [FSY00]. Expanding [Jor71, Kra15a, Kra15b, Sch41]. Expansion [BG91, Dir42b, GPS86, Jor66, Rij66, Csa81, Fin00, IM95]. Expansions [Din58, BDM81a, GFG01, Szm01]. expectation [YS03]. Experiences [Ano55, L.88]. Experiment [Pei36, Zic84]. Experimental [Ano34a, Dic64, DuM28, MBS02]. experimenter [Kri87]. experiments [KPW75]. Explicit [Cor70]. exponential [JT10]. exponentials [CPV10]. Expression [Lam69, Lam70]. Expressions [BBR85, Nov74, Wal76, Car92, Cor70]. Extend [Ano34a]. extended [BG90, Dir64e, IK91, MHN98, Pru73]. extensible [Dir62c]. Extension [CV90, AT94]. Extensions [All03]. external [AD87, Dar90, HM99a, Rou84, SV89b, SV89a, SC91, SC92b, SC92a, SV93, ZQ90]. extra [Car01, Kha86, KPW86]. Extraordinary [Dir90c, Gib19, Dir88a, Dir88b, Dir90d]. extrapolation [CY91].

F.R.S [Dir37c]. F.R.S. [DP86]. Faces [Bri18, RdO07, RdO16]. factorization [Cra85]. Factors [Din58]. Fail [GLR14]. faite [Dir48e]. Fantastic [WD06]. Farmelo [Cas10, Dys10, RS10, Sti13]. Fast [IKRT89a, IKRT91]. Faster [zXZ06]. Father [Dir03a]. FD [FR86]. FDG [FR86]. FDH [FR86]. fears [Dir69c]. feet [Pol87a]. fejlödésük [Dir78a]. Feldgleichungen [EM33a]. feldtheoretische [Lan30]. félé [Neu33]. Fermi [CT67a, Csa81, Csa92, DuM28, GG51, MHN98, Rij66, Tu91, YG77, BDM81a, BDM81b, BDM84, Bel82, Bel94, BBR03, BG91, CT67b, Cra88, Din58, Doa91, Faz15, Fer86, FA10, Fuk14, FR86, GFG01, Gau93, Goa95, GZDA01, Kap40, Kos86, Löw64, Mac98, MS38, MR51, Mil07, MN97, MKK05, Moh07, NK93, NM01, Pei72, Pic89, Rho50, Sag91, SS11, Tu91, WBD70]. fermion [Fei02, ZQ90]. Fermions [FK08, Bel85, Col12, MZF^{+}12]. Feynman [And88, Bus10, CY91, Bro05, DR93, FBD05, II08, Jac25, KY93, Sch94a, Wüt11a, Wüt11b]. fidelity [AME14]. Field [Dir50a, Dir55d, Dir83b, Dun12, Har28, Inf53a, Ish90, KK10, KP83, Pau43, Sch69, Sch91a, Suc67, AD87, BN72, Bor85, CDK90, Dar90, Dir31b, Dir39a, Dir51a, Dir55b, Dir55e, Dir59c, Dir62a, Dir62b, Dir62e, Dir64d, Dir64e, Dir66b, Dir68d, Dir87d, Ehb22, Ehb23, EM33a, Gre77, HD77, Inf61, IU86, Jos72, Kur87a, LMV91, Lan30, Lan05a, Mal75a, Mal75b, MS87, Oev88, Ogu96, Olv86, RU81, Ran87, Rou84, Sal87a, Sal87b, Tam03, TCC97, YI90, ZQ90, dG80]. field-theoretical [Lan30]. Fields [BSW88, Fab22, SBB^{+}62, AW95, AFG75, BCK97, Can98, Car92, Car94, CH02, Dir25a, Dir49a, Dir62d, Hau72, HM99a, Ker90, LW78, Leh91, PR86, Reu99, SM91, SV89b, SV89a, SC91, SC92b, SC92a, SV93, Vil95, Yan78b]. fifth [Dor82]. Fifty [Kae48, RZ80]. Fig [Wil36]. find [Lub82]. Finding [CMO00, Car01]. Findings [Ano30]. Fine [BS67, Bek86, FK07, MWF07, SCPA04, WFC^{+}99, AG03, BBSF10, Car01, CKYY07, Han99, MG87, Sta12]. Fine-Structure [BS67, Bek86, FK07, MWF07, SCPA04, BBSF10, Car01]. ‘Fine-Tuning [Sta12]. Finite [QGW01, Dir62e, Sal87a, Sal87b, SWO76]. First [Ano33, Dat95, Ell91, GMW80, SDEW90, Stu18]. First-Row [Dat95].

Fisica [RCC^{+}06, Wei77]. Five [Sch31b]. Fixation [Dir59d]. fixed [BB85]. Fizika [Med90, Koz90]. fizikai [Dir78a]. fizike [BDH70]. fiziki [Dir83d]. Florida [BB03, CMO00, Hal78, KPS^{+}78, Per80, CH16, KP83, Lan87b, Oak00]. flow [Bun52a]. fluid [IV85]. Flux [DP03, Lub82]. Flux-across-surfaces [DP03]. Fock [Ish90, Dat95, DN96, VIK98]. Fock-CI [DN96]. Fois [Ano05, Kle05]. Foldy [Hos89]. Folgerungen [Jor66, HE35]. Following [UL31, Dir28a, Jär64]. fonction [Sch72]. Forbidden [GPS86]. Forces [Ell91, Dir73c, Dir73d, SDEW90, SHD91, Zic88]. Form [BBR85, Dir51a, Dir58d, Olv86]. Formalism [GG02, Rob66, Suc67, Ant69a, Ant69b, Kam86, Pru73, Wot91, vEdG85, vSJLEdJG86]. Formation [Jor49, MS87]. former [Phi87]. Forms [Dir49b, LC98, Tha88]. formula [Doa91, Mor85]. formulas [Wig01]. Formulation [Fab22, GG02, Mar65, MR82b, Dir55a, GH75, RM96, Szm98, Szm99, Tjo75]. fortalt [BR64, RRK^{+}64]. Fossil [Mau76]. Foundation [Jos72, Mat87, MR82d, MR82e]. Foundations [Dir64b, Dir66a, GL23, Dir78d, Med89]. Four [Ano75, HDW75, DW09, IU86, Lég95, Zas89, Sch31a]. four-dimensional [IU86, Lég95]. Fourier [CK90, CK91, Lig87]. Fourth [Jär64]. fractionating [Dir42c]. Fragments [Kra19]. frames [UK82]. Framework [Dun12]. France [Ano33]. free [BT90, Dir25d]. French [Bal85, BI34, Bor49, Bou01, CCJ^{+}34, CT12, Dav98, Dir31a, Dir34c, Dir39a, Dir49c, Dir52c, Kle05, PS32a, PS32b, Sch72, dB32b, dB32a, dB33, dB34, dB51, dB52b, dB52a]. frequencies [Rob85]. frequency [Dir24b]. Friedmann [Sri83]. friends [BR64, RRK^{+}64, Roz67]. frontier [BBC^{+}94, Pet85, Pet88]. frontiers [KPS^{+}78]. FRS [Dal87b]. fuerzas [SHD91]. Function [AT99, JP73, Lam69, Lam70, Rij66, BMP14, Car92, Deb13, Hos83, Hos87, Hyl84, IM95, JT10, Lut07, Rod80, Sch72, SS11, WY85]. Functional [Nil03b, All03, Csa81, SS11]. Functionals [Csa92]. Functions [Din58, Mac98, MS38, Nov74, ROB64, Rho50, Tau61b, Wal76, AW95, BDM81a, BDM81b, BDM84, Bel91, Dir55c, Fer86, GZDA01, GGS04, IP56, IP57, Kos86, Lig87, LW75, Mal75c, RS87, SWO76, ZQ90]. Fundamental [Dir25e, Dir67a, Dir71c, Dir73b, Dir77d, Edd46, Ell91, ES10, IPM73, Kur73, Mar84, Per67, SV95, WBR76, Alp73, CGK72, DPI71, DSW08, Dir84c, Dir84e, DOV02, Dys72, DFW99, IC04, Kha86, KPW86, LMSV08, Lee95, MR82c, Nil03a, Nor86, Pet85, Pet88, SDEW90, SHD91, SV90, Uza03, Dir78a, Str71]. fundamentales [SHD91]. Furry [Ano34a]. Further [Sta67, IP57]. fusion [Rei72a, Rei72b]. Future [Dir84b, Peg77, Pet11].

G [Dar91, L.88, Stu88]. Gables [DPI71, IPM73, Kur73, KPS^{+}78, KP83, Per80, Kur87b]. galaxies [Hoy87]. Galileo [Buc80]. Gallery [Pai00a]. Games [Tau63c]. Gamma [Din58, IKRT89a, Lee95, Pod97]. Gamma-Matrices [IKRT89a]. Gamow [Alp73, GG76, Har01, Kra91b, Kra91a, Rei72a, Rei72b]. gaps [NP96]. Gas [DuM28, Fer26]. Gases [Fer26]. Gauge [BSW88, Dir55a, GB90, Zic84, ZDF^{+}84, Ara93, BV85, Dir50c, FG87, Kam86, Pon05, RM96, Yan78b, Dir52c]. Gauge-invariant [Dir55a]. Gaussian [QGW01, Sag91]. geb [Deh36]. gegenwärtigen [Jor59b]. geh [Ben34]. gehaltenen [HSD34a, Wil36]. Gell [MHN98]. Gell-Mann [MHN98]. Genauigkeit [Neu33]. General [Ano62, Das93, Dor82, GL92, Hov96, Inf34a, Kil76, NM84, Ruf82, Sch31a, Sch31b, CV90, Fuk14, LC98, SMI83, UB83, Ant69a, Dir43a, Dir75g, Dir96, Tra92, Wil93b]. générales [Dir53c]. generalised [Lig87, NK93]. Generalization [Yan78b, BCK97, UK82]. Generalized [AT99, Dir50b, Dir58b, FR86, GZDA01, Lan01, BBR03, DW09, Fei02, Gau93, NM01, Pic89, Riv94, Sag91]. generates [Nas97]. Generation [Hür22]. generators [Nas84]. Genesis [Wüt11b, Kra81b]. Genius [Bri95a, Dys10, Pai00a, Far16]. geodesic [KM84]. geology [FG22]. Geometric [KR92, Ham84, Kru94, Ran87]. Geometrical [Ano00, Jun97, Dir74e]. Geometry [Tau62, Ati98]. Geophysical [Jor62a, SV90]. Geophysics [Jor59a, Kra15a, Kra16, Goe16]. Geophysik [Jor59a]. George [Alp73, Dys93, GG76, Har01, Rei72a, Rei72b]. Georges [Dir68e, Dir72e]. German [Bos24, Dir27e, Dir28f, Dir28d, Dir28e, Dir30b, Dir53a, Dir60b, Dir60a, Dir77a, Dir80b, Dir84e, Dir87b, EM33b, EM33a, Far16, Fer26, Fis10a, Fis12a, Haa34, Hei34b, Hei34a, HSD34a, HE35, Jär64, JW28, Jor37, Jor52, Jor55, Jor59a, Jor59b, Jor66, Lan30, Neu33, Pau25, Pei35, Sch38, Sei36, Smo93, Som34, SM35, Som36, Tak31, vN28, vW02]. Germany [Wat91]. Gesetz [Bos24]. Get [Ano33]. Gets [Ano33, Ano34b]. Gleichung [EM33b, SM35]. Gleichungen [EM33a]. Global [BF12, KM12, KK10, FSS98]. God [MB11b]. Goes [Ano33]. Gold [Dir52a]. golden [Meh72, Meh87b]. Gordon [GS79, CV02, SBZ90, VC02, de 89]. Goudsmit [Lak96]. GPS [Ano12]. gradient [Csa81, Dir24a, PSA94]. gradient-holonomic [PSA94]. gradients [MHN98]. Graduate [Suc67, Bus10]. Graham [Dys10, Cas10, RS10, Sti13]. grand [DNRV03]. Grant [Ano33]. graphs [BT90]. Graßmann [Pet11, Hor11, PKKL09]. Grassmannian [Rum86]. Gravitation [BD61, Dic59, Dir78g, Dir79d, Dir05b, Jor71, ZdSS91, Dir58d, Dir59d, Dir79f, Dir79m, Jor62b, LT62, SS98, Dir79g, Dir80e]. Gravitational [Dir72f, Dir75b, Dir79e, Dir83c, JRF06, Mal75b, Rob83, Sch91a, SBB^{+}62, VKA21, Wil93a, WW86, Cho81, Dir59c, Dir62b, Dir62d, Dir62e, Dir64e, Dir68d, Dir78c, Hal78, Inf61, Jor66, Lev80, Ros83, Sch99, Web87, Dir60b, Dir60a, Dir87b, Mal75a]. Gravitationshypothese [Jor66]. Gravitationswellen [Dir60b, Dir60a, Dir87b]. Gravitatsiya [Dir05b]. Gravitochemical [JRF06]. Gravity [BSW88, Gam67a, Gam67c, Gam67b, Goe16, Kra15b, Wes80, BP95, Bel85, Can98, Kra15a, Kra16, SS72, Jor52, Jor55]. gravomagnetic [LBNZ98]. Great [Far03, Hen89, L.88, Ram89, Sch88, Stu88, Wüt11a, Kae48, KW87, Pic08, vW02]. Greatest [Gun15, Hür22, Gab13, Wal09]. Green [Dys93, Hos83, Hos87, Hyl84, IM95, WY85]. Grenzen [Neu33]. Grosse [vW02]. Grossmann [Ruf82]. Group [DSP03, Far01, Tau61b, Wig39, WY84, Dir45c, Dir63b, Dir72a, KM84, Lég95, Mal75c, NdlC08, RL68, RRK^{+}64, Ten68, Wig89]. Groups [Wey31b, Wey50, Pau25]. Grund [Deh36, Haa34]. Grundlagen [Jor52, Jor55]. Gruppentheorie [Wey31a]. GTO [Dat95]. Guide [Rod19]. Guy [Lan87a].

H [Dar96, Dir61a, Hov96, War96]. H. [Dir52a, Dir64a]. Haas [Deh36]. Half [Ano18]. Half-light [Ano18]. Half-matter [Ano18]. halos [Hoy87]. Hamilton [AW17, Caw69, PP71]. Hamiltonian [CY91, Dir50b, Dir51a, Dir58b, Dir58d, Dir59d, Dir64c, DEŠ92, EKI94, JM99, Oev88, Ogu96, Olv86, Rou84, Sei99, TM85, YZ08]. Hamiltonians [NP96]. hans [BR64]. harmonics [Yan78a]. Hartree [Dat95, DN96]. harvest [Bra09]. határai [Neu33]. Hawking [KN03, Pic08]. HD [IRV^{+}08]. Heating [JRF06]. Heisenberg [Ano51, Ben34, Deh36, Haa34, LW34, Lak96, Wil36, Ano33, Ano34a, Ano75, Bok04, BR87, Dir28a, Dir64a, Dir76b, Haa34, HDW75, Nis32, Smo93, vW02]. Held [Meh73, CCJ^{+}34, Hal78, Jef87, KPS^{+}78, KP83, Ruf82, Sal69, Zic88]. Helge [Bro91, Dar92, Dre90, Sch91b, Goe16, Hen91]. hemisphere [DFPP43, PFD43]. hemispheroid [DFPP43, PFD43]. heritage [Esp08]. Hermann [PKKL09]. Hermite [Doa91]. Hess [Ano51]. Hestenes [Ham84]. Hidden [Cas10, Dys10, Far09b, RS10, Sti13, Bla11, Far16, Lee95]. hidrógeno [Bun52b]. Hierarchy [CV02, DW09, YZ08, YL09]. Higgs [AFG75, Tol98, YS03]. High [AME14, BtGW95, Fer86, IKRT89b, IPM73, Rij66, DPI71, Dir70a, Dir71d, KPW75, KPS^{+}78, NP96, Per80, KMPD85]. High- [BtGW95]. high-energy [Dir70a, Dir71d, KPW75, KPS^{+}78, NP96, Per80, KMPD85]. High-fidelity [AME14]. High-precision [Fer86]. higher [Kru94]. higher-order [Kru94]. Hilbert [Tol98, Dir70c, Dir74g, Dir91a, Pru73]. Hintertreppe [Fis10a, Fis12a]. Hirshfeld [Tiw12]. Hirzel [Ben34, Wil36]. Historical [GL23, Deb13, Dir75c, HE82, Kra81a, Kra06b]. Histories [Pei97]. History [Dir87c, KN03, Kra82, Kra19, UM86, Wei77, BN13, Fra04, Kra15a, KHFA67, Nor65, Nor90, Oak00, Pai72]. Hold [Ano18, Dir36a]. Holders [Ano24a, Ano24b]. hole [FSY00, Pei35]. Holes [Mon95]. holonomic [PSA94]. Homogeneous [Dir33a, Inf53a]. honor [KMPD85, LW78, SH99]. Honoring [Dys74]. Honour [Ble86, Sch85, Tor87, BVV84, Sno73, dBDW^{+}84]. Hooft [CK78]. Hopes [Dir69c]. host [Hal78]. Hulthén [IS10a]. Human [Bus10, II08, Lip87]. hundred [GMW80, Rob15]. Hungarian [Dir78a, Neu33]. hybrid [MZF^{+}12]. Hydrodynamics [Tau63c]. Hydrogen [Dir26e, Dir67c, Lak96, Rig02, Bel91, Bun52b, Dat04, DH32a, MR10]. hydrogen-like [DH32a]. Hydrogenic [Tiw12]. hyperbolic [CV90, Köp06]. Hypergraphs [RRS06]. hypersurfaces [Tra92]. Hypothese [Jor59a]. Hypothesen [Jor59b]. hypotheses [Jor59b]. Hypothesis [Bos24, Jor62a, Jor71, Tel73, Dir25a, Jor62b, Jor66, Dir74b, Dir75d, Dir78c, Dir79g, Jor59a].

Ia [FA10]. Idea [Dir79b, Car01]. ideal [Fer26]. idealen [Fer26]. Ideas [Dir78g, Dir69a, Dir73e, MS87, dB52b]. idées [dB52b]. identical [Fer24, Mil07]. identici [Fer24]. Identities [Nas84, Kat00]. idöbeni [Dir78a]. ignited [Gre05]. ihre [Lan30, Som36]. II [Wig96d, Ant69b, Dir28c, Dir52d, Dir72c, Dir88a, Dir03b, Hos83, Kim80b, Kox97, Lam70, LRK80, MB90, Nis35b, PS32b, PFD43, PSA94, Tau61b]. III [BMP14, Din58, Dir54c, Dir88a, Dir04, MB91, Tau63a]. Illuminating [Bat07, HW06]. illus [Dre90]. im [Jär64, Pau25, Pei35]. imaginäre [Sch38]. imaginary [Sch38]. implementability [FG87]. implementation [IKRT91]. implications [Mil97]. Important [Ano30]. Improved [SBK^{+}08]. impulses [Nil03b]. IMT [MKK05]. inadequacies [Dir87d]. inception [Ehb22, Ehb23]. including [Ano65, IS10a, Sno73]. Inclusion [RM96]. inclusive [Ano10]. Incomplete [Goa95, Pon05]. increased [Car01]. Indefinite [Hei72]. independent [Dic59]. Infeld [Dir52b, Dir60d]. Infinite [NdlC08, Dir34a]. Infinite-dimensional [NdlC08]. Infinitesimal [Ker83]. Influence [Dir76b, Dir80d, Kur87a]. informal [CH16]. Information [Plo15, DSK72, RS10]. Inhomogeneous [Wig39, Wig89]. initial [CG07, Ehb22, Ehb23]. Innocence [Stu18]. Inquiry [Gib19]. institut [CCJ^{+}34]. Institute [Ano34b, CCJ^{+}34, F^{+}00]. Instruments [L.88]. Insulator [FK08]. integrability [SOS84]. integrable [PSA94, YZ08, YL09]. Integral [Goa95, JP73, MD88, Rho50, Fuk14, IT84, Jac25, KY93, Lin10, MN97, MKK05, Moh07, NM01, OS81, Rum86, SS72]. Integrals [BBR85, BG91, CT67b, Din58, ROB64, BBR03, CT67a, Dir25b, Doa91, FH65, FHS10, FR86, GFG01, Gau93, NK93, Pic89, Sag91, SWO76, Tal04, WBD70]. Integration [BOS96, Csa92, BO92, Sei99]. Integrator [BCJT16]. Interacting [Dir62d, Bel85, Rou84]. Interaction [KK10, BSW98, DEŠ89, DEŠ92]. Interactions [Kur73, BK76, BJ03, CV90, DPI71, Dir86, IPM73, LC98, Web87, Zic84]. interdisciplinary [F^{+}00]. Interest [L.88, K.L88]. interfaces [dG97]. internal [MB90]. International [BCEP94, Meh73, Ruf82, Sal69, Wat91, Zic88, Ano10, Car01, CCJ^{+}34, RCC^{+}06]. Interpretación [Bun52b]. Interpretation [Dir27b, Pei36, Dir30d, Dir42e, Dir68b, Lan30, Lan05a, Pol87c, Ran87, vEdG85, Bun52b]. intertwining [Wil98]. Interview [Ano94, BP79a]. intrinsèque [Bou01]. Intrinsic [RL68, Bou01]. intrinsic-particle [Bou01]. Introduction [Csa81, Haa34, dB52b, vSJLEdJG86]. Invariance [BF12, DPP97, KM12, Rob85, WB78, ZDF^{+}84, Cra93, Dir25b, DPP42, FA01, MSB92]. Invariant [Lég95, Dir55a, Hos82, Hos83]. inventory [KHFA67]. Inverse [Cor70, HJKS91, Jun97, Leo80]. inversion [Cra85]. Investigation [Dir26e, Dir67c]. investigations [CH16]. Invited [DPI71]. involutional [Kim80a, Kim80b]. Ions [VIK98]. ISBN [Ble86, Hen89, Hen91, Ryc17, Sch85, War96]. isometry [KM84]. isotopes [Dir41]. Israel [UM86]. ISSAC [Wat91]. Italian [Fer24, Maj37]. Italy [Dys74, Meh73, RCC^{+}06, Zic88]. iterative [Ata89]. IV [Dir05b, Far01, Tau62]. Ivancevic [Bus10, Bus10].

J [Dar91, Fal97, L.88, Sch85, Szm99, Fal97]. J. [Dir71a]. J.-P [Sch85]. Jacobi [AW17, PP71]. Jagdish [Dys74]. James [Ryc17, Dir75e]. Jan [KPS^{+}78, Per80]. January [DPI71, IPM73, Kur73, KP83]. jauge [Dir52c]. Jean [Ble86, Tor87]. Jean-Pierre [Ble86, Tor87]. jepohe [Dir90d]. Jerusalem [HE82]. John [Dir37c, F^{+}00, Jef87, Stu88, Réd05, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, vN96]. Johnson [DJ95]. joint [F^{+}00]. Jones [L.88]. Jordan [Kra15b, Sch99]. Journey [FR13]. journeys [WD06]. juin [LT62]. Jülich [F^{+}00]. July [Ruf82, Wat91]. June [Pei92, Sal69].

Kähler [PT86, Reu99]. Kahn [Kos86]. Kalinga [Aur71]. Kálnay [Hin77]. Kaluza [GT85a, GT85b, GT86, Mar84]. Kapitza [Bat00, Bat07, DRFS80]. Kargon [L.88]. kart [Wil36]. Kasner [Sri89]. Kelvin [L.88]. Kepler [Cor86, DJ95, IU86, YI90]. kernel [MD88, Rod80]. Kerr [EF77, KM92]. Ket [Rob66, D’A02, Pru73]. kets [Jau72]. Key [Ano18]. Killing [BCK97]. Kilmister [Hen89]. Kinetic [Ano30, Csa92]. Kinetic-Energy [Csa92]. Kinks [Sky71, Leo80]. Klein [CV02, GS79, GT85a, GT85b, GT86, Mar84, SBZ90, Som36, VC02, de 89]. Klein’schen [Som36]. kleinsten [Fis10a, Fis12a]. Knowledge [Ano34a]. Komplexstruktur [Pau25]. koordináta [Neu33]. koordináta-mérés [Neu33]. Koordinatenmessungs [Neu33]. Koordinatenmessungs-Genauigkeit [Neu33]. Kosmologie [Jor52, Jor55]. kosmologischen [Jor59b]. Kosmologiya [Dir05b]. Kostant [Pen99]. Kragh [Dre90, Goe16, Hen91, Bro91, Dar92, Sch91b]. kreds [BR64, RRK^{+}64]. Kursunoglu [Hen89, L.88, Ram89, Sch88, Stu88]. Kvantovaya [Dir02, Dir03b, Dir04]. kvantovoi [Dir60c, Dir90b]. kvantovoj [Dir68a, Dir99, Med89]. Kvantovoĭ [Dir32a]. kwanty [Inf33].

Laboratory [SBK^{+}08, Dic59]. Lagrangian [Dir33b, Dir05a, OP95]. Lagrangians [RM95]. Laguerre [MR10]. Lanczos [BCEP94, GH94]. language [Ehb22, Ehb23]. Large [Alp73, Ata89, Dir74b, Dir75d, Dir78c, Dir79g]. Large- [Ata89]. last [Alp73, Rec97]. later [Esp08]. Lattice [HJL00, MKS^{+}13, F^{+}00]. laureate [Ano65]. Law [Bor49, Bos24, Ano34a]. Laws [And88, Irv83, Bou85, D’A02, FW87, GS79, Pic08, Rob85]. Lax [PSA94]. lead [Rob85]. Leap [Bus10, Fis10a, Fis12a, II08]. Leben [Far16]. lecture [Dir60b, HSD34a, Pei92]. Lectures [Dir35b, Dir55b, Dir64d, Dir64f, Dir66b, Dir67f, Dir01, Dir02, Dir05b, IPM73, L.88, Nis32, Nis35a, Nis35b, Sch69, Suc67, Ano65, BDH70, DPI71, DHS78, FW87, Dir68a, Dir99, Ell91, SDEW90, dG97, And88]. led [Bel82]. Lee [WP96]. Legacy [Goe16, RCC^{+}06, Kra16, McC04]. Leipzig [Ben34, Deh36, Wil36]. Lekcii [Dir68a, BDH70]. Lektsii [Dir99, Dir02, Dir05b]. L’électron [Som34, dB34]. Lemaître [Dir68e, Dir72e]. lengths [ES10]. lensing [LBNZ98]. Leopold [Dir52b]. Let [Kra17]. Letter [Aur71, Jac25]. letters [Réd05]. Level [Bus10, Tiw12]. Levinson [Kla90]. Library [CMO00]. Lichtgeschwindigkeit [Pei35, Tak31]. Lichtquantenhypothese [Bos24]. Lie [DW09, FZ91, ML93, Pen99, UB83]. Life [Cas10, Dys10, Far09b, Gun15, RS10, Sti13, BDH^{+}89, Bla11, BR64, Dal87b, Dir83a, Far16, Gre05, Kem87, Pai98, RRK^{+}64, Roz67, BDH70]. Life-Changing [Gun15]. lifts [Oev88]. Light [Bos24, CR12, Fin00, Irv83, Lee23, Ano18, Dav98, KD33, Kra06b, Lee21, Lou83, Pei35, Tak31]. Light-cone [Fin00]. Light-Quantum [Bos24]. Like [MRB75, DH32a, IS10a, IS10b, LC98, WP96]. Limit [BCJT16, CP68, MWF07, Sch06, SCPA04, All03]. Limitations [Ano34a]. Limits [FA10, MWF07, Neu33, SCPA04, WBR76, Rob85]. line [Gre77, HJKS91]. linear [GS84]. Lines [MWF07, SBK^{+}08, SCPA04]. Lippmann [DJ95]. List [Meh87c]. Literaturberichte [Som34]. Literature [Som34]. Little [Dir80f]. liv [BR64, RRK^{+}64]. livre [Ano05]. Lobachevskian [Ker90]. Local [BF12, Hes73, KM12, SV95, Cra93, FGM^{+}96, IV85, IV87, IK91, Jin91, MB90, dG80]. localizable [Dir48c]. Localized [Rad96]. Löchertheorie [Pei35]. Lodge [Ano30]. Logic [Dir54b, Tau61a]. logical [Med89]. logicheskie [Med89]. London [Aur71]. Long [Dir73c, Dir73d]. Lorentz [sC78, Coh70, DPP42, Dir45a, Dir45c, Dir53b, DPP97, Eeg80, GLMR94, GPS86, Rob85, RL68, San79, Son99, Ten68, Wig39, Wig89]. Lorentz-transformations [Eeg80]. Lorentzian [IK91]. loss [SWO76]. Lost [Aug95]. Lösung [SM35]. Louis [Ble86, Bou01, Sch85, Tor87, BVV84, Dav98, dBDW^{+}84]. Love [Gun15]. Low [Ano08, MWF07, SCPA04, BtGW95]. low-energy [BtGW95]. Lucasian [Ano24c, KN03]. lumière [Dav98]. luminaries [Bri95a]. Lyman [SBK^{+}08].

M [Ano85, Ano10, Anoxx, Ben34, Bro06, BP79a, Cha73, CH16, CMO00, Dal87b, Dal95, Dar96, Dav36, Deb13, Dir37c, Dir40, Dir80a, Dys87, Eli87, Eps35, Ett31, Fis10b, Fis12b, Got11, Hal78, Hei30a, Hei30b, HK93, Hov96, Jär64, Kil76, Kri87, KMPD85, LW78, Lan87a, LJ31, MS87, Med90, Meh72, Meh87b, Meh87c, Oak00, Pau30, Pol58, Sch69, Sch94b, Smo84, Suc67, Tem35, War96, Wil36, Wri16, Wil36]. Mach [BP95, BD61, Dic61, Dir61a, HND^{+}62, PD62]. Machian [Son99]. MACSYMA [TM85]. made [Haw11, Sch94a]. magic [Wil03]. Magnetic [CT83, Dir48a, Dir77e, For63, God87, Hoy87, Inf53a, Kol67, MC70, MKS^{+}13, Pic81, Sch06, Ama68, AC72, BK76, BB85, Beh85, BM70, BG90, CK78, CH02, Cor86, Dir25a, Dir48d, Dir76c, Dir84d, Gre77, Hau72, LMV91, Lub75, MG87, Ogu96, Ros83, Rou84, Som34, TCC97, Vil95, YI90, dB33, dB34, Dir74f]. magnétique [dB33, dB34, Som34]. Magnitnii [Dir74f]. Main [Ano34a]. Majorana [RCC^{+}06, Ano08, ACM10, Esp08, ES10, Esp12, EAW15, Fio06, FK08, Gar84, GG11, MRB75, MZF^{+}12, zXZ06]. Majorana-neutrino-mass [Gar84]. make [Dir81a]. makers [CM86]. making [Bel99]. Man [Cas10, Dys10, Far09b, Kib98, RS10, Sti13, Bla11, Far09a, Far16, God98]. manifold [MB90, MB91, Mar88]. manifolds [GNH78]. Mann [MHN98]. Many [Bri18, Kat00, Dir29b, Dir30d, DFW99, RdO07, RdO16]. many-electron [Dir30d, DFW99]. Many-particle [Kat00]. MAPLE [Vul03]. Marcel [Ruf82]. Mass [HN71, IRV^{+}08, SBK^{+}08, Car94, CKYY07, Gar84, IS10b, KY01, KBN83]. masses [Mal75b, Mal75a]. massive [AW17, Dar90, KL92, PR86]. massless [Bor85, Nil82]. matemática [Bas08]. materials [WBSB14]. Materiewellen [Deh36, Haa34]. materja [Inf33]. Math [Szm99]. Mathematical [GG02, CH16, Dir78d, EAW15, Mur71, vEdG85, vSJLEdJG86, Mar65]. mathematically [Bel91]. Mathematics [Dir82e, Ano24c, Bue05, Dir39c, Dir82a, Dir85, FF91, Pen97, KN03]. Mathisson [Dir40]. Matrices [Goo55, IKRT89a, Lee95, Mor85, Pod97]. Matrix [KMW86, MR82b, Dir30d, Hos87, IKRT91, NM84, Szm98, Szm99, Wig01, WY84, ZQ90]. Matter [Ano30, HND^{+}62, Plo15, Ano18, Bat07, Dir77b, DRF17, Inf33, Inf34b, Haa34]. matters [Rei72a, Rei72b]. Maurice [Ano52, Ano83a, Ano83b, DP86, Hal85, Hen89, KW87, Pei85, Ram89, Sch88, Stu88, Wig88, Dir74d, Kra06a]. Max [Fis10a, Fis12a, Ano52, Fis10a, Fis12a, Gre05]. Max-Planck-Medaille [Ano52]. maximum [LR79, LRK80]. Maxwell [Bri18, BCK97, BR97, Can98, Cha72, DK89, Das93, Das96, Faz15, GS79, HD77, KBN83, LU31, LCH10, PR86, Rad96, RA11, RdO07, RdO16, Sch41, SM03]. May [Ano18, Car01]. Mayer [PS32a, PS32b]. McDougall [Fuk14]. Me [Ber09]. Mean [KK10, dB32a]. Meaning [Jor59a, Smo93, Dir50c, Som36]. Means [LR59, Eeg80, Fer86, IKRT89b]. meant [Kem87]. Measure [LR59, Nak00, Nor65, Nor90, Zas89]. Measurement [Neu33, Hal78, Pet85, Pet88]. measurements [Rob85]. measures [Nak97]. mecánica [Dir54a, Dir67b, Dir68c, Dir69e]. mécanique [Dir30h, Dir31a]. mechanical [Lan30, Lan05a]. Mechanics [DL82, Dir25e, Dir26a, Dir26d, Dir26e, Dir30b, Dir30f, Dir32a, Dir35c, Dir47b, Dir54d, Dir55d, Dir58c, Dir59e, Dir67a, Dir67b, Dir67e, Dir67c, Dir68c, Dir69e, Dir81d, Dir03c, Dir05a, Dys07a, Fey48, GG02, GGK02, GL23, Har28, HSD34b, Mar65, MR82b, Nis32, Sei36, Tau61a, Wey31b, Wey50, vdW67, Ant69a, Ant69b, Bok04, CY91, Dir26f, Dir26c, Dir27a, Dir27e, Dir29a, Dir29b, Dir31a, Dir32b, Dir33b, Dir33c, Dir37a, Dir37d, Dir39b, Dir42e, Dir45b, Dir51d, Dir55b, Dir64b, Dir64c, Dir64f, Dir66a, Dir67f, Dir68a, Dir69a, Dir72d, Dir73f, Dir74c, Dir77d, Dir82b, Dir99, Dir01, Dir09, DR93, Fei02, FH65, FHS10, FK70, Got11, Haa34, Jär64, JT10, Mat87, MR82a, MR82c, Meh87a, MR87, MR00, MS68, Pol87c, Tjo75, vEdG85, vdW68, vdW07, Eps35, LJ31, Pol58, Tem35, Hei30a]. Mechanics [Dav36, Hei30b, Pau30, Bro06, Dra23, Kra13]. Mechanics. [Ett31]. mechanique [Dir53c]. mechnanics [Dir54a]. Medaille [Ano52]. Medal [Ano85, Ano13, Ano24a]. medarbejdere [BR64, RRK^{+}64]. méecanique [Dir90a]. Meeting [CCJ^{+}34, Pei87a, Ruf82, Wig87a, Wig96a, Hal78]. Mehra [Dys74]. mekhanike [Dir68a, Dir99]. Mekhaniki [Dir32a, Dir60c]. Mem [Dir61b]. Memo [BBD^{+}46]. Memoirs [Dir05b]. Memorial [And88, Ell91, Pei87a, SDEW90, dG97, Dir71a, FW87, Haw98, Pei92, Rei72a, Rei72b, Wig87a, Wig96a]. Memories [Dir88a, Dir88b, Dir90d, Kir97]. memory [Ano85]. men [Sch94a]. Mensch [Far16]. mérés [Neu33]. Merit [Ano24b]. Merwe [Ble86, Sch85, Tor87]. Mesotron [Kap40]. meteorites [PD62]. Meteorology [Tau63c]. Method [BCJT16, Pau43, Tau49, BO92, Cra88, Fuk14, GL92, Inf53b, LCH10, NM01, RSK99, SM35]. Méthodes [Dir53c]. Methods [Dir69d, Dir82d, Dir89, Har28, Dir30a, Dir41, Dir64c, Dir70b, MKK05, RS87]. Metric [Lee23, Hei72, HM99b]. metrics [IK91, SC92b]. Metrological [IC04]. Miami [KPS^{+}78, KP83, Sta87]. microwave [Han99, YS03]. Mills [LH86, RTV99, Tol98, Lan01, MB90]. Mind [Hür22, II08, WD06, Sno73, Bus10]. minds [Pic08, Wal09]. Minkowski [GLMR94, KLMW88, Pod97, RK96]. Miramare [Meh73]. mirror [HW96]. Mixed [Lub75]. Mode [DJ95]. Model [GD10, Lee23, Mon95, Das96, Dir62c, ERW75, Gab13, KM04, OS81, RM95, WP96, Wot91]. Modeled [Csa92]. Modelers [Lak96]. Models [Dir74b, Dir81c, Lak96, All03, Len14]. moderation [LT04]. Modern [BBBM90, Far03, Gib19, HSD34b, L.88, Bra09, Inf34b, Nor65, Nor90, ZdSS91, HSD34a]. moderne [Ben34, HSD34a, Wil36]. modification [IP57]. Modifications [MR82b]. modified [Bel85, IP56]. Molecular [DN96, McW73, QGW01]. Molecules [FK07, BBSF10, ERW75, EG84]. Möller [Ano67]. Moment [Roz20, BK76, BB85, Beh85, Ogu96, Rou84]. momentum [Dat04, DMH85, Nak00]. monatomic [Fer26]. monobromide [BBSF10]. Monografii [Dir02]. monograph [Bus10]. Monographs [Dir02, Sch85]. monopol [Dir74f]. Monopole [Ble76, Che77, Fer49, HSZ76, Kol67, MC70, Pic81, RZ80, BB85, BG90, CK78, Coq85, Cor86, Dir77e, Dir78e, Fre95, Gre77, IU86, Kra81a, LMV91, Lub75, Lub82, MG87, NdlC08, Oli98, Pos84, TCC97, Vil95, Yan78b, Yan78a, YI90]. Monopoles [For63, Hen83, MKS^{+}13, Sch06, CT83, Dir74f, Dir76c, Fre95, God87, Hoy87, LBNZ98, Ros83]. Moon [Dir80g]. moons [RA11]. most [EM33a, Haw11]. motion [Dir42c, Dir64e, Dir69b, Dir70a, Dir71d, Dir74a]. motions [AW95]. moving [Bet83]. moyennes [dB32a]. MR2807927 [Bla11]. multicomponent [YZ08]. Multiconfigurational [VIK98]. Multiplication [Dir43a, Dir42d, PFD43]. Multiscale [BCJT16]. multivectorial [KRR90]. Must [Ano30, Far03]. My [Dir83a, Dir03a, HC87, Dir87a, Kem87]. Mystery [Ano30, FR13]. Mystic [Cas10, Dys10, Far09b, RS10, Bla11]. mystifying [Lut07].

N [Hen89, L.88, Ram89, Sch88, Stu88]. Nach [Dir60b, Sei36]. näherungsweisen [SM35]. Named [Ano33, BBC^{+}94]. nanowire [MZF^{+}12]. Nation [Ano33]. Natural [McW73, EM33a, LT04]. Nature [Dir63a, Dir87c, Gam68, HN71, Kra19, Wil58, Alp73, Bar02, Dir71b, Dir73a, Dir74e, Meh73, Zic88, Dys74]. natürlichsten [EM33a]. Naturwissenschaft [Smo93]. nauchnych [Koz90]. Nauchnye [Dir03b, Dir04, Dir05b]. Nauchnykh [Dir02, Dir03b, Dir04, Dir05b]. nauki [Inf33]. Near [Dat95]. Negative [Ano30]. neobychajnoj [Dir90d]. Neobychaĭnoĭ [Dir90c]. net [Sel94]. Neuberger [HJL00]. Neue [Ben34]. Neumann [F^{+}00, Mar65, Réd05, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Tjo75, vN96]. Neurons [Kir97]. neutral [BB85, Csa81, Ogu96, SV93]. Neutrino [GLR14, Gar84, KM84, PR86, Gar84]. neutrino-electron [Gar84]. Neutrinos [Fio06, zXZ06, Ano08, CGKS94, CT12, Fra04, HM99a, CT12]. Neutron [Dir42d, Dir43a, JRF06, LT04, VKA21, PFD43]. Newman [EF77, KM92]. Newton [BP95, KN03, LBNZ98, WW86]. Niels [BR64, Dir64g, Dir67d, Kra17, RRK^{+}64, Roz67]. No [EKI94]. No-Pair [EKI94]. Nobel [HSD34a, Ano33, Ano34b, Ano65, Ano24d, Bri95a, DS83, Gre05, HSD34b, LW34, Pau45]. Nobelpreis [Ano51]. Nobelpreises [HSD34a, Wil36]. Nodes [Dir26a]. Non [Fey48, FSY00, Har28, Dir42a, Dir42d, Dir55c, FZ91, MB90, MB91, Mar88]. Non-Coulomb [Har28]. Non-existence [FSY00]. non-Lie [FZ91]. non-orthogonal [Dir55c]. Non-Relativistic [Fey48]. non-Riemannian [MB90, MB91, Mar88]. non-scattering [Dir42a]. non-uniform [Dir42d]. noncompact [BT90]. noncovariant [CY91]. nondiagonal [SC92b]. Nondispersive [Hil92, SBZ90]. nonexistence [Zas89]. nonextensive [JT10]. nonextensive-statistical-mechanics [JT10]. nonfactorizable [HM99b]. Nonlinear [Bou08, Ker83, KK10, AKV83, FZ91, PSA94, RU81, SES82, SOS84, Tak79]. Nonrelativistic [BCJT16]. nonstandard [LW75, Nak97]. Nonunitary [yRSBSV99]. Nordström [FSY00]. norm [Tam03]. Normal [DJ95, Tha88]. Normal-Mode [DJ95]. North [BCEP94]. Notation [Ano67, Dir39b]. Note [Dir24b, Dir24c, Dir30c, Dir30d, Dir55c, Wal76, Tjo75]. Notes [AG13, Pei42, RS10, Bel82]. nothing [Wea16, Ryc17]. novel [SBZ90]. November [Hal78, Lan87a]. Nowe [Inf33]. noyaux [CCJ^{+}34, Sch72]. Nuclear [Irv83, Mau76, Mla98, Stu18, Wig96d, AG13, Kae48, RC13]. Nuclei [CR12, CCJ^{+}34, Sch72]. nucleosynthesis [Kha86, KPW86, YS03]. Nucleus [FR13]. null [KM84]. number [DOV02]. Numbers [Alp73, Dir74b, Dir75d, Dir78c, Dir79g, Bar02, D’A02, Hes75, PP71]. Numerical [Csa92, HJL00, Pic89, Sei99, Tau63b, AME14, AKV83, LCH10, NK93, SM91, SFVW95, SG11, F^{+}00]. Numerology [Gam68]. NUT [CV02, LBNZ98].

O [Ble86, Sch85, Tor87, Bel82, Dir90c, Dir90d]. Obey [Kap40]. Obituary [Bin00, DP86, Dir40, Dir61b, Dir75e, Pei85]. Objects [Gam67d]. oblate [DFPP43, PFD43]. Observables [Hes75, Cor07, Hes73]. observational [Uza03]. Observations [VKA21, FA10]. observed [Dic59]. Obtain [Ano34a]. October [CCJ^{+}34, DP86, Far01, RCC^{+}06]. octobre [CCJ^{+}34]. octonions [Köp06, MB91]. odd [Paĭ90]. odyssey [SG11]. off [Wig01]. off-diagonal [Wig01]. Oklo [LT04]. Old [Dic59, PB68, Zic88, Zic88]. Oliver [Ano30]. OM [Dal87b]. Omega [Bar02]. onde [dB32b]. One [Gib19, Gun15, Kir97, NP96, Wal09, AKV83, Dar90]. One-dimensional [NP96, AKV83]. ongoing [Das96]. ontology [Wri16]. Open [Bok04, KL92, VC02]. Operator [DJ95, HJL00, Hin77, Neu00, Szm98, Szm99, Ara93, BT90, Cla94, Dir37d, FG87, FSS98, Fei02, HW90, KMW86, KLMW88, Lan01, Mil98, Pen99, Pod97, Tam03, Tha88, Tol98, Tra92, de 89]. Operators [GZM05, Tau63a, All03, BILP03, CV02, DEŠ89, FGM^{+}96, Hes75, KY01, KM84, KLMW88, Kla90, LR79, LRK80, Olv86, Wil98, WY84, Zor80, de 91, Tau61b]. Oppenheimer [Ano34a, Dir71a]. Optical [CR12]. optics [Bat07]. Orbis [Per80, KPS^{+}78, KP83]. Orbit [Löw64]. orbitals [Tal04]. Orbiting [RA11]. orbits [Hal87]. Order [Mac98, Rho50, VIK98, Hos87, Hos89, Kru94, Ano24b]. Orders [CT67b, CT67a]. organized [Ruf82, Sal69]. Organizing [Dra23]. Origin [Dir83b, Cor70]. Original [Ano34a]. Origins [Moy81b]. orthogonal [Dir55c, Vil90b]. orthogonally [KM84]. oscillator [Cra93, DST92, LMV91, MSB92, de 91]. Osnovy [Dir32a, Med89]. Other [L.88, K.L88, Tau62, Bri95a, Rei72a, Rei72b, Sha87]. Our [Bek86, Dir71b]. Overlap [Neu00, Tal04]. Owen [Ryc17]. Own [Bek86]. Oxford [RS10, Ano24d].

P [And88, Bel73, Ben34, Dal87b, Dar96, Dav36, Dir37c, Eps35, Ett31, Hal78, Hei30a, Hen89, Hov96, Jär64, Kil76, KMPD85, L.88, LW78, LJ31, Pol58, Ram89, Sch69, Sch85, Sch88, Stu88, Suc67, Tem35, War96, Wil36, Ryc17]. P. [Ano85, BP79a, Cha73, Dal95, Dar96, Dir80a, Eli87, Got11, Hei30b, HK93, Jär64, Kri87, MS87, Med90, Meh72, Meh87b, Meh87c, Pau30, Sch94b, Smo84, Wri16]. P/b [Sch85]. packet [SBZ90]. Padé [Rod80]. pages [RS10]. Pair [EKI94, KM84, EKI94]. Palais [Dir48e]. Pantheon [Lee21]. paper [Dit15, Pei42]. Papers [Ano75, Anoxx, Dir03b, Dir04, Dir05b, HDW75, Sch58, Sch03, vN96, CH16, DPI71, DP97, Dor82, GG76, Haw11, Oak00, Wig96d]. para [JM99]. para-supersymmetry [JM99]. Parameter [Som36, KM84]. parameters [DW09, Som36]. Part [MR87, Dir43a, Har28, PFD43, Wig96d]. Particle [BH83a, BH83b, Dir24c, Dir73h, Roz20, Ser12, AW17, BB85, Bou01, BG90, Bro77, BH82, CY91, Dir64e, DP03, FA01, IT84, IT88, Kat00, LCH10, Nil82, OS81, Ogu96, Rou84, SMI83, SS98, Son99, dB51]. Particles [And88, Ano18, Dir62e, Cor70, Dar90, Dir74a, Dir86, FW87, Fis10a, Fis12a, Kos86, KP83, Mil07, PB68, RS84, Riv94, SV93, Sri83, dB52a]. particular [CGKS94, MN97]. particule [Bou01]. particules [dB51, dB52a]. partir [BI34]. partnership [BR87]. Pascual [Kra15b]. Passion [Jon08]. past [Pet11]. Path [Nak97, Nak00, OS81, PS79, Bel82, FH65, FHS10, IT84, Jac25, KY93, Len14, Rum86, SS72, SH99, Zas89]. path-integral [SS72]. path-space [Zas89]. pathways [Inf33]. Paul [Bla11, Ble86, Cas10, Dar91, Dir74d, Dys10, Far16, Hen89, Kra06a, L.88, Ram89, RS10, Sch85, Sch88, Sti13, Stu88, Tor87, Wig88, Wig96a, Ano52, Ano83a, Ano83b, Ano10, Anoxx, BvdM83, BVV84, Bok08, Bri95b, BR87, Bro06, Cas85, CH16, CMO00, Cri18, DP86, Dal87a, Deb13, Dir87a, Dys87, Far09a, Far09b, Far16, Fis10b, Fis12b, God98, Hal85, Kem87, Koz90, Kra13, KW87, K.L88, Lan87a, Lan87b, McC04, MB11a, Mot87, Oak00, Pai98, Pai00b, Pei85, Sno73, Tay87b, Wal09, Wig87b, Wig96c, dBDW^{+}84, Kib98]. Pauli [CV02, Dit15, GH75, Hor11, JW28, Sch99]. Paulische [JW28]. PC [IKRT91]. peace [Kra17]. pedagogy [BN13]. Peierls [DP97]. perfect [IV85]. perfection [Dir79m]. Periodic [Tau61b, FSY00, NP96]. Perlman [Dir64a]. permeate [BMP14]. personal [Oak00]. perspective [Dir75c, Kra06b]. perspectives [HE82]. Perturbation [Kis98, Löw64, Zor80, Dir55c]. Peter [L.88]. Phase [Lin10, Nil82]. Phase-integral [Lin10]. phenomena [Dir30c]. phenomenological [EAW15]. Philosophy [UM86, Dir37c]. Photo [Ano83a, Ano83b, DH32a]. Photo-electric [DH32a]. Photon [MRB75]. Phys [Szm99]. Physical [Dir27b, Dir37c, Dir87c, Dor82, GL23, Tel48, D’A02, Dic59, Dir42e, Dir68b, Dir84c, Dir84e, ES10, Jor37, Pet85, Pet88, DSK72]. physiciens [CT12]. Physicist [Dir63a, Dys74, Hen89, Ram89, Sch88, Stu88, Dir73a, Dir83a, Gre05, KW87, Meh73, L.88]. Physicists [Ano33, Gun15, Hür22, Ser12, vW02, CT12, Fre87]. Physics [And88, Ano34b, BDH70, BH83a, BH83b, Bun82, Dir76a, Dir76b, Dir84b, Dys93, Gam85, Hal78, HSD34a, Hen84, HK93, Irv83, Koz90, Kra82, Kur73, McW73, Med90, Meh73, Mla98, Pau45, RCC^{+}06, Rob85, Ruf82, Ryc17, Sch85, Stu18, Suc67, Zic88, AG13, Ano65, BN13, BDH^{+}89, Bra09, BH82, BP79b, CM86, Dir39c, Dir53a, Dir69d, Dir70a, Dir70b, Dir71c, Dir71d, Dir77a, DHS78, Dir80d, Dir81b, Dir82d, Dir89, Dor82, EAW15, FF91, FW87, Gam66a, Gam66b, Gam72, Haw11, IPM73, Kae48, KHFA67, KPW75, KPS^{+}78, KMPD85, McC04, Meh72, Meh75, Meh87b, Per80, RC13, Rec97, Sal69, SS03, Str11, Wea16, Wei77, Zic00, CCJ^{+}34, L.88, LW34, Wig96d, Str71]. Physik [Ano51, Dir53a, Dir77a]. physikalischen [Jor37]. Physiker [vW02]. physique [CCJ^{+}34]. Picture [Dir63a]. Pictures [Dir64a]. piece [Wil03]. Pierre [Ble86, Tor87]. pilot [Col12, dB52b]. pilot-wave [Col12, dB52b]. pilote [dB52b]. Pinpointing [Ano12]. pioneers [DR93]. Planar [CH02]. Planck [Ano52, BF12, Fis10a, Fis12a, Ano12, Bos24, Cra88, Fis10a, Fis12a, Inf61, Inf63, KM12, MR82d, MR82e, Noe73, Peg77, WB78]. Plancks [Bos24]. plane [BN72, DK89, Kim80a, Mur71]. plane-wave [BN72]. play [MB11b]. Playing [Pai87]. plus [DSP03]. Pluto [RA11]. Poet [Ano33]. Poincaré [Dir72a]. point [All03, Dir48b, Ric03]. point-circle [All03]. Poisson [FLS76, Lan72]. Pol′′ [Koz90]. Poles [Dir48a, Ama68, AC72, Dir48d, Dir84d]. Polish [Inf33]. politics [Sch99]. polya [Dir90b]. Polygamma [Din58]. Polymath [Har01, Hen89]. polynomials [MR10]. Pontifical [Dir80c]. Pontificia [Buc80]. pontosságának [Neu33]. Portrait [Pai00a]. portraits [Far01]. Position [BF12, Hin77, KM12, Dic59]. Positioning [BF12, KM12]. Positive [Dir71f, Dir72c, Dir73h, Dir71e, Gam33, Pei34]. Positive-Energy [Dir71f, Dir72c, Dir73h, Dir71e]. Positron [Dir75a, HE06, Mon95, Pas12, CDK90, Dir34a, Dir34c, Dir34b, Dir84a, Far10, Gol05, Hei34b, Hei34a, HE35, Maj37, BBC^{+}94]. positrone [Maj37]. Positrons [Dir34d, Hei34b, Dir65b, Dir91b, Hei34a, HE35]. possibility [Mal75a, Mal75b]. Possible [Han99, JRF06, Dir71d, Vul03]. Possibly [Can98]. Post [Ano34b]. posteriori [ABR06]. posters [Ano02]. Postscript [Sch94c, Mit10]. potential [Ara93, BC78, DEŠ92, DSP03, GS84, Hyl84, IS10a, IS10b, IU86, KY93, yRSBSV99, RU81, SMI83, Vil95, YI90, de 89]. potentials [BCK97, BM70, Cor70, Cri76, Fin00, SCSB93]. power [SMI83]. pozitrona [Dir34b]. Pp [Hen89, War96, Dre90, Dys10, Suc67]. Precisely [Cor07]. Precision [Csa92, Fer86]. predict [Far10]. predictable [Cor07]. Prediction [Dir78h, GLR14, Pas12, Gol05]. Prehistoric [Irv83]. Prehistory [Dir80f]. Preis [Ben34, Ble86, Deh36, Wil36]. Prejudice [Dir76a]. prejudices [Dir77a]. Preliminary [Dir26e, Dir67c]. Presence [MC70, AD87, Fin00, Hau72, HM99a, SV93, SS98, Vil95, YI90]. Present [Dir83c, Dir59f]. presentation [Ano65]. presented [Dor82]. preservation [Oak00]. Press [Ble86, Dre90, Hen89, Hen91, RS10, Ryc17, Sch85, Suc67, War96]. Presymplectic [GNH78]. Pretty [Dir82a, Dir82e, Dir85]. previous [Dat04]. Primordial [Noe73, Kha86, KPW86]. principe [Bor49]. principes [Dir31a, Dir90a]. Principios [Dir67b, Dir68c, Dir69e]. Principle [BD61, Dic61, HND^{+}62, Plo15, BP95, CGK72, Dir24b, Dir74a, LR79, LRK80, PD62, Dir61a]. Principles [Bro06, Dav36, Dir30b, Dir30f, Dir35c, Dir47b, Dir58c, Dir59e, Dir67b, Dir67e, Dir68c, Dir69e, Dir81d, Dir03c, Dra23, Eps35, Ett31, Hei30b, LJ31, Plo15, Pol58, Tem35, Dir31a, Dir09, Hei30a, Jär64, Jor52, Jor55, Kra13, Pau30]. Printsipy [Dir60c]. Prinzipien [Dir30b]. Priscilla [BBC^{+}94]. Prism [UM86]. privedshem [Bel82]. Prize [HSD34a, Pau45, Ano24d, Dir71a, Ano33, Ano34b, Aur71, DS83]. prize-winners [Ano24d]. Prizes [LW34]. probabilistic [Leh91]. Probability [Kos86]. Probe [IRV^{+}08]. Problem [DJ95, Cha72, Cor70, Cor86, Dir30d, Dir33c, Dir80g, IU86, NdlC08, YI90, ZdSS91, Sta12]. problèmes [Dir30h]. Problems [DH32b, JP73, Nis32, All04, Ant69a, Ant69b, ABR06, BM98, Dir71c, PB68, Str71]. procedure [PSA94]. Proceedings [Dys74, KP83, Zic88, Hal78, Kur73, KPS^{+}78, Per80, Sal69, Wat91, BB03, BCEP94, Ruf82]. processes [Car94, Dir36a]. product [Mor85]. Production [Sch06]. Prof. [Ano34b, Dir37c, Dir61b]. professional [Bus10]. Professor [Dal87b, Fre87, HC87]. Professors [KN03, Nis35a, Nis35b, Ano24c, LW34]. Profs [Ano33]. program [BDM81b, BDM84]. prohibition [JW28]. Projections [RK96]. Projective [Sch91a, TVvN34]. Proof [Bar81]. propagating [CGKS94]. Propagation [IT84]. propagator [AW95, IJ02, Rum86]. propagators [RR93]. Properties [Ara93, Goo55, Beh85, CCJ^{+}34, GZDA01]. property [Mur71]. propriétés [CCJ^{+}34]. Proton [Dys67, Mon95, SBK^{+}08, CKYY07, Dir30g]. proton-electron [CKYY07]. Proton-to-Electron [SBK^{+}08]. Protons [Dir30i, Dir30e, Dir30j]. Proximity [FK08]. Pseudospectral [BCJT16]. pseudospheres [Cla94]. pseudospin [IS10b]. Psi [Din58]. pure [Ric03]. Puti [Dir83d, Bel82]. Putting [Wil93b].

QED [Sch94a]. quadrature [MKK05, Sag91]. quanta [Bor49, Inf33, Inf34b, Lub82]. Quantelung [Fer26]. Quantengenies [Far16]. Quantenmechanik [Deh36, Dir27e, Haa34, Dir30b, Wey31a]. Quantensprung [Fis10a, Fis12a]. Quantentheorie [Dir28f, Dir28d, Dir28e]. quantification [Dir49c]. quantique [BI34, Dir30h, Dir31a, Dir53c, Dir90a]. quantisation [Her18]. Quantised [Dir31b]. Quantities [WBR76]. Quantization [Pau43, Riv94, AT94, BV85, CDK90, Dar90, Dir49c, Dir68d, Fer24, Fer26, OS81, OP95, PSA94, SS72, YI90]. quantized [BSW98]. quantizzazione [Fer24]. Quantum [Ano75, BVV84, Bel73, Bel99, Bos24, Bro05, Bro06, Dav36, DL82, Dir25e, Dir26a, Dir26d, Dir26e, Dir27b, Dir27d, Dir29b, Dir30b, Dir30f, Dir32a, DH32b, Dir35c, Dir43b, Dir47b, Dir48c, Dir54d, Dir55d, Dir58c, Dir59e, Dir65a, Dir67a, Dir67b, Dir67e, Dir67c, Dir68c, Dir69e, Dir81d, Dir82f, Dir83b, DPP97, Dir02, Dir03c, Dir03b, Dir04, Dir05a, Dra23, Dun12, Dys07a, Dys10, Eps35, Ett31, Fey48, FH65, FBD05, FHS10, GG02, GGK02, Gam85, GL23, Hei30a, Hei30b, HSD34b, HDW75, Hür22, II08, Jon08, LJ31, Mar65, Nis32, Plo15, Pol58, RS10, Sch69, Sch85, Sch94b, Sch58, Sch03, Suc67, Tau61a, Tem35, Wey31b, Wey50, dBDW^{+}84, vdW67, Ant69a, Ant69b, BN13, BP95, Bok04, BI34, Bor49, Bro77, Col12, Dir25b, Dir26b, Dir26f, Dir26c, Dir27e, Dir27c, Dir28f, Dir28d]. quantum [Dir28e, Dir28b, Dir28c, Dir29a, Dir31a, DFP32, Dir32b, Dir33b, Dir33c, Dir35b, Dir37a, Dir37d, Dir39b, DPP42, Dir42e, Dir45b, Dir46, Dir47a, Dir49a, Dir51d, Dir55a, Dir55b, Dir57b, Dir62a, Dir64b, Dir64c, Dir64d, Dir64f, Dir66a, Dir66b, Dir67f, Dir68a, Dir68b, Dir69a, Dir71a, Dir72d, Dir73f, Dir73g, Dir74c, Dir77d, Dir78d, Dir82b, Dir87d, Dir99, Dir01, Dir09, DR93, Ehb22, Ehb23, Far16, Fei02, Fis10a, Fis12a, F^{+}00, Gam66a, Gam66b, Gam72, Got11, Gre05, Haa34, Haw11, Jär64, Jos72, Kra13, KHFA67, Len14, Lip87, Lou83, Mat87, MS87, Med89, MR82a, MR82c, MR82d, MR82e, Meh87a, MR00, MB11b, Pau30, Pod97, Pol87c, SW72, Sel94, Str11, Tjo75, vEdG85, vdW68, vdW07, Ble86, Dir54a, Tor87, Bus10]. quantumelektrodynamischen [Jär64]. quark [Dir77c]. Quasars [MWF07, SCPA04]. quasiclassical [YI90]. Quasistellar [Gam67d]. quaternions [Dir45a, MB91]. Quelques [Dir30h, Dir48e, dB33]. Quest [Ble86, Sch85, Tor87, BVV84, dBDW^{+}84]. question [BP79b]. quintessence [AG03].

R [Dar96, L.88]. R. [Dir61a, War96]. R.S. [Dir61b]. Radial [BBR85, ROB64, SWO76, IJ02, IM95, Lin10, SFVW95, SG11]. radiating [Dir38a]. Radiation [Dir27d, Noe73, Dir25c, Ehb22, Ehb23, Web87]. Radium [Kae48]. Raleigh [BCEP94]. random [RR93]. range [BC78, Dir73c, Dir73d]. Rank [Kir97]. Rank-One [Kir97]. Rapports [CCJ^{+}34]. Rate [Dir43a, PFD43]. Ratio [SBK^{+}08, CKYY07]. Rational [CT67a, CT67b, SS11, D’A02]. Ray [Sei36]. Rayleigh [WY85]. Razmyshleniya [Dir05b]. reactor [LT04]. Reactors [Irv83, Mau76]. Realism [Pas12]. realities [WD06]. Reality [RS10]. really [Fra04, Nor86]. realm [Kae48]. recalls [DRFS80]. reception [Ehb22, Ehb23, Lub75, MR82c]. Recollections [Dir77f, Dir87e, Dir90c, Lan87b]. red [Inf63]. REDUCE [Vul03]. reduced [Hyl84, Lég95, MHN98, OP95]. reduces [IU86]. reducible [Vil90a]. reducing [GL92]. Reduction [Nil82, FZ91, Hos89, LH86, PSA94]. réécriture [Dav98]. reexamined [GT86]. reference [UK82]. reflection [KD33]. Reflections [RK96]. reformulation [Dir60d]. Regime [BCJT16]. Reissner [FSY00]. relation [Bok04, Dir39c, Dir51d, Pau25, Wig72b]. Relations [UL31, Bel85]. relationships [Lan05b]. relativiste [Dir53c]. relativistes [LT62]. Relativistic [BD61, Dir32b, Dir36b, Dir55d, Dir71f, Dir72c, Dir77h, Dir79h, Dir79i, Dir79j, Dir79k, EKI94, Fey48, Hos87, Sei36, AW17, Das96, Dir49b, Dir55b, Dir59f, Dir62a, Dir71e, Dir77g, Dir77i, DFW99, Gol05, Lip87, PP71, WY85, Kra81b]. relativistischen [Sei36]. Relativity [Ano62, Dic64, Dir24c, Hov96, Inf34a, Ruf82, Sch31a, Sch31b, Hor11, Rob15, TVvN34, DN96, Dir26f, Dir72d, Dir73f, Dir75g, Dir82c, Dir96, Kil76, Wil93b, Wil05]. Relaxation [Col12]. release [Dir42a]. relics [Lee95]. relyativistskoi [Dir59f]. Remark [Goa97]. remarkable [Dir63b]. Remarks [CK78, Wig96b, vN28, dB33, Hei34b, Hei34a, Kim80a, Kim80b]. remarques [dB33]. Remembering [Wig87b, Wig96c, Wig88]. Remembrances [GJEF10, Sha87]. reminiscence [Pol87b]. Reminiscences [Hen89, Jef87, KW87, Mot87, Eli87, L.88, Sch88, Stu88, Ram89]. Renormalization [BJ03, Dir81a]. Replace [GH94]. Reply [Dir61a, Dir64a, Dir52a, Dir52b]. report [KHFA67, Som34]. Reports [HSD34b, Dir28a, CCJ^{+}34]. Representation [AW95, JP73, CPV10, D’A02, Dat04, Dir63b, Hos82, Hos83, Rum86]. Representations [Bor85, DJ95, Wig39, Dir45c, JT10, Lee95, yRSBSV99, NdlC08, Wig89]. requirements [Dir84c, Dir84e]. Research [BN13, Dic59, Bus10, Sha87]. Resolvent [Tam03]. resonances [BB85]. Resonant [Bou08]. Results [Ano34a, Dat95, WP96]. retrieval [DSK72]. Retrospect [BBBM90]. reversal [Dir37d]. Review [And88, Bel73, Ble86, Bri18, Bro91, Cas10, Dar96, Dav36, Dre90, Dys74, Dys10, Ell91, Eps35, Ett31, Goe16, Hen89, Hen91, Hov96, Kib98, Kil76, L.88, LJ31, Pau30, Pol58, Ram89, Ryc17, Sch69, Sch85, Sch88, Sch91b, Sti13, Stu88, Suc67, Tem35, Tor87, War96, Bla11, Her18]. Revised [Ano34a]. revisited [Rug88]. Revolution [Ano18, Kae48, Bel99, CM86, Gre05, Kle05, Ano05, Kle05]. rewriting [Dav98]. Richard [Hov96, And88]. Riemannian [Dir59a, MB90, MB91, Mar88]. Right [Wil93b]. ring [ML93]. Rings [Tau63a]. rise [MR82d, MR82e, MR87]. Robert [Dir71a, L.88]. Robertson [KL90, KL92, VP90, Zec96]. Rockies [Van72]. role [Dir53a]. Röntgenstrahlen [Sei36]. roots [PKKL09]. rotation [NdlC08]. rotational [IV85, IV87, IK91, Jin91]. route [Bel94]. Row [Dat95]. Royaumont [LT62]. Run [zXZ06]. runaway [CKYY07]. Russian [Koz90, Ano33, Bel82, BDH70, Dir32a, Dir34b, Dir59f, Dir68a, Dir74f, Dir99, Dir02, Dir03b, Dir04, Dir90c, Dir90d, Dir05b, Med89, Med90]. Rutherford [Lak96]. RW [Lee23].

S [Ben34, Dar92, Deh36, Dir64a, Sch85, Sch91b, Wil36]. Salam [Bel73, Ell91]. Sallhofer [Lak96]. satisfied [Nas84]. Savant [Ano30]. Says [Ano30]. Sb [Dir90d]. sbornik [Koz90]. scalability [RSK99]. Scalar [Cri76, DST92, DSP03, Fin00, KBN83, Vil95]. Scale [IRV^{+}08, DNRV03]. Scales [WBR76]. scattering [BC78, BG90, Car94, Cor70, Dir25d, Dir26f, Dir42a, Gar84, HJKS91, Jun97, Tha88, WY85]. scholarly [Ano10]. Scholars [Ano34a]. Schönen [Smo93]. School [Bou01, Suc67, Zic88]. schoolroom [Hor11]. Schrödinger [Ano51, Ben34, Deh36, Dir61b, Haa34, Hen89, LW34, Lak96, SM35, Szm99, Wil36, BC78, DL82, Dir64a, DS83, GH75, Haa34, Lam87, LC98, MR10, MR87, Nak97, RA11, SM91, SFVW95, SM35, Szm98, de 89]. Schrödinger- [SM35]. Schrödinger-like [LC98]. Schroedinger [Ano33]. Schucking [SH99]. Schwerkraft [Jor52, Jor55]. Schwinger [AW95, FG87, Sch94a, Wot91]. Sciama [Bin00]. Science [Dir79c, Dir79b, Far03, F^{+}00, GGK02, Gib19, Jon08, Pai00a, Smo93, Suc67, UM86, BR87, Dir37c, Gre05, GM80, Inf33, Inf34b, Pic08]. Scientiae [KPS^{+}78, KP83, Per80]. scientiarum [Buc80]. Scientific [Bro91, Dir02, Dir03b, Dir04, Dir05b, Dre90, Har01, Hen91, Rod19, Sch91b, Sch94c, Alp73, BBC^{+}94, DP97, Dar92, Dir68e, Dir72e, Haw11, Kra90, Meh72, Meh87b, Wig96d]. Scientists [Ano30, Bri95a, Str11]. scoops [Ano13]. Scope [Bus10, Tiw12]. Sea [Kol67, BSW98, Fin00, FG22, Wri16]. Search [Kol67, Sch06, WFC^{+}99, CH16]. Searching [Esp12]. secolo [Wei77]. Second [Roz20, Ruf82, VIK98, Bor49, CM86, Hos87, Hos89, Bor49, Dir49c, Stu18]. Second-Order [VIK98, Hos87, Hos89]. seconde [Dir49c]. secrets [Bar02, HW06]. Secular [Noe73]. Sediment [Kol67]. Seems [GLR14]. seen [Roz67]. Seiberg [Fre95]. Selected [DP97, Sch58, Sch03, Réd05]. selection [Rec97]. Self [BSW88, EG84, Ish90, MHN98, Roh60, Dir42c, KY01, LR79, YL09, Zor80]. self-adjoint [Zor80]. self-adjointness [KY01, LR79]. Self-Consistent [Ish90, EG84, MHN98, YL09]. Self-Consistently [BSW88]. Self-Energy [Roh60]. self-fractionating [Dir42c]. selfadjointness [LRK80]. seltsamste [Far16]. Semi [EM33a, EM33b]. Semi-Vektoren [EM33a, EM33b]. Semiclassical [SM03, NP96]. semiconductor [MZF^{+}12]. semilinear [EV97]. Semivectors [EM33b, EM33a]. sense [Dir81a]. Sensitivity [FK07, BBSF10]. Separability [IV85]. Separable [Rüd84]. separates [IV87, Jin91]. Separation [IK91, KM84, PT86, VP90, Vil90b, VC02, Dir41, KMW86, SV89b, SC91, SC92b, SC92a, Löw64]. Sept [Dys74, Ano05, Kle05]. September [DHS78, Meh73, Pet11]. septième [CCJ^{+}34]. Series [KM92, Bel91, Moh07]. Session [Dir80c, ZDF^{+}84]. Set [Ano34a, QGW01, KL92]. Sets [Dat95, QGW01, Sch06, Tau61a, KLMW88]. setting [Kru94]. seven [Kle05]. seven-times [Kle05]. Seventh [CCJ^{+}34, Far01]. seventieth [Cha73]. seventy [Esp08]. Several [Smo87]. Shankland [Pei36]. Shape [Dir43a, GPS86, FA01, PFD43]. Shaped [Gib19]. Share [Ano33]. Sharp [CG07]. shearfree [KM84]. shell [Csa81, Dir42d, DEŠ89, DEŠ92]. Shift [de 91, Inf63]. shine [Kra17]. Shook [Gam85, Gam66a, Gam66b, Gam72, Haw11]. Short [Ano33, BC78, Deb13]. Short-Story [Ano33]. Sicily [Zic88]. side [Dal87a]. Signatures [MZF^{+}12]. Significance [Dic64]. Silent [Dys10]. silicon [BBSF10]. Similarity [SES82, Wil98]. simmetrica [Maj37]. simple [Rum86]. Sinc [AT07]. Sinc-based [AT07]. since [CT83, Meh75]. Single [Hür22]. singular [All04, Ara93]. singularities [BK10, Dir31b]. Sir [Ano30]. sistemi [Fer24]. Sitter [Dir35a, Dir63b, vN40]. situation [Dir59f]. six [Ten68]. six-dimensional [Ten68]. Sixty [FR13]. size [Dir62e, Kha86, KPW86]. sketch [Dal87b]. Skew [GZM05]. Skew-Symmetric [GZM05]. Skyrmion [MKS^{+}13]. Slater [ERW75, EG84, Kos86, Tal04]. Small [Bou08]. smallest [Fis10a, Fis12a]. so-called [Gol05]. Sobranie [Dir02, Dir03b, Dir04, Dir05b]. Society [GGK02, Gib19, Dor82]. Sociology [UM86]. Soft [dG97]. Solid [Dir43a, PFD43]. Solitary [KK10]. Soliton [Tak79, YL09]. Solitons [GPS86, RS84]. Solution [BK76, GG51, LR59, SMI83, Sri89, TCC97, AME14, Ata89, BC78, BN72, Das93, FA01, Lin10, SM91, SFVW95, SM35, Tu91, Vil90a, VP90, Vil90b, Vil95, VC02, Won90, dB52b]. Solutions [BSW88, DMH85, GB13, IS10b, MR51, Sta67, BM98, CGKS94, DK89, Das93, EF77, FM75, FSY00, Hau72, Hil92, HM99b, HM99a, IS10a, IV87, Jin91, KM92, KL90, KL92, KBN83, PR86, Por95, Rad96, RA11, RTV99, SBZ90, SV89a, SV93, SES82, Tak79, Wil91]. Solvay [CCJ^{+}34, Far01, Meh75, CCJ^{+}34]. Solving [Ano30, SG11, Tau49, TKS19]. Some [Ano55, Dir73g, Eli87, ES10, Jor71, KBN83, L.88, K.L88, Paĭ90, PR86, Phi87, ROB64, RS87, WBD70, Jär64, dB33, vN28]. Sommerfeld [Lak96, DuM28, MR82d, MR82e]. Sort [Ser12]. sostoyanie [Dir59f]. source [ABR06]. Sources [KHFA67, vdW67, vdW68, vdW07, YL09]. sous [CCJ^{+}34]. Soviet [Fre87]. Sovremennoe [Dir59f]. sozdaniyu [Dir90b]. Space [Ble86, Fey48, RK96, Sch85, Tor87, vN40, BO92, BMS02b, BVV84, Das93, Dat04, Dir35a, Dir36c, Dir59a, Dir70c, Dir73e, Dir74e, Dir74g, DFW99, EF77, GLMR94, Hei72, IT88, IKRT89b, IV85, IV87, JM99, Jin91, Jor52, Jor55, KM92, KM84, KLMW88, Kis98, KL92, Lég95, LBNZ98, MB90, Nak97, Nak00, Nas86, Nil82, PT86, Pru73, Ran87, Ric03, Rüd84, Sri89, Vil90a, VP90, Zas89, Zec96, dBDW^{+}84]. spaces [BILP03, FA01, KL90, OP95, Pod97]. spacetimes [Vul03]. Spaltung [EM33a]. Spanish [Bun52b, Dir54a, Dir67b, Dir68c, Dir69e, SHD91]. spatially [Das93, IS10b]. spatially-dependent [IS10b]. Special [Tau49, ZDF^{+}84, Das96, Hor11]. specified [Nas97]. Spectra [MWF07, Sla29, SCPA04, LBNZ98, Pau25]. Spectral [Beh85, BK10, Leo80]. Spectroscopic [Ano67]. spectroscopy [SMI83]. Spectrum [Mil98, Gin09, MD88, Ogu96]. speech [Dir71a]. speeches [Ano65]. Speed [Lee23, IKRT89b, Kra06b, Lee21, Pei35, Tak31]. Spektren [Pau25]. sphere [Dir42b, Dir42d]. spheres [Cla94]. spherically [DEŠ89, DEŠ92, PT86]. spheroid [DFPP43, PFD43]. Spin [Löw64, vN28, Dar90, Dir75c, FA01, IS10b, Ogu96, Pen97, Seb19, dB51, dB52a]. spin- [FA01, dB51, dB52a]. spin-one [Dar90]. Spin-Orbit [Löw64]. spinning [AW17, Riv94]. Spinor [EM33a, LU31, Dir62d, Eeg80, Hal87, Nas97, RL68, Wig01]. Spinor-Gleichungen [EM33a]. spinorial [NM84]. Spinors [Dir70c, Dir74g, RK96, Tra92, Lég95, NM84, Nas86, RR93]. spins [Kat00]. Spirit [Gib19]. Splitting [EM33a]. St [Dir37c, Dir90d, Jef87]. Stability [Bou08, Roh60]. staircase [Fis10a, Fis12a]. Stand [Jor59b]. standard [Gab13]. Standing [Bou08, KD33]. Stars [JRF06, Jor49, VKA21, Kra17]. State [BB03, CH16, CMO00, Dir83c, Hal78, Lan87b, Oak00, Cor70, Hei72, IS10a, Jor59b]. Statement [Dir33c]. States [GPS86, BB85, RU81, ZQ90]. Stat’i [Dir03b, Dir04, Dir05b]. static [KL92]. statistical [Dir25c, Dir29a, Dir41, JT10]. Statistics [Kap40, Bel82, Bel94, Pei72]. statistike [Bel82]. status [Uza03]. stellar [Dir25d, Kra17]. Stellung [Dir53a]. Step [Nov74, Wal76]. Steven [And88]. still [Kae48]. Stockholm [HSD34a, Wil36, WD06]. Stoner [Fuk14]. Storia [Wei77]. Story [Ano33, Bel73, Gam85, Jon08, Gam66a, Gam66b, Gam72]. Stoßvorgänge [Dir27e]. Strange [Ryc17, Wea16]. Strangest [Cas10, Dys10, Far09b, RS10, Sti13, Bla11, Far16]. streams [Dir77b]. stress [Dir55e]. string [Che77]. strongly [Ara93]. Structure [BS67, Bek86, CCJ^{+}34, FK07, MS68, MWF07, SCPA04, WFC^{+}99, AG03, BBSF10, Car01, CKYY07, Csa81, Han99, Jär64, MG87, Pau25, PSA94, YZ08, dG80, CCJ^{+}34]. Structures [GZM05]. Struggle [Wüt11a]. Struktur [Jär64]. student [Bus10, Phi87]. Studies [Ble86, KPS^{+}78, KP83, Löw64, Sch85, Tor87, UM86, BVV84, dBDW^{+}84]. Study [Ano34b, DN96, Löw64, Pas12, AKV83, Fis10a, Fis12a, Kra81a, LR79, LRK80, RSK99]. stuff [Haw11]. Sturmian [Hos87]. subelectron [Dir77c]. subgroups [Dir72a, Lég95]. submanifolds [Bet83]. Subnuclear [Zic88]. sulla [Fer24]. Summary [Str71, Ano30]. summation [Bel91]. super [Nil82]. superalgebra [KR92]. superalgebras [HW90, Pen99]. Superconducting [FK08]. superconductor [MZF^{+}12]. superluminal [Sri83]. supernova [FA10]. Superstring [WW86]. Supersymmetric [Tiw12, DST92, Fei02, OS81]. Supersymmetry [DJ95, JM99]. supervisor [Sha87]. support [Car01]. Surface [FK08]. surfaces [DP03]. surrounded [Dir42d]. SWU [Ber09]. Symbolic [D’A02, Wat91]. Symmetric [BSW88, FK70, GZM05, DEŠ89, DEŠ92, PT86]. symmetrical [Dar90, MD88, Maj37]. Symmetries [GM80, Ker83, LMV91, SES81, ZDF^{+}84, BB85, Dir73c, Dir73d, KMW86, Pei92]. Symmetry [Dir55f, Ant69a, Ant69b, BR97, FA01, IS10b, IV85, IV87, IK91, Jin91, KM84, Lee95, Vil90b, Ant69b]. Symplectic [BNB94, Reu99]. Symposium [Meh73, Sal69, Wat91, BB03, Dys74, HE82, LW78, Sno73, Wig96b]. synthesis [SBZ90]. System [BF12, KM12, All04, HJKS91, IKRT89b, KY93, LCH10, MD88, SM03]. Systematic [SBB^{+}62]. Systems [GZM05, AT07, Dir29b, Dir48c, Fer24, GB90, Oev88, PSA94, Rüd84, RTV99, Sei99, TM85, YG77].

T [Bus10, CK78, Dir02, Dir03b, Dir04, Dir05b]. T. [Dir52a]. tale [BMS02a]. talk [Bri95a]. Tallahassee [BB03, Hal78, LW78, Wig87a]. Taub [CV02]. Taylor [L.88, Stu88, Dar91]. team [Car01]. techniques [RL68]. Teilchen [Fis10a, Fis12a]. temperature [Dir24a, PD62]. Ten [Jon08]. tensor [Bet83, Dar90, DMH85, Dir55e, IS10a, Lan05b, RM96, SS98, WY84]. tenu [CCJ^{+}34]. Teoria [Maj37, Bun52b]. teorii [Dir59f, Dir90b, Med89]. Teoriya [Dir34b, Dir02, Dir03b, Dir04]. Term [Sch91a, FG87, KY01, Wot91]. Terms [Löw64, AW95, ABR06, Cor70, Mal75c, MHN98]. Test [BF12, Dir79l, KM12, Rob83, Wil93b, Dir80b]. Tests [Rob85, Ano08, MBS02]. tetrad [Can98, Nas97, RM95]. tetrads [UB83]. Tevatron [Sch06]. textbook [Tiw12]. textbooks [BN13]. Thales [Lak96]. théeorie [Dir48e]. Their [Sch06, Cor70, Dir73b, Dir78a, Dir86, Dys72, LC98, Oak00, Som36, Uza03]. them [Pic08]. Theorem [RRS06, DP03, Kla90, Sem93, Bar81]. theorems [Cra85]. theoretic [WY84]. Theoretical [Dic64, KPS^{+}78, KP83, L.88, Meh73, Ruf82, Sal69, Dir69d, Dir70b, Dir82d, Dir89, EAW15, Jor52, Jor55, Lan30, Lan05a, Meh72, Meh87b, Uza03]. theoretischen [Jor52, Jor55]. Theorie [Hei34b, Hei34a, HE35, Lan30, Neu33, vN28, Dir34c, Bor49, Bou01, Dav98, Dir39a, PS32a, PS32b, dB32a, dB33, dB34, dB51, dB52b, dB52a, Som36]. Theorien [Deh36, Haa34]. Theories [BSW88, KPW75, Mar84, McA90, WW86, BM98, GH75, Haa34, Sal87a, Sal87b, Szm98, Szm99, LT62]. theorist [Ehb22, Ehb23]. Theory [Ano30, Ano34a, BD61, Bro05, Dir27d, Dir30i, Dir30j, DH32b, Dir34d, Dir43a, Dir55d, Dir65b, Dir75g, Dir75a, Dir76c, Dir79d, Dir82f, Dir83b, Dir83c, Dir91b, Dir96, DPP97, Dir02, Dir03b, Dir04, Dra23, DuM28, Dun12, Dys07b, EKI94, FBD05, Gam85, HSD34a, HE06, HL38, Hov96, Inf34a, Inf53a, Kil76, Kim80a, Kim80b, Kra81b, Löw64, Mar65, Neu33, Plo15, Sch91a, Sch31a, Sch31b, Sla29, Tau61a, Tau61b, Tau63b, Wey31b, Wey50, Wie29, vN28, AW17, BNB94, Bor49, Bor52, Bou85, Bou01, Bun52b, BV85, Caw69, Che77, Col12, Dav98, Dir26c, Dir27c, Dir28f, Dir28d, Dir28e, Dir28b, Dir28c, Dir34a, Dir38a, Dir41, DPP42, Dir48c, Dir48d, Dir48b, Dir49a, Dir50a, Dir51c, Dir52d, Dir54c, Dir55b, Dir58d, Dir59d, Dir59f, Dir62a, Dir64d, Dir66b, Dir71a, Dir73g, Dir78d, Dir78f, Dir79f, Dir80h, Dir84c, Dir84d]. theory [Dir84e, Dir87d, Dir79m, DSP03, Edd46, Ehb22, Ehb23, GB90, Gam33, Gam66a, Gam66b, Gam72, Gol05, GNH78, Hei34b, Hei34a, HE35, Hes73, Hes75, Inf63, IU86, Jos72, Kam86, Kis98, Kru94, KP83, Kur87a, Lan30, Lan05a, LW78, Lip87, Lou83, Maj37, MHN98, MS87, Med89, MR82d, MR82e, Oev88, Olv86, Pai72, PS32a, PS32b, Pei34, Pei35, RM96, SW72, SS11, Som36, Tha88, Tra92, Zic84, dB32a, dB33, dB34, dB52b, dB52a, Dir34c, Dir34b, Dir39a, Dir79g, Dir80e, Har28, Sch69, Suc67, Tau63c, dB51, Bel73]. There [Ano30, Dir51b, Dir52a, Dir52b, Fra04]. thermodynamic [BSW98]. Thermodynamics [Bor49]. thermodynamique [Bor49]. Thesis [Bro05, FBD05]. Things [MB11a]. Thinking [Ano00, Dir87a]. third [Zic88]. Thirties [Hen84, Kra82]. Thirty [Gam66a, Gam66b, Gam72, Gam85]. Thomas [MR51, Csa81, Csa92, Dir30c, GG51, MHN98, Rij66, Tu91, YG77]. Thomson [Lak96, Fal97]. Thought [Dir79a, Paĭ90]. Thoughts [Dir05b]. Three [Ano33, HSD34b, Kam86, Wig01, BSW98, BV85, FA01, IU86]. Three-dimensional [Kam86, BV85, FA01, IU86]. three-wave [BSW98]. tidy [Mil97]. tighter [Rob85]. Time [AG03, BS67, BCJT16, Bek86, Ble86, CKYY07, CP68, DNRV03, Dir79l, Dys67, FK07, Gam67a, GT85a, GT85b, GT86, IRV^{+}08, JRF06, Kha86, KPW86, Mar84, MWF07, SBK^{+}08, Sch85, SV90, SV95, SCPA04, Tor87, VKA21, WB78, WFC^{+}99, Wes80, Wil58, WBR76, WW86, AME14, Alp73, BO92, BMS02b, BVV84, BBSF10, Car01, Das93, DSW08, Dir53b, Dir73b, Dir73e, Dir74e, Dir78a, Dir80b, Dys72, DFW99, FA10, Fey48, FSY00, Han99, IT88, JM99, Jin91, KM92, KLMW88, LT04, LMSV08, Lee21, Lev80, Mal75a, Mal75b, Nas86, Nil03a, RSK99, Sri89, Wig72b, YS03, Zas89, Zec96, dBDW^{+}84]. time-dependent [AME14, RSK99]. time-energy [Wig72b]. time-periodic [FSY00]. time-variation [Nil03a]. Times [Ano30, ES10, IV85, IV87, KM84, Kle05, PT86, Rüd84, VP90]. Todorov [CY91]. Together [MB11a]. told [BR64, RRK^{+}64]. Tomonaga [Sch94a]. Too [Hür22]. tool [Ehb22, Ehb23, TKS19]. Topics [Tau62]. Topological [FK08]. Topology [AFG75]. torsion [UB83, Vul03]. trace [Eeg80, IKRT91]. Traces [IKRT89a, PKKL09]. Tragedy [Jon08]. Transform [YYH^{+}01, Leo80]. transformation [CV90, Dir53b, NM84]. transformations [BC69, Dir45a, Dir50c, Dir52c, Eeg80, FG87, GB90, Kim80a, Kim80b, Pon05]. Transformed [Dat04]. Transforming [AT99]. Transforms [LR59]. transitions [Rob85]. transitive [KM84]. Translations [Ano75, HDW75]. Travelling [Rod19]. Travels [Van72]. treasury [FF91]. Treatment [HJL00, WY84]. trends [LW78, Dor82]. trial [RS87]. Trialogue [DOV02]. Tributes [K.L88, Stu88, Tay87b, Dar91, L.88]. Trieste [Dys74, Meh73, Ruf82, Sal69, BDH70]. trio [Ano13]. trip [WD06]. triumph [Gab13]. Trudov [Dir02, Dir03b, Dir04, Dir05b, Koz90]. truly [Alp73]. Tuning [Sta12]. twentieth [CM86, Wei77]. twentieth-century [CM86]. twenty [Zic88]. twenty-third [Zic88]. twistorial [Car94]. Two [LC98, Moh07, SCSB93, BMS02a, CV90, CY91, FA01, GL92, IT88, KM84, Kis98, MHN98, RSK99, Sem93, Ano34a]. Two-body [LC98, SCSB93, CV90, GL92, Sem93]. two-dimensional [FA01, IT88, Kis98, RSK99]. two-parameter [KM84]. Tying [MB11a]. Type [RRS06, Ser12, EM33a, FA10, Hos89, MD88, MB90, MS95]. Typus [EM33a].

U [Ryc17, Haa34, Wil36, Ber09, Dir42b]. U-sphere [Dir42b]. UF [DN96]. unbounded [Kae48]. uncertainty [Wig72b]. Uncovered [Hür22]. undergraduates [Tiw12]. Underlie [RK96]. understanding [Dir71b]. unificaciön [SHD91]. Unification [Ell91, RM95, SDEW90, DNRV03, SHD91]. Unified [GG02, Mar65, Sch91a, Kur87a, Tjo75]. Uniform [Dir43a, RRS06, Dir42d, PFD43]. Uniformly [BCJT16]. Unifying [DSW08]. Unit [Nov74, Wal76, Sch38]. Unit-Step [Nov74, Wal76]. unitaire [PS32a, PS32b]. Unitary [Dir45c, FG87, Wig39, PS32a, PS32b, Wig89]. Units [McW73, SH59]. Universe [Bus10, Jor49, RS10, Bar02, Car01, Dir81c, HW06, II08, LMSV08, Nor65, Nor90, VC02, Sch41, Sri83]. Università [RCC^{+}06]. University [Ble86, Dre90, F^{+}00, Hal78, Hen89, Hen91, KPS^{+}78, KP83, RS10, Sch85, Suc67, War96, BB03, Ano24c, CH16, CMO00, KN03, Lan87b, Oak00, Sta87]. unknown [Dir84e, ES10, PKKL09]. Unorthodoxy [DL82]. Unravelling [FR13]. Unwinding [MKS^{+}13]. Upper [Sch06]. USA [BB03]. Use [Sta12, Dir55c, Dir71d, Inf53b]. used [Dir69b, Dir70a]. Using [AT99, Cra88, Csa92, QGW01, VKA21, Bel91, CPV10, JT10, Lee21, MB90, MB91, Mar88, RS87, Sei99, Vul03, Wot91].

V [Bus10, L.88, BDH70, Tau63b]. vacuum [Dir57b, Dir58a, FG22, Pei34, Pei35, Wri16, YS03, Dir59b]. Vaidya [CGKS94]. Vakuum [Dir59b, Dir57a, Pei35]. valeurs [dB32a]. Value [JP73, YS03]. Values [SBK^{+}08, Cor70, dB32a]. vanishing [DMH85]. Variability [Dir72f, Gam67d, Sch91a, CGK72, Cho81, IC04]. Variable [Rob83, KY01, Sch99, SC91]. Variables [AT99, Bun52a, Dir33a, Dir37a, IK91, KMW86, KM84, SV89b, SC92b, SC92a, VP90, Vil90b, VC02]. Variation [CP68, Dir75f, Dir82f, Dys67, Dys78a, Dys78b, FK07, IRV^{+}08, JRF06, Mar84, MWF07, Noe73, SBK^{+}08, SV95, SCPA04, VKA21, WFC^{+}99, WBR76, WW86, AG03, BBSF10, CKYY07, Dir78c, Dir80g, Dys72, DFW99, FA10, GT85a, GT85b, GT86, Han99, Kha86, KPW86, LT04, LMSV08, Lee21, Mal75a, Mal75b, Nil03a, Peg77, SV90, Uza03, YS03]. Variations [BMS02b, DSW08, DNRV03, Hal78]. Various [DH32b]. Vary [BS67, Dir75b, Alp73]. Varying [AWW^{+}81, BM98, KM04, Kra15b, Kra16, Kra19, Lee23, Mag15, BMS02a, Dir78b, Kra06b, MBS02, Goe16]. Varying- [BM98]. Vechernie [BDH70]. vector [SV89b, SV89a]. vectors [Nas86]. Vedral [RS10]. Veka [Med90, Koz90]. Vektoren [EM33a, EM33b]. velocity [Dic59, Rob85]. venner [BR64, RRK^{+}64]. verb [Haa34]. verborgene [Far16]. Verfahren [SM35]. Verlag [Ben34, Wil36]. Verlagsgesellschaft [Deh36]. verm [Haa34]. versatility [Dir64g, Dir67d]. Version [CO03]. versus [Ano00]. VI [Tau63c]. via [LW75, PSA94]. Viable [Lee23]. vibrates [FG22]. Vienna [MR87]. View [Dir87c, Kri87, Mil07]. Vigier [Ble86, Sch85, Tor87]. Vignette [Ano94]. Vii [Sch85, QGW01]. VIII [Löw64]. Vindications [Moy81c]. Virial [Sem93]. virke [BR64, RRK^{+}64]. Visa [Ano54, Ano55]. visit [DHS78]. Vlatko [RS10]. vodka [Mag15]. Void [Wea16, Ryc17]. Vol [Dir02, Dir04, Dir05b, Wig96d]. volume [Rei72a, Rei72b, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, UM86]. vom [EM33a]. Voreingenommenheit [Dir77a]. Vortrag [Dir60b]. Vorträge [Wil36, HSD34a]. Vospominanija [Dir90d]. Vospominaniya [Dir90c, Dir05b]. vs [Ano08].

W [Ben34, Wil36]. walk [RR93]. Walker [KL90, KL92, VP90, Zec96]. Wars [Stu18]. Was [Wil93b, Lan87a]. watershed [RC13]. Wave [Dir36c, Dir71f, Dir72c, Dir77h, Dir79h, Dir79i, Dir79j, Eva04, Har28, Sei36, VKA21, ZQ90, AD87, BSW98, BN72, Bel91, BI34, Bro77, Col12, DK89, Dir27a, Dir35a, Dir36b, Dir55c, Dir59a, Dir71e, Dir77g, Dir77i, Dir79k, Lan30, Lan05a, MR87, SM91, SFVW95, SBZ90, SS98, dB32b, dB52b]. wave-mechanical [Lan30]. wave-particle [Bro77]. Wavelet [YYH^{+}01]. Waves [Bou08, KK10, Dir60b, Dir60a, Dir87b, Haa34, KD33, Kim80a, Kim80b, Mur71]. Way [Dra23, Pai87, Pei87b]. weak [Web87]. Weatherall [Ryc17]. Weinberg [And88]. Weird [Str11]. Weisskopf [Ano75, Dit15, HDW75]. Weizsäcker [Tu91]. Wellenmechanik [Sei36]. wellenmechanische [Lan30]. Weltall [Jor52, Jor55]. Weltkonstanten [Jor37]. Werner [vW02, BR87, SBK^{+}08, vW02]. wes [Haa34]. Wess [Wot91]. Weyl [Ara93, Mal75c, VP90]. Wheeler [CY91]. which [IV87, Jin91]. Who [Lan87a, Gre05, Sch94a]. Wigner [Hen89, L.88, Sch85, Sch88, Stu88, Szm99, BVV84, Bel73, Ble86, Ram89, Szm98, Wig96d, dBDW^{+}84, Tor87]. Winner [Ano34b]. winners [Ano24d]. without [Che77, Dir65a, Mag15]. Witten [Fre95]. wood [Dir65a]. words [Phi87]. work [Dir68e, Dir72e, God98, Hei34b, Meh72, Pai98, Pet11, Roz67, Kib98]. worker [Bus10]. workers [RRK^{+}64]. Works [Dar96, Dir02, Dir03b, Dir04, Dir05b, Hov96, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, War96, BR64, Dal95, Koz90, Meh87c, RRK^{+}64, Wig96d]. workshop [F^{+}00, Hal78]. World [Hür22, Stu18, Dir55f, FF91, Gre05, Haw11, Inf34b, Jor37]. Wouthuysen [Hos89]. Wouthuysen-type [Hos89]. Write [Ano33]. Wuppertal [F^{+}00].

x [Dre90, Hen89, RS10, Sei36]. X-Ray [Sei36]. Xe [EKI94]. XLIX [Sto94]. xviii [Hen89]. XX [Koz90, Med90, Wei77]. xxiv [War96].

Yale [Ryc17]. Yang [RTV99, Tol98, Lan01, LH86, MB90]. Yano [BCK97]. Year [FR13, RC13, KMPD85]. Years [Dir82c, Gam85, Mla98, Esp08, Gam66a, Gam66b, Gam72, GMW80, Kae48, Rec97, Rob15, RZ80]. Yeshiva [Suc67]. York [Dre90, Suc67]. Yukawa [Tol98].

Zametki [Bel82]. Zealand [DHS78]. Zeeman [Dat04, Kox97]. Zeilinger [Fis10a, Fis12a]. Zeittest [Dir80b]. zero [KBN83, Lev80]. Zhizn [BDH70]. Zitterbewegung [Dir73h, IT88]. Züge [Jär64]. zum [Fis10a, Fis12a, Jor59b]. Zumino [Wot91]. zur [Hei34b, Hei34a, SM35, vN28, Dir28f, Dir28e, Fer26, Sei36]. Zurich [MR87]. Zusammenhang [Pau25].

References

Araya:2006:PEE

[ABR06] Rodolfo Araya, Edwin Behrens, and Rodolfo Rodríguez. A posteriori error estimates for elliptic problems with Dirac delta source terms. Numerische Mathematik, 105(2):193–216, December 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amaldi:1972:DMP

[AC72] E. Amaldi and N. Cabibbo. On the Dirac magnetic poles. In Salam and Wigner [SW72], pages 183–212. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Arimondo:2010:CEM

[ACM10] E. Arimondo, Charles W. Clark, and W. C. Martin. Colloquium: Ettore Majorana and the birth of autoionization. Reviews of Modern Physics, 82(3):1947–1958, July 2010. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://adsabs.harvard.edu/abs/2010RvMP...
82.1947A;   http://link.aps.org/doi/10.1103/RevModPhys.
82.1947;   http://rmp.aps.org/abstract/RMP/v82/i3/p1947_
1.

Abounadi:1987:DWE

[AD87] Jinane Abounadi and Teymour Darkhosh. The Dirac wave equation in the presence of an external field. Journal of Mathematical Physics, 28(7):1505–1507, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arafune:1975:THF

[AFG75] J. Arafune, P. G. O. Freund, and C. J. Goebel. Topology of Higgs fields. Journal of Mathematical Physics, 16(2):433–437, February 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v16/i2/p433_s1.

Anchordoqui:2003:TVF

[AG03] Luis Anchordoqui and Haim Goldberg. Time variation of the fine structure constant driven by quintessence. Physical Review D (Particles and Fields), 68(8):083513, October 15, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.68.083513.

Al-Ghazi:2013:NNP

[AG13] Muthana Al-Ghazi. Notes on nuclear physics in 1932. Physics Today, 66(11):11, November 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [RC13].

Alvarez:1983:NSN

[AKV83] A. Alvarez, Pen Yu Kuo, and Luis Vázquez. The numerical study of a nonlinear one-dimensional Dirac equation. Applied Mathematics and Computation, 13(1–2):1–15, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Allahverdiev:2003:EDF

[All03] B. P. Allahverdiev. Extensions, dilations and functional models of discrete Dirac operators in limit point-circle cases. IMA Journal of Applied Mathematics, 68(3):251–267, June 2003. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic). URL

http://www3.oup.co.uk/imamat/hdb/Volume_68/Issue_03/
hxg016.sgm.abs.html;   http://www3.oup.co.uk/imamat/
hdb/Volume_68/Issue_03/pdf/hxg016.pdf.

Allahverdiev:2004:DEP

[All04] Bilender P. Allahverdiev. Dissipative eigenvalue problems for a singular Dirac system. Applied Mathematics and Computation, 152(1):127–139, April 26, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alpher:1973:LNC

[Alp73] Ralph A. Alpher. Large numbers, cosmology and Gamow: Are the fundamental constants of nature truly constant, or do they vary with time? An account of George Gamow’s last scientific enquiry. American Scientist, 61(1):52–58, January/February 1973. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://adsabs.harvard.

edu/abs/1973AmSci..61...52A;   http://www.jstor.org/
stable/27843563.

Amaldi:1968:DMP

[Ama68] E. Amaldi. On the Dirac magnetic poles. In Puppi and Bernardini [PB68], page ?? LCCN QC721 .O4 1968. A volume dedicated to Gilberto Bernardini in his sixtieth birthday.

Almquist:2014:HFN

[AME14] Martin Almquist, Ken Mattsson, and Tomas Edvinsson. Highfidelity numerical solution of the time-dependent Dirac equation. Journal of Computational Physics, 262(??):86–103, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008462.

Andersen:1988:BRR

[And88] Per H. Andersen. Book review: Richard P. Feynman and Steven Weinberg, Elementary Particles and the Laws of Physics: The 1986 Dirac Memorial Lectures. Physics Today, 41(4):96, April 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1930:SAN

[Ano30] Anonymous. Scientists acclaim new atom theory: Sir Oliver Lodge says Dr. Dirac has important contribution to solving mystery of matter. Summary of his findings. Cambridge savant says there must be times when kinetic energy of electron is negative. New York Times, ??(??):11, September 10, 1930. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1933:NPG

[Ano33] Anonymous. Nobel Prize goes to Bunin, Russian; short-story write and poet, exile in France, is first of his nation to get award; three physicists named; profs. Dirac and Schroedinger share for 1933 — Heisenberg gets the 1932 grant. New York Times, ??(??):23, November 10, 1933. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/100827281.

Anonymous:1934:ETD

[Ano34a] Anonymous. Electron theory of Dirac revised: California scholars dispose of main difficulties but obtain original results. Extend Heisenberg law. Two limitations to experimental knowledge set up by Oppenheimer and Furry. Electron theory of Dirac revised. New York Times, ??(??):n1, February 18, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1934:PDA

[Ano34b] Anonymous. Prof. Dirac appointed: Nobel Prize winner in physics gets advance study institute post. New York Times, ??(??):19, March 16, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1951:NPH

[Ano51] Anonymous. Der Nobelpreis für Physik: Heisenberg–Schrödinger 1932 Dirac 1933/Chadwick 1935/1936 Hess–Anderson. Physikalische Blätter, 7(9):404–409, September 1951. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19510070904/abstract.

Anonymous:1952:MPM

[Ano52] Anonymous. Max-Planck-Medaille an Paul Adrien Maurice Dirac. Physikalische Blätter, 8(11):512–513, 1952. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1954:DDV

[Ano54] Anonymous. Dirac denied visa. Physics Today, 7(7):7, July 1954. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1955:SRV

[Ano55] Anonymous. Some recent visa experiences. Bulletin of the Atomic Scientists, 11(10):368–371, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:RDG

[Ano62] Anonymous, editor. Recent Developments in General Relativity. Pergamon Press, London, UK, 1962. 472 pp. LCCN QC 6 .R295. This book is dedicated to Leopold Infeld in connection with his 60th birthday.

Anonymous:1965:NLI

[Ano65] Anonymous, editor. Nobel lectures: including presentation speeches and laureates’ biographies: physics 1922–1941. Elsevier, Amsterdam, The Netherlands, 1965. x + 456 pp. LCCN QC71

.P455 1965. URL http://nobelprize.org/nobel_prizes/
physics/laureates/1922/bohr-lecture.html;   http://nobelprize.
org/nobel_prizes/physics/laureates/1922/bohr-lecture.
pdf.

Anonymous:1967:CDD

[Ano67] Anonymous. Correspondence of Dennison, Dirac, the Spectroscopic Notation Committe and Möller. ????, American Philosophical Society, Philadelphia, PA, USA, 1967. ?? pp.

Anonymous:1975:DGC

[Ano75] Anonymous. Quantum electrodynamics: Translations of four papers by Heisenberg, Dirac, and Weisskopf. Technical Report RL-75-073, Science and Technology Facilities Council ePublication Archive, Chilbolton, Daresbury, and Rutherford Appleton Laboratories, UK, 1975. URL http://epubs.cclrc.ac.

uk/work-details?w=21197.

Anonymous:1983:PPAa

[Ano83a] Anonymous. Photo of Paul Adrien Maurice Dirac. Foundations of Physics, 13(2):ii, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889478.

Anonymous:1983:PPAb

[Ano83b] Anonymous. Photo of Paul Adrien Maurice Dirac. Foundations of Physics, 13(3):i, March 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01906177.

Anonymous:1985:MEM

[Ano85] Anonymous. Medal established in memory of P. A. M. Dirac. Physics Today, 38(8):64, August 1985. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

scitation.aip.org/content/aip/magazine/physicstoday/
article/38/8/10.1063/1.2814664.

Anonymous:1994:VID

[Ano94] Anonymous. Vignette: Interview with Dirac. Science, 266 (5192):1889, December 16, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:2000:GVA

[Ano00] Anonymous. Geometrical versus algebraical thinking. Physics in Perspective (PIP), 2(4):453, December 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2002:DP

[Ano02] Anonymous. Dirac posters. Web site, 2002. URL http://

web.archive.org/web/20040107032703/http://education.
iop.org/schools/supteach/dirac.html.

Anonymous:2005:CLB

[Ano05] Anonymous. Critique de livre: Il était sept Fois La révolution: Albert Einstein et les autres. La Recherche, ??(386):11–??, May 1, 2005. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (electronic). URL http:/

/www.larecherche.fr/idees/livres/il-etait-sept-fois-
revolution-albert-einstein-autres-01-05-2005-73924.

Anonymous:2008:DVM

[Ano08] Anonymous. Dirac vs Majorana neutrinos: Low energy tests. AIP Conference Proceedings, 72:21–23, 2008. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic),

1935-0465. URL http://proceedings.aip.org/resource/2/
apcpcs/72/1/21_1.

Anonymous:2010:PPP

[Ano10] Anonymous. Paul A. M. Dirac international scholarly certificates: 1914–1979 (inclusive). Paul A. M. Dirac Papers, 1788–

1999., 2010. URL http://purl.fcla.edu/fsu/lib/digcoll/
dirac/dirac-papers/certificates.

Anonymous:2012:PPC

[Ano12] Anonymous. Pinpointing Planck’s constant with GPS. Web site,

March 15, 2012. URL http://physics.aps.org/synopsis-
for/10.1103/PhysRevLett.108.110801. See [KM12].

Anonymous:2013:ETS

[Ano13] Anonymous. Emeritus trio scoops the 2013 Dirac Medal. Physics World, 26(8):??, August 8, 2013. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworld.com/cws/article/news/2013/
aug/08/emeritus-trio-scoops-the-2013-dirac-medal.

Anonymous:2018:NHL

[Ano18] Anonymous. New half-light half-matter particles may hold the key to a computing revolution. R&D Magazine, ??(??):??, Oc-

tober 11, 2018. URL https://www.rdmag.com/news/2018/10/
new-half-light-half-matter-particles-may-hold-key-computing-
revolution.

Anonymous:2024:HCM

[Ano24a] Anonymous. Holders of the Copley Medal (1731–2020). In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2024.

Anonymous:2024:HOM

[Ano24b] Anonymous. Holders of the Order of Merit (1902–2020). In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2024.

Anonymous:2024:LPM

[Ano24c] Anonymous. Lucasian professors of mathematics at Cambridge University (1664–2020). In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2024.

Anonymous:2024:NPW

[Ano24d] Anonymous. Nobel prize-winners in the Oxford DNB. In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2024.

Anonymous:20xx:PMD

[Anoxx] Anonymous. Paul A. M. Dirac papers, 1788–1999. Web archive,

20xx. URL http://purl.fcla.edu/fsu/lib/digcoll/dirac/

dirac-papers/findingaid. Shelved at Paul A. M. Dirac Science Library.

Antoine:1969:DFSa

[Ant69a] J. P. Antoine. Dirac formalism and symmetry problems in quantum mechanics. I. General Dirac formalism. Journal of Mathematical Physics, 10(1):53–69, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoine:1969:DFSb

[Ant69b] J.-P. Antoine. Dirac formalism and symmetry problems in quantum mechanics. II. Symmetry problems. Journal of Mathematical Physics, 10(12):2276–2290, December 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arai:1993:PDW

[Ara93] Asao Arai. Properties of the Dirac–Weyl operator with a strongly singular gauge potential. Journal of Mathematical Physics, 34(3):915–935, March 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashtekar:1994:AED

[AT94] Abhay Ashtekar and Ranjeet S. Tate. An algebraic extension of Dirac quantization: examples. Journal of Mathematical Physics, 35(12):6434–6470, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Au:1999:TVU

[AT99] Chi Au and Judy Tam. Transforming variables using the Dirac generalized function. The American Statistician, 53(3):270–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annaby:2007:SBC

[AT07] M. H. Annaby and M. M. Tharwat. Sinc-based computations of eigenvalues of Dirac systems. BIT Numerical Mathematics, 47(4):699–713, December 2007. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=47&issue=4&spage=699.

Atag:1989:LIS

[Ata89] Satilmis Atag. Large-N iterative solution of the Dirac equation. Journal of Mathematical Physics, 30(3):696–699, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atiyah:1998:DEG

[Ati98] Michael F. Atiyah. The Dirac equation and geometry. In Goddard [God98], pages 108–124. ISBN 0-521-58382-9 (hardcover), 0-521-01953-2 (paperback). LCCN QC16.D57

P38 1998. URL http://www.loc.gov/catdir/description/
cam028/97022443.html;   http://www.loc.gov/catdir/toc/
cam021/97022443.html.

Augenstein:1995:LAD

[Aug95] Bruno W. Augenstein. A lost alternative to Dirac’s equation. Physics Today, 48(5):86, May 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Aurum:1971:LLD

[Aur71] Aurum. Letter from London: Development and the Kalinga Prize. Bulletin of the Atomic Scientists, 27(3):24, 26–27, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albeverio:1995:RPS

[AW95] S. Albeverio and Zheng Dong Wang. Representation of the propagator and Schwinger functions of Dirac fields in terms of Brownian motions. Journal of Mathematical Physics, 36(10): 5207–5216, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alves:2017:RHJ

[AW17] Waldyr Alves, jun.Rodrigues and Samuel A. Wainer. The relativistic Hamilton–Jacobi equation for a massive, charged and spinning particle, its equivalent Dirac equation and the de Broglie–Bohm theory. Adv. Appl. Clifford Algebr., 27(2):1779– 1799, 2017. ISSN 0188-7009 (print), 1661-4909 (electronic).

Alpher:1981:MV

[AWW^{+}81] Ralph A. Alpher, Paul S. Wesson, Arthur R. Wagner, Paul Wesson, Grote Reber, Harry L. Shipman, R. D. Davies, Paul Wesson, Herbert Malamud, and Halton Arp. More on varying G. Physics Today, 34(3):11–78, March 1981. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/10.1063/1.2914462. See

[Wes80].

Balibar:1985:BEE

[Bal85] Françoise Balibar. Bohr entre Einstein et Dirac. (French) [Bohr between Einstein and Dirac]. Revue d’Histoire des Sciences, 38(3–4):293–307, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic).

URL http://www.persee.fr/web/revues/home/prescript/
article/rhs_0151-4105_1985_num_38_3_4008.

Barlow:1981:APD

[Bar81] Penelope Barlow. An alternative proof to Dirac’s Theorem. Two-Year College Mathematics Journal, 12(1):57–58, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/00494925.1981.11972496.

Barrow:2002:CNA

[Bar02] John D. Barrow. The constants of nature: from Alpha to Omega—the numbers that encode the deepest secrets of the universe. Pantheon Books, New York, NY, USA, 2002. ISBN 0-375-42221-8. xv + 352 pp. LCCN QC39 .B37 2002. URL http://

www.loc.gov/catdir/bios/random051/2002075975.html;
http://www.loc.gov/catdir/description/random042/2002075975.
html;   http://www.loc.gov/catdir/samples/random051/
2002075975.html;   http://www.loc.gov/catdir/toc/fy035/
2002075975.html.

BaselgaMoreno:2008:DBM

[Bas08] Sergio Baselga Moreno. Dirac: la belleza matemática, volume 34 of La matemática en sus personajes. Nivola, Tres Cantos, Spain, 2008. ISBN 84-96566-76-5. 171 pp. LCCN ????

Batelaan:2000:KDE

[Bat00] H. Batelaan. The Kapitza–Dirac effect. Contemporary Physics, 41(6):369–381, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Batelaan:2007:CIK

[Bat07] H. Batelaan. Colloquium: Illuminating the Kapitza–Dirac effect with electron matter optics. Reviews of Modern Physics, 79(3):929–941, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.929;   http://rmp.
aps.org/abstract/RMP/v79/i3/p929_1.

Barut:1985:BSR

[BB85] A. O. Barut and A. J. Bracken. Bound states, resonances, and symmetries of a neutral Dirac particle with anomalous magnetic moment, coupled to a fixed monopole. Journal of Mathematical Physics, 26(6):1390–1396, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baer:2003:PDC

[BB03] Howard Baer and Alexander Belyaev, editors. Proceedings of the Dirac centennial symposium: Florida State university, Tallahassee, USA, 6–7 December 2002. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-412-X. LCCN QC174.45.A1 D57 2003.

Bertotti:1990:MCR

[BBBM90] B. Bertotti, R. Balbinot, S. Bergia, and A. Messina, editors. Modern Cosmology in Retrospect. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37213-5. xx + 426 pp. LCCN QB981 .M774 1990. URL http://www.loc.

gov/catdir/description/cam024/90041803.html;   http:
//www.loc.gov/catdir/toc/cam028/90041803.html.

Bartusiak:1994:PNP

[BBC^{+}94] Marcia Bartusiak, Barbara Burke, Andrew Chaikin, Addison Greenwood, T. A. Heppenheimer, Michelle Hoffman, David Holzman, Elizabeth J. Maggio, and Anne Simon Moffat. A Positron named Priscilla: scientific discovery at the frontier. National Academy Press, Washington, DC, USA, 1994. ISBN 0-309-04893-1. viii + 348 pp. LCCN Q180.55.D57 P67 1994.

URL http://www.nap.edu/catalog.php?record_id=2110.

Blackett:1946:MAE

[BBD^{+}46] P. M. S. Blackett, M. Born, P. I. Dee, P. A. M. Dirac, N. Feather, E. A. Guggenheim, H. S. W. Massey, P. B. Moon, N. F. Mott, M. L. E. Oliphant, F. A. Paneth, R. E. Peierls, M. H. L. Pryce, F. E. Simon, Sir George Thompson, O. R. Frisch, and H. W. B. Skinner. Memo to the UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(12):6–8, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bessis:1985:CFE

[BBR85] N. Bessis, G. Bessis, and D. Roux. Closed-form expressions for the Dirac–Coulomb radial r^{t} integrals. Physical Review A (Atomic, Molecular, and Optical Physics), 32(??):2044–2050, 1985. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Bhagat:2003:EGB

[BBR03] Vikram Bhagat, Ranjan Bhattacharya, and Dhiranjan Roy. On the evaluation of generalized Bose–Einstein and Fermi– Dirac integrals. Computer Physics Communications, 155(1): 7–20, September 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503002947.

Beloy:2010:EST

[BBSF10] K. Beloy, A. Borschevsky, P. Schwerdtfeger, and V. V. Flambaum. Enhanced sensitivity to the time variation of the fine-structure constant and m_{p}/m_{e} in diatomic molecules: a closer examination of silicon monobromide. Physical Review A (Atomic, Molecular, and Optical Physics), 82(2):022106, August 2010. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/doi/

10.1103/PhysRevA.82.022106.

Bracken:1969:CTD

[BC69] A. J. Bracken and H. A. Cohen. On canonical SO(4, 1) transformations of the Dirac equation. Journal of Mathematical Physics, 10(11):2024–2032, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banerjee:1978:ESS

[BC78] S. N. Banerjee and S. N. Chakravorty. On the exact scattering solution of the Schrödinger and Dirac equations with a short range potential. Journal of Mathematical Physics, 19(12):2481– 2484, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Bao:2016:UAM

[BCJT16] Weizhu Bao, Yongyong Cai, Xiaowei Jia, and Qinglin Tang. A uniformly accurate multiscale time integrator pseudospectral method for the Dirac equation in the nonrelativistic limit regime. SIAM Journal on Numerical Analysis, 54(3):1785–1812, ???? 2016. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Benn:1997:DPM

[BCK97] I. M. Benn, Philip Charlton, and Jonathan Kress. Debye potentials for Maxwell and Dirac fields from a generalization of the Killing–Yano equation. Journal of Mathematical Physics, 38(9):4504–4527, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brans:1961:MPR

[BD61] C. Brans and R. H. Dicke. Mach’s principle and a relativistic theory of gravitation. Physical Review, 124(3):925–935, November 1961. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.124.925.

Bethe:1970:ZVF

[BDH70] H. A. Bethe, P. A. M. Dirac, and V. Heisenberg. Zhizn’ v fizike. Vechernie lekcii v Trieste. (Russian) [Life in physics: Evening lectures at Trieste]. Uspekhi Fizicheskikh Nauk, 102 (10):279–312, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/ articles/1970/10/f/. From a Life of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Bethe:1989:LP

[BDH^{+}89] Hans A. (Hans Albrecht) Bethe, P. A. M. Dirac, W. Heisenberg, E. P. Wigner, O. Klein, L. D. Landau, and E. M. Lifshitz, editors. From a life of physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0937-7. ix + 92 pp. LCCN QC71 .F74 1989. URL

http://adsabs.harvard.edu/abs/1989liph.book.....B.

Banuelos:1981:AEF

[BDM81a] Alicia Ba nuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. Analytical expansions for Fermi–Dirac functions. Journal of Mathematical Physics, 22(3):452–455, March 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banuelos:1981:PCF

[BDM81b] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. A program for computing the Fermi–Dirac functions. Computer Physics Communications, 21(3):315–322, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900126.

Banuelos:1984:PCF

[BDM84] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. A program for computing the Fermi–Dirac functions. Computer Physics Communications, 35(1–3):C–670, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828246.

Behncke:1985:SPD

[Beh85] Horst Behncke. Spectral properties of the Dirac equation with anomalous magnetic moment. Journal of Mathematical Physics, 26(10):2556–2559, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bekenstein:1986:FSC

[Bek86] Jacob D. Bekenstein. The fine-structure constant: From Eddington’s time to our own. In Ullmann-Margalit [UM86], pages 209–224. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. LCCN Q174 .B67

vol. 95 Q175. URL http://www.springerlink.com/content/
978-94-009-4566-1;   https://link.springer.com/book/
10.1007/978-94-009-4566-1.

Bell:1973:BRD

[Bel73] John S. Bell. Book review: The Dirac story Aspects of Quantum Theory, by Abdus Salam and E. P. Wigner. Science, 180 (4089):943–944, June 1, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1736041.

Belloni:1982:ZPP

[Bel82] L. Belloni. Zametki o puti, privedshem E. Fermi k statistike Fermi–Diraka. (Russian) [Notes on the path that led Fermi to Fermi–Dirac statistics]. Uspekhi Fizicheskikh Nauk, 136(1):167– 175, January 1982. CODEN UFNAAG. ISSN 0042-1294 (print),

1996-6652 (electronic). URL http://ufn.ru/ru/articles/
1982/1/f/.

Belinfante:1985:CDM

[Bel85] Frederik J. Belinfante. On the covariance of Dirac’s modified canonical commutation relations for fermions interacting with gravity and bosons. Journal of Mathematical Physics, 26(11): 2827–2835, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belinfante:1991:CDA

[Bel91] F. J. Belinfante. Computing Dirac’s atomic hydrogen wave functions of the continuum, using summation of mathematically divergent series. Computers in Physics, 5(3):319–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.
168410.

Belloni:1994:FRF

[Bel94] Lanfranco Belloni. On Fermi’s route to Fermi–Dirac statistics. European Journal of Physics, 15(3):102–109, May 1994. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL

http://iopscience.iop.org/0143-0807/15/3/002.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13 .B45 1999. URL http://www.loc.

gov/catdir/description/uchi052/99035499.htm;   http:
//www.loc.gov/catdir/toc/uchi052/99035499.htm.

Bennewitz:1934:NBM

[Ben34] ?. Bennewitz. Neue Bücher: Die moderne Atomtheorie. W. Heisenberg; E. Schrödinger; P. A. M. Dirac. Verlag S. Hirzel, Leipzig 1934. Preis geh. RM. 2,50. Angewandte Chemie, 47(37): 655, 1934. CODEN ANCEAD. ISSN 1521-3757.

Bernstein:2009:SUM

[Ber09] Jeremy Bernstein. SWU for U and me. arXiv.org, ??(??):1–65,

June 13, 2009. URL https://arxiv.org/abs/0906.2505.

Betounes:1983:DTD

[Bet83] David E. Betounes. Dirac tensor distributions for moving submanifolds of R^{n}. Journal of Mathematical Physics, 24(11):2566– 2572, November 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berengut:2012:CGP

[BF12] J. C. Berengut and V. V. Flambaum. Comment on “Global Positioning System Test of the Local Position Invariance of Planck’s Constant”. Physical Review Letters, 109(6):068901:1, August 2012. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://journals.aps.org/prl/
abstract/10.1103/PhysRevLett.109.068901. See [KM12].

Bowlin:1990:CSC

[BG90] James B. Bowlin and Alfred S. Goldhaber. Classical scattering of a charged particle on an extended magnetic monopole. Journal of Mathematical Physics, 31(9):2305–2313, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v31/i9/p2305_s1.

Boersma:1991:AEC

[BG91] J. Boersma and M. L. Glasser. Asymptotic expansion of a class of Fermi–Dirac integrals. SIAM Journal on Mathematical Analysis, 22(3):810–820, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1982:BEP

[BH82] Laurie M. Brown and Lillian Hoddeson. The birth of elementaryparticle physics. Physics Today, 35(4):36–43, April 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v35/i4/p36_s1. This is a much-abridged version of [BH83a].

Brown:1983:BEP

[BH83a] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. In The Birth of Particle Physics [BH83b], pages 3–36. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Brown:1983:BPP

[BH83b] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Born:1934:DEO

[BI34] Max Born and Leopold Infeld. Deduction de l’équation d’ondes de Dirac à partir de l’électrodynamique quantique. (French) [Deduction of the Dirac wave equation from quantum electrodynamics]. Comptes Rendus Hebdomadaires des Séances de l’Academie des Sciences, 199:1596–1598, November–December 1934. ISSN 0001-4036.

Balachandran:2003:DOC

[BILP03] A. P. Balachandran, Giorgio Immirzi, Joohan Lee, and Peter Presnajder. Dirac operators on coset spaces. Journal of Mathematical Physics, 44(10):4713–4735, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Binney:2000:ODS

[Bin00] James Binney. Obituary: Dennis Sciama 1926–1999. Physics World, 13(2):11, February 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

stacks.iop.org/2058-7058/13/i=2/a=13.

Brzezniak:2003:RCI

[BJ03] Zdzislaw Brzezniak and Brian Jefferies. Renormalization of Coulomb interactions for the 1D Dirac equation. Journal of Mathematical Physics, 44(4):1638–1659, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barut:1976:SDE

[BK76] A. O. Barut and J. Kraus. Solution of the Dirac equation with Coulomb and magnetic moment interactions. Journal of Mathematical Physics, 17(4):506–508, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bairamov:2010:EDD

[BK10] Elgiz Bairamov and Turhan Koprubasi. Eigenparameter dependent discrete Dirac equations with spectral singularities. Applied Mathematics and Computation, 215(12):4216–4220, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Blank:2011:SMH

[Bla11] Brian E. Blank. The strangest man: the hidden life of Paul Dirac, mystic of the atom [book review of MR2807927]. Notices of the American Mathematical Society, 58(9):1278–1284, October 2011. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bleick:1976:M

[Ble76] Willard E. Bleick. Monopole in 1927. Physics Today, 29(8): 68, August 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bleyer:1986:BRA

[Ble86] Ulrich Bleyer. Book review: Asim O. Barut, Alwyn Van der Merwe and Jean-Pierre Vigier (eds.): Quantum, Space and Time — The Quest Continues. Studies and Essays in Honour of Louis de Broglie, Paul Dirac and Eugene Wigner. Cambridge University Press, Cambridge 1984, ISBN 0-521-31911-0, Preis: £25. Astronomische Nachrichten, 307(3), 1986. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Berrondo:1970:DDE

[BM70] M. Berrondo and H. V. McIntosh. Degeneracy of the Dirac equation with electric and magnetic Coulomb potentials. Journal of Mathematical Physics, 11(1):125–140, January 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barrow:1998:VTS

[BM98] John D. Barrow and João Magueijo. Varying-α theories and solutions to the cosmological problems. Physics Letters B, 443(1?4):104–110, 1998. CODEN PYLBAJ. ISSN 0370-2693 (print), 1873-2445 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0370269398012945.

Benham:2014:CPI

[BMP14] Richard Benham, Chris Mortensen, and Graham Priest. Chunk and permeate III: the Dirac delta function. Synthese, 191 (13):3057–3062, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/s11229-014-0473-7;
 http://link.springer.com/content/pdf/10.1007/s11229-
014-0473-7.pdf.

Barrow:2002:CTT

[BMS02a] John D. Barrow, João Magueijo, and Håvard B. Sandvik. A cosmological tale of two varying constants. Physics Letters B, 541(3–4):201–210, 2002. CODEN PYLBAJ. ISSN 0370-2693 (print), 1873-2445 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0370269302022335.

Barrow:2002:VAS

[BMS02b] John D. Barrow, João Magueijo, and Håvard Bunes Sandvik. Variations of alpha in space and time. Physical Review D (Particles and Fields), 66(4):043515, August 15, 2002. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.66.043515.

Beers:1972:ASD

[BN72] B. Beers and H. H. Nickle. Algebraic solution for a Dirac electron in a plane-wave electromagnetic field. Journal of Mathematical Physics, 13(10):1592–1595, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Badino:2013:RPH

[BN13] Massimiliano Badino and Jaume Navarro. Research and pedagogy: a history of quantum physics through its textbooks, volume 2 of Max Planck research library for the history and development of knowledge. Studies. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5871-X. 302 pp. LCCN QC173.98.

URL http://www.edition-open-access.de/studies/2/.

Barcelos-Neto:1994:SAD

[BNB94] J. Barcelos-Neto and N. R. F. Braga. Symplectic analysis of a Dirac constrained theory. Journal of Mathematical Physics, 35 (7):3497–3503, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagrov:1992:NMI

[BO92] V. G. Bagrov and V. V. Obukhov. New method of integration for the Dirac equation on a curved space–time. Journal of Mathematical Physics, 33(6):2279–2289, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bokulich:2004:OCD

[Bok04] Alisa Bokulich. Open or closed? Dirac, Heisenberg, and the relation between classical and quantum mechanics. Studies in History and Philosophy of Modern Physics, 35(3):377–396, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000139.

Bokulich:2008:PDE

[Bok08] Alisa Bokulich. Paul Dirac and the Einstein–Bohr debate. Perspectives on Science, 16(1):103–114, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL

http://muse.jhu.edu/journals/perspectives_on_science/
v016/16.1bokulich.html;   http://muse.jhu.edu/journals/
perspectives_on_science/v016/16.1bokulich.pdf;   http:
//www.mitpressjournals.org/doi/abs/10.1162/posc.2008.
16.1.103.

Born:1949:SPT

[Bor49] Max Born. Le second principe de la thermodynamique déduit de la théorie des quanta. (French) [The Second Law of Thermodynamics derived from quantum theory]. Annales de l’Institut Henri Poincaré, 11(1):47, 1949. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic).

Born:1952:DNT

[Bor52] Max Born. Dirac’s new theory of the electron. Nature, 169 (4313):1105, June 28, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Borek:1985:RCA

[Bor85] R. Borek. Representations of the current algebra of a charged massless Dirac field. Journal of Mathematical Physics, 26(2): 339–344, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bose:1924:PGL

[Bos24] Satyendra Nath Bose. Plancks Gesetz und Lichtquantenhypothese. (German) [Planck’s law and the light-quantum hypothesis]. Zeitschrift für Physik, 26(1):178–181, ???? 1924. CODEN

ZEPYAA. ISSN 0044-3328. URL http://link.springer.com/
article/10.1007/BF01327326.

Bagrov:1996:IED

[BOS96] V. G. Bagrov, V. V. Obukhov, and A. G. Sakhapov. Integration of the Einstein–Dirac equations. Journal of Mathematical Physics, 37(11):5599–5610, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boudet:1985:CLD

[Bou85] Roger Boudet. Conservation laws in the Dirac theory. Journal of Mathematical Physics, 26(4):718–724, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boudet:2001:TIP

[Bou01] Roger Boudet. La théorie intrinsèque de la particule de Dirac et “l’École Louis de Broglie”. (French) [The intrinsic-particle theory of Dirac and ‘the School of Louis de Broglie’]. Annales de la Fondation Louis de Broglie, 26(1-3):95–114, 2001. ISSN 0182-4295 (print), 2108-6397 (electronic).

Boussaid:2008:ASS

[Bou08] Nabile Boussaid. On the asymptotic stability of small nonlinear Dirac standing waves in a resonant case. SIAM Journal on Mathematical Analysis, 40(4):1621–1670, ???? 2008. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buckley:1979:IPM

[BP79a] P. Buckley and F. D. Peat. Interview with P. A. M. Dirac. In Buckley and Peat [BP79b], pages 34–40. ISBN 0-7100-

0313-7. LCCN QC24.5. URL http://www.fdavidpeat.com/
interviews/dirac.htm.

Buckley:1979:QPC

[BP79b] Paul Buckley and F. David Peat, editors. A question of physics: conversations in physics and biology. Routledge and Kegan Paul, London, UK, 1979. ISBN 0-7100-0313-7. x + 159 pp. LCCN QC24.5.

Barbour:1995:MPN

[BP95] Julian B. Barbour and Herbert Pfister, editors. Mach’s principle: from Newton’s bucket to quantum gravity, volume 6 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1995. ISBN 0-8176-3823-7, 3-7643-3823-7 (paperback). vii + 536 pp. LCCN QC137 .M33 1995.

URL http://www.gbv.de/dms/bowker/toc/9780817638238.
pdf;   http://www.zentralblattmath.org/zmath/en/search/
?an=0846.01008.

Bohr:1964:NBH

[BR64] Niels Bohr and S. (Stefan) Rozental, editors. Niels Bohr: hans liv og virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: his life and works told by a circle of friends and colleagues]. J. H. Schultz, København, Danmark, 1964. 341 pp. LCCN QC16.B63 N5. English translation in [Roz67].

Brown:1987:PDW

[BR87] Laurie M. Brown and Helmut Rechenberg. Paul Dirac and Werner Heisenberg—a partnership in science. In Kurşunoğlu and Wigner [KW87], pages 117–162. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://adsabs.harvard.

edu/abs/1987ragp.book.....D;   http://www.loc.gov/catdir/
description/cam023/86033409.html;   http://www.loc.gov/
catdir/toc/cam028/86033409.html.

Booth:1997:DME

[BR97] H. S. Booth and C. J. Radford. The Dirac–Maxwell equations with cylindrical symmetry. Journal of Mathematical Physics, 38 (3):1257–1268, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Brian:1995:GTC

[Bri95a] Denis Brian. Genius talk: conversations with Nobel scientists and other luminaries. Plenum Press, New York, NY, USA; London, UK, 1995. ISBN 0-306-45089-5. xii + 423 pp. LCCN fa1941.

Brian:1995:PD

[Bri95b] Denis Brian. Paul Dirac. In Genius talk: conversations with Nobel scientists and other luminaries [Bri95a], pages 61–72. ISBN 0-306-45089-5. LCCN fa1941.

Bridges:2018:BRM

[Bri18] Thomas J. Bridges. Book review: The Many Faces of Maxwell, Dirac and Einstein Equations: A Clifford Bundle Approach. Contemporary Physics, 59(3):323, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bromberg:1977:DQE

[Bro77] J. Bromberg. Dirac’s quantum electrodynamics and the waveparticle equivalence. In Weiner [Wei77], pages 147–157. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Brown:1991:BRBb

[Bro91] Laurie M. Brown. Book review: Dirac: A Scientific Biography by Helge Kragh. Isis, 82(4):769–770, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233386.

Brown:2005:FTN

[Bro05] Laurie M. Brown. Feynman’s Thesis — a New Approach To Quantum Theory. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2005. ISBN 981-256-366-0 (print), 981-256-763-1 (e-book). 142 pp. LCCN QC174.12

.F49 2005. URL http://public.eblib.com/EBLPublic/
PublicView.do?ptiID=244532;   http://www.worldscibooks.
com/promotion/feynman.html#5852.

Brown:2006:PMD

[Bro06] Laurie M. Brown. Paul A. M. Dirac’s The Principles of Quantum Mechanics. Physics in Perspective (PIP), 8(4):381– 407, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-006-0276-4;   http://www.
springerlink.com/content/n5354k7513351301/.

Bahcall:1967:DFS

[BS67] John N. Bahcall and Maarten Schmidt. Does the fine-structure constant vary with cosmic time? Physical Review Letters, 19 (22):1294–1295, November 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

link.aps.org/doi/10.1103/PhysRevLett.19.1294.

Baekler:1988:CSS

[BSW88] P. Baekler, M. Seitz, and V. Winkelmann. Cylindrically symmetric solutions of self-consistently coupled Dirac fields in gauge theories of gravity. Classical and quantum gravity, 5(??):479– 490, 1988. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Beals:1998:DST

[BSW98] Richard Beals, David Sattinger, and Eric Williams. A Dirac sea and thermodynamic equilibrium for the quantized three-wave interaction. Journal of Mathematical Physics, 39(1):1–29, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bulla:1990:FDO

[BT90] W. Bulla and T. Trenkler. The free Dirac operator on compact and noncompact graphs. Journal of Mathematical Physics, 31 (5):1157–1163, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blancarte:1995:HLE

[BtGW95] Herminio Blancarte, Benoî t Grébert, and Ricardo Weder. Highand low-energy estimates for the Dirac equation. Journal of Mathematical Physics, 36(3):991–1015, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bucciarelli:1980:EGC

[Buc80] B. Bucciarelli, editor. Einstein, Galileo: commemoration d’Albert Einstein, 1979, Pontificia academia scientiarum, Collana Scienza e fede. Pontificia Accademia delle scienze, Libreria editrice vaticana, Vatican City, 1980. URL http://books.

google.com/books?id=pnpkYgEACAAJ.

Bueno:2005:DDM

[Bue05] Otávio Bueno. Dirac and the dispensability of mathematics. Studies in History and Philosophy of Modern Physics, 36(3):465–490, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219805000377.

Buneman:1952:CFE

[Bun52a] O. Buneman. Circulation in the flow of electricity: Dirac’s new variables. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 215(1122):346–352, December 1952. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Bunge:1952:IAH

[Bun52b] Mario Augusto Bunge. Interpretación del átomo de hidrógeno en la teoría de Dirac. (Spanish) [Interpretation of the hydrogen atom in the Dirac theory]. Doctor en ciencias físicas, Facultad de Ciencias Exactas y Naturales, Universidad de Buenos Aires, Buenos Aires, Argentina, 1952. 121 pp.

URL http://catalogosuba.sisbi.uba.ar/vufind/Record/
20160331012339142.

Bunge:1982:CBP

[Bun82] Mario Bunge. Is chemistry a branch of physics? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):209–223, ???? 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170620.

Bussey:2010:QLD

[Bus10] Dr. P. J. Bussey. Quantum Leap: From Dirac and Feynman, Across the Universe to Human Body and Mind, by V. Ivancevic and T. Ivancevic, Scope: monograph collection. Level: graduate student and professional research worker. Contemporary Physics, 51(2):185–186, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burnel:1985:DQT

[BV85] A. Burnel and M. Van Der Rest-Jaspers. Dirac quantization of a three-dimensional gauge theory. Journal of Mathematical Physics, 26(12):3155–3159, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barut:1983:PDH

[BvdM83] Asim O. Barut and Alwyn van der Merwe. Paul Dirac on his eightieth birthday. Foundations of Physics, 13(2):187– 188, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889479.

Barut:1984:QST

[BVV84] A. O. (Asim Orhan) Barut, Alwyn Van der Merwe, and Jean-Pierre Vigier, editors. Quantum, space, and time–the quest continues: studies and essays in honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-31911-0 (paperback). vii + 659 pp. LCCN QC174.125 .Q376 1984.

Canarutto:1998:PDT

[Can98] Daniel Canarutto. Possibly degenerate tetrad gravity and Maxwell–Dirac fields. Journal of Mathematical Physics, 39(9): 4814–4823, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cardoso:1992:BFE

[Car92] J. G. Cardoso. Bessel function expressions for Dirac fields. Journal of Mathematical Physics, 33(3):1109–1113, March 1992. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cardoso:1994:TDP

[Car94] J. G. Cardoso. A twistorial description of the processes of mass scattering of Dirac fields. Journal of Mathematical Physics, 35 (1):294–309, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carilli:2001:WCC

[Car01] Chris L. Carilli. When constants are not constant: The recent finding by an international team of astronomers that the finestructure constant has increased over time may support the idea that the universe has extra dimensions. Physics World, 14(10): 26–27, October 2001. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.org/

2058-7058/14/i=10/a=27.

Casimir:1985:PD

[Cas85] Hendrik B. G. Casimir. Paul Dirac 1902–1984. Naturwissenschaftliche Rundschau, 38(6):219–223, ???? 1985. CODEN NARSAC. ISSN 0028-1050.

Cassidy:2010:BRG

[Cas10] David C. Cassidy. Book review: Graham Farmelo, The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom. Isis, 101(3):661, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/657209.

Cawley:1969:HDT

[Caw69] R. G. Cawley. Hamilton–Dirac theory of Hamilton’s equations. Journal of Mathematical Physics, 10(5):928–932, May 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Cho:1990:CAQ

[CDK90] H. T. Cho, Adel Diek, and R. Kantowski. A Clifford algebra quantization of Dirac’s electron positron field. Journal of Mathematical Physics, 31(9):2192–2200, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carter:2007:SEE

[CG07] Rebecca Carter and Michael B. Giles. Sharp error estimates for discretizations of the 1D convection–diffusion equation with Dirac initial data. IMA Journal of Numerical Analysis, 27(2):406–425, April 2007. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http://

imajna.oxfordjournals.org/cgi/content/abstract/27/2/
406;   http://imajna.oxfordjournals.org/cgi/reprint/
27/2/406.

Chechev:1972:DPV

[CGK72] V. P. Chechev, L. E. Gurevich, and Ya. M. Kramarovsky. Dirac’s principle and variability of fundamental constants. Physics Letters B, 42(2):261–263, November 27, 1972. CODEN PYLBAJ. ISSN 0370-2693 (print), 1873-2445 (electronic).

Christodoulakis:1994:ESD

[CGKS94] T. Christodoulakis, T. Grammenos, E. Korfiatis, and V. C. Spanos. Exact solutions to the Dirac equation for neutrinos propagating in a particular Vaidya background. Journal of Mathematical Physics, 35(5):2430–2437, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiang:2002:PDE

[CH02] Chun-Ming Chiang and Choon-Lin Ho. Planar Dirac electron in Coulomb and magnetic fields: a Bethe Ansatz approach. Journal of Mathematical Physics, 43(1):43–51, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clark:2016:PMD

[CH16] Kathleen M. Clark and Emmet P. Harrington. The Paul A. M. Dirac papers at Florida State University: a search for informal mathematical investigations. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):205– 214, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17498430.2016.1175143.

Chadam:1972:CPC

[Cha72] J. M. Chadam. On the Cauchy problem for the coupled Maxwell–Dirac equations. Journal of Mathematical Physics, 13 (5):597–604, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chandrasekhar:1973:PMD

[Cha73] S. Chandrasekhar. P. A. M. Dirac on his seventieth birthday. Contemporary Physics, 14(4):389–394, 1973. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cheng:1977:MTS

[Che77] Kuo-Shung Cheng. Monopole theory without string. Journal of Mathematical Physics, 18(4):746–749, April 1977. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v18/i4/p746_s1.

Chow:1981:VGC

[Cho81] T. L. Chow. The variability of the gravitational constant. Lettere al Nuovo Cimento Series 2, 31(4):119–120, 1981. CODEN LNUCAE. ISSN 0375-930X (print), 1827-613X (electronic).

Chernavskii:1978:RAB

[CK78] D. S. Chernavskii and R. Kerner. Remarks on the asymptotic behaviour of the ’t Hooft’s magnetic monopole. Journal of Mathematical Physics, 19(1):287–291, January 1978. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v19/i1/p287_s1.

Cohen:1990:ADD

[CK90] Stuart B. Cohen and Ivan N. Krischner. Approximating the Dirac distribution for Fourier analysis. Journal of Computational Physics, 89(1):255, July 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/002199919090141M.

Cohen:1991:ADD

[CK91] Stuart B. Cohen and Ivan N. Kirschner. Approximating the Dirac distribution for Fourier analysis. Journal of Computational Physics, 93(2):312–324, April 1991. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/002199919190185N.

Cardot:2004:BDE

[CKPT04] Hervé Cardot, Ja-Yong Koo, Heon Jin Park, and Alain Trubuil. Boosting Diracs for electrophoresis. Journal of Computational and Graphical Statistics, 13(3):659–673, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1198/106186004X2615.

Chiba:2007:TVP

[CKYY07] Takeshi Chiba, Tatsuo Kobayashi, Masahide Yamaguchi, and Jun’ichi Yokoyama. Time variation of the proton-electron mass ratio and the fine structure constant with a runaway dilaton. Physical Review D (Particles and Fields), 75(4):043516, February 15, 2007. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.75.043516.

Clarkson:1994:EED

[Cla94] Eric Clarkson. Eigenvalues and eigenfunctions of the Dirac operator on spheres and pseudospheres. Journal of Mathematical Physics, 35(5):2064–2073, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Close:2010:A

[Clo10] Frank E. Close. Antimatter. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-955016-6 (hardcover), 0-19-957887-7 (paperback). x + 166 pp. LCCN QC173.3 .C56 2010.

Crease:1986:SCM

[CM86] Robert P. Crease and Charles C. Mann. The second creation: makers of the revolution in twentieth-century physics. MacMillan Publishing Company, New York, NY, USA, 1986. ISBN 0-02-521440-3. xi + 480 pp. LCCN QC7 .C74 1986.

Conaway:2000:PMD

[CMO00] Charles Wm. Conaway, Joseph R. McElrath, Jr., and Billie B. Oakes. The Paul A. M. Dirac Collection: a finding aid to his collection in the Paul A. M. Dirac Library at Florida State University. Report, Florida State University, Tallahassee, FL 32306-4140, 2000. 65 pp. With the technical assistance of John R. Nemmers and Burton H. Altman.

Carvalho:2003:NVD

[CO03] Marcelo Carvalho and Alexandre L. Oliveira. A new version of the Dirac’s Æther and its cosmological applications. Foundations of Physics Letters, 16(3):255–263, June 2003. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cohn:1970:CLD

[Coh70] J. Cohn. Considerations on the Lorentz–Dirac equation. Journal of Mathematical Physics, 11(1):294–295, January 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colin:2012:RQE

[Col12] Samuel Colin. Relaxation to quantum equilibrium for Dirac fermions in the de Broglie–Bohm pilot-wave theory. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 468(2140):1116–1135, 2012. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Coquereaux:1985:DCD

[Coq85] Robert Coquereaux. Differential characters: the Dirac monopole as an example. Journal of Mathematical Physics, 26(12):3176– 3179, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corbella:1970:ISP

[Cor70] O. D. Corbella. Inverse scattering problem for Dirac particles. Explicit expressions for the values of the potentials and their derivatives at the origin in terms of the scattering and boundstate data. Journal of Mathematical Physics, 11(5):1695–1713, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cordani:1986:KPM

[Cor86] Bruno Cordani. Kepler problem with a magnetic monopole. Journal of Mathematical Physics, 27(12):2920–2921, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v27/i12/p2920_s1.

Cordes:2007:PPD

[Cor07] H. O. (Heinz Otto) Cordes. Precisely predictable Dirac observables, volume 154 of Fundamental theories of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-5168-9 (hardcover), 90-481-7299-3 (softcover). xix + 268 pp. LCCN

QC174.45 .C65 2007. URL http://www.loc.gov/catdir/
enhancements/fy0824/2007425740-d.html;   http://www.
loc.gov/catdir/toc/fy0712/2007425740.html.

Chitre:1968:LVT

[CP68] S. M. Chitre and Yash Pal. Limit on variation of e^{2} with time. Physical Review Letters, 20(6):278–279, February 5, 1968. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.20.278.

Chevreuil:2010:CAD

[CPV10] A. Chevreuil, A. Plastino, and C. Vignat. On a conjecture about Dirac’s delta representation using q-exponentials. Journal of Mathematical Physics, 51(9):093502, September 2010. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i9/p093502_s1.

Clark:2012:LAN

[CR12] Charles W. Clark and Joseph Reader. Light, atoms and nuclei: The optical discovery of deuterium. Optics and Photonics News, 23(5):36–41, May 2012. CODEN OPPHEL. ISSN 1047-6938 (print), 1541-3721 (electronic). URL https://www.osa-opn.

org/home/articles/volume_23/issue_5/features/light,_
atoms_and_nuclei_the_optical_discovery_of/. See also

[RC13].

Crawford:1985:ADB

[Cra85] J. P. Crawford. On the algebra of Dirac bispinor densities: factorization and inversion theorems. Journal of Mathematical Physics, 26(7):1439–1441, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crawford:1988:UEM

[Cra88] Frank S. Crawford. Using Einstein’s method to derive both the Planck and Fermi–Dirac distributions. American Journal of Physics, 56(10):883–885, October 1988. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

ajp.aapt.org/resource/1/ajpias/v56/i10/p883_s1.

Crawford:1993:DOL

[Cra93] James P. Crawford. The Dirac oscillator and local automorphism invariance. Journal of Mathematical Physics, 34(10): 4428–4435, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Critchfield:1976:SPD

[Cri76] C. L. Critchfield. Scalar potentials in the Dirac equation. Journal of Mathematical Physics, 17(2):261–266, February 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crilly:2018:WBP

[Cri18] Tony Crilly. What became of Paul Dirac’s classmate? BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):179–188, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/17498430.2018.
1518508.

Csavinszky:1981:ISS

[Csa81] P. Csavinszky. Introduction of the shell structure with gradient expansion corrections into the Thomas–Fermi–Dirac energy density functional for neutral atoms. International Journal of Quantum Chemistry, 20(S15):387–391, March 8–14, 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Csavinszky:1992:PNI

[Csa92] P. Csavinszky. Precision in the numerical integration of the Thomas–Fermi–Dirac kinetic-energy and exchange-energy functionals using a modeled electron density. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 26(??):371–??, 1992. CODEN IJQSDI. ISSN 0161-3642.

Cody:1967:CRC

[CT67a] W. J. Cody and Henry C. Thacher, Jr. Corrigendum: “Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2, and 3/2”. Mathematics of Computation, 21(99): 525, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.jstor.org/stable/

2003289. See [CT67b].

Cody:1967:RCA

[CT67b] W. J. Cody and Henry C. Thacher, Jr. Rational Chebyshev approximations for Fermi–Dirac integrals of orders −\frac{1}{2}, \frac{1}{2} and \frac{3}{2}. Mathematics of Computation, 21(97):30–40, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/stable/2003468. See

also [CT67a].

Cabrera:1983:MME

[CT83] Blas Cabrera and W. Peter Trower. Magnetic monopoles: Evidence since the Dirac conjecture. Foundations of Physics, 13(2): 195–215, February 1983. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889481.

Cohen-Tannoudji:2012:NDP

[CT12] Gilles Cohen-Tannoudji. Les neutrinos défient les physiciens. (French) [Neutrinos challenge physicists]. La Recherche, ??(466): 38, June 19, 2012. CODEN RCCHBV. ISSN 0029-5671 (print),

1625-9955 (electronic). URL http://www.larecherche.fr/
physique-du-xxie-siecle/neutrinos-defient-physiciens-
01-07-2012-91376.

Crater:1990:ETB

[CV90] H. W. Crater and P. Van Alstine. Extension of two-body Dirac equations to general covariant interactions through a hyperbolic transformation. Journal of Mathematical Physics, 31(8):1998– 2014, August 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cotaescu:2002:HDP

[CV02] Ion I. Cotaescu and Mihai Visinescu. Hierarchy of Dirac, Pauli, and Klein–Gordon conserved operators in Taub–NUT background. Journal of Mathematical Physics, 43(6):2978–2987, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crater:1991:CEN

[CY91] Horace Crater and Dujiu Yang. A covariant extrapolation of the noncovariant two particle Wheeler–Feynman Hamiltonian from the Todorov equation and Dirac’s constraint mechanics. Journal of Mathematical Physics, 32(9):2374–2394, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DAgostino:2002:RND

[D’A02] Salvo D’Agostino. From rational numbers to Dirac’s bra and ket: Symbolic representation of physical laws. Physics in Perspective (PIP), 4(2):216–229, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-002-8364-6.

Dalitz:1987:ASP

[Dal87a] R. H. Dalitz. Another side to Paul Dirac. In Kurşunoğlu and Wigner [KW87], pages 69–92. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Dalitz:1987:BSL

[Dal87b] R. H. Dalitz. A biographical sketch of the life of Professor P. A. M. Dirac, OM, FRS. In Taylor [Tay87b], pages 3–28. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Dalitz:1995:CWP

[Dal95] Richard H. Dalitz, editor. The collected works of P. A. M. Dirac, 1924–1948. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-36231-8. xxiiv + 1310 pp. LCCN QC21.2 .D57 1995.

URL http://adsabs.harvard.edu/abs/1995cwpm.book...
..D;   http://www.loc.gov/catdir/description/cam026/
95001010.html;   http://www.loc.gov/catdir/enhancements/

fy1006/95001010-t.html. Edited and with a preface by R. H. Dalitz.

Darkhosh:1990:DQM

[Dar90] Teymour Darkhosh. Dirac quantization of massive spin-one particles in an external symmetrical tensor field. Journal of Mathematical Physics, 31(9):2201–2203, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Darrigol:1991:AOBa

[Dar91] Olivier Darrigol. Analyses d’ouvrages: Tributes to Paul Dirac par J. G. Taylor. Revue d’Histoire des Sciences, 44(2):258, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23633155.

Darrigol:1992:AOBc

[Dar92] Olivier Darrigol. Analyses d’ouvrages: Dirac : A scientific biography par Helge S. Kragh. Revue d’Histoire des Sciences, 45 (2–3):380–381, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633014.

Darrigol:1996:BRP

[Dar96] Olivier Darrigol. Book review: P. A. M. Dirac and R. H. Dalitz: The Collected Works of P. A. M. Dirac: 1924–1948. Isis, 87(4): 747, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
235248.

Das:1993:GSM

[Das93] A. Das. General solutions of Maxwell–Dirac equations in (1+1)- dimensional space–time and a spatially confined solution. Journal of Mathematical Physics, 34(9):3986–3999, September 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:1996:OBB

[Das96] A. Das. An ongoing big bang model in the special relativistic Maxwell–Dirac equations. Journal of Mathematical Physics, 37 (5):2253–2259, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Datta:1995:NDH

[Dat95] S. N. Datta. Near Dirac–Hartree–Fock results for first-row atoms calculated with GTO basis sets. International Journal of Quantum Chemistry, 56(2):91–??, 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Datta:2004:TDE

[Dat04] Sambhu N. Datta. Transformed Dirac equation for the hydrogen atom, comparison with previous approaches in momentum space, and the anomalous Zeeman effect in momentum representation. International Journal of Quantum Chemistry, 96(1): 42–55, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davis:1936:BRBc

[Dav36] H. T. Davis. Book review: The Principles of Quantum Mechanics by P. A. M. Dirac. Isis, 25(2):493–496, September 1936. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/225395.

Daviau:1998:ATL

[Dav98] Claude Daviau. Application à la théorie de la lumière de Louis de Broglie d’une réécriture de l’équation de Dirac. (French) [Application to Louis de Broglie’s theory of light in a rewriting of the Dirac equation]. Annales de la Fondation Louis de Broglie, 23(3-4):121–127, 1998. ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1932:DVM

[dB32a] Louis de Broglie. Sur les densités de valeurs moyennes dans la théorie de Dirac. (French) [On the densities of mean values in Dirac’s theory]. Comptes rendus de l’Académie des sciences, Paris, 194(??):1062–??, ???? 1932. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:AEE

[dB32b] Louis de Broglie. Sur une analogie entre l’électron de Dirac et l’onde électromagnétique. (French) [On an analogy between Dirac’s electron and the electromagnetic wave]. Comptes rendus de l’Académie des sciences, Paris, 195(??):536–??, ???? 1932. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1933:QRT

[dB33] Louis de Broglie. Quelques remarques sur la théorie de l’électron magnétique de Dirac. (French) [Some remarks on the theory of Dirac’s magnetic electron]. Archives Sc. physiques (5), 15(??): 465–483, 1933.

deBroglie:1934:LMT

[dB34] Louis de Broglie. L’électron magnétique (théorie de Dirac). (French) [The magnetic electron (Dirac theory)]. Hermann et cie, Paris, France, 1934. vii + 315 pp. LCCN ????

URL http://hooke.lib.cam.ac.uk/cgi-bin/bib_seek.cgi?
cat=ul&bib=4782869.

deBroglie:1951:TPS

[dB51] Louis de Broglie. La théorie des particules de spin 1/2 (Électrons de Dirac). (French) [Theory of spin-1/2 particle (Dirac electrons)]. Gauthier-Villars et cie, Paris, France, 1951. ???? pp. LCCN ????

deBroglie:1952:TPS

[dB52a] Louis de Broglie. La théorie des particules de spin 1/2: électrons de Dirac. (French) [The theory of spin-1/2 particles: Dirac electrons]. Gauthier-Villars et cie, Paris, France, 1952. 164 pp. LCCN ????

deBroglie:1952:IIO

[dB52b] Louis de Broglie. Sur l’introduction des idées d’onde-pilote et de double solution dans la théorie de l’électron de Dirac. (French) [On the introduction of pilot-wave and double-solution ideas in Dirac’s electron theory]. Comptes rendus de l’Académie des sciences, Paris, 235(??):557–??, ???? 1952. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1984:QST

[dBDW^{+}84] Louis de Broglie, P. A. M. (Paul Adrien Maurice) Dirac, Eugene Paul Wigner, A. O. (Asim Orhan) Barut, Alwyn Van der Merwe, and Jean-Pierre Vigier, editors. Quantum, space, and time — the quest continues: studies and essays in honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-31911-0 (paperback). vii + 659 pp. LCCN QC174.125 .Q376 1984.

deLange:1989:OAS

[de 89] O. L. de Lange. An operator analysis for the Schrödinger, Klein– Gordon, and Dirac equations with a Coulomb potential. Journal of Mathematical Physics, 30(4):858–866, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLange:1991:SOD

[de 91] O. L. de Lange. Shift operators for a Dirac oscillator. Journal of Mathematical Physics, 32(5):1296–1300, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Debnath:2013:SBP

[Deb13] Lokenath Debnath. A short biography of Paul A. M. Dirac and historical development of Dirac delta function. International Journal of Mathematical Education in Science and Technology, 44(8):1201–1223, December 2013. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Dehlinger:1936:BMQ

[Deh36] U. Dehlinger. Buchbesprechung: Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von De Broglie, Schrödinger, Heisenberg und Dirac. 4. und 5. Aufl. Von Arthur Haas. Leipzig 1934, Akademische Verlagsgesellschaft. 299 S., 7 Abb. Preis geb. RM 7.80. Zeitschrift für Elektrochemie und angewandte physikalische Chemie, 42(1):43, 1936. ISSN 0005-9021.

Dittrich:1989:DOS

[DEŠ89] J. Dittrich, P. Exner, and P. Šeba. Dirac operators with a spherically symmetric δ-shell interaction. Journal of Mathematical Physics, 30(12):2875–2882, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dittrich:1992:DHC

[DEŠ92] J. Dittrich, P. Exner, and P. Šeba. Dirac Hamiltonian with Coulomb potential and spherically symmetric shell contact interaction. Journal of Mathematical Physics, 33(6):2207–2214, June 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dirac:1932:QE

[DFP32] P. A. M. Dirac, V. A. Fock, and Boris Podolsky. On quantum electrodynamics. Physikalische Zeitschrift der Sowjetunion, 2 (??):468–479, 1932. CODEN PHZSAL. ISSN 0369-9811. Russian translation in [Dir03b, Vol. II, pp. 409ff]. Reprinted in [Sch58].

Dirac:1943:AOS

[DFPP43] P. A. M. Dirac, Klaus Fuchs, Rudolf Peierls, and P. Preston. Application to the oblate spheroid hemisphere and oblate hemispheroid. Report number MSD 5, part II., 1943.

Dzuba:1999:CRE

[DFW99] V. A. Dzuba, V. V. Flambaum, and J. K. Webb. Calculations of the relativistic effects in many-electron atoms and space–time variation of fundamental constants. Physical Review A (Atomic, Molecular, and Optical Physics), 59(1):230–237, January 1999. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-

4446, 1538-4519. URL http://link.aps.org/doi/10.1103/
PhysRevA.59.230.

deFalco:1980:LSE

[dG80] Diego de Falco and Francesco Guerra. On the local structure of the Euclidean Dirac field. Journal of Mathematical Physics, 21 (5):1111–1114, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGennes:1997:SID

[dG97] Pierre-Gilles de Gennes. Soft interfaces: the 1994 Dirac Memorial Lectures. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56417-4 (hardback). x + 117 pp. LCCN QC173.4.I57 G46 1997. URL http://cup.es/aus/

catalogue/catalogue.asp?isbn=9780511822322;   http:/
/www.loc.gov/catdir/description/cam028/96045963.html;
 http://www.loc.gov/catdir/toc/cam021/96045963.html.

With an introduction by Sir Sam Edwards.

Dirac:1932:PEA

[DH32a] P. A. M. Dirac and J. W. Harding. Photo-electric absorption in hydrogen-like atoms. Mathematical proceedings of the Cambridge Philosophical Society, 28(2):209–218, ???? 1932. CODEN PCPSA4. ISSN 0008-1981.

Dirac:1932:VPQ

[DH32b] P. A. M. Dirac and W. Heisenberg. Various Problems in Quantum Theory. Kei Mekai, Tokyo, Japan, 1932. ???? pp. LCCN ???? Based on notes of lectures held in Tokyo on September 2–7, 1929. Translated by Y. Nishina, M. Kotani, and T. Inui.

Dirac:1978:DP

[DHS78] Paul A. M. Dirac, H. Hora, and J. R. Shepanski. Directions in physics: lectures delivered during a visit to Australia and New Zealand, August–September 1975. Wiley-Interscience, New York, NY, USA, 1978. ISBN 0-471-02997-1. ix + 95

+ 1 pp. LCCN QC71 .D55. URL http://adsabs.harvard.
edu/abs/1978dpld.book.....D;   http://hooke.lib.cam.
ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1206117. Fore-

word by Sir Mark Oliphant. Lecture 1: The Development of Quantum Mechanics. Lecture 2: Quantum Electrodynamics. Lecture 3: Magnetic Monopoles. Lecture 4: A Positive Energy Relativistic Wave Equation. Lecture 5: Cosmological and Gravitational Constants.

Dicke:1959:NRO

[Dic59] R. H. Dicke. New research on old gravitation: Are the observed physical constants independent of the position, epoch, and velocity of the laboratory? Science, 129(3349):621–624, March 6, 1959. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/129/3349/621.

Dicke:1961:DCM

[Dic61] R. H. Dicke. Dirac’s cosmology and Mach’s principle. Nature, 192(4801):440–441, November 4, 1961. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v192/n4801/abs/
192440a0.html;   http://www.nature.com/nature/journal/
v192/n4801/pdf/192440a0.pdf.

Dicke:1964:TSE

[Dic64] Robert H. (Robert Henry) Dicke, editor. The Theoretical Significance of Experimental Relativity. Documents on modern physics. Gordon and Breach, New York, NY, USA, 1964. xii + 153 pp. LCCN QC6 .D476.

Dingle:1958:AEC

[Din58] R. B. Dingle. Asymptotic expansions and converging factors. III. gamma, psi and polygamma functions, and Fermi–Dirac and Bose–Einstein integrals. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 244(1239): 484–490, April 22, 1958. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/100264.

Dirac:1924:DUT

[Dir24a] P. A. M. Dirac. Dissociation under a temperature gradient. Mathematical proceedings of the Cambridge Philosophical Society, 22(2):132–137, ???? 1924. CODEN PCPSA4. ISSN 0008-1981.

Dirac:1924:NDP

[Dir24b] P. A. M. Dirac. Note on the Doppler principle and Bohr’s frequency condition. Mathematical proceedings of the Cambridge Philosophical Society, 22(3):432–433, May 19, 1924. CODEN PCPSA4. ISSN 0008-1981.

Dirac:1924:NRD

[Dir24c] P. A. M. Dirac. Note on the relativity dynamics of a particle. Philosophical Magazine, 47(282):1158–1159, June 1924. CODEN PHMAA4. ISSN 0031-8086. Russian translation in [Dir03b, Vol. II, pp. 16ff].

Dirac:1925:AHM

[Dir25a] P. A. M. Dirac. The adiabatic hypothesis for magnetic fields. Mathematical proceedings of the Cambridge Philosophical Society, 23(1):69–72, ???? 1925. CODEN PCPSA4. ISSN 0008-1981.

Dirac:1925:AIQ

[Dir25b] P. A. M. Dirac. The adiabatic invariance of the quantum integrals. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 107(744):725–734, April 1, 1925. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/94420.

Dirac:1925:CSE

[Dir25c] P. A. M. Dirac. The conditions for statistical equilibrium between atoms, electrons and radiation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 106(739):581–596, November 1, 1925. CODEN PRLAAZ.

ISSN 0080-4630. URL http://www.jstor.org/stable/94294.

Dirac:1925:ECS

[Dir25d] P. A. M. Dirac. The effect of Compton scattering by free electrons in a stellar atmosphere. Monthly Notices of the Royal Astronomical Society, 85(??):825–832, June 1925. CODEN MN-RAA4. ISSN 0035-8711 (print), 1365-2966 (electronic). URL

http://adsabs.harvard.edu/full/1925MNRAS..85..825D.

Dirac:1925:FEQ

[Dir25e] P. A. M. Dirac. The fundamental equations of quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 109(752):642–653, December 1, 1925. CODEN PRLAAZ. ISSN 0080-4630. URL http:// www.jstor.org/stable/94441. Reprinted in [Dir67a]. Russian translation in [Dir77d].

Dirac:1926:ENQ

[Dir26a] P. A. M. Dirac. The elimination of the nodes in quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 111(757):281–305, May 1, 1926. CODEN PRLAAZ. ISSN 0080-4630. URL http://www. jstor.org/stable/94909. Russian translation in [Dir03b, Vol. II, pp. 91ff].

Dirac:1926:QA

[Dir26b] P. A. M. Dirac. On quantum algebra. Mathematical proceedings of the Cambridge Philosophical Society, 23(4):412–418, 1926. CODEN MPCPCO. ISSN 0305-0041, 1469-8064. Russian translation in [Dir03b, Vol. II, pp. 140ff].

Dirac:1926:TQM

[Dir26c] P. A. M. Dirac. On the theory of quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 112(762):661–667, October 1, 1926. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor. org/stable/94692. Russian translation in [Dir03b, Vol. II, pp. 147ff].

Dirac:1926:QM

[Dir26d] P. A. M. Dirac. Quantum Mechanics. Ph.D. thesis, Cambridge University, Cambridge, UK, June 17, 1926. According to [Far09b, p. 101], this is the first thesis to be submitted anywhere on the subject of quantum mechanics.

Dirac:1926:QMP

[Dir26e] P. A. M. Dirac. Quantum mechanics and a preliminary investigation of the hydrogen atom. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 110 (755):561–579, March 1, 1926. CODEN PRLAAZ. ISSN 0080-

4630. URL http://www.jstor.org/stable/94410. Reprinted

in [Dir67c]. Russian translation in [Dir03b, Vol. II, pp. 72ff].

Dirac:1926:RQM

[Dir26f] P. A. M. Dirac. Relativity quantum mechanics with an application to Compton scattering. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 111 (758):405–423, June 2, 1926. CODEN PRLAAZ. ISSN 0080-

4630. URL http://www.jstor.org/stable/94709. Russian

translation in [Dir03b, Vol. II, pp. 117ff].

Dirac:1927:CEW

[Dir27a] P. A. M. Dirac. The Compton effect in wave mechanics. Mathematical proceedings of the Cambridge Philosophical Society, 23 (5):500–507, 1927. CODEN MPCPCO. ISSN 0305-0041, 1469-8064. Russian translation in [Dir03b, Vol. II, pp. 163ff].

Dirac:1927:PIQ

[Dir27b] P. A. M. Dirac. The physical interpretation of the quantum dynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 113(765):621–641, January 1, 1927. CODEN PRLAAZ. ISSN 0080-4630. URL http:// www.jstor.org/stable/94646. Russian translation in [Dir03b, Vol. II, pp. 171ff].

Dirac:1927:QTD

[Dir27c] P. A. M. Dirac. The quantum theory of dispersion. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 114(769):710–728, May 2, 1927. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/94735.

Dirac:1927:QTE

[Dir27d] P. A. M. Dirac. The quantum theory of the emission and absorption of radiation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 114(767): 243–265, March 1, 1927. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/94746. Russian transla-

tion in [Dir03b, Vol. II, pp. 285ff].

Dirac:1927:QS

[Dir27e] P. A. M. Dirac. Über die Quantenmechanik der Stoßvorgänge. (German) [On the quantum mechanics of collisions]. Zeitschrift für Physik, 44(??):585–595, 1927. CODEN ZEPYAA. ISSN 0044-3328. Russian translation in [Dir03b, Vol. II, pp. 191ff].

Dirac:1928:DFR

[Dir28a] P. A. M. Dirac. Discussion following reports of Bohr and Heisenberg. In ????, editor, Electrons and photons (5th Conseil de physique de l’Institut International de Physique, Solvay, 24–29 October 1927, Bruxelles), pages 258–263. Gauthier-Villars et cie, Paris, France, 1928. LCCN ????

Dirac:1928:QTEa

[Dir28b] P. A. M. Dirac. The quantum theory of the electron. I. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 117(778):610–624, February 1, 1928. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/94981.

Dirac:1928:QTEb

[Dir28c] P. A. M. Dirac. The quantum theory of the electron. II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 118(779):351–361, March 1, 1928. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/94909.

Dirac:1928:QEGa

[Dir28d] P. A. M. Dirac. Über der Quantentheorie des Elektrons. (German) [On the quantum theory of electrons]. Physikalische Zeitschrift, 29(??):561–563, ???? 1928. CODEN PHZTAO. ISSN 0369-982X.

Dirac:1928:QEGb

[Dir28e] P. A. M. Dirac. Zur quantentheorie des elektrons. (German) [On the quantum theory of electrons]. In Hans Falkenhagen, editor, Leipziger Vorträge 1928: Quantentheorie und Chemie, pages 85– 94. Verlag von S. Hirzel, Leipzig, Germany, 1928. LCCN ???? Lecture given in June 1928; manuscript translated to German by A. Eucken.

Dirac:1928:QE

[Dir28f] P. A. M. Dirac. Zur Quantentheorie des Elektrons. (German) [Toward a quantum theory of the electron]. Leipziger Vorträge, ??(??):85–94, ???? 1928.

Dirac:1929:BSQ

[Dir29a] P. A. M. Dirac. The basis of statistical quantum mechanics. Mathematical proceedings of the Cambridge Philosophical Society, 25(1):62–66, 1929. CODEN MPCPCO. ISSN 0305-0041, 1469-8064. Russian translation in [Dir03b, Vol. II, pp. 208ff].

Dirac:1929:QMM

[Dir29b] P. A. M. Dirac. Quantum mechanics of many electron systems. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 123(792):714–733, April 6, 1929. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor. org/stable/95222. Russian translation in [Dir03b, Vol. II, pp. 213ff].

Dirac:1930:AM

[Dir30a] P. A. M. Dirac. Approximate methods. Manuscript in English held at Churchill College Archives (Cambridge, England). This is the text of chapter XIA of the Russian edition of [Dir30f] (M.-L. GTTI, 1932; translated by M. P. Bronshtein, ed. D. D. Iwanenko), pp. 243-257, with the note ’Author’s addition to the Russian edition’ [note from [DP86, p. 179, ref. 31]]., 1930.

Dirac:1930:PQG

[Dir30b] P. A. M. Dirac. Die Prinzipien der Quantenmechanik. (German) [The Principles of Quantum Mechanics]. Verlag von S. Hirzel, Leipzig, Germany, 1930. xi + 274 pp. LCCN ???? Translated by Werner Bloch from the English original.

Dirac:1930:NEP

[Dir30c] P. A. M. Dirac. Note on exchange phenomena in the Thomas atom. Mathematical proceedings of the Cambridge Philosophical Society, 26(3):376–385, 1930. CODEN MPCPCO. ISSN 0305-0041, 1469-8064. Russian translation in [Dir03b, Vol. II, pp. 268ff].

Dirac:1930:NID

[Dir30d] P. A. M. Dirac. Note on the interpretation of the density matrix in the many-electron problem. Mathematical proceedings of the Cambridge Philosophical Society, 27(2):240–243, ???? 1930. CODEN PCPSA4. ISSN 0008-1981. Russian translation in [Dir03b, Vol. II, pp. 279ff].

Dirac:1930:AEP

[Dir30e] P. A. M. Dirac. On the annihilation of electrons and protons. Mathematical proceedings of the Cambridge Philosophical Society, 26(3):361–375, 1930. CODEN MPCPCO. ISSN 0305-0041, 1469-8064.

Dirac:1930:PQM

[Dir30f] P. A. M. Dirac. The Principles of Quantum Mechanics. International series of monographs on physics (Oxford, England). Clarendon Press, Oxford, UK, 1930. x + 257 pp. LCCN QC174.12 .D57 1930. On page 96 of this book, Dirac introduces the notation h̄ = h/(2π), where “h̄ is a new universal constant having the dimensions of action.” That symbol is actually a letter in the Serbian alphabet (Unicode character U045B, Cyrillic small letter tshe), but it is unknown to this bibliographer whether Dirac knew that. It seems more likely that he simply added a stroke across h as a convenient indication of its division by 2π.

Dirac:1930:P

[Dir30g] P. A. M. Dirac. The proton. Nature, 126(3181):605–606, October 18, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v126/n3181/pdf/126605a0.pdf.

Dirac:1930:QPM

[Dir30h] P. A. M. Dirac. Quelques problèmes de mécanique quantique. Annales de l’Institut Henri Poincaré, 1(4):357–400, 1930.

CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (elec-

tronic). URL http://www.numdam.org/item?id=AIHP_1930_

_1_4_357_0. Baed on lectures given at the Institut Henri Poincaré, 13, 14, 19, and 20 December, 1929.

Dirac:1930:TEPa

[Dir30i] P. A. M. Dirac. A theory of electrons and protons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 126(801):360–365, January 1, 1930. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/95359.

Dirac:1930:TEPb

[Dir30j] P. A. M. Dirac. Theory of electrons and protons. Uspekhi Fizicheskikh Nauk, ??(9):581–591, September 1930. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1930/9/a/.

Dirac:1931:PMQ

[Dir31a] P. A. M. Dirac. Les principes de la mécanique quantique. (French) [The principles of quantum mechanics], volume 21. Les Presses Universitaires de France, Paris, France, 1931. viii +

314 pp. URL http://longstreet.typepad.com/books/2011/
03/first-french-edition-of-diracs-quantum-mechanics-

1931.html. Recueil des conférences-rapports de documentation sur la physique. Translated by Al. Proca and J. Ullmo.

Dirac:1931:QSE

[Dir31b] P. A. M. Dirac. Quantised singularities in the electromagnetic field. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 133(821):60–72, September 1, 1931. CODEN PRLAAZ. ISSN 0080-4630. URL http:// www.jstor.org/stable/95639. Russian translation in [Dir03b, Vol. II, pp. 388–??].

Dirac:1932:OKM

[Dir32a] P. A. M. Dirac. Osnovy Kvantovoĭ Mekhaniki. (Russian) [The Elements of Quantum Mechanics]. Gostekhizdat, Moscow and Leningrad, USSR, 1932. ???? pp. LCCN ???? See also [Dir02, Vol. I, pp. 303ff]. Contains author’s supplement to the Russian edition: Priblizhennye metody (Approximate methods).

Dirac:1932:RQM

[Dir32b] P. A. M. Dirac. Relativistic quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 136(829):453–464, May 2, 1932. CODEN PRLAAZ.

ISSN 0080-4630. URL http://www.jstor.org/stable/95782.

Russian translation in [Dir03b, Vol. II, pp. 399ff].

Dirac:1933:HVC

[Dir33a] P. A. M. Dirac. Homogeneous variables in classical dynamics. Mathematical proceedings of the Cambridge Philosophical Society, 29(3):389–400, ???? 1933. CODEN PCPSA4. ISSN 0008-1981. Reprinted in [Sch58]. Russian translation in [Dir03b, Vol. II, pp. 461ff].

Dirac:1933:LQM

[Dir33b] P. A. M. Dirac. The Lagrangian in quantum mechanics. Physikalische Zeitschrift der Sowjetunion, 3(1):64–72, ???? 1933. CODEN PHZSAL. ISSN 0369-9811. URL http://

ebooks.worldscinet.com/ISBN/9789812567635/9789812567635_

0003.html. Reprinted in [Sch58, 312–320], [Dir05a], and [FBD05, 111–119].

Dirac:1933:SPQ

[Dir33c] P. A. M. Dirac. Statement of a problem in quantum mechanics. Journal of the London Mathematical Society, 8(??):274–277, October 1933. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Dirac:1934:DID

[Dir34a] P. A. M. Dirac. Discussion of the infinite distribution of electrons in the theory of the positron. Mathematical proceedings of the Cambridge Philosophical Society, 30(2):150–163, ???? 1934. CODEN PCPSA4. ISSN 0008-1981.

Dirac:1934:TPR

[Dir34b] P. A. M. Dirac. Teoriya pozitrona. (Russian) [Theory of the positron]. In M. P. Bronstein et al., editors, Atomnoe yadro: Sbornik dokladov i vsesoyoznoi yadernoi konferentsii [Leningrad, 24–30 September 1933], pages 129–144. Izdatelstvo, Akademii Nauk SSR, Leningrad, USSR, 1934. LCCN ????

Dirac:1934:TDP

[Dir34c] P. A. M. Dirac. Théorie du positron. (French) [Theory of the positron]. In Cockcroft et al. [CCJ^{+}34], pages 203–230. LCCN ???? Address given at the Solvay Conference, October 22–29, 1933. Russian translation in [Dir03b, Vol. II, pp. 431ff].

Dirac:1934:TEP

[Dir34d] P. A. M. Dirac. Theory of Electrons and Positrons. A. Norstedt och Sönner, Stockholm, Sweden, 1934. 6 pp. LCCN ????

Dirac:1935:EWE

[Dir35a] P. A. M. Dirac. The electron wave equation in de Sitter space. Annals of Mathematics (2), 36(3):657–669, 1935. CO-DEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (elec-

tronic). URL http://www.jstor.org/stable/1968649;
http://www.jstor.org/stable/pdfplus/1968649.pdf.

Dirac:1935:LQE

[Dir35b] P. A. M. Dirac. Lectures on quantum electrodynamics: 1934– 1935. Report, Institute for Advanced Study, Princeton, NJ, USA, 1935. 71 pp. Notes by Boris Podolsky (first semester) and Nathan Rosen (second semester).

Dirac:1935:PQM

[Dir35c] P. A. M. Dirac. The Principles of Quantum Mechanics, volume 27 of International series of monographs on physics (Oxford, England). Clarendon Press, Oxford, UK, second edition, 1935. xi + 300 pp.

Dirac:1936:DCE

[Dir36a] P. A. M. Dirac. Does conservation of energy hold in atomic processes? Nature, 137(3460):298–299, February 22, 1936. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v137/n3460/
pdf/137298a0.pdf. See also [Pei36].

Dirac:1936:RWE

[Dir36b] P. A. M. Dirac. Relativistic wave equations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 155(886):447–459, July 1, 1936. CODEN PRLAAZ.

ISSN 0080-4630. URL http://www.jstor.org/stable/96758.

Dirac:1936:WEC

[Dir36c] P. A. M. Dirac. Wave equations in conformal space. Annals of Mathematics (2), 37(2):429–442, 1936. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://

www.jstor.org/stable/1968455.

Dirac:1937:CVQ

[Dir37a] P. A. M. Dirac. Complex variables in quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 160(900):48–59, May 1, 1937. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/96854.

Dirac:1937:CC

[Dir37b] P. A. M. Dirac. The cosmological constants. Nature, 139 (3512):323, February 20, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

adsabs.harvard.edu/abs/1937Natur.139..323D;   http:/
/www.nature.com/nature/journal/v139/n3512/pdf/139323a0.

pdf. See related work in [Dir38b, Tel48, Wil58, Dys67, Gam67a, Gam67d, Gam68, Alp73].

Dirac:1937:PSP

[Dir37c] P. A. M. Dirac. Physical science and philosophy: Prof. P. A. M. Dirac, F.R.S., St. John’s College, Cambridge. Nature, 139(3528): 1001–1002, June 12, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v139/n3528/pdf/1391001a0.pdf. A reply to

Dr. H. Dingle.

Dirac:1937:ROQ

[Dir37d] P. A. M. Dirac. The reversal operator in quantum mechanics. Izv. Akad. Nauk. SSSR, 4–5(??):569–575 (English), 576–582 (Russian), ???? 1937.

Dirac:1938:CTR

[Dir38a] P. A. M. Dirac. Classical theory of radiating electrons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 167(929):148–169, August 5, 1938. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/97128.

Dirac:1938:NBC

[Dir38b] P. A. M. Dirac. A new basis for cosmology. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 165(921):199–208, April 5, 1938. CODEN PRLAAZ. ISSN 0080-4630. URL http://

adsabs.harvard.edu/abs/1938RSPSA.165..199D;   http:/
/rspa.royalsocietypublishing.org/content/165/921/199;
 http://www.jstor.org/stable/97105. See related work

in [Dir37b, Tel48, Wil58, Dys67, Gam67a, Gam67d, Gam68, Alp73].

Dirac:1939:TED

[Dir39a] P. A. M. Dirac. La théorie de l’électron et du champ électromagnétique. (French) [Theory of the electron and the electromagnetic field]. Annales de l’Institut Henri Poincaré, 9 (2):13–49, 1939. CODEN AIHPA2. ISSN 0365-320x (print),

2400-4855 (electronic). URL http://archive.numdam.org/
article/AIHP_1939__9_2_13_0.pdf.

Dirac:1939:NNQ

[Dir39b] P. A. M. Dirac. A new notation for quantum mechanics. Mathematical proceedings of the Cambridge Philosophical Society, 35 (3):416–418, ???? 1939. CODEN PCPSA4. ISSN 0008-1981. This paper introduces the Dirac bra and ket notation.

Dirac:1939:RBM

[Dir39c] P. A. M. Dirac. The relation between mathematics and physics. Proceedings of the Royal Society of Edinburgh, 59(??):122–129, ???? 1939. CODEN PRSEAE. ISSN 0370-1646. James Scott Prize Lecture, February 6, 1939.

Dirac:1940:ODM

[Dir40] P. A. M. Dirac. Obituary: Dr. M. Mathisson. Nature, 146(3706): 613, November 9, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v146/n3706/pdf/146613b0.pdf.

Dirac:1941:TSI

[Dir41] P. A. M. Dirac. The theory of the separation of isotopes by statistical methods. British Report, 1941.

Dirac:1942:EEE

[Dir42a] P. A. M. Dirac. Estimate of the efficiency of energy release with a non-scattering container. Report number MSD 4., 1942.

Dirac:1942:EUS

[Dir42b] P. A. M. Dirac. Expansion of U-sphere enclosed in a container. Report number MSD 2., 1942.

Dirac:1942:MSF

[Dir42c] P. A. M. Dirac. The motion in a self-fractionating centrifuge. Technical report, HMSO, London, UK, 1942. Report number BR 42, May 1942. Declassified 1946 as Report BDDA 7.

Dirac:1942:NMS

[Dir42d] P. A. M. Dirac. Neutron multiplication in a sphere of uniform density surrounded by a shell of non-uniform density. Report number MSD 3., 1942.

Dirac:1942:PIQ

[Dir42e] P. A. M. Dirac. The physical interpretation of quantum mechanics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 180(980):1–40, March 18, 1942. CODEN PRLAAZ. ISSN 0080-4630. URL http://www. jstor.org/stable/97777. Bakerian Lecture, June 19, 1941. Russian translation in [Dir03b, Vol. II, pp. 587ff].

Dirac:1943:ARN

[Dir43a] P. A. M. Dirac. Approximate rate of neutron multiplication for a solid of arbitrary shape and uniform density. Part 1. General theory. British Declassified Document Association, number 33. Report number MSD 5, part I., 1943. URL http://discovery.

nationalarchives.gov.uk/details/r/C1978994.

Dirac:1943:QE

[Dir43b] Paul A. M. Dirac. Quantum electrodynamics. Communications of the Dublin Institute for Advanced Studies. Series A, Theoretical physics, 1943(1):1–36, 1943. ISSN 0070-7414.

Dirac:1945:AQL

[Dir45a] P. A. M. Dirac. Application of quaternions to Lorentz transformations. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 50(??):261–270, 1945.

CODEN PRIAAK. ISSN 0035-8975. URL http://www.jstor. org/stable/20520646. Russian translation in [Dir03b, Vol. II, pp. 553ff].

Dirac:1945:ABC

[Dir45b] P. A. M. Dirac. On the analogy between classical and quantum mechanics. Reviews of Modern Physics, 17(2–3):195–199, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.17.195.

Dirac:1945:URL

[Dir45c] P. A. M. Dirac. Unitary representations of the Lorentz group. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 183(994):284–295, February 22, 1945. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/97721.

Dirac:1946:DQE

[Dir46] Paul A. M. Dirac. Developments in quantum electrodynamics. Communications of the Dublin Institute for Advanced Studies. Series A, 1946(1):3–33, 1946. ISSN 0070-7414.

Dirac:1947:DQE

[Dir47a] P. A. M. Dirac. The difficulties in quantum electrodynamics. In ????, editor, Report of the International Conference on Fundamental Particles and Low Temperature, July 1946, volume 1, pages 10–14. Physical Society, London, UK, 1947. LCCN ????

Dirac:1947:PQM

[Dir47b] P. A. M. Dirac. The Principles of Quantum Mechanics, volume 27 of International series of monographs on physics (Oxford, England). Clarendon Press, Oxford, UK, third edition, 1947. xii + 311 pp. URL http://adsabs.harvard.

edu/abs/1947pqm..book.....D;   http://hooke.lib.cam.
ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=203807.

Dirac:1948:MP

[Dir48a] P. A. M. Dirac. Magnetic poles. Physics Today, 1(7):26, November 1948. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Dirac:1948:TPE

[Dir48b] P. A. M. Dirac. On the theory of point electrons. Philosophical Magazine, 39(??):31–34, 1948. CODEN PHMAA4. ISSN 0031-8086.

Dirac:1948:QTL

[Dir48c] P. A. M. Dirac. Quantum theory of localizable dynamical systems. Physical Review (2), 73(8):1092–1103, May 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.1092.

Dirac:1948:TMP

[Dir48d] P. A. M. Dirac. The theory of magnetic poles. Physical Review (2), 74(7):817–830, October 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.74.817.

Dirac:1948:QDT

[Dir48e] Paul A. M. Dirac. Quelques déeveloppements sur la théeorie atomique: conféerence faite au Palais de la Déecouverte le 6 déecembre 1945. Université de Paris, Paris, France, 1948. 15 pp. LCCN ????

Dirac:1949:DTF

[Dir49a] P. A. M. Dirac. The dynamical theory of fields, classical and quantum. Lecture notes, Canadian Mathematical Congress: second summer seminar held at the University of British Columbia, 1949, Vancouver, BC, Canada, 1949. 114 pp.

Dirac:1949:FRD

[Dir49b] P. A. M. Dirac. Forms of relativistic dynamics. Reviews of Modern Physics, 21(3):392–399, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.392.

Dirac:1949:SQ

[Dir49c] P. A. M. Dirac. La seconde quantification. (French) [Second quantization]. Annales de l’Institut Henri Poincaré, 11 (1):15–47, 1949. CODEN AIHPA2. ISSN 0365-320x (print),

2400-4855 (electronic). URL http://archive.numdam.org/
article/AIHP_1949__11_1_15_0.pdf.

Dirac:1950:FT

[Dir50a] P. A. M. Dirac. Field theory. In Sir Erneset William Titterton, editor, Proceedings of the Harwell Nuclear Physics Conference, September 1950, pages 114–115. Ministry of Supply, Harwell,

UK, 1950. LCCN ???? URL http://www.abaa.org/books/
251337932.html.

Dirac:1950:GHD

[Dir50b] P. A. M. Dirac. Generalized Hamiltonian dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):129–148, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1950:NMG

[Dir50c] P. A. M. Dirac. A new meaning for gauge transformations in electrodynamics. Il Nuovo Cimento (9), 7(??):925–938, 1950. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Dirac:1951:HFF

[Dir51a] P. A. M. Dirac. The Hamiltonian form of field dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):1–23, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1951:TA

[Dir51b] P. A. M. Dirac. Is there an æther? Nature, 168(4282):906– 907, November 24, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v168/n4282/abs/168906a0.html.

Dirac:1951:NCT

[Dir51c] P. A. M. Dirac. A new classical theory of electrons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 209(1098):291–296, November 7, 1951. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/98908.

Dirac:1951:RCQ

[Dir51d] P. A. M. Dirac. The relation of classical to quantum mechanics. In Proceedings of the Second Canadian Mathematics Congress, Vancouver, 1949, pages 10–31. University of Toronto Press, Toronto, ON, Canada, 1951.

Dirac:1952:TARa

[Dir52a] P. A. M. Dirac. Is there an æther?: Reply to H. Bondi and T. Gold. Nature, 169(4291):146, January 26, 1952. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v169/n4291/
abs/169146b0.html.

Dirac:1952:TARb

[Dir52b] P. A. M. Dirac. Is there an æther?: Reply to Leopold Infeld. Nature, 169(4304):702, April 26, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v169/n4304/
abs/169702b0.html.

Dirac:1952:TJE

[Dir52c] P. A. M. Dirac. Les transformations de jauge en électrodynamique. (French) [Gauge transformations in electrodynamics]. Annales de l’Institut Henri Poincaré, 13(1):1–42, 1952. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic). URL

http://archive.numdam.org/article/AIHP_1952__13_1_1_
0.pdf.

Dirac:1952:NCT

[Dir52d] P. A. M. Dirac. A new classical theory of electrons. II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 212(??):330–339, May 7, 1952. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/98900.

Dirac:1953:SAP

[Dir53a] P. A. M. Dirac. Die Stellung des Aethers in der Physik. (German) [The role of æthers in physics]. Naturwissenschaftliche Rundschau, 6(??):441–446, ???? 1953. CODEN NARSAC. ISSN 0028-1050. Lindau Lecture, July 1, 1953.

Dirac:1953:LTA

[Dir53b] P. A. M. Dirac. The Lorentz transformation and absolute time. Physica, 19(??):888–896, 1953. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic). Proceedings of Lorentz– Kamerlingh Onnes Memorial Conference. Leiden, June 22–26,

Dirac:1953:MGM

[Dir53c] P. A. M. Dirac. Méthodes générales de la mechanique quantique relativiste. In Particules Fondamentales et Noyaux (Paris 24–29 avril 1950), volume 38 of Colloques internationaux du Centre national de la recherche scientifique, pages 27–45. Éditions du Centre national de la recherche scientifique (CNRS), Paris, France, 1953. LCCN ???? URL http:/

/gateway-bayern.de/BV024322243;   http://ipac.ub.tu-
berlin.de/cgi-bin/katalog-link?u0980940.

Dirac:1954:MCE

[Dir54a] P. A. M. Dirac. La mecánica cuántico y el éter. (Spanish) [Quantum mechnanics and the aether]. Ciencia e Investigacion, 10 (??):399–405, ???? 1954. CODEN CIBAAH. ISSN 0009-6733.

Dirac:1954:LB

[Dir54b] P. A. M. Dirac. Logic or beauty? The Scientific Monthly, 79(4): 268–269, October 1954. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/21106.

Dirac:1954:NCT

[Dir54c] P. A. M. Dirac. A new classical theory of electrons. III. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 223(1155):438–445, May 20, 1954. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/99450.

Dirac:1954:QMA

[Dir54d] P. A. M. Dirac. Quantum mechanics and the æther. The Scientific Monthly, 78(3):142–146, March 1954. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://

www.jstor.org/stable/20945.

Dirac:1955:GIF

[Dir55a] P. A. M. Dirac. Gauge-invariant formulation of quantum electrodynamics. Canadian Journal of Physics = Journal canadien de physique, 33(11):650–660, November 1955. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Dirac:1955:LQM

[Dir55b] P. A. M. Dirac. Lectures on quantum mechanics and relativistic field theory. Technical Report 1, Tata Institute of Fundamental Research, Bombay, India (??), 1955.

Dirac:1955:NUN

[Dir55c] P. A. M. Dirac. Note on the use of non-orthogonal wave functions in perturbation calculations. Canadian Journal of Physics = Journal canadien de physique, 33(12):709–712, December 1955. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Dirac:1955:QMR

[Dir55d] P. A. M. Dirac. Quantum mechanics and relativistic field theory. Listed in [CMO00, p. 48] without publication information., 1955.

Dirac:1955:STF

[Dir55e] P. A. M. Dirac. The stress tensor in field dynamics. Il Nuovo Cimento (10), 1(??):16–36, 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Dirac:1955:SAW

[Dir55f] P. A. M. Dirac. Symmetry in the atomic world. Journal of Scientific and Industrial Research, 14A(??):153–155, ???? 1955. CO-DEN JSIRAC. ISSN 0022-4456 (print), 0975-1084 (electronic). Address given at the Indian Science Congress, January 1955.

Dirac:1957:EV

[Dir57a] P. A. M. Dirac. Elektroni i vakuum. Vseoiuz. o-vo po rasprostraneniiu polit. i nauch. znanii, 8(37):15–??, ???? 1957.

Dirac:1957:VQE

[Dir57b] P. A. M. Dirac. The vacuum in quantum electrodynamics. Nuovo Cimento, Suppl., X. Ser., 6(supplemento):322–339, 1957. CO-DEN NUCUAF. ISSN 0550-3868.

Dirac:1958:EV

[Dir58a] P. A. M. Dirac. Electrons and vacuum. Fizika v Skole, 2(??): 20–32, ???? 1958. Lecture given in Moscow on 9 October 1956.

Dirac:1958:GHD

[Dir58b] P. A. M. Dirac. Generalized Hamiltonian dynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246(1246):326–332, August 19, 1958. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/100496.

Dirac:1958:PQM

[Dir58c] P. A. M. Dirac. The Principles of Quantum Mechanics. International Series of Monographs on Physics. Clarendon Press, Oxford, UK, fourth edition, 1958. xii + 312 pp.

Dirac:1958:TGH

[Dir58d] P. A. M. Dirac. The theory of gravitation in Hamiltonian form. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 246(1246):333–343, August 19, 1958. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/100497.

Dirac:1959:EWE

[Dir59a] P. A. M. Dirac. The electron wave equation in Riemannian space. In B. Kockel, W. Macke, and A. Papapetrou, editors, Max-Planck-Festschrift 1958, pages 339–344. VEB Deutscher Verlag der Wissenschaften, Berlin, East Germany, 1959.

Dirac:1959:EVR

[Dir59b] P. A. M. Dirac. Elektrony a vakuum. (Czech) [Vacuum electrons]. Prokroky Matematiky Fysiky a Astronomie, 4(??):309– 317, ???? 1959.

Dirac:1959:EGF

[Dir59c] P. A. M. Dirac. Energy of the gravitational field. Physical Review Letters, 2(8):368–371, April 1959. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/10.1103/PhysRevLett.2.368.

Dirac:1959:FCH

[Dir59d] P. A. M. Dirac. Fixation of coordinates in the Hamiltonian theory of gravitation. Physical Review (2), 114(3):924–930, May 1959. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.114.924.

Dirac:1959:PQM

[Dir59e] P. A. M. Dirac. The Principles of Quantum Mechanics. International Series of Monographs on Physics. Clarendon Press, Oxford, UK, fourth (reprinted) edition, 1959. xii + 312 pp.

Dirac:1959:SSR

[Dir59f] P. A. M. Dirac. Sovremennoe sostoyanie relyativistskoi teorii electrona. (Russian) [The present situation in the relativistic theory of electrons]. Trudy Instituta Istorii Estestvoznanija i Techniki, Akademiya Nauk SSSR, 22(??):32–33, ???? 1959. CO-DEN TIETA4. ISSN 0568-5826.

Dirac:1960:GGGb

[Dir60a] P. A. M. Dirac. Gravitationswellen (German) [Gravitational waves]. Naturwissenschaftliche Rundschau, 13(??):165–168, ???? 1960. CODEN NARSAC. ISSN 0028-1050. Unabridged version of Lindau lecture given at the 9th Lindau Meeting, June 29–July 3, 1960.

Dirac:1960:GGGa

[Dir60b] P. A. M. Dirac. Gravitationswellen: Nach einem Vortrag (German) [Gravitational waves: After a lecture]. Physikalische Blätter, 16(7):364–366, July 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/phbl.19600160703/

abstract. Abridged version of Lindau lecture given at the 9th Lindau Meeting, June 29–July 3, 1960.

Dirac:1960:PKM

[Dir60c] P. A. M. Dirac. Printsipy kvantovoi mekhaniki. Translated from the 4th English ed. by Yu. N. Demkov and G. F. Drukarev; edited and with a preface by V. A. Fok. Gosudarstv. Izdat. Fis.- Mat. Lit., Moscow, 1960. 434 pp.

Dirac:1960:RBI

[Dir60d] P. A. M. Dirac. A reformulation of the Born–Infeld electrodynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 257(1288):32–43, August 23, 1960. CODEN PRLAAZ. ISSN 0080-4630. URL http://

www.jstor.org/stable/2413797.

Dirac:1961:DCM

[Dir61a] P. A. M. Dirac. Dirac’s cosmology and Mach’s Principle [reply to R. H. Dicke]. Nature, 192(4801):441, November 4, 1961. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://adsabs.harvard.edu/abs/1961Natur.192..441D.

Dirac:1961:OPE

[Dir61b] P. A. M. Dirac. Obituary: Prof. Erwin Schrödinger, For. Mem. R.S. Nature, 189(4762):355–356, February 4, 1961. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v189/n4762/
pdf/189355a0.pdf.

Dirac:1962:CQF

[Dir62a] P. A. M. Dirac. The conditions for a quantum field theory to be relativistic. Reviews of Modern Physics, 34(4):592–596, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.34.592.

Dirac:1962:EGF

[Dir62b] P. A. M. Dirac. The energy of the gravitational field. In Lichnerowicz and Tonnelat [LT62], pages 385–394. LCCN ????

Dirac:1962:EME

[Dir62c] P. A. M. Dirac. An extensible model of the electron. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 268(1332):57–67, June 19, 1962. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/2414316.

Dirac:1962:IGS

[Dir62d] P. A. M. Dirac. Interacting gravitational and spinor fields. In Anonymous [Ano62], pages 191–200. LCCN QC 6

.R295. URL http://adsabs.harvard.edu/abs/1962rdgr.

book..191D. This book is dedicated to Leopold Infeld in connection with his 60th birthday.

Dirac:1962:PFS

[Dir62e] P. A. M. Dirac. Particles of finite size in the gravitational field. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 270(1342):354–356, November 27, 1962. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/2414537.

Dirac:1963:EPP

[Dir63a] P. A. M. Dirac. The evolution of the physicist’s picture of nature. Scientific American, 208(5):45–53, May 1963. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v208/n5/pdf/scientificamerican0563-45.pdf.

Dirac:1963:RRS

[Dir63b] P. A. M. Dirac. A remarkable representation of the 3 + 2 de Sitter group. Journal of Mathematical Physics, 4(??):901–909, 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v4/i7/p901_s1.

Dirac:1964:ESH

[Dir64a] P. A. M. Dirac. Equivalence of the Schrödinger and Heisenberg pictures: [reply to H. S. Perlman]. Nature, 204(4960):772, November 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v204/n4960/abs/204772a0.html.

Dirac:1964:FQM

[Dir64b] P. A. M. Dirac. Foundations of quantum mechanics. Nature, 203 (4941):115–116, July 11, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v203/n4941/pdf/203115a0.pdf.

Dirac:1964:HMQ

[Dir64c] P. A. M. Dirac. Hamiltonian methods and quantum mechanics. Proceedings of the Royal Irish Academy, Section A, 63(??):49– 59 (1964), 1964. CODEN PRIAAK. ISSN 0035-8975. Larmor Lecture.

Dirac:1964:LQF

[Dir64d] P. A. M. Dirac. Lectures on quantum field theory, volume 2 of Monographs series. Belfer Graduate School of Science, Yeshiva University, New York, NY, USA, 1964. v + 87 pp. LCCN

QC174.45 .D57 1964. URL http://hooke.lib.cam.ac.uk/
cgi-bin/bib_seek.cgi?cat=ul&bib=338668.

Dirac:1964:MEP

[Dir64e] P. A. M. Dirac. The motion of an extended particle in the gravitational field. In Léopold Infeld, editor, Proceedings on Theory of Gravitation: Conference in Warszawa and Jabłonna, 25–31 July, 1962, pages 163–171 (discussion 171–175). Gauthier-Villars et cie, Paris, France, 1964. LCCN ????

Dirac:1964:LQM

[Dir64f] Paul A. M. Dirac. Lectures on quantum mechanics, volume 2 of Belfer Graduate School of Science Monographs Series. Belfer Graduate School of Science, New York, 1964. v + 87 pp. Reprinted in [Dir67f].

Dirac:1964:NBA

[Dir64g] Paul A. M. Dirac. Niels Bohrs alsidighed. (Danish) [Niels Bohr’s versatility]. In Bohr and Rozental [BR64], pages 297–300. LCCN QC16.B63 N5. English translation in [Roz67].

Dirac:1965:QED

[Dir65a] P. A. M. Dirac. Quantum electrodynamics without dead wood. Physical Review, 139(3B):684–690, August 1965. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Dirac:1965:TEP

[Dir65b] P. A. M. Dirac. Theory of electrons and positrons. In Anonymous [Ano65], pages 320–225. LCCN QC71 .P455

1965. URL http://nobelprize.org/nobel_prizes/physics/
laureates/1922/bohr-lecture.html;   http://nobelprize.
org/nobel_prizes/physics/laureates/1922/bohr-lecture.

pdf. Nobel Prize in Physics lecture, December 12, 1933.

Dirac:1966:FQM

[Dir66a] P. A. M. Dirac. Foundations of quantum mechanics. In A. Gelbart, editor, Proceedings: Some Recent Advances in the Basic Sciences, volume 1, pages 1–10. Belfer Graduate School of Science, Yeshiva University, New York, NY, USA, 1966. LCCN ????

Dirac:1966:LQF

[Dir66b] P. A. M. Dirac. Lectures on quantum field theory, volume 3 of Monographs series. Belfer Graduate School of Science, Yeshiva University, New York, NY, USA, 1966. viii + 151 pp. LCCN ????

URL http://hooke.lib.cam.ac.uk/cgi-bin/bib_seek.cgi?
cat=ul&bib=35554.

Dirac:1967:FEQ

[Dir67a] P. A. M. Dirac. The fundamental equations of quantum mechanics. In van der Waerden [vdW67], pages 307–320. LCCN QC174.12 S655. Reprint of [Dir25e].

Dirac:1967:PMC

[Dir67b] P. A. M. Dirac. Principios de mecánica cuántica. (Spanish) [The Principles of Quantum Mechanics]. Editorial Ariel, Barcelona, Spain, 1967. 330 pp. LCCN ???? URL http://adsabs.

harvard.edu/abs/1967pdmc.book.....D. Translated by An-

tonio Montes from English original.

Dirac:1967:QMP

[Dir67c] P. A. M. Dirac. Quantum mechanics and a preliminary investigation of the hydrogen atom. In van der Waerden [vdW67], pages 417–427. LCCN QC174.12 S655. Reprint of [Dir26e].

Dirac:1967:VNB

[Dir67d] P. A. M. Dirac. The versatility of Niels Bohr. In Rozental [Roz67], pages 306–309. LCCN QC16.B63 N53.

Dirac:1967:PQM

[Dir67e] P. A. M. (Paul Adrien Maurice) Dirac. The Principles of Quantum Mechanics, volume 27 of International series of monographs on physics (Oxford, England). Clarendon Press, Oxford, UK, fourth revised edition, 1967. ISBN 0-19-852011-5 (paperback). xii + 314 pp. LCCN ????

Dirac:1967:LQM

[Dir67f] Paul A. M. Dirac. Lectures on quantum mechanics, volume 2 of Belfer Graduate School of Science Monographs Series. Belfer Graduate School of Science, New York, 1967. v + 87 pp. Second printing of the 1964 original.

Dirac:1968:LPK

[Dir68a] P. A. M. Dirac. Lekcii po kvantovoj mekhanike. (Russian) [Lectures on quantum mechanics]. Mir, Moscow, USSR, 1968. ???? pp. LCCN ????

Dirac:1968:PIQ

[Dir68b] P. A. M. Dirac. The physical interpretation of quantum electrodynamics. Commentarii — Pontificia Academia Scientiarum, 2 (13):1–12, ???? 1968. CODEN CPOSAF. ISSN 0554-6648.

Dirac:1968:PMC

[Dir68c] P. A. M. Dirac. Principios de mecánica cuántica. (Spanish) [The Principles of Quantum Mechanics]. Ariel, Barcelona, Spain, 1968. 330 pp. LCCN ???? Translated by Antonio Montes from English original.

Dirac:1969:QGF

[Dir68d] P. A. M. Dirac. The quantization of the gravitational field. In Salam [Sal69], pages 539–543. LCCN ????

Dirac:1968:SWG

[Dir68e] P. A. M. Dirac. The scientific work of Georges Lemaître. Pontificiae Acad. Sci., Scripta Varia, 2(11):1–20, 1968. CODEN PASVAE. ISSN 0377-9971.

Dirac:1969:BIQ

[Dir69a] P. A. M. Dirac. The basic ideas of quantum mechanics. A course of lectures … January–March, 1969. Notes taken by G. D. Kaiser and R. M. Williams, Coral Gables, FL, 1969.

URL http://hooke.lib.cam.ac.uk/cgi-bin/bib_seek.cgi?
cat=ul&bib=3532179.

Dirac:1969:CEM

[Dir69b] P. A. M. Dirac. Can equations of motion be used? In Timm Gudehus and Geoffrey Kaiser, editors, Coral Gables Conference on Fundamental Interactions at High Energy, pages 1–18. Gordon and Breach, New York, NY, USA, 1969. LCCN ????

Dirac:1969:HF

[Dir69c] P. A. M. Dirac. Hopes and fears. Eureka, 32(??):2–4, October 1969.

Dirac:1969:MTP

[Dir69d] P. A. M. Dirac. Methods in theoretical physics. International Atomic Energy Agency Bulletin, ??(??):21–28, ???? 1969. CO-DEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). Reprinted from [BDH^{+}89].

Dirac:1969:PMC

[Dir69e] P. A. M. Dirac. Principios de mecánica cuántica. (Spanish) [The Principles of Quantum Mechanics]. Ariel, Barcelona, Spain, 1969. 330 pp. LCCN ???? Translated by Antonio Montes from English original.

Dirac:1970:CEM

[Dir70a] P. A. M. Dirac. Can equations of motion be used in high-energy physics? Physics Today, 23(4):29–31, April 1970. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). From the abstract: “A theory with mathematical beauty is more likely to be correct than an ugly one that fits some experimental data.”.

Dirac:1970:MTP

[Dir70b] P. A. M. Dirac. The methods of theoretical physics. Uspekhi Fizicheskikh Nauk, 102(10):291, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

URL http://ufn.ru/ru/articles/1970/10/j/. From a Life

of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Dirac:1970:SHS

[Dir70c] P. A. M. Dirac. Spinors in Hilbert space. Report ????, Center for Theoretical Studies, University of Miami, Coral Gables, FL, USA, 1970. 100 pp. A series of lectures given in Miami in 1969 and revised in 1970. Published in [Dir74g].

Dirac:1971:DQT

[Dir71a] P. A. M. Dirac. The development of quantum theory: J. Robert Oppenheimer memorial prize acceptance speech. Gordon and Breach Science Publishers, New York, NY, USA, 1971. 66 pp. LCCN ????

Dirac:1971:EOU

[Dir71b] P. A. M. Dirac. The evolution of our understanding of nature. Report CTS-LN-71-1, Center for Theoretical Studies, University of Miami, Coral Gables, FL, USA, 1971. 22 pp.

Dirac:1971:FPP

[Dir71c] P. A. M. Dirac. Fundamental problems in physics. Lecture given at 21st Lindau Meeting, June 28–July 2, 1971, 1971.

Dirac:1971:IPU

[Dir71d] P. A. M. Dirac. Is it possible to use equations of motion in highenergy physics? Uspekhi Fizicheskikh Nauk, 103(1):121–126, January 1971. CODEN UFNAAG. ISSN 0042-1294 (print),

1996-6652 (electronic). URL http://ufn.ru/ru/articles/
1971/1/d/.

Dirac:1971:PERa

[Dir71e] P. A. M. Dirac. A positive-energy relativistic wave equation. In Dal Cin et al. [DPI71], pages 1–11. LCCN ????

Dirac:1971:PERb

[Dir71f] P. A. M. Dirac. A positive-energy relativistic wave equation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 332(1551):435–445, May 4, 1971. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/77762.

Dirac:1972:DSP

[Dir72a] P. A. M. Dirac. Discrete subgroups of the Poincaré group. In V. I. Ritus, editor, Problemi teoretichestoi fiziki. (Russian) [Problems of theoretical physics] (in memory of Igor E. Tamm), pages 45–51, 484, 491 (1 plate). Nauka, Moscow, Russia, 1972.

Dirac:1972:EC

[Dir72b] P. A. M. Dirac. Evolutionary cosmology. Commentarii — Pontificia Academia Scientiarum, 2(46):1–15, ???? 1972. CODEN CPOSAF. ISSN 0554-6648.

Dirac:1972:PER

[Dir72c] P. A. M. Dirac. A positive-energy relativistic wave equation. II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 328(1572):1–7, May 2, 1972. CO-DEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/78078.

Dirac:1972:RQM

[Dir72d] P. A. M. Dirac. Relativity and quantum mechanics. Fields and Quanta, 3(??):139–164, ???? 1972. CODEN FDQTAB. ISSN 0046-3744.

Dirac:1972:SWG

[Dir72e] P. A. M. Dirac. The scientific work of Georges Lemaître. Pontificiae Acad. Sci., Scripta Varia, 36:67–83, 1972. CODEN PASVAE. ISSN 0377-9971. URL http://adsabs.harvard.

edu/abs/1972PASSV..36...67D;   http://www.uclouvain.

be/en-316446.html. L’Académie Pontificale des Sciences en mémoire de son second Président Georges Lemaître.

Dirac:1972:VGC

[Dir72f] P. A. M. Dirac. The variability of the gravitational constant. In Reines [Rei72a], pages 56–59. ISBN 0-87081-025-1.

LCCN QC780 .C65. URL http://adsabs.harvard.edu/abs/
1972cht..conf...56D.

Dirac:1973:DPC

[Dir73a] P. A. M. Dirac. Development of the physicist’s conception of nature. In Mehra [Meh73], pages 1–14. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.

springer.com/us/book/9789027703453.

Dirac:1973:FCT

[Dir73b] P. A. M. Dirac. Fundamental constants and their development in time. In Mehra [Meh73], pages 45–54. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.

springer.com/us/book/9789027703453.

Dirac:1973:LRFa

[Dir73c] P. A. M. Dirac. Long range forces and broken symmetries. In Behram Kurşunoğlu, editor, Fundamental Interactions in Physics: proceedings of the Coral Gables Conference on Fundamental Interactions, January 22–26, 1973, volume 2 of Studies in the natural sciences, pages 1–17. Plenum Press, New York, NY, USA; London, UK, 1973. ISBN 0-306-36902-8. LCCN QC793.9 .C67 1973.

Dirac:1973:LRFb

[Dir73d] P. A. M. Dirac. Long range forces and broken symmetries. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 333(1595):403–418, June 26, 1973. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/78370.

Dirac:1973:NIS

[Dir73e] P. A. M. Dirac. New ideas of space and time. Die Naturwissenschaften, 60(12):529–531, December 1973. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1973NW.....

60..529D. Lecture held at the Lindau Nobelpreisträger-Tagung, July 2, 1973.

Dirac:1973:RQM

[Dir73f] P. A. M. Dirac. Relativity and quantum mechanics. In E. Sudarshan and Y. Ne’eman, editors, The Past Decade in Particle Theory, pages 747–772. Gordon and Breach Science Publishers, New York, NY, USA, 1973. ISBN 0-677-12010-9. LCCN QC793 .P37. First published in [Dir72d].

Dirac:1973:SED

[Dir73g] P. A. M. Dirac. Some of the early developments of quantum theory. In Behram Kurşunoğlu and Arnold Perlmutter, editors, Impact of Basic Research on Technology, pages 1–13. Plenum Press, New York, NY, USA; London, UK, 1973. ISBN 0-306-36901-X. LCCN QC30 .K82.

Dirac:1973:ZNP

[Dir73h] P. A. M. Dirac. Zitterbewegung of the new positive-energy particle. In Iverson et al. [IPM73], pages 354–364. ISBN 0-306-36903-6. LCCN QC793.9 .C68 1972. URL http://

adsabs.harvard.edu/abs/1973fipa.conf..354D;   http:/
/www.osti.gov/energycitations/product.biblio.jsp?osti_

id=4181820. A volume dedicated to P. A. M. Dirac on the occasion of his seventieth birthday.

Dirac:1974:APM

[Dir74a] P. A. M. Dirac. An action principle for the motion of particles. General relativity and gravitation, 5(6):741–748, 1974. CODEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic). URL

http://adsabs.harvard.edu/abs/1974GReGr...5..741D.

Dirac:1974:CML

[Dir74b] P. A. M. Dirac. Cosmological models and the Large Numbers Hypothesis. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 338(1615):439–446, July 16, 1974. CODEN PRLAAZ. ISSN 0080-4630. URL http://

adsabs.harvard.edu/abs/1974RSPSA.338..439D;   http:/
/www.jstor.org/stable/78591. See [Rob83, Rob85, Jor59a].

Dirac:1974:DQM

[Dir74c] P. A. M. Dirac. The development of quantum mechanics. Contributi del Centro Linceo Interdisciplinare di scienze mathematiche e loro applicazioni, 4(??):1–11, 1974. CODEN CCLADS. ISSN 0391-8041.

Dirac:1974:DPA

[Dir74d] P. A. M. Dirac. Dirac, Paul Adrien Maurice. In Edgardo Macorini, editor, Scienziati e Tecnologici Contemporanei, page ?? Arnoldo Mondadori, Milan, Italy, 1974. ISBN ???? LCCN ????

Dirac:1974:GNS

[Dir74e] P. A. M. Dirac. The geometrical nature of space and time. In Stephan L. Mintz, Laurence Mittag, and Susan M. Widmayer, editors, Fundamental Theories in Physics, volume 5 of Studies in the natural sciences, pages 1–18. Plenum Press, New York, NY, USA; London, UK, 1974. ISBN 0-306-36905-2. LCCN QC793.9 .O72 1974.

Dirac:1974:MMR

[Dir74f] P. A. M. Dirac. Magnitnii monopol. (Russian) [Magnetic monopoles]. In A. M. Prokhorov, editor, Bolshaya Sovetskaya Entsiklopediya, volume 15, pages 186–187 (columns 546–548). Izdatelstvo ’Sovetskaya Entsiklopediya’, Moscow, USSR, 1974. ISBN ???? LCCN ????

Dirac:1974:SHS

[Dir74g] Paul Adrien Maurice Dirac. Spinors in Hilbert space. Plenum Press, New York, NY, USA; London, UK, 1974. ISBN 0-306-30798-7. vi + 91 pp. LCCN QA433 .D57. A series of lectures given in Miami in 1969 and revised in 1970.

Dirac:1975:TP

[Dir75a] P. A. M. Dirac. 5. theory of the positron. In Mehra [Meh75], chapter 8, pages 218–220. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Dirac:1975:DGC

[Dir75b] P. A. M. Dirac. Does the gravitational constant vary? Commentarii — Pontificia Academia Scientiarum, 3(7):1–7, ???? 1975. CODEN CPOSAF. ISSN 0554-6648.

Dirac:1975:HPS

[Dir75c] P. A. M. Dirac. An historical perspective on spin. In J. B. Roberts, editor, Proceedings of Summer Study on High-Energy Physics with Polarized Beams, July 22–26, 1974, Ann Arbor, MI, USA, volume ANL/HEP 75-02, pages 1–11. Argonne National Laboratory, Argonne, IL, USA, 1975. ISBN ???? LCCN ????

Dirac:1975:LNH

[Dir75d] P. A. M. Dirac. The Large Numbers Hypothesis and its consequences. In Kurşunoğlu et al. [KPW75], pages 443–455. ISBN 0-306-36909-5. LCCN QC793 .O7 1975. URL http://adsabs.

harvard.edu/abs/1975tehe.conf..443D.

Dirac:1975:JC

[Dir75e] P. A. M. Dirac. Obituary: James Chadwick. Commentarii — Pontificia Academia Scientiarum, 3(6):1–5, ???? 1975. CODEN CPOSAF. ISSN 0554-6648.

Dirac:1975:V

[Dir75f] P. A. M. Dirac. Variation of G. Nature, 254(5497):273, March 20, 1975. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://adsabs.harvard.edu/abs/
1975Natur.254..273D;   http://www.nature.com/nature/
journal/v254/n5497/abs/254273a0.html.

Dirac:1975:GTR

[Dir75g] P. A. M. (Paul Adrien Maurice) Dirac. General Theory of Relativity. Wiley, New York, NY, USA, 1975. ISBN 0-471-21575-9. viii + 69 pp. LCCN QC173.6 .D57.

Dirac:1976:BPP

[Dir76a] P. A. M. Dirac. Belief and prejudice in physics. Chemiker-Zeitung, 100(12):541–542, ???? 1976. ISSN 0009-2894.

Dirac:1976:HIP

[Dir76b] P. A. M. Dirac. Heisenberg’s influence on physics. Commentarii — Pontificia Academia Scientiarum, 3(14):1–15, ???? 1976. CODEN CPOSAF. ISSN 0554-6648.

Dirac:1976:TMM

[Dir76c] P. A. M. Dirac. Theory of magnetic monopoles. In Arnold Perlmutter, editor, New Pathways in High-Energy Physics: Magnetic charge and other fundamental approaches. Proceedings of Orbis Scientiae 1976 held by the Center for Theoretical Studies, University of Miami, Coral Gables, Florida, January 19–22, 1976, volume 10 of Studies in the natural sciences, pages 1–14. Plenum Press, New York, NY, USA; London, UK, 1976. ISBN 0-306-36910-9 (vol. 1). LCCN QC793 .O7 1976.

Dirac:1977:AVP

[Dir77a] P. A. M. Dirac. Annahmen und Voreingenommenheit in der Physik. (German) [Assumptions and prejudices in physics]. Naturwissenschaftliche Rundschau, 30(??):429–432, ???? 1977. CODEN NARSAC. ISSN 0028-1050. Lindau lecture 1976.

Dirac:1977:DSM

[Dir77b] P. A. M. Dirac. The dynamics of streams of matter. In Behram Kurşunoğlu, Arnold Perlmutter, and Linda F. Scott, editors, Deeper Pathways in High Energy Physics: Proceedings of a Conference Held by the Center for Theoretical Studies, University of Miami, Coral Gables, Florida, January 17–21, 1977, volume 12 of Studies in the natural sciences, pages 1–11. Plenum Press, New York, NY, USA; London, UK, 1977. ISBN 0-306-36912-5. LCCN QC793 .O7 1977.

Dirac:1977:ESQ

[Dir77c] P. A. M. Dirac. Ehrenhaft, the subelectron and the quark. In Weiner [Wei77], pages 290–293. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Dirac:1977:FEQ

[Dir77d] P. A. M. Dirac. Fundamental equations of quantum mechanics. Uspekhi Fizicheskikh Nauk, 122(8):611–621, August 1977. CO-DEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

URL http://ufn.ru/ru/articles/1977/8/e/.

Dirac:1977:MM

[Dir77e] P. A. M. Dirac. Magnetic monopole. In Aleksandr M. Prochorov and Jean Paradise, editors, Great Soviet Encyclopedia, page ?? MacMillan Publishing Company, New York, NY, USA, third edition, 1977. ISBN ???? LCCN ????

Dirac:1977:REE

[Dir77f] P. A. M. Dirac. Recollections of an exciting era. In Weiner [Wei77], pages 109–146. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Dirac:1977:REWb

[Dir77g] P. A. M. Dirac. The relativistic electron wave electron. In L. (Livia) Jenik and I. Montvay, editors, Proceedings of the 1977 European Conference on Particle Physics, Budapest, Hungary, 4–9 July 1977, pages 15–34. Central Research Institute for Physics, Budapest, Hungary, 1977. ISBN 963-371-349-8. LCCN QC770 .E9 1977; QC793 +E97 1977. Reprinted in [Dir79h, Dir79i, Dir79j].

Dirac:1977:REWa

[Dir77h] P. A. M. Dirac. The relativistic electron wave equation. Europhysics News, 8(10):1–4, ???? 1977. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). Abridged version of [Dir77g] presented at the European Conference on Particle Physics, Budapest, July 4–9, 1977.

Dirac:1977:RWE

[Dir77i] P. A. M. Dirac. The relativistic wave equation of electron. Fizikai Szemle (Budapest), 27(??):443–450, ???? 1977. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Dirac:1978:AFA

[Dir78a] P. A. M. Dirac. Alapveto fizikai állandôk és idöbeni fejlödésük. (Hungarian) [Fundamental constants and their development in time]. Fizikai Szemle (Budapest), 28(??):201–204, ???? 1978. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1978FizSz.

.28..201D. Abridged Hungarian version of [Dir73b].

Dirac:1978:CV

[Dir78b] P. A. M. Dirac. Consequences of varying G. In J. E. Lannutti and P. K. Williams, editors, Current trends in the theory of fields (Symposium in honor of P. A. M. Dirac, Florida State University, Tallahassee, FL, April 6–7, 1978), volume 48 of Particles and Fields Subseries, pages 169–174. American Institute of Physics, Woodbury, NY, USA, 1978. LCCN QC793.3.F5 C87

1978. URL http://adsabs.harvard.edu/abs/1978AIPC...
48..169D;   http://link.aip.org/link/?APCPCS/48/169/
1.

Dirac:1978:LNH

[Dir78c] P. A. M. Dirac. The Large Numbers Hypothesis and the cosmological variation of the gravitational constant. In Halpern [Hal78], pages 3–20. ISBN 0-8130-0610-4. LCCN QC178

.O4. URL http://adsabs.harvard.edu/abs/1978omcv.conf.
...3D.

Dirac:1978:MFQ

[Dir78d] P. A. M. Dirac. The mathematical foundations of quantum theory. In A. R. Marlow, editor, Mathematical foundations of quantum theory (Proceedings of a Conference, Loyola University, New Orleans, LA, 1977), pages 1–8. Academic Press, New York, USA, 1978.

Dirac:1978:MC

[Dir78e] P. A. M. Dirac. The monopole concept. International Journal of Theoretical Physics, 17(4):235–247, 1978. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Dirac:1978:NAC

[Dir78f] P. A. M. Dirac. New approach to cosmological theory. In Kurşunoğlu et al. [KPS^{+}78], pages 1–16. ISBN 0-306-40037-5. LCCN QC793 .O7 1978.

Dirac:1978:NIA

[Dir78g] P. A. M. Dirac. New ideas about gravitation and cosmology. Commentarii — Pontificia Academia Scientiarum, 3(24):1–10, ???? 1978. CODEN CPOSAF. ISSN 0554-6648.

Dirac:1978:PA

[Dir78h] P. A. M. Dirac. The prediction of antimatter. Technical report, University of Michigan, Ann Arbor, Ann Arbor, MI, USA, 1978. 25 pp. First H. R. Crane Lecture given on April 17, 1978 at the University of Michigan, Ann Arbor.

Dirac:1979:BET

[Dir79a] P. Dirac. Beauty of Einstein’s thought. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 101(6):46–47, ???? 1979. CODEN MEENAH. ISSN 0025-6501 (print), 1943-5649 (electronic).

Dirac:1979:EDI

[Dir79b] P. Dirac. Einstein, definite idea about science. Folia Humanistica, 17(196):199–201, ???? 1979. ISSN 0015-5594.

Dirac:1979:ECC

[Dir79c] P. Dirac. Einstein, or a certain conception of science. Chimia, 33(9):346–347, ???? 1979. ISSN 0009-4293.

Dirac:1979:DET

[Dir79d] P. A. M. Dirac. Developments of Einstein’s theory of gravitation. In Perlmutter and Scott [PS79], pages 1–13. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (ebook). LCCN QC19.2-20.85; QC178 .O72 1979. URL http:/

/adsabs.harvard.edu/abs/1979opae.conf....1D;   http:
//link.springer.com/chapter/10.1007/978-1-4684-3596-
2_1/;   http://link.springer.com/content/pdf/bfm:978-
1-4684-3596-2/1.pdf.

Dirac:1979:DGC

[Dir79e] P. A. M. Dirac. Does the gravitational constant change? Chemiker-Zeitung, 103(10):331–332, ???? 1979. ISSN 0009-2894.

Dirac:1979:EET

[Dir79f] P. A. M. Dirac. The excellence of Einstein’s theory of gravitation. Impact of Science on Society [Unesco], 29(1):11–14, 1979. CODEN ISSOA8. ISSN 0019-2872. Article adapted from talk at the UNESCO Symposium at Ulm, September 19, 1978. Reprinted in [Dir80e].

Dirac:1979:LNH

[Dir79g] P. A. M. Dirac. The Large Numbers Hypothesis and the Einstein Theory of Gravitation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 365(1720):

19–30, February 19, 1979. CODEN PRLAAZ. ISSN 0080-

4630. URL http://adsabs.harvard.edu/abs/1979RSPSA.
365...19D;   http://www.jstor.org/stable/79806. See

[Jor59a, Rob83, Rob85].

Dirac:1979:REWa

[Dir79h] P. A. M. Dirac. The relativistic electron wave equation. Fizikai Szemle (Budapest), 27(??):443–450, ???? 1979. CO-DEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). Reprint of [Dir77g].

Dirac:1979:REWb

[Dir79i] P. A. M. Dirac. The relativistic electron wave equation. Soviet Physics. Uspekhi, 22(??):648–653, ???? 1979. CODEN SOP-UAP. ISSN 0038-5670.

Dirac:1979:REWc

[Dir79j] P. A. M. Dirac. The relativistic electron wave equation. Soviet Physics. Uspekhi, 22(??):648–653, ???? 1979. CODEN SOP-UAP. ISSN 0038-5670. Reprint of [Dir77g].

Dirac:1979:RWE

[Dir79k] P. A. M. Dirac. Relativistic wave equation of the electron. Uspekhi Fizicheskikh Nauk, 129(4):681–691, April 1979. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1979/8/e/. Also published in

EPS Conf.1977:0015 (QCD161:I48:1977:V.1).

Dirac:1979:TT

[Dir79l] P. A. M. Dirac. The test of time. The Unesco Courier, 32(5): 17–23, May 1979. ISSN 1993-8616. Abridged version of lecture given at the UNESCO Symposium in Ulm, September 1978. German translation in [Dir80b].

Dirak:1979:PET

[Dir79m] Pol A. M. Dirak. The perfection of Einstein’s theory of gravitation. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 22(55)(3):216–218, 1979. CODEN FMBMAC. ISSN 0015-3265. Translated from the French by D. Vačov.

Dirac:1980:APM

[Dir80a] P. A. M. Dirac. Address of P. A. M. Dirac. Science, 207(4436): 1161–1162, March 14, 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://adsabs.harvard.

edu/abs/1979Sci...207.1161D.

Dirac:1980:ZGT

[Dir80b] P. A. M. Dirac. Der Zeittest. (German) [The test of time]. Naturwissenschaftliche Rundschau, 33(??):353–356, ???? 1980. CODEN NARSAC. ISSN 0028-1050. German translation of [Dir79l].

Dirac:1980:ESP

[Dir80c] P. A. M. Dirac. Einstein session of the Pontifical Academy. Science, 207(4436):1161–1162, ???? 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Dirac:1980:EIP

[Dir80d] P. A. M. Dirac. Einstein’s influence on physics. In Bucciarelli

[Buc80], pages 19–23. URL http://books.google.com/books?
id=pnpkYgEACAAJ.

Dirac:1980:EET

[Dir80e] P. A. M. Dirac. Excellence of Einstein’s Theory of Gravitation. In Goldsmith et al. [GMW80], pages 41–46. ISBN 0-08-025019-X. LCCN QC16.E5 E52 1980. Reprint of [Dir79f].

Dirac:1980:LP

[Dir80f] P. A. M. Dirac. A little ‘prehistory’. The Old Cothamian, page 9, 1980.

Dirac:1980:VPM

[Dir80g] P. A. M. Dirac. The variation of G and the problem of the Moon. In Perlmutter [Per80], pages 1–7. ISBN 0-306-40565-2. LCCN

QC793.28 .O72 1980. URL http://adsabs.harvard.edu/abs/
1980rdhe.conf....1D.

Dirac:1980:WWB

[Dir80h] P. A. M. Dirac. Why we believe in the Einstein theory. In Gruber and Millman [GM80], pages 1–11. ISBN 0-306-40541-5. LCCN Q172.5.S95 S92.

Dirac:1981:DRM

[Dir81a] P. A. M. Dirac. Does renormalization make sense? In Dennis W. Duke and Jeff F. Owens, editors, Perturbative Quantum Chromodynamics: Conference held on the Florida State University Campus during March 25–28, 1981, volume 74 of AIP Conference Proceedings, pages 129–130. American Institute of Physics, Woodbury, NY, USA, October 1, 1981. CODEN APCPCS. ISBN 0-88318-173-8. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QC793.3.Q35 P46. URL http://

link.aip.org/link/?APCPCS/74/129/1.

Dirac:1981:EDP

[Dir81b] P. A. M. Dirac. Einstein and the development of physics. In Colette M. Kinnon, A. N. Kholodilin, and J. G. Richardson, editors, The Impact of Modern Scientific Ideas on Society: in commemoration of Einstein: papers presented at the Unesco symposium on the impact of modern scientific ideas on society, Munich–Ulm, 18–20 September 1978, and the addresses delivered on the occasion of Unesco’s celebration of the hundredth anniversary of Einstein’s birth, Paris, 9 May 1979, pages 13–23. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1238-7. LCCN Q175.4 .I54.

Dirac:1981:MU

[Dir81c] P. A. M. Dirac. Models of the universe. In Behram Kurşunoğlu and Arnold Perlmutter, editors, Gauge Theories, Massive Neutrinos, and Proton Decay: Proceedings of Orbis Scientiae 1981, Coral Gables, Florida, January 19–22, 1981, volume 18 of Studies in the natural sciences, pages 1–9. Plenum Press, New York, NY, USA; London, UK, 1981. ISBN 0-306-40821-x. LCCN QC793.3.F5 O7 1981.

Dirac:1981:PQM

[Dir81d] P. A. M. (Paul Adrien Maurice) Dirac. The Principles of Quantum Mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised edition, 1981. ISBN 0-19-852011-5. xii + 314 pp. LCCN QC6 .D55 1981M.

Dirac:1981:PM

[Dir82a] P. A. M. Dirac. Pretty mathematics. International Journal of Theoretical Physics, 21(8–9):603–605, 1981–1982. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). Contribution to symposium at Loyola University, New Orleans, May 1981, in honor of Dirac’s 80th birthday.

Dirac:1982:AQM

[Dir82b] P. A. M. Dirac. The arrival of quantum mechanics. In Dorobantu [Dor82], pages 185–190. LCCN QC1 .E87 1981.

Dirac:1982:EYR

[Dir82c] P. A. M. Dirac. The early years of Relativity. In Holton and Elkana [HE82], pages 79–90. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Dirac:1982:MTP

[Dir82d] P. A. M. Dirac. Methods in theoretical physics. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 24(57)(4):257–263 (1983), 1982. CODEN FMBMAC. ISSN 0015-3265. Translated from the English by D. Vachov.

Dirac:1982:PM

[Dir82e] P. A. M. Dirac. Pretty mathematics. International Journal of Theoretical Physics, 21(8–9):603–605, ???? 1982. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Dirac:1982:VQT

[Dir82f] P. A. M. Dirac. The variation of G and the quantum theory. In Ruffini [Ruf82], pages 1–6. ISBN 0-444-86357-5 (set). LCCN

QC173.6 .M37 1979. URL http://adsabs.harvard.edu/abs/

1982mgm..conf....1D. Paper presented at 1979 conference.

Dirac:1983:MLP

[Dir83a] P. A. M. Dirac. My life as a physicist. In Antonino Zichichi, editor, The Unity of the Fundamental Interactions: proceedings of the nineteenth Course of the International School of Subnuclear Physics, held July 31–August 11, 1981, in Erice, Trapani, Sicily, volume 19 of The subnuclear series, pages 733–749. Plenum Press, New York, NY, USA; London, UK, 1983. ISBN 0-306-41242-X. LCCN QC794.8.H5 I577 1981.

Dirac:1983:OQF

[Dir83b] P. A. M. Dirac. The origin of quantum field theory. In Brown and Hoddeson [BH83b], pages 39–55. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983. Paper presented at May 1980 conference.

Dirac:1983:PSG

[Dir83c] P. A. M. Dirac. The present state of gravitational theory. In Kuršunoğlu and Perlmutter [KP83], pages 1–11. ISBN 0-306-41345-0. LCCN QC793.3.F5 O7 1982. Paper presented at 1982 conference.

Dirak:1983:PF

[Dir83d] P. A. M. Dirak. Puti fiziki. Ènergoizdat, Moscow, USSR, 1983. 88 pp. Translated from the English by N. Ya. Smorodinskaya, Translation edited and with a preface and an afterword by Ya. A. Smorodinskiĭ.

Dirac:1984:BP

[Dir84a] P. A. M. Dirac. Blackett and the positron. In Hendry [Hen84], pages 61–62. ISBN 0-85274-761-6. LCCN QC51.G72 C352 1984.

Dirac:1984:FAP

[Dir84b] P. A. M. Dirac. The future of atomic physics. International Journal of Theoretical Physics, 23(8):677–681, ???? 1984. CO-DEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Dirac:1984:RFP

[Dir84c] P. A. M. Dirac. The requirements of fundamental physical theory. European Journal of Physics, 5(??):65–67, ???? 1984. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). Lindau lecture 1982 (the last that Dirac gave there).

Dirac:1984:TMP

[Dir84d] P. A. M. Dirac. The theory of magnetic poles. In Rebbi and Soliani [RS84], pages 485–498. ISBN 9971-966-42-5, 9971-966-43-3 (paperback). LCCN QC174.26.W28 S635 1984. Reprint of [Dir48d].

Dirac:1984:UGR

[Dir84e] P. A. M. Dirac. [unknown]. (German) [The requirements of fundamental physical theory]. Naturwissenschaftliche Rundschau, 30(??):429–431, ???? 1984. CODEN NARSAC. ISSN 0028-1050.

Dirak:1985:PM

[Dir85] P. A. M. Dirak. Pretty mathematics. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 27(60)(4):357–358, 1985. CODEN FMBMAC. ISSN 0015-3265. Translated from the English by B. Penkov.

Dirac:1986:EPT

[Dir86] P. A. M. Dirac. Elementary particles and their interactions. Manuscript held by the Pauli Library at CERN, Geneva, Switzerland [DP86, Ref. 196]., 1986.

Dirac:1987:TMD

[Dir87a] Margit Dirac. Thinking of my darling Paul. In Kurşunoğlu and Wigner [KW87], pages 3–8. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Dirac:1987:GGG

[Dir87b] P. A. M. Dirac. Gravitationswellen (German) [Gravitational waves]. In R. Schmid and W. Wessinger, editors, Nobel Prize Winners, pages 232–235. ????, ????, 1987. ISBN ???? LCCN ???? Reprint of [Dir60b, Dir60a].

Dirac:1987:HPV

[Dir87c] P. A. M. Dirac. The history of the physical view of nature. In W. Greiner and G. Wolschine, editors, Elementary Matter, Vacuum and Fields, pages 18–27. ????, ????, 1987. ISBN ???? LCCN ????

Dirac:1987:IQF

[Dir87d] P. A. M. Dirac. The inadequacies of quantum field theory. In Kurşunoğlu and Wigner [KW87], pages 194–

198. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Dirac:1987:REE

[Dir87e] P. A. M. Dirac. Recollections of an exciting era. Uspekhi Fizicheskikh Nauk, 153(9):105–134, September 1987. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1987/9/c/.

Dirac:1988:MEE

[Dir88a] P. A. M. Dirac. Memories of an extraordinary epoch. II, III. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 30(63)(3):195–209, 1988.

CODEN FMBMAC. ISSN 0015-3265. Translated from the English by D. Vachov.

Dirak:1988:MEE

[Dir88b] P. A. M. Dirak. Memories of an extraordinary epoch. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 30(63)(2):127–133, 1988. CODEN FMBMAC. ISSN 0015-3265. Translated from the English by D. Vachov.

Dirac:1989:MTP

[Dir89] P. A. M. Dirac. Methods in theoretical physics. In Bethe et al. [BDH^{+}89], pages 19–30. ISBN 9971-5-0937-7. LCCN QC71

.F74 1989. URL http://adsabs.harvard.edu/abs/1989liph.

book.....B. Lecture given at the International Symposium on Contemporary Physics. Triest, the International Centre of Theoretical Physics, June 1968. Vienna: IAEA. (Special Suppl. of IAEA Bulletin).

Dirac:1990:PMQ

[Dir90a] Paul A. M. Dirac. Les principes de la méecanique quantique. J. Gabay, Sceaux, France, 1990. ISBN 2-87647-071-3. viii + 314 pp. LCCN ???? Revised and enlarged version of [Dir31a].

Dirak:1990:KSK

[Dir90b] P. A. M. Dirak. K sozdaniyu kvantovoi teorii polya, volume 7 of Biblioteka Teoreticheskoĭ Fiziki [Library of Theoretical Physics]. Nauka, Moscow, Russia, 1990. ISBN 5-02-014024-4. 368 pp. Osnovnye stati 1925–1958 godov. [Major papers, 1925–1958], Translated from the English and French by A. B. Kozhevnikov, V. P. Pavlov, M. K. Polivanov and V. P. Shelest, Edited and with an introduction by B. V. Medvedev.

Dirak:1990:VNE

[Dir90c] P. A. M. Dirak. Vospominaniya o Neobychaĭnoĭ Epokhe (Russian). [Recollections of the Extraordinary Epoch]. Nauka, Moscow, Russia, 1990. ISBN 5-02-014344-8. 208 pp. Translated from the English by N. Ya. Smorodinskaya, Edited by Ya. A. Smorodinskiĭ.

Dirak:1990:VNJ

[Dir90d] Poul’ Andrei Moris Dirak. Vospominanija o neobychajnoj jepohe Sb. st. (Russian) [Memories of an extraordinary epoch]. Nauka, Moscow, Russia, 1990. ISBN 5-02-014344-8. 205 + 2 pp. Russian translation by N. Ya. Smorodinskaya of [?].

Dirac:1991:H

[Dir91a] Paul A. M. Dirac. Hilbert. In Ferris and Fadiman [FF91], page 604. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Dirac:1991:TEP

[Dir91b] Paul A. M. Dirac. Theory of electrons and positrons. In Ferris and Fadiman [FF91], pages 80–85. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Dirac:1996:GTR

[Dir96] P. A. M. (Paul Adrien Maurice) Dirac. General Theory of Relativity. Princeton landmarks in mathematics and physics; Princeton paperbacks. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01146-X (paperback). viii + 71 pp. LCCN QC173.6 .D57 1996. URL http://

adsabs.harvard.edu/abs/1996gtr..book.....D;   http:/
/press.princeton.edu/titles/5813.html;   http://www.
loc.gov/catdir/description/prin021/95046196.html;   http:
//www.loc.gov/catdir/toc/prin031/95046196.html.

Dirac:1999:LQM

[Dir99] P. A. M. Dirac. Lektsii po kvantovoj mekhanike. (Russian) [Lectures on the quantum mechanics]. Nauchno-Izdatel’skij Tsentr “Regulyarnaya i Khaoticheskaya Dinamika, Izhevsk, USSR, 1999. 148 pp.

Dirac:2001:LQM

[Dir01] P. A. M. (Paul Adrien Maurice) Dirac. Lectures on quantum mechanics. Dover, New York, NY, USA, 2001. ISBN 0-486-41713-1 (paperback). v + 87 pp. LCCN QC174.125

.D55 2001. URL http://www.loc.gov/catdir/description/
dover032/00065608.html.

Dirac:2002:SNT

[Dir02] P. A. M. Dirac. Sobranie Nauchnykh Trudov T. I. Kvantovaya Teoriya (Monografii, Lektsii). (Russian) [Collected Scientific Works. Vol. I Quantum Theory (Monographs, Lectures)]. Klassiki Nauki. Fizmatlit, Moscow, Russia, 2002. ISBN 5-9221-0201-X. 698 pp. LCCN ???? Edited by A. D. Sukhanov.

Dirac:2003:MF

[Dir03a] Monica Dirac. My father. In Baer and Belyaev [BB03], pages 39–42. ISBN 981-238-412-X. LCCN QC174.45.A1 D57 2003.

Dirac:2003:SNT

[Dir03b] P. A. M. Dirac. Sobranie Nauchnykh Trudov T. II. Kvantovaya Teoriya (Nauchnye Stat’i: 1924–1947). (Russian) [Collected Scientific Works. Quantum Theory (Scientific Papers: 1924–1947)]. Klassiki Nauki. Fizmatlit, Moscow, Russia, 2003. ISBN 5-9221-0381-4. 846 pp. LCCN ???? Edited by A. D. Sukhanov.

Dirac:2003:PQM

[Dir03c] Paul Adrien Maurice Dirac. The Principles of Quantum Mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised reprinted edition, 2003. ISBN 0-19-852011-5. xii + 314 pp. LCCN QC6

.D55 2003. URL 04; 2;   http://www.gbv.de/dms/bowker/
toc/9780198520115..

Dirac:2004:SNT

[Dir04] P. A. M. Dirac. Sobranie Nauchnykh Trudov T. III. Kvantovaya Teoriya (Nauchnye Stat’i: 1948–1984). (Russian) [Collected Scientific Works Vol. III. Quantum Theory (Scientific Papers: 1948–1984)]. Klassiki Nauki. Fizmatlit, Moscow, Russia, 2004. ISBN 5-9221-0503-5. 718 pp. LCCN ???? Edited by A. D. Sukhanov.

Dirac:2005:LQM

[Dir05a] P. A. M. Dirac. The Lagrangian in quantum mechanics. In Feynman’s Thesis — a New Approach To Quantum Theory [Bro05], page ?? ISBN 981-256-366-0 (print), 981-256-763-1 (e-book).

LCCN QC174.12 .F49 2005. URL http://public.eblib.com/
EBLPublic/PublicView.do?ptiID=244532;   http://www.worldscibooks.
com/promotion/feynman.html#5852.

Dirak:2005:SNT

[Dir05b] P. A. M. Dirak. Sobranie Nauchnykh Trudov T. IV. Gravitatsiya i Kosmologiya. Vospominaniya i Razmyshleniya. (Lektsii, Nauchnye Stat’i: 1937–1984). (Russian) [Collected Scientific Works Vol. IV. Gravitation and Cosmology. Memoirs and Thoughts (Lectures, Scientific Papers: 1937–1984)]. Klassiki Nauki. Fizmatlit, Moscow, Russia, 2005. ISBN 5-9221-0613-9. 784 pp. LCCN ???? Edited by A. D. Sukhanov.

Dirac:2009:PQM

[Dir09] Paul Adrien Maurice Dirac. The principles of quantum mechanics, volume 27 of The international series of monographs on physics. Clarendon Press, Oxford, UK, fourth revised reprinted edition, 2009. ISBN 0-19-852011-5 (paperback). xii + 314 pp. LCCN QC6 .D55 2009.

Dittrich:2015:PWA

[Dit15] Walter Dittrich. On the Pauli–Weisskopf anti-Dirac paper. European Physical Journal H, 40(2):261–278, March 2015. CO-DEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2015-60006-1.

Dahl:1995:DKP

[DJ95] J. P. Dahl and T. Jørgensen. On the Dirac–Kepler problem: The Johnson–Lippmann operator, supersymmetry, and normalmode representations. International Journal of Quantum Chemistry, 53(2):161–??, 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Das:1989:CEP

[DK89] A. Das and D. Kay. A class of exact plane wave solutions of the Maxwell–Dirac equations. Journal of Mathematical Physics, 30 (10):2280–2284, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeMaria:1982:SDU

[DL82] M. De Maria and F. La Teana. Schrödinger’s and Dirac’s unorthodoxy in quantum mechanics. Fundamenta Scientiae, 3(??): 129–148, ???? 1982.

Dimakis:1985:SEC

[DMH85] A. Dimakis and F. Müller-Hoissen. Solutions of the Einstein– Cartan–Dirac equations with vanishing energy-momentum tensor. Journal of Mathematical Physics, 26(5):1040–1048, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeJong:1996:RCU

[DN96] W. A. De Jong and W. C. Nieuwpoort. Relativity and the chemistry of UF_{6}: a molecular Dirac–Hartree–Fock-CI study. International Journal of Quantum Chemistry, 58(2):203–216, ???? 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-

461X (electronic). URL http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=60476.

Dine:2003:TVS

[DNRV03] Michael Dine, Yosef Nir, Guy Raz, and Tomer Volansky. Time variations in the scale of grand unification. Physical Review D (Particles and Fields), 67(1):015009, January 1, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.67.015009.

Doan:1991:CFD

[Doa91] Khanh Bui Doan. A computation of Fermi–Dirac integrals by asymptotics and the Hermite corrector formula. Applied Mathematics and Computation, 41(1 (part I)):61–68, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dorobantu:1982:TPP

[Dor82] I. A. Dorobantu, editor. Trends in physics, 1981: papers presented at the fifth General Conference of the European Physical Society. Central Institute of Physics, Bucharest, Romania, 1982. LCCN QC1 .E87 1981.

Duff:2002:TNF

[DOV02] M. J. Duff, L. B. Okun, and G. Veneziano. Trialogue on the number of fundamental constants. Journal of High Energy Physics, 0203(023):1–31, 2002. CODEN JHEPAB. ISSN

1126-6708. URL http://arxiv.org/pdf/physics/0110060;
http://iopscience.iop.org/1126-6708/2002/03/023/.

Dalitz:1986:OPA

[DP86] Richard Henry Dalitz and Sir Rudolf Ernst Peierls. Obituary: Paul Adrien Maurice Dirac. 8 August 1902–20 October 1984. elected F.R.S. 1930. Biographical Memoirs of Fellows of the Royal Society, 32(??):137–185, December 1986. CODEN BM-FRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL

http://www.jstor.org/stable/770111.

Dalitz:1997:SSP

[DP97] Richard Henry Dalitz and Sir Rudolf Peierls, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (ebook). xxiii + 805 pp. LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book....
.D;   https://ui.adsabs.harvard.edu/#abs/1997ssps.book.
....P;   https://ui.adsabs.harvard.edu/#abs/1997WSSP.
..19.....D;   https://www.worldscientific.com/worldscibooks/
10.1142/3128.

Durr:2003:FAS

[DP03] D. Dürr and P. Pickl. Flux-across-surfaces theorem for a Dirac particle. Journal of Mathematical Physics, 44(2):423–465, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DalCin:1971:IPL

[DPI71] Mario Dal Cin, Arnold Perlmutter, and Geoffrey J. Iverson, editors. Invited papers: lectures from the Coral Gables conference on fundamental interactions at high energy, January 20–22, 1971, volume 3 of Tracts in mathematics and natural sciences. Gordon and Breach Science Publishers, New York, NY, USA, 1971. LCCN ????

Dirac:1942:LIQ

[DPP42] P. A. M. Dirac, Rudolf Peierls, and Maurice H. L. Pryce. On Lorentz invariance in the quantum theory. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 38(2):193–200, January 1942. CODEN PCPSA4.

ISSN 0008-1981. URL https://ui.adsabs.harvard.edu/#
abs/1942PCPS...38..193D.

Dirac:1997:LIQ

[DPP97] P. A. M. Dirac, R. Peierls, and M. H. L. Pryce. On Lorentz invariance in the quantum theory. In Dalitz and Peierls [DP97], pages 321–328. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2

.P42 1997. URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book..321D.

Dutt:1993:DFP

[DR93] Ranabir Dutt and Asim K. Ray, editors. Dirac and Feynman: pioneers in quantum mechanics. Wiley Eastern, New Delhi, India, 1993. ISBN 81-224-0493-6. viii + 214 pp. LCCN QC174.26.W28 D39 1993.

Drago:2023:DBE

[Dra23] Antonino Drago. Dirac’s book The Principles of Quantum Mechanics as an alternative way of organizing a theory. Foundations of Science, 28(2):551–574, June 2023. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL https://

link.springer.com/article/10.1007/s10699-022-09835-3.

Dresden:1990:BRD

[Dre90] M. Dresden. Book review: Dirac. A Scientific Biography. Helge Kragh. Cambridge University Press, New York, 1990. x, 389 pp., illus. $44.50. Science, 249(4971):937, August 24, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Duplantier:2017:DM

[DRF17] Bertrand Duplantier, Vincent Rivasseau, and Jean-Nöel Fuchs, editors. Dirac matter, volume 71 of Progress in mathematical physics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2017. ISBN 3-319-32535-3 (print), 3-319-32536-1 (e-book). ISSN 1544-9998. LCCN QC174.45; QA1-939.

URL http://lib.myilibrary.com?id=990350;   http://
link.springer.com/10.1007/978-3-319-32536-1;   http:
//link.springer.com/openurl?genre=book\%26isbn=978-3-
319-32535-4;   http://public.eblib.com/choice/PublicFullRecord.
aspx?p=4791271.

Dirac:1980:DRK

[DRFS80] P. A. M. Dirac, Joseph Rotblat, Bernard T. Feld, and Grace Marmor Spruch. Dirac recalls Kapitza. Physics Today, 33 (5):81–82, May 15, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Dirac:1983:ANP

[DS83] Paul Dirac and Erwin Schrödinger. 50th anniversary of the Nobel Prize for Schrödinger and Dirac. Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 25(58)(4):277–297, 1983. CODEN FMBMAC. ISSN 0015-3265. Translated from the French and German by D. Vachov.

Davies:1972:DDI

[DSK72] Alan R. Davies, Kenneth Smith, and K. L. Kwok. Dirac, dynamic information retrieval of atomic codes: 1. Physical design criteria. Computer Physics Communications, 3(4):277–295, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557290077X.

Dong:2003:GTA

[DSP03] Shi-Hai Dong, Guo-Hua Sun, and Dusan Popov. Group theory approach to the Dirac equation with a Coulomb plus scalar potential in D + 1 dimensions. Journal of Mathematical Physics, 44(10):4467–4479, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dixit:1992:SDO

[DST92] V. V. Dixit, T. S. Santhanam, and W. D. Thacker. A supersymmetric Dirac oscillator with scalar coupling. Journal of Mathematical Physics, 33(3):1114–1117, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dent:2008:UCR

[DSW08] Thomas Dent, Steffen Stern, and Christof Wetterich. Unifying cosmological and recent time variations of fundamental couplings. Physical Review D (Particles and Fields), 78(10):103518, November 15, 2008. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.78.103518.

DuMond:1928:ECS

[DuM28]  Jesse W. M. DuMond.  Experimental confirmation for

Sommerfeld–Fermi–Dirac degenerate gas theory of conduction electrons. Science, 68(1767):452, November 9, 1928. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/1652417;   http://
www.sciencemag.org/content/68/1767/452.1.full.pdf.

Duncan:2012:CFQ

[Dun12] Anthony Duncan. The Conceptual Framework of Quantum Field Theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-957326-3. x + 782 pp. LCCN

QC174.45 .D86 2012. URL http://www.loc.gov/catdir/
enhancements/fy1306/2012472641-b.html;   http://www.
loc.gov/catdir/enhancements/fy1306/2012472641-d.html;
 http://www.loc.gov/catdir/enhancements/fy1306/2012472641-
t.html.

Dong:2009:LAC

[DW09] Huan-He Dong and Xin-Zeng Wang. A Lie algebra containing four parameters for the generalized Dirac hierarchy. Applied Mathematics and Computation, 215(2):459–463, September 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dyson:1967:TVC

[Dys67] Freeman J. Dyson. Time variation of the charge of the proton. Physical Review Letters, 19(22):1291–1293, November 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.19.1291.

Dyson:1972:FCT

[Dys72] F. J. Dyson. The fundamental constants and their time variation. In Salam and Wigner [SW72], pages 213–236. ISBN 0-521-08600-0. LCCN QC174.1 .A85 1972. URL http://hooke.lib.

cam.ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;
http://www.loc.gov/catdir/enhancements/fy1001/72075298-
d.html;   http://www.loc.gov/catdir/enhancements/fy1001/
72075298-t.html.

Dyson:1974:BRH

[Dys74] Freeman Dyson. Book review: Honoring Dirac: The Physicist’s Conception of Nature, by Jagdish Mehra. Proceedings of a symposium, Trieste, Italy, Sept. 1972. Science, 185(4157):1160– 1161, September 27, 1974. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1738373.

Dyson:1978:VC

[Dys78a] F. J. Dyson. Variation of constants. In Lannutti and Williams [LW78], pages 163–167. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

Dyson:1978:VCa

[Dys78b] F. J. Dyson. Variation of constants. In Lannutti and Williams [LW78], pages 163–167. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

Dyson:1987:PMD

[Dys87] Freeman J. Dyson. Paul A. M. Dirac. American Philosophical Society Year Book for 1986, pages 100–104, 1987. CODEN YAPSAL. ISSN 0065-9762. Obituary notice.

Dyson:1993:GGP

[Dys93] Freeman Dyson. George Green and physics. Physics World, 6(8):33–38, August 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL iopscience.iop.

org/pwa/full/pwa-pdf/6/8/phwv6i8a28.pdf.

Dyson:2007:AQM

[Dys07a] Freeman Dyson. Advanced Quantum Mechanics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-270-661-5 (paperback), 981-270-

622-4. xiv + 220 pp. URL http://www.gbv.de/dms/bowker/
toc/9789812706614.pdf;   http://www.worldscientific.
com/worldscibooks/10.1142/6427. Transcribed and with a

preface by David Derbes.

Dyson:2007:DT

[Dys07b] Freeman J. Dyson. The Dirac theory. In Advanced Quantum Mechanics [Dys07a], pages 5–30. ISBN 981-270-661-5 (paperback),

981-270-622-4. URL http://www.worldscientific.com/doi/
abs/10.1142/9789812708212_0002. Transcribed and with a

preface by David Derbes.

Dyson:2010:BRS

[Dys10] Freeman J. Dyson. Book review: Silent quantum genius: The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom, by Graham Farmelo. Basic Books, 539 pp., $29.95. New York Review of Books, 57(3):20–??, February 25, 2010. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2010/feb/25/silent-quantum-genius/. See response [GD10].

Esposito:2015:PEM

[EAW15] Salvatore Esposito, E. (Evgeny) Akhmedov, and Frank Wilczek. The physics of Ettore Majorana: phenomenological, theoretical, and mathematical. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-04402-2 (hardcover), 1-316-19108-7 (PDF ebook). xi + 382 pp. LCCN QC19.6 .E87 2015.

Eddington:1946:FT

[Edd46] Sir Arthur Eddington. Fundamental theory. Cambridge University Press, Cambridge, UK, 1946. viii + 292 pp. LCCN ????

Eeg:1980:CDS

[Eeg80] J. O. Eeg. Calculations of Dirac-spinor amplitudes by means of complex Lorentz-transformations and trace calculus. Journal of Mathematical Physics, 21(1):170–174, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Einstein:1977:CSD

[EF77] S. Einstein and R. Finkelstein. A class of solutions of the Dirac equation in the Kerr–Newman space. Journal of Mathematical Physics, 18(4):664–671, April 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ellis:1984:SCD

[EG84] D. E. Ellis and G. L. Goodman. Self-consistent Dirac–Slater calculations for molecules and embedded clusters. International Journal of Quantum Chemistry, 25(1):185–200, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Erdogan:2021:WLD

[EGT21] M. Burak Erdoğan, William R. Green, and Ebru Toprak. What is … the Dirac equation? Notices of the American Mathematical Society, 68(10):1782–1785, 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ehberger:2022:LDT

[Ehb22] Markus Ehberger. “The language of Dirac’s theory of radiation”: the inception and initial reception of a tool for the quantum field theorist. Archive for History of Exact Sciences, 76(6): 531–571, November 2022. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.springer.

com/article/10.1007/s00407-022-00293-8. See correction

[Ehb23].

Ehberger:2023:CLD

[Ehb23] Markus Ehberger. Correction to: “The language of Dirac’s theory of radiation”: the inception and initial reception of a tool for the quantum field theorist. Archive for History of Exact Sciences, 77(1):121–122, January 2023. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL https://link.

springer.com/article/10.1007/s00407-022-00300-y. See

[Ehb22].

Eliav:1994:RCC

[EKI94] E. Eliav, U. Kaldor, and Y. Ishikawa. Relativistic coupled cluster theory based on the No-Pair Dirac–Coulomb–Breit Hamiltonian: Relativistic pair correlation energies of the Xe atom. International Journal of Quantum Chemistry. Symposium, 28 (??):205–??, 1994. CODEN IJQSAF. ISSN 0538-821X.

Eliezer:1987:SRP

[Eli87] C. J. Eliezer. Some reminiscences of P. A. M. Dirac. In Taylor [Tay87b], pages 58–60. ISBN 0-85274-480-3. LCCN QC16.D57

T75 1987. US$10.00. URL http://www.loc.gov/catdir/
enhancements/fy0745/87153334-d.html. Based on the papers

presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Ellis:1991:BRU

[Ell91] John Ellis. Book review: Unification of Fundamental Forces: The First of the 1988 Dirac Memorial Lectures, by Abdus Salam. Physics Today, 44(7):59–60, July 1991. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Einstein:1933:SNF

[EM33a] Albert Einstein and Walter Mayer. Spaltung der natürlichsten Feldgleichungen für Semi-Vektoren in Spinor-Gleichungen vom Diracschen Typus. (German) [Splitting the most natural field equations for semivectors into spinor equations of the Dirac type]. Proc. Acad. Sci. Amst., 36 (part 2)(??):615–619, 1933.

Einstein:1933:DGS

[EM33b] Albert Einstein and Walther Mayer. Dirac-Gleichung für Semi-Vektoren. (German) [Dirac equations for semivectors]. Proc. Acad. Sci. Amst., 36 (part 2)(??):497–502, 1933.

Eden:1972:DC

[EP72] R. J. Eden and J. C. Polkinghorne. Dirac in Cambridge. In Salam and Wigner [SW72], pages 1–5. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Epstein:1935:BRP

[Eps35] Paul S. Epstein. Book review: Principles of Quantum Mechanics, by P. A. M. Dirac. Science, 81(2113):640–641, June 28, 1935. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1658937.

Ellis:1975:ADS

[ERW75] D. E. Ellis, A. Rosén, and P. F. Walch. Applications of the Dirac–Slater model to molecules. International Journal of Quantum Chemistry, 9(S9):351–358, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Esposito:2010:FTL

[ES10] Salvatore Esposito and Giovanni Salesi. Fundamental times, lengths and physical constants: Some unknown contributions by Ettore Majorana. Annalen der Physik (1900), 522(7):456–466, July 2010. ISSN 0003-3804 (print), 1521-3889 (electronic).

Esposito:2008:EMH

[Esp08] Salvatore Esposito. Ettore Majorana and his heritage seventy years later. Annalen der Physik (1900), 17(5):302– 318, May 2008. ISSN 0003-3804 (print), 1521-3889 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2008AnP..
.520..302E;   http://onlinelibrary.wiley.com/doi/10.
1002/andp.200810296/.

Esposito:2012:SED

[Esp12] Salvatore Esposito. Searching for an equation: Dirac, Majorana and the others. Annals of Physics, 327(6):1617–1644, June 2012. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0003491612000346.

Ettlinger:1931:BRP

[Ett31] H. J. Ettlinger. Book review: The Principles of Quantum Mechanics., by P. A. M. Dirac. American Mathematical Monthly, 38(9):524, November 1931. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2300785.

Escobedo:1997:SDE

[EV97] M. Escobedo and L. Vega. A semilinear Dirac equation in H^{s}(R^{3}) for s > 1. SIAM Journal on Mathematical Analysis, 28 (2):338–362, March 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://epubs.siam.org/

sam-bin/dbq/article/28301.

Evans:2004:DDE

[Eva04] M. W. Evans. Derivation of Dirac’s equation from the Evans wave equation. Foundations of Physics Letters, 17(2):149–166, April 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Frommer:2000:NCL

[F^{+}00] Andreas Frommer et al., editors. Numerical challenges in lattice quantum chromodynamics: joint interdisciplinary workshop of John von Neumann Institute for Computing, Jülich, and Institute of Applied Computer Science, Wuppertal University, August 1999, volume 15 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67732-1. ISSN 1439-7358. LCCN QC793.3.G38 N86 2000.

Fakhri:2001:ESD

[FA01] H. Fakhri and N. Abbasi. Exact solution of the Dirac equation for a spin-(1/2 ) charged particle in two-dimensional and three-dimensional Euclidean spaces with shape invariance symmetry. Journal of Mathematical Physics, 42(6):2416–2437, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrero:2010:LTV

[FA10] Alejandro Ferrero and Brett Altschul. Limits on the time variation of the Fermi constant G_{F} based on type Ia supernova observations. Physical Review D (Particles and Fields), 82(12): 123002, December 15, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.82.123002.

Fabbri:2022:BBF

[Fab22] Luca Fabbri. de Broglie–Bohm formulation of Dirac fields. Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

link.springer.com/article/10.1007/s10701-022-00641-2.

Falconer:1997:JJT

[Fal97] Isobel Falconer. J J Thomson and the discovery of the electron. Physics Education, 32(4):226–231, July 1997. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL

http://iopscience.iop.org/0031-9120/32/4/015.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Farmelo:2003:IMB

[Far03] Graham Farmelo, editor. It Must Be Beautiful: Great Equations of Modern Science. Granta, London, UK, 2003. ISBN 1-86207-555-7. xviii + 284 pp. LCCN Q125 .I88 2003.

Farmelo:2009:PDM

[Far09a] Graham Farmelo. Paul Dirac, a man apart. Physics Today, 62 (11):46–50, November 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Farmelo:2009:SMH

[Far09b] Graham Farmelo. The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom. Basic Books, New York, NY, USA, 2009. ISBN 0-465-01827-0. 539 + 8 pp. LCCN QC16.D57; QC16.D57 F37 2009. US$29.95. URL http://bnreview.

barnesandnoble.com/t5/The-Thinking-Read/The-Strangest-
Man-The-Hidden-Life-of-Paul-Dirac-Mystic-of-the/ba-p/
1243.

Farmelo:2010:DDP

[Far10] Graham Farmelo. Did Dirac predict the positron? Contemporary Physics, 51(2):97–101, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Farmelo:2016:SMV

[Far16] Graham Farmelo. Der seltsamste Mensch: das verborgene Leben des Quantengenies Paul Dirac. (German) [The strangest man: the hidden life of the quantum genius Paul Dirac]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-49949-5. xxiii + 614 pp. LCCN QC16.D57 F3715 2016. Translation to German by Reimara Rössler of [Far09b].

Fazarinc:2015:FDB

[Faz15] Zvonko Fazarinc. Fermi–Dirac, Bose–Einstein, Maxwell– Boltzmann, and computers. Computer Applications in Engineering Education, 23(5):746–759, September 2015. CODEN CAPEED. ISSN 1061-3773 (print), 1099-0542 (electronic).

URL https://dblp.org/db/journals/caee/caee23.html#
Fazarinc15;   https://www.wikidata.org/entity/Q60032924.

Feynman:2005:FTN

[FBD05] Richard P. (Richard Phillips) Feynman, Laurie M. Brown, and P. A. M. (Paul Adrien Maurice) Dirac. Feynman’s Thesis: a New Approach to Quantum Theory. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2005. ISBN 981-256-366-0, 981-256-380-6 (paperback). xxii + 119 pp. LCCN QC174.12 .F48 2005.

Feinberg:2002:DOF

[Fei02] Joshua Feinberg. The Dirac operator in a fermion bag background in 1 + 1 dimensions and generalized supersymmetric quantum mechanics. Journal of Mathematical Physics, 43(8): 3927–3936, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fermi:1924:CSQ

[Fer24] Enrico Fermi. Considerazioni sulla quantizzazione dei sistemi che contengono degli elementi identici. (Italian) [Considerations on the quantization of systems that contain identical elements]. Il Nuovo Cimento (8), 1(1):145–152, December 1924. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121

(electronic). URL http://www.springerlink.com/content/
97203r302q8w7unj/.

Fermi:1926:QIE

[Fer26] Enrico Fermi. Zur Quantelung des idealen einatomigen Gases. (German) [On the quantization of an ideal monatomic gas]. Zeitschrift für Physik, 36(11–12):902–912, November 1926. CO-

DEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/k763270092273181/fulltext.pdf.

Fermi:1949:DM

[Fer49] Enrico Fermi. The Dirac monopole. Report VPI-EPP-5-86, ????, ????, 1949.

FernandezVelicia:1986:HPA

[Fer86] F. J. Fernández Velicia. High-precision analytic approximations for the Fermi–Dirac functions by means of elementary functions. Physical Review A (3), 34(5):4387–4395, 1986. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Feynman:1948:STA

[Fey48] R. P. Feynman. Space–time approach to non-relativistic quantum mechanics. Reviews of Modern Physics, 20(2):367–387, April 1, 1948. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Falkensteiner:1987:UIG

[FG87] P. Falkensteiner and H. Grosse. Unitary implementability of gauge transformations for the Dirac operator and the Schwinger term. Journal of Mathematical Physics, 28(4):850–854, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Furlan:2022:EVA

[FG22] Stefano Furlan and Rocco Gaudenzi. The earth vibrates with analogies: the Dirac sea and the geology of the vacuum. Studies in History and Philosophy of Science Part A, 93(??):163– 174, June 2022. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368122000590.

Falomir:1996:DDO

[FGM^{+}96] H. Falomir, R. E. Gamboa Saraví, M. A. Muschietti, E. M. Santangelo, and J. E. Solomin. Determinants of Dirac operators with local boundary conditions. Journal of Mathematical Physics, 37(11):5805–5819, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feynman:1965:QMP

[FH65] Richard P. (Richard Phillips) Feynman and Albert R. Hibbs. Quantum mechanics and path integrals. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1965. xiv + 365 pp. LCCN QC174.1 .F39.

Feynman:2010:QMP

[FHS10] Richard P. (Richard Phillips) Feynman, Albert R. Hibbs, and Daniel F. Styer. Quantum mechanics and path integrals. Dover, New York, NY, USA, 2010. ISBN 0-486-47722-3. xii + 371 pp. LCCN QC174.12 .F484 2010. URL http://www.loc.gov/

catdir/enhancements/fy1006/2010004550-d.html.

Finster:2000:LCE

[Fin00] Felix Finster. Light-cone expansion of the Dirac sea in the presence of chiral and scalar potentials. Journal of Mathematical Physics, 41(10):6689–6746, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiorini:2006:NDM

[Fio06] E. Fiorini. Neutrinos: Dirac or Majorana? In Rapisarda et al. [RCC^{+}06], page ?? LCCN ???? URL http://pos.sissa.it/

archive/conferences/037/007/EMC2006_007.pdf.

Fischer:2010:HQE

[Fis10a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2010:PMD

[Fis10b] Ernst Peter Fischer. Paul A. M. Dirac (1902–1984). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis10a], pages 199–208. ISBN 3-7766-2643-7. LCCN ????

Fischer:2012:HQE

[Fis12a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fischer:2012:PMD

[Fis12b] Ernst Peter Fischer. Paul A. M. Dirac (1902–1984). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis12a], pages 199–208. ISBN 3-596-19406-7. LCCN ????

Franke:1970:SDB

[FK70] W. H. Franke and A. J. Kalnay. Symmetric Dirac bracket in classical mechanics. Journal of Mathematical Physics, 11(5): 1729–1734, May 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flambaum:2007:EST

[FK07] V. V. Flambaum and M. G. Kozlov. Enhanced sensitivity to the time variation of the fine-structure constant and m_{p}/m_{e} in diatomic molecules. Physical Review Letters, 99(15):150801, October 12, 2007. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/doi/

10.1103/PhysRevLett.99.150801.

Fu:2008:SPE

[FK08] Liang Fu and C. L. Kane. Superconducting proximity effect and Majorana fermions at the surface of a topological insulator. Physical Review Letters, 100(9):096407, March 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.100.096407.

Flato:1976:DPB

[FLS76] M. Flato, A. Lichnerowicz, and D. Sternheimer. Deformations of Poisson brackets, Dirac brackets and applications. Journal of Mathematical Physics, 17(9):1754–1762, September 1976. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Felber:1975:NCE

[FM75] F. S. Felber and J. H. Marburger. New class of exact solutions of the Dirac equation. Journal of Mathematical Physics, 16(10): 2089–2092, October 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ford:1963:MM

[For63] Kenneth W. Ford. Magnetic monopoles. Scientific American, 209(6):122–131, December 1963. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.nature.com/scientificamerican/journal/v209/n6/
pdf/scientificamerican1263-122.pdf.

Fullerton:1986:GFD

[FR86] L. W. Fullerton and G. A. Rinker. Generalized Fermi– Dirac integrals — FD, FDG, FDH. Computer Physics Communications, 39(2):181–185, February/March 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465586901293.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Franklin:2004:TRN

[Fra04] Allan Franklin. Are there really neutrinos?: an evidential history. Perseus Publishers, Cambridge, MA, USA, 2004. ISBN 0-7382-0265-7, 0-8133-4128-0 (paperback). ix + 371 pp. LCCN

QC793.5.N42 F73 2004. URL http://www.loc.gov/catdir/
enhancements/fy0837/2008530710-d.html;   http://www.
loc.gov/catdir/toc/fy0805/2008530710.html.

Frenkel:1987:PDS

[Fre87] V. Ya. Frenkel. Professor Dirac and Soviet physicists. Uspekhi Fizicheskikh Nauk, 153(9):173–186, September 1987. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1987/9/f/. English translation

in Sov. Phys. Usp. 30 816–822 (1987).

Freund:1995:DMS

[Fre95] Peter G. O. Freund. Dirac monopoles and the Seiberg–Witten monopole equations. Journal of Mathematical Physics, 36(6): 2673–2674, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falomir:1998:DOD

[FSS98] H. Falomir, R. E. Gamboa Saraví, and E. M. Santangelo. Dirac operator on a disk with global boundary conditions. Journal of Mathematical Physics, 39(1):532–544, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2000:NET

[FSY00] Felix Finster, Joel Smoller, and Shing-Tung Yau. Non-existence of time-periodic solutions of the Dirac equation in a Reissner– Nordström black hole background. Journal of Mathematical Physics, 41(4):2173–2194, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fukushima:2014:CGI

[Fuk14] Toshio Fukushima. Computation of a general integral of Fermi–Dirac distribution by McDougall–Stoner method. Applied Mathematics and Computation, 238(??):485–510, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S009630031400561X.

Feynman:1987:EPL

[FW87] Richard P. (Richard Phillips) Feynman and Steven Weinberg. Elementary particles and the laws of physics: the 1986 Dirac memorial lectures. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34000-4 (hardcover), 0-521-65862-4 (paperback). x + 110 pp. LCCN QC793.28

.F49 1987. URL http://www.loc.gov/catdir/description/
cam023/87026362.html;   http://www.loc.gov/catdir/toc/
cam023/87026362.html.

Fushchich:1991:NLR

[FZ91] W. I. Fushchich and R. Z. Zhdanov. On the non-Lie reduction of the nonlinear Dirac equation. Journal of Mathematical Physics, 32(12):3488–3490, December 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gabrielse:2013:SMG

[Gab13] Gerald Gabrielse. The standard model’s greatest triumph. Physics Today, 66(12):64–65, December 2013. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Gamow:1933:TDE

[Gam33] George Gamow. The theory of Dirac electrons and positive. Sorena, 8(??):25–30, ???? 1933. CODEN ???? ISSN ????

Gamow:1966:TYSa

[Gam66a] George Gamow. Thirty years that shook physics; the story of quantum theory. Doubleday, Garden City, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3 1966b.

Gamow:1966:TYSb

[Gam66b] George Gamow. Thirty years that shook physics; the story of quantum theory, volume S45 of Science study series. Anchor Books, New York, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3.

Gamow:1967:DGC

[Gam67a] George Gamow. Does gravity change with time? Proceedings of the National Academy of Sciences of the United States of America, 57(2):187–193, February 15, 1967. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http:/

/adsabs.harvard.edu/abs/1967PNAS...57..187G;   http:
//www.jstor.org/stable/pdfplus/57929.pdf.

Gamow:1967:EGC

[Gam67b] George Gamow. Electricity, gravity, and cosmology. Physical Review Letters, 19(13):759–761, September 25, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://adsabs.harvard.edu/abs/1967PhRvL..
19..759G;   http://prl.aps.org/abstract/PRL/v19/i13/

p759_1. See erratum [Gam67c].

Gamow:1967:EEG

[Gam67c] George Gamow. Erratum: “Electricity, Gravity, and Cosmology”. Physical Review Letters, 19(17):1000, October 23, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://adsabs.harvard.edu/abs/
1967PhRvL..19.1000G. See [Gam67b].

Gamow:1967:VEC

[Gam67d] George Gamow. Variability of elementary charge and quasistellar objects. Physical Review Letters, 19(16):913–914, October 16, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://prl.aps.org/abstract/
PRL/v19/i16/p913_1.

Gamow:1968:NCN

[Gam68] George Gamow. Numerology of the constants of nature. Proceedings of the National Academy of Sciences of the United States of America, 59(2):313–318, February 15, 1968. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/pdfplus/58638.pdf.

Gamow:1972:TYS

[Gam72] George Gamow. Thirty years that shook physics; the story of quantum theory, volume 38 of Science study series. Heinemann Educational, London, UK, 1972. ISBN 0-435-55071-3. viii + 224 + 12 pp. LCCN QC173.98 .G35 1972.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12

.G35 1985. US$4.95. URL http://www.loc.gov/catdir/
description/dover032/85006797.html.

Garavaglia:1984:DMN

[Gar84] Theodore Garavaglia. Dirac- and Majorana-neutrino-mass effects in neutrino-electron elastic scattering. Physical Review D (Particles and Fields), 29(3):387–392, February 1984. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500.

Gautschi:1993:CGF

[Gau93] Walter Gautschi. On the computation of generalized Fermi– Dirac and Bose–Einstein integrals. Computer Physics Communications, 74(2):233–238, February 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/001046559390093R.

Galvao:1990:GTD

[GB90] Carlos A. P. Galvão and João Batista T. Boechat. Gauge transformations in Dirac theory of constrained systems. Journal of Mathematical Physics, 31(2):448–451, February 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gitman:2013:DES

[GB13] Dmitry Gitman and Vladislav G. Bagrov. The Dirac Equation and its Solutions, volume 4 of De Gruyter Studies in Mathematical Physics. Walter de Gruyter, Berlin, Germany, 2013. ISBN 3-11-026329-7. ???? pp. LCCN ???? URL http://www.

degruyter.com/view/product/177851.

Gerward:1997:DEC

[GC97] Leif Gerward and Christopher Cousins. The discovery of the electron: a centenary. Physics Education, 32(4):219–225, July 1997. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552

(electronic). URL http://iopscience.iop.org/0031-9120/
32/4/014/.

Gluck:2010:DMA

[GD10] Malcolm Gluck and Freeman J. Dyson. Dirac’s model airplane. New York Review of Books, 57(5):46, March 25, 2010. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/2010/
mar/25/diracs-model-airplane/. In response to [Dys10].

Galiautdinov:2002:CCD

[GF02] Andrei A. Galiautdinov and David R. Finkelstein. Chronon corrections to the Dirac equation. Journal of Mathematical Physics, 43(10):4741–4752, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garoni:2001:CAE

[GFG01] T. M. Garoni, N. E. Frankel, and M. L. Glasser. Complete asymptotic expansions of the Fermi–Dirac integrals F_{p}(η) = 1/Γ(p+1) ∫_{0}^{∞}[ϵ^{p}/(1+e^{ϵ−η})]dϵ. Journal of Mathematical Physics, 42(4):1860–1868, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gombas:1951:STF

[GG51] P. Gombás and R. Gáspár. Solution of the Thomas–Fermi– Dirac equation. Nature, 168(4264):122, July 21, 1951. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v168/n4264/
pdf/168122a0.pdf.

Gamow:1976:GGB

[GG76] George Gamow and Barbara Gamow. George Gamow and Barbara Gamow papers, 1915–1975, 1976. URL http:

//hdl.loc.gov/loc.mss/eadmss.ms010191;   http://hdl.
loc.gov/loc.mss/eadmss.ms010191.3.

Gadella:2002:UMF

[GG02] M. Gadella and F. Gómez. A unified mathematical formalism for the Dirac formulation of quantum mechanics. Foundations of Physics, 32(6):815–869, June 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1023/A%3A1016069311589.

Goldhaber:2011:CDM

[GG11] Alfred Scharff Goldhaber and Maurice Goldhaber. Clarifying Dirac and Majorana distinctions. Physics Today, 64(7):12, July 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/
1/phtoad/v64/i7/p12_s3.

Galison:2002:QMS

[GGK02] Peter Galison, Michael Gordin, and David Kaiser, editors. Quantum Mechanics: Science and Society. Routledge, London, UK, 2002. ISBN 1-136-70972-X. 433 (est.) pp. LCCN ????

Guilpin:2004:AAD

[GGS04] Christian Guilpin, Jacques Gacougnolle, and Yvan Simon. The ϵ-algorithm allows to detect Dirac delta functions. Applied Numerical Mathematics: Transactions of IMACS, 48(1):27–40, January 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gurtler:1975:CFD

[GH75] R. Gurtler and D. Hestenes. Consistency in the formulation of the Dirac, Pauli, and Schrödinger theories. Journal of Mathematical Physics, 16(3):573–584, March 1975. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gsponer:1994:LER

[GH94] André A. Gsponer and Jean-Pierre Hurni. Lanczos’s equation to replace Dirac’s equation? In Brown et al. [BCEP94], pages 509–512. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Ginoux:2009:DS

[Gin09] Nicolas Ginoux. The Dirac spectrum, volume 1976 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-01569-7 (paperback), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 156 pp. LCCN QA614.95

.G56 2009. URL http://www.springerlink.com/openurl.
asp?genre=issue&issn=0075-8434&1976.

Gilinsky:2010:RD

[GJEF10] Victor Gilinsky, Ted Jacobson, Ron Edge, and Graham Farmelo. Remembrances of Dirac. Physics Today, 63(5):59–60, May 2010.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Galeao:1992:GMR

[GL92] A. P. Galeão and P. Leal Ferreira. General method for reducing the two-body Dirac equation. Journal of Mathematical Physics, 33(7):2618–2625, July 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Gaftoi:1994:LDE

[GLMR94] Violeta Gaftoi, José L. López Bonilla, Jesús Morales, and Marco A. Rosales. The Lorentz–Dirac equation in Minkowski space. Journal of Mathematical Physics, 35(7):3482–3489, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerra:2014:WEC

[GLR14] Francesco Guerra, Matteo Leone, and Nadia Robotti. When energy conservation seems to fail: The prediction of the neutrino. Science & Education (Springer), 23(6):1339–1359, June 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Gruber:1980:SS

[GM80] Bruno Gruber and Richard S. Millman, editors. Symmetries in science. Plenum Press, New York, NY, USA; London, UK, 1980. ISBN 0-306-40541-5. LCCN Q172.5.S95 S92.

Goldsmith:1980:EFH

[GMW80] Maurice Goldsmith, Alan L. (Alan Lindsay) Mackay, and James Woudhuysen, editors. Einstein, the first hundred years. Pergamon, New York, NY, USA, 1980. ISBN 0-08-025019-X. xiii + 200 pp. LCCN QC16.E5 E52 1980.

Gotay:1978:PMD

[GNH78] Mark J. Gotay, James M. Nester, and George Hinds. Presymplectic manifolds and the Dirac–Bergmann theory of constraints. Journal of Mathematical Physics, 19(11):2388–2399, November 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goano:1995:ACC

[Goa95] Michele Goano. Algorithm 745: Computation of the complete and incomplete Fermi–Dirac integral. ACM Transactions on Mathematical Software, 21(3):221–232, September 1995. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/citations/journals/toms/
1995-21-3/p221-goano/. See remark [Goa97].

Goano:1997:RA7

[Goa97] Michele Goano. Remark on Algorithm 745. ACM Transactions on Mathematical Software, 23(2):295, June 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Goa95].

Goddard:1987:MM

[God87] Peter Goddard. Magnetic monopoles. In Taylor [Tay87b], pages 104–113. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Goddard:1998:PDM

[God98] Peter Goddard, editor. Paul Dirac: the man and his work. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58382-9 (hardcover), 0-521-01953-2 (paperback). xv + 124 pp. LCCN QC16.D57 P38 1998. URL http://www.loc.

gov/catdir/description/cam028/97022443.html;   http:
//www.loc.gov/catdir/toc/cam021/97022443.html.

Goenner:2016:BRH

[Goe16] Hubert Goenner. Book review: Helge Kragh. Varying Gravity: Dirac’s Legacy in Cosmology and Geophysics. Isis, 107(4):885– 887, December 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goldman:2005:DRE

[Gol05] Josh D. Goldman. Dirac’s relativistic electron theory and the so-called prediction of the positron. M.Sc. dissertation, Imperial College of Science and Technology, London, UK, 2005. ???? pp.

Good:1955:PDM

[Goo55] R. H. Good, Jr. Properties of the Dirac matrices. Reviews of Modern Physics, 27(2):187–211, April 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.27.187;
http://rmp.aps.org/abstract/RMP/v27/i2/p187_1.

Gottfried:2011:PMD

[Got11] Kurt Gottfried. P. A. M. Dirac and the discovery of quantum mechanics. American Journal of Physics, 79(3):261–266, March 2011. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v79/i3/p261_s1.

Guinea:1986:SSC

[GPS86] F. Guinea, R. E. Peierls, and R. Schrieffer. Shape of solitons in classically forbidden states: “Lorentz expansion”. Physica Scripta, 33:282–283, March 1986. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1986PhyS...33..282G.

Greub:1977:CLB

[Gre77] Werner H. Greub. Complex line bundles and the magnetic field of a monopole. Lecture Notes in Mathematics, 570:350– 354, 1977. CODEN LNMAA2. ISBN 3-540-08068-6 (print), 3-540-37498-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0087790/.

Greenspan:2005:ECW

[Gre05] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Grujic:2000:ED

[Gru00] Petar Grujic. Einstein 1, Dirac 1. Physics World, 13(5):19– 20, May 2000. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://physicsworldarchive.
iop.org/full/pwa-pdf/13/5/phwv13i5a19.pdf.

Glassey:1979:CLC

[GS79] R. T. Glassey and W. A. Strauss. Conservation laws for the classical Maxwell–Dirac and Klein–Gordon–Dirac equations. Journal of Mathematical Physics, 20(3):454–458, March 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Glasser:1984:DEL

[GS84] M. L. Glasser and N. Shawagfeh. Dirac equation for a linear potential. Journal of Mathematical Physics, 25(8):2533–2537, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gleiser:1985:ETV

[GT85a] M. Gleiser and J. G. Taylor. Erratum: Time variation of coupling constants in Kaluza–Klein cosmologies. Physical Review D (Particles and Fields), 32(12):3337, December 15, 1985. CO-DEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic),

1538-4500 (CD-ROM). URL http://link.aps.org/doi/10.
1103/PhysRevD.32.3337.2. See [GT85b].

Gleiser:1985:TVC

[GT85b] M. Gleiser and J. G. Taylor. Time variation of coupling constants in Kaluza–Klein cosmologies. Physical Review D (Particles and Fields), 31(8):1904–1910, April 15, 1985. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/

PhysRevD.31.1904. See corrigendum [GT85a].

Gleiser:1986:TVC

[GT86] M. Gleiser and J. G. Taylor. Time variation of coupling constants in Kaluza–Klein cosmologies reexamined. Physical Review D (Particles and Fields), 33(2):570–571, January 15, 1986. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.33.570.

Gunderman:2015:LCL

[Gun15] Richard Gunderman. The life-changing love of one of the 20th century’s greatest physicists. The Conversation, ??(??):??, De-

cember 8, 2015. URL https://getpocket.com/explore/item/
the-life-changing-love-of-one-of-the-20th-century-s-
greatest-physicists.

Gong:2001:GFD

[GZDA01] Zhigang Gong, Ladislav Zejda, Werner Däppen, and Josep M. Aparicio. Generalized Fermi–Dirac functions and derivatives: properties and evaluation. Computer Physics Communications, 136(3):294–309, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550100145X.

Gorrec:2005:DSB

[GZM05] Y. Le Gorrec, H. Zwart, and B. Maschke. Dirac structures and boundary control systems associated with skew-symmetric differential operators. SIAM Journal on Control and Optimization, 44(5):1864–1892, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/61167.

Haas:1934:MQE

[Haa34] Arthur Haas. Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. 4. u. 5. verb. u. abermals wes. verm. Aufl. (German) [Matter waves and quantum mechanics. An introduction based on the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Akad. Verlagsges., Leipzig, Germany, 1934. viii + 299 + 7 pp.

Halpern:1978:MCV

[Hal78] Leopold Halpern, editor. On the measurement of cosmological variations of the gravitational constant: proceedings of the workshop meeting held November 12–14, 1975, at the Department of Physics, Florida State University, Tallahassee: host, P. A. M. Dirac, Monograph publishing on demand: Imprint series. University Presses of Florida, Gainesville, FL, USA, 1978. ISBN 0-8130-0610-4. LCCN QC178 .O4.

Halpern:1985:PAM

[Hal85] Leopold Halpern. Paul Adrien Maurice Dirac (1902–1984). Foundations of Physics, 15(3):257–259, March 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00737316.

Hall:1987:DSO

[Hal87] Richard L. Hall. Dirac spinor orbits. Journal of Mathematical Physics, 28(2):457–462, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamilton:1984:DEH

[Ham84] J. Dwayne Hamilton. The Dirac equation and Hestenes’ geometric algebra. Journal of Mathematical Physics, 25(6):1823–1832, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hannestad:1999:PCT

[Han99] Steen Hannestad. Possible constraints on the time variation of the fine structure constant from cosmic microwave background data. Physical Review D (Particles and Fields), 60(2):023515, June 15, 1999. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.60.023515.

Hartree:1928:WMA

[Har28] D. R. Hartree. The wave mechanics of an atom with a noncoulomb central field. Part I. Theory and methods. Mathematical proceedings of the Cambridge Philosophical Society, 24(1):89–110, ???? 1928. CODEN PCPSA4. ISSN 0008-1981. Page 91 of this paper introduces atomic units for length (a_{H} = h^{2}/(4π^{2}me^{2}), the radius of first Bohr orbit of hydrogen), charge (e, the electron charge), and mass (m, electron mass). From these are derived the unit of action (h/(2π) ), the unit of energy (e^{2}/a, twice the ionization energy of hydrogen), and the unit of time (1/(4πcR)). The value R is the Rydberg constant.

Harper:2001:AGG

[Har01] Eamon Harper. In appreciation: George Gamow: Scientific amateur and polymath. Physics in Perspective (PIP), 3(3):335–372, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://adsabs.harvard.edu/
abs/2001PhP.....3..335H;   http://link.springer.com/
article/10.1007/PL00000536;   http://www.springerlink.
com/content/m0htmt9ww9b9jjv1/;   http://www.springerlink.
com/openurl.asp?genre=journal&issn=1422-6944.

Hautot:1972:ASD

[Hau72] A. Hautot. About the solutions of Dirac’s equations in the presence of new magnetic fields. Journal of Mathematical Physics, 13(5):710–714, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hawking:1998:DMA

[Haw98] Stephen Hawking. Dirac memorial address. In Goddard [God98], pages xiii–?? ISBN 0-521-58382-9 (hardcover), 0-521-01953-2 (paperback). LCCN QC16.D57 P38 1998. URL http://

www.loc.gov/catdir/description/cam028/97022443.html;
 http://www.loc.gov/catdir/toc/cam021/97022443.html.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Harish-Chandra:1987:MAP

[HC87] Harish-Chandra. My association with Professor Dirac. In Kurşunoğlu and Wigner [KW87], pages 34–36. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

Hamilton:1977:CDE

[HD77] M. J. Hamilton and A. Das. On the combined Dirac–Einstein– Maxwell field equations. Journal of Mathematical Physics, 18 (10):2026–2030, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heisenberg:1975:QET

[HDW75] W. Heisenberg, Paul A. M. Dirac, and V. Weisskopf. Quantum electrodynamics: Translations of four papers by Heisenberg, Dirac, and Weisskopf. Technical Report RL-75-073, Science and Technology Facilities Council ePublication Archive, Chilbolton, Daresbury, and Rutherford Appleton Laboratories, UK, 1975.

URL http://epubs.stfc.ac.uk/work-details?w=21197;
http://purl.org/oai/oai:epubs.cclrc.ac.uk:work/21197.

Heisenberg:1935:FDT

[HE35] Werner Heisenberg and Hans Euler. Folgerungen aus der Diracschen Theorie des Positrons. (German) [Consequences of Dirac’s theory of the positron]. Zeitschrift für Physik, 98(11–12):714– 732, ???? 1935. CODEN ZEPYAA. ISSN ???? URL http:

//www.springerlink.com/content/p710710041x81t32/. En-

glish translation in [HE06].

Holton:1982:AEH

[HE82] Gerald James Holton and Yehuda Elkana, editors. Albert Einstein, historical and cultural perspectives: the centennial symposium in Jerusalem. Princeton University Press, Princeton, NJ, USA, 1982. ISBN 0-691-08299-5. LCCN QC16.E5 J48 1979. US$27.50.

Heisenberg:2006:CDT

[HE06] Werner Heisenberg and Hans Euler. Consequences of Dirac’s theory of the positron. ArXiv Physics e-prints, May 4,

2006. URL http://arxiv.org/pdf/physics/0605038. English

translation by W. Korolevski and H. Kleinert.

Heisenberg:1930:BBP

[Hei30a] Werner Heisenberg. Besprechung: The principles of Quantum mechanics. Von P. A. M. Dirac. Metallwirtschaft, 9(??):988, ???? 1930.

Heisenberg:1930:BPM

[Hei30b] Werner Heisenberg. Buchbesprechung: P. A. M. Dirac The Principles of Quantum Mechanics. Metallwirtschaft, 9(??):988–??, ???? 1930.

Heisenberg:1934:BDT

[Hei34a] Werner Heisenberg. Bemerkungen zur Diracschen Theorie des Positrons. (German) [Remarks on the Dirac theory of the positron]. Zeitschrift für Physik, 90(3–4):209–231, ???? 1934. CODEN ZEPYAA. ISSN ???? URL http://www.

springerlink.com/content/vr14wp0778r73025/. See correc-

tion [Hei34b].

Heisenberg:1934:BAB

[Hei34b] Werner Heisenberg. Berichtigung zu der Arbeit: “Bemerkungen zur Diracschen Theorie des Positrons”. (German) [Correction on the work: “Remarks on the Dirac theory of the positron”]. Zeitschrift für Physik, 92(9–10):692, ???? 1934. CO-

DEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/l0710nm9813n7622/. See [Hei34a].

Heisenberg:1972:IMS

[Hei72] W. Heisenberg. Indefinite metric in state space. In Salam and Wigner [SW72], pages 129–136. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Hendry:1983:MBD

[Hen83] John Hendry. Monopoles before Dirac. Studies in History and Philosophy of Science Part A, 14(1):81–87, March 1983. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368183900079.

Hendry:1984:CPT

[Hen84] John Hendry, editor. Cambridge Physics in the Thirties. Adam Hilger Ltd., Bristol, UK, 1984. ISBN 0-85274-761-6. 209 pp. LCCN QC51.G72 C352 1984.

Hendry:1989:BRB

[Hen89] John Hendry. Book review: Behram N. Kurşunoğlu & Eugene P. Wigner (eds). Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac. Cambridge: Cambridge University Press, 1987. Pp. xviii + 297. ISBN 0-521-34013-6. £30.00. C.W. Kilmister (ed.). Schrödinger: Centenary Celebration of a Polymath. Cambridge: Cambridge University Press, 1987. Pp. x + 253. ISBN 0-521-34017-9. £30.00. British Journal for the History of Science, 22(1):87–88, March 1989. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026687.

Hendry:1991:BRH

[Hen91] John Hendry. Book review: Helge Kragh. Dirac: A Scientific Biography. Cambridge: Cambridge University Press, 1990. ISBN 0-521-38089-8. £35. British Journal for the History of Science, 24(1):116, March 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4027031.

Heras:2018:DQC

[Her18] Ricardo Heras. Dirac quantisation condition: a comprehensive review. Contemporary Physics, 59(4):331–355, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hestenes:1973:LOD

[Hes73] D. Hestenes. Local observables in the Dirac theory. Journal of Mathematical Physics, 14(7):893–905, July 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hestenes:1975:OOC

[Hes75] D. Hestenes. Observables, operators, and complex numbers in the Dirac theory. Journal of Mathematical Physics, 16(3):556– 572, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hillion:1992:NSD

[Hil92] Pierre Hillion. Nondispersive solutions of the Dirac equation. Journal of Mathematical Physics, 33(5):1822–1830, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hindmarsh:1977:BKP

[Hin77] J. L. Hindmarsh. On the Bunge–Kálnay position operator for the Dirac electron. International Journal of Theoretical Physics, 16:443–445, 1977. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Hinton:1991:ISL

[HJKS91] D. B. Hinton, A. K. Jordan, M. Klaus, and J. K. Shaw. Inverse scattering on the line for a Dirac system. Journal of Mathematical Physics, 32(11):3015–3030, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Hernandez:2000:NTN

[HJL00] P. Hernandez, K. Jansen, and L. Lellouch. A numerical treatment of Neuberger’s lattice Dirac operator. In Frommer et al. [F^{+}00], pages 29–39. ISBN 3-540-67732-1. ISSN 1439-7358. LCCN QC793.3.G38 N86 2000.

Hovis:1993:PMD

[HK93] R. Corby Hovis and Helge Kragh. P. A. M. Dirac and the beauty of physics. Scientific American, 268(5):104–?? (Intl. ed. 62–??), May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hill:1938:DET

[HL38] E. L. Hill and R. Landshoff. The Dirac electron theory. Reviews of Modern Physics, 10(2):87–132, April 1938. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
10.87;   http://rmp.aps.org/abstract/RMP/v10/i2/p87_
1.

Hounkonnou:1999:ESDb

[HM99a] M. N. Hounkonnou and J. E. B. Mendy. Exact solutions of Dirac equation for neutrinos in presence of external fields. Journal of Mathematical Physics, 40(9):4240–4254, September 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:1999:ESDa

[HM99b] M. N. Hounkonnou and J. E. B. Mendy. Exact solutions of the Dirac equation in a nonfactorizable metric. Journal of Mathematical Physics, 40(8):3827–3842, August 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoyle:1971:NM

[HN71] F. Hoyle and J. V. Narlikar. On the nature of mass. Nature, 233(5314):41–44, September 3, 1971. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v233/n5314/pdf/
233041a0.pdf.

Hoyle:1962:MPC

[HND^{+}62] F. Hoyle, J. V. Narlikar, P. A. M. Dirac, H. Bondi, R. Schlegel, W. Davidson, and J. L. Synge. Mach’s principle and the creation of matter [and discussion]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 270 (1342):334–341, November 27, 1962. CODEN PRLAAZ. ISSN

0080-4630. URL http://www.jstor.org/stable/2414534.

Horn:2011:GPD

[Hor11] Martin Erik Horn. Grassmann, Pauli, Dirac: special relativity in the schoolroom. In Petsche [Pet11], pages 435–452. ISBN 3-0346-0404-1 (print), 3-0346-0405-X (e-book). LCCN

QA29.G73 F76 2011. URL http://www.springerlink.com/
content/978-3-0346-0405-5.

Hostler:1982:IRD

[Hos82] Levere C. Hostler. An SL(2, C)-invariant representation of the Dirac equation. Journal of Mathematical Physics, 23(6):1179– 1184, June 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hostler:1983:IRD

[Hos83] Levere C. Hostler. An SL(2, C)-invariant representation of the Dirac equation. II. Coulomb Green’s function. Journal of Mathematical Physics, 24(9):2366–2373, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hostler:1987:RCS

[Hos87] Levere Hostler. Relativistic Coulomb Sturmian matrix elements and the Coulomb Green’s function of the second-order Dirac equation. Journal of Mathematical Physics, 28(12):2984–2989, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hostler:1989:FWT

[Hos89] Levere Hostler. Foldy–Wouthuysen-type reduction of the second-order Dirac equation. Journal of Mathematical Physics, 30(7):1621–1630, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hovis:1996:BRC

[Hov96] R. Corby Hovis. Book review: The Collected Works of P. A. M. Dirac, 1924–1948 and General Theory of Relativity Richard H. Dalitz and P. A. M. Dirac. Physics Today, 49(9):84–88, September 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hoyle:1987:MMH

[Hoy87] F. Hoyle. Magnetic monopoles and the halos of galaxies. In Kurşunoğlu and Wigner [KW87], pages 174–193. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

Hoyle:1992:AD

[Hoy92] F. Hoyle. The achievement of Dirac. Notes and Records of the Royal Society of London, 46(1):183–187, January 1992. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Heisenberg:1934:MAE

[HSD34a] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Die moderne Atomtheorie: die bei der entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. (German) [Modern Atomic Theory: the 1933 Nobel Prize in Physics lecture in Stockholm]. Verlag von S. Hirzel, Leipzig, Germany, 1934. 2 + 45 + 1 pp. LCCN QC173 .H38.

Heisenberg:1934:MQM

[HSD34b] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Modern Quantum Mechanics: Three Nobel Reports. Technico-Theoretical State Press, Leningrad and Moscow, USSR, 1934. ???? pp.

Hagen:1976:MD

[HSZ76] C. R. Hagen, Julian Schwinger, and Daniel Zwanziger. Monopole debate. Physics Today, 29(4):15–??, April 1976. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v29/i4/p15.

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hannabuss:1990:CSA

[HW90] K. C. Hannabuss and W. R. Weiss. Corepresentations of superalgebras and an antilinear Dirac operator. Journal of Mathematical Physics, 31(10):2520–2524, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hey:1996:EM

[HW96] Anthony J. G. Hey and Patrick Walters. Einstein’s mirror. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43504-8 (hardcover), 0-521-43532-3 (paperback). xii + 291 pp. LCCN QC173.55 .H49 1996. URL http://www.

zentralblattmath.org/zmath/en/search/?an=0968.83002.

Halpern:2006:BNU

[HW06] Paul Halpern and Paul S. Wesson. Brave new universe: illuminating the darkest secrets of the cosmos. Joseph Henry Press, Washington, DC, USA, 2006. ISBN 0-309-10137-9 (hardcover), 0-309-65823-3 (PDFs). viii + 264 + 8 pp. LCCN QB981 .H248 2006. URL http://www.

loc.gov/catdir/toc/ecip068/2006004464.html;   http:/
/www.nap.edu/catalog.php?record_id=11636.

Hylton:1984:RDG

[Hyl84] D. J. Hylton. The reduced Dirac Green function for the Coulomb potential. Journal of Mathematical Physics, 25(4):1125–1132, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ignatiev:2004:MCV

[IC04] A. Yu. Ignatiev and B. J. Carson. Metrological constraints on the variability of the fundamental constants e, h̄, and c. Physics Letters A, 331(6):361–365, November 1, 2004. CODEN PY-LAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Ivancevic:2008:QLD

[II08] Vladimir G. Ivancevic and Tijana T. Ivancevic. Quantum leap: from Dirac and Feynman, across the universe, to human body and mind. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008. ISBN 981-281-927-4, 981-281-928-2 (e-book). xvi + 839 pp. LCCN QC174.12 .I927 2008.

Ichinose:2002:PRD

[IJ02] Takashi Ichinose and Brian Jefferies. The propagator of the radial Dirac equation. Journal of Mathematical Physics, 43(8): 3963–3983, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iyer:1991:SVD

[IK91] B. R. Iyer and N. Kamran. Separation of variables for the Dirac equation in an extended class of Lorentzian metrics with local rotational symmetry. Journal of Mathematical Physics, 32(9): 2497–2502, September 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ilyin:1989:FAC

[IKRT89a] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodioniov, and A. Yu. Taranov. Fast algorithm for calculation of Dirac’s gamma-matrices traces. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(4):15–24, October 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Ilyin:1989:HSD

[IKRT89b] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodioniov, and A. Yu. Taranov. High speed Dirac algebra calculations in a space of arbitrary dimension by means of a computer algebra system. Computer Physics Communications, 57(1-3):505–506, December 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ilyin:1991:PIF

[IKRT91] V. A. Ilyin, A. P. Kryukov, A. Ya. Rodionov, and A. Yu. Taranov. PC implementation of fast Dirac matrix trace calculations. In Watt [Wat91], pages 456–457. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991. URL http://www.acm.org:80/

pubs/citations/proceedings/issac/120694/p456-ilyin/.

Indelicato:1995:AED

[IM95] P. Indelicato and Peter J. Mohr. Asymptotic expansion of the Dirac–Coulomb radial Green’s function. Journal of Mathematical Physics, 36(2):714–724, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Infeld:1933:NDN

[Inf33] Leopold Infeld. Nowe drogi nauki: kwanty i materja (Polish) [New pathways of science: quanta and matter], volume 2 of Z Dziedziny Nauki i Techniki. nakł. Mathesis Polskiej, Warszawa, Poland, 1933. x + 284 pp. LCCN ????

Infeld:1934:DEG

[Inf34a] Leopold Infeld. Dirac’s equation in the general relativity theory. Acta Physica Polonica, 3(??):1–14, ???? 1934. CODEN APPOAK. ISSN 0001-673X.

Infeld:1934:WMS

[Inf34b] Leopold Infeld. The world of modern science: matter and quanta. G. P. Putnam’s Sons, New York, NY, USA, 1934. 287 pp. LCCN ???? Translation by Louis Infield of [Inf33]. Introduction by Albert Einstein.

Infeld:1953:ECH

[Inf53a] Leopold Infeld. An electronic cloud in a homogeneous electric and magnetic field according to Dirac’s theory. Bulletin de l’Académie polonaise des sciences. Classe troisième, 1(??): 99–104, ???? 1953. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1953:UAM

[Inf53b] Leopold Infeld. On the use of an approximation method in Dirac’s electrodynamics. Bulletin de l’Académie polonaise des sciences. Classe troisième, 1:18–22, 1953. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1961:PCC

[Inf61] Leopold Infeld. Is Planck’s constant a constant in a gravitational field? Bulletin de l’Académie Polonaise des Sciences, Série des Sciences, Mathématiques, Astronomiques et Physiques, 9:617– 620, 1961. CODEN BAPMAM. ISSN 0001-4117.

Infeld:1962:PCT

[Inf63] Leopold Infeld. Planck’s constant and the theory of the red shift. Zeitschrift für Physik, 171:34–43, 1962–1963. CODEN ZEPYAA. ISSN 0170-9739 (print), 1431-5858 (electronic).

Infeld:1956:MDF

[IP56] Leopold Infeld and Jerzy Plebański. On modified Dirac δfunctions. Bulletin de l’Académie polonaise des sciences. Classe troisième, 4:687–691, 1956. CODEN BPSMA6. ISSN 0554-5897.

Infeld:1957:FMD

[IP57] Leopold Infeld and Jerzy Plebański. On a further modification of Dirac’s δ-functions. Bulletin de l’Académie polonaise des sciences. Classe troisième, 5:51–54, vi, 1957. CODEN BPSMA6. ISSN 0554-5897.

Iverson:1973:FIP

[IPM73] Geoffrey J. Iverson, Arnold Perlmutter, and Stephan L. Mintz, editors. Fundamental interactions in physics and astrophysics: Lectures from the 1972 Coral Gables Conference on Fundamental Interactions at High Energy, January 19–21, 1972, volume 3 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1973. ISBN 0-306-36903-6. LCCN QC793.9 .C68 1972. A volume dedicated to P. A. M. Dirac on the occasion of his seventieth birthday.

Irvine:1983:CLP

[Irv83] J. M. Irvine. The constancy of the laws of physics in the light of prehistoric nuclear reactors. Contemporary Physics, 24(5):427–437, September 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00107518308210695.

Ivanov:2008:HPD

[IRV^{+}08] T. I. Ivanov, M. Roudjane, M. O. Vieitez, C. A. de Lange, W.-Ü L. Tchang-Brillet, and W. Ubachs. HD as a probe for detecting mass variation on a cosmological time scale. Physical Review Letters, 100(9):093007, March 7, 2008. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.100.093007.

Ikhdair:2010:ABS

[IS10a] Sameer M. Ikhdair and Ramazan Sever. Approximate bound state solutions of Dirac equation with Hulthén potential including Coulomb-like tensor potential. Applied Mathematics and Computation, 216(3):911–923, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ikhdair:2010:SSD

[IS10b] Sameer M. Ikhdair and Ramazan Sever. Solutions of the spatially-dependent mass Dirac equation with the spin and pseudospin symmetry for the Coulomb-like potential. Applied Mathematics and Computation, 216(2):545–555, March 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ishikawa:1990:ADF

[Ish90] Y. Ishikawa. Atomic Dirac–Fock–Breit self-consistent field calculations. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 24(??):383–??, 1990. CODEN IJQSDI. ISSN 0161-3642.

Ichinose:1984:PDP

[IT84] Takashi Ichinose and Hiroshi Tamura. Propagation of a Dirac particle. A path integral approach. Journal of Mathematical Physics, 25(6):1810–1819, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ichinose:1988:ZDP

[IT88] Takashi Ichinose and Hiroshi Tamura. The Zitterbewegung of a Dirac particle in two-dimensional space–time. Journal of Mathematical Physics, 29(1):103–109, January 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:1986:FDC

[IU86] Toshihiro Iwai and Yoshio Uwano. The four-dimensional conformal Kepler problem reduces to the three-dimensional Kepler problem with a centrifugal potential and Dirac’s monopole field. Classical theory. Journal of Mathematical Physics, 27(6):1523– 1529, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iyer:1985:SDE

[IV85] B. R. Iyer and C. V. Vishveshwara. Separability of the Dirac equation in a class of perfect fluid space–times with local rotational symmetry. Journal of Mathematical Physics, 26(5):1034– 1039, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iyer:1987:ESS

[IV87] B. R. Iyer and C. V. Vishveshwara. Exact solutions for space– times with local rotational symmetry in which the Dirac equation separates. Journal of Mathematical Physics, 28(6):1377– 1381, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jacob:1998:A

[Jac98] Maurice Jacob. Antimatter. In Goddard [God98], pages 46– 87. ISBN 0-521-58382-9 (hardcover), 0-521-01953-2 (paperback). LCCN QC16.D57 P38 1998. URL http://www.loc.

gov/catdir/description/cam028/97022443.html;   http:
//www.loc.gov/catdir/toc/cam021/97022443.html.

Jacobson:2025:FLP

[Jac25] Ted Jacobson. Feynman 1947 letter on path integral for the Dirac equation. European Physical Journal H, 50(1):??, December 2025. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL https://link.springer.com/article/10.
1140/epjh/s13129-025-00090-1.

Jarnefelt:1964:EZA

[Jär64] G. Järnefelt. Einige Züge der abstrakten quantumelektrodynamischen Struktur im Anschluss an das Buch von P. A. M. Dirac, “The principles of quantum mechanics” (Fourth edition 1958). (German) [Some consequences of the abstract quantum electrodynamic structure following the book of P. A. M. Dirac, “The Principles of Quantum Mechanics” (fourth edition 1958)]. Publications of the Astronomical Observatory, Helsinki, Finland 104, Astronomical Observatory, Helsinki, Finland, 1964. ?? pp.

Jauch:1972:BK

[Jau72] J. M. Jauch. On bras and kets. In Salam and Wigner [SW72], pages 137–167. ISBN 0-521-08600-0. LCCN

QC174.1 .A85 1972.  URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

JefferysSwirles:1987:RDH

[Jef87] Lady Bertha Jefferys (Swirles). Reminiscences at the dinner held at St John’s College. In Taylor [Tay87b], pages 38–39. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Jing:1991:ESE

[Jin91] Ji Liang Jing. Exact solutions of Einstein’s equations for space– time with local rotational symmetry in which the Dirac equation separates. Journal of Mathematical Physics, 32(5):1334–1336, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:1999:DPS

[JM99] M. A. Jafarizadeh and S. K. Moayedi. Degeneracy and parasupersymmetry of Dirac Hamiltonian in (2 + 1)-space–time. Journal of Mathematical Physics, 40(9):4274–4289, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Jordan:1937:PWG

[Jor37] Pascual Jordan. Die physikalischen Weltkonstanten. (German) [The physical world constants]. Die Naturwissenschaften, 25 (??):513–516, ???? 1937. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL https://www.

digizeitschriften.de/dms/img/?PID=PPN34557155X_0025%
7CLOG_0173.

Jordan:1949:FSD

[Jor49] Pascual Jordan. Formation of the stars and development of the universe. Nature, 164(4172):637–640, October 15, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Jordan:1952:SWG

[Jor52] Pascual Jordan. Schwerkraft und Weltall; Grundlagen der theoretischen Kosmologie. (German) [Gravity and space: basic principles of theoretical cosmology], volume 107 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1952. viii + 207 pp. LCCN QC6 .J662.

Jordan:1955:SWG

[Jor55] Pascual Jordan. Schwerkraft und Weltall; Grundlagen der theoretischen Kosmologie. (German) [Gravity and space: basic principles of theoretical cosmology]. Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1955. xi + 277 + 13 pp. LCCN QC6 .J662.

Jordan:1959:BDH

[Jor59a] Pascual Jordan. Die Bedeutung der Diracschen Hypothese für die Geophysik, etc.. (German) [The meaning of the Dirac Hypothesis for geophysics, etc.]. Akademie der Wissenschaften und der Literatur. Abhandlungen der mathematischnaturwissenschaftlichen Klasse, 9(??):27–??, ???? 1959. CO-DEN AWLMA9. ISSN 0002-2993.

Jordan:1959:GSD

[Jor59b] Pascual Jordan. Zum gegenwärtigen Stand der Diracschen kosmologischen Hypothesen. (German) [On the current state of Dirac’s cosmological hypotheses]. Zeitschrift für Physik, 157 (1):112–121, February 1959. CODEN ZEPYAA. ISSN 0044-

3328. URL https://link.springer.com/article/10.1007/
BF01375155.

Jordan:1962:GCD

[Jor62a] P. Jordan. Geophysical consequences of Dirac’s hypothesis. Reviews of Modern Physics, 34(4):596–600, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.34.596;   http://rmp.aps.org/
abstract/RMP/v34/i4/p596_1.

Jordan:1962:ECD

[Jor62b] Pascual Jordan. Empirical confirmation of Dirac’s hypothesis of diminishing gravitation. In Anonymous [Ano62], pages 283–288. LCCN QC 6 .R295. This book is dedicated to Leopold Infeld in connection with his 60th birthday.

Jordan:1966:EEF

[Jor66] Pascual Jordan. Die Expansion der Erde: Folgerungen aus d. Diracschen Gravitationshypothese. (German) [The expansion of the Earth: conclusions from Dirac’s gravitational hypothesis]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1966. ISBN 3-663-00255-1 (paperback), 3-663-02168-8 (e-book). xiv + 182 + 47 pp.

Jordan:1971:EES

[Jor71] Pascual Jordan, editor. The Expanding Earth; Some Consequences of Dirac’s Gravitation Hypothesis, volume 37 of International series of monographs in natural philosophy. Pergamon, New York, NY, USA, 1971. ISBN 0-08-015827-7 (paperback), 0-08-015827-7 (e-book), 1-4831-5587-0. xv + 202

pp. URL http://www.sciencedirect.com/science/book/

9780080158273. Edited by Arthur Beer. With the collaboration of J. B. Hutchings and T. R. Stoeckley.

Jost:1972:FQF

[Jos72] R. Jost. Foundation of quantum field theory. In Salam and Wigner [SW72], pages 69–85. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Johnson:1973:IRD

[JP73] Edward R. Johnson and Raymond Parnes. An integral representation of the Dirac delta function for axisymmetric boundary value problems. SIAM Journal on Applied Mathematics, 24 (4):539–544, June 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jofre:2006:CPT

[JRF06] Paula Jofré, Andreas Reisenegger, and Rodrigo Fernández. Constraining a possible time variation of the gravitational constant through “gravitochemical heating” of neutron stars. Physical Review Letters, 97(13):131102, September 29, 2006. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.97.131102.

Jauregui:2010:NRD

[JT10] M. Jauregui and C. Tsallis. New representations of π and Dirac delta using the nonextensive-statistical-mechanics q-exponential function. Journal of Mathematical Physics, 51(6):063304, June 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v51/i6/p063304_s1.

Jung:1997:GAI

[Jun97] Wolf Jung. Geometrical approach to inverse scattering for the Dirac equation. Journal of Mathematical Physics, 38(1):39–48, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jordan:1928:PAG

[JW28] P. Jordan and E. Wigner. Über das Paulische Äquivalenzverbot. (German) [On Pauli’s equivalence prohibition]. Zeitschrift für Physik, 47(9–10):631–651, ???? 1928. CODEN ZEPYAA.

ISSN ???? URL http://www.springerlink.com/content/
hx1t32272451437h.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kamimura:1986:TDG

[Kam86] Kiyoshi Kamimura. Three-dimensional gauge theory in Dirac formalism. Journal of Mathematical Physics, 27(8):2137–2139, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kapur:1940:DMO

[Kap40] P. L. Kapur. Does the mesotron obey Bose–Einstein or Fermi– Dirac statistics? Nature, 145(3663):69, January 13, 1940. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v145/n3663/
pdf/145069a0.pdf.

Katriel:2000:MPD

[Kat00] Jacob Katriel. Many-particle Dirac identities for arbitrary elementary spins. International Journal of Quantum Chemistry, 78(6):407–411, ???? 2000. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract/72502335/START;
 http://www3.interscience.wiley.com/cgi-bin/fulltext/
72502335/FILE?TPL=ftx_start;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=72502335&PLACEBO=IE.pdf.

Krori:1983:SES

[KBN83] K. D. Krori, R. Bhattacharya, and D. Nandy. Some exact solutions of Einstein–Dirac–Maxwell zero mass scalar equations. Journal of Mathematical Physics, 24(6):1574–1576, June 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kapitza:1933:RES

[KD33] P. L. Kapitza and P. A. M. Dirac. The reflection of electrons from standing light waves. Mathematical proceedings of the Cambridge Philosophical Society, 29(2):297–300, ???? 1933. CODEN PCPSA4. ISSN 0008-1981.

Kemmer:1987:WPD

[Kem87] Nicholas Kemmer. What Paul Dirac meant in my life. In Kurşunoğlu and Wigner [KW87], pages 37–42. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

Kersten:1983:ISC

[Ker83] Paul H. M. Kersten. Infinitesimal symmetries and conserved currents for nonlinear Dirac equations. Journal of Mathematical Physics, 24(9):2374–2376, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerimov:1990:LDF

[Ker90] G. A. Kerimov. Lobachevskian Dirac fields. Journal of Mathematical Physics, 31(7):1745–1754, July 1990. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:1986:CTV

[Kha86] Pushpa Khare. Comment on “Time variation of fundamental constants, primordial nucleosynthesis, and the size of extra dimensions.”. Physical Review D (Particles and Fields), 34(6):1936–1937, September 15, 1986. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.

34.1936. See [KPW86].

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kibble:1998:BRB

[Kib98] T. W. B. Kibble. Book review: Paul Dirac: The Man and his Work. European Journal of Physics, 19(3):??, 1998. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL http://stacks.iop.org/0143-0807/19/i=3/a=017. Re-

view of [God98].

Kilmister:1976:BRG

[Kil76] C. W. Kilmister. Book review: General Theory of Relativity, by P. A. M. Dirac. Mathematical Gazette, 60(413):239, October 1976. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3617516.

Kim:1980:TITa

[Kim80a] Shoon K. Kim. Theory of involutional transformations applied to the Dirac theory of the electron. I. Remarks on the Dirac plane waves. Journal of Mathematical Physics, 21(8):2282–2285, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:1980:TITb

[Kim80b] Shoon K. Kim. Theory of involutional transformations applied to the Dirac theory of the electron. II. Remarks on the Dirac– Coulomb waves. Journal of Mathematical Physics, 21(8):2286– 2290, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirby:1997:MNR

[Kir97] Kevin G. Kirby. Of memories, neurons, and rank-one corrections. College Mathematics Journal, 28(1):2–19, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07468342.1997.11973823. This paper contains an ex-

tensive tutorial on the Dirac bra and ket notation.

Kiselev:1998:PTD

[Kis98] O. M. Kiselev. Perturbation theory for the Dirac equation in two-dimensional space. Journal of Mathematical Physics, 39(4): 2333–2345, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Komech:2010:GAS

[KK10] Alexander Komech and Andrew Komech. Global attraction to solitary waves for a nonlinear Dirac equation with mean field interaction. SIAM Journal on Mathematical Analysis, 42(6): 2944–2964, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://epubs.siam.org/

sima/resource/1/sjmaah/v42/i6/p2944_s1.

L:1988:SOB

[K.L88] K.L. Some other books of interest: Tributes to Paul Dirac. Science, 241(4870):1239, September 2, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Kovalyov:1990:DER

[KL90] M. Kovalyov and M. Légaré. The Dirac equation in Robertson– Walker spaces: a class of solutions. Journal of Mathematical Physics, 31(1):191–198, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovalyov:1992:CSS

[KL92] M. Kovalyov and M. Légaré. A complete set of solutions to the massive Dirac equation in static open Robertson–Walker space. Journal of Mathematical Physics, 33(3):1118–1126, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klaus:1990:LTD

[Kla90] Martin Klaus. On the Levinson theorem for Dirac operators. Journal of Mathematical Physics, 31(1):182–190, January 1990.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klein:2005:ESF

[Kle05] Etienne Klein. Il était sept fois la révolution: Albert Einstein et les autres .... (French) [The seven-times revolution: Albert Einstein and the others]. Flammarion, Paris, France, 2005. ISBN 2-08-210343-9. 237 pp. LCCN QC16.E5.

Kamran:1988:CCS

[KLMW88] N. Kamran, M. Légaré, R. G. McLenaghan, and P. Winternitz. The classification of complete sets of operators commuting with the Dirac operator in Minkowski space–time. Journal of Mathematical Physics, 29(2):403–411, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamran:1984:SVS

[KM84] N. Kamran and R. G. McLenaghan. Separation of variables and symmetry operators for the neutrino and Dirac equations in the space–times admitting a two-parameter abelian orthogonally transitive isometry group and a pair of shearfree geodesic null congruences. Journal of Mathematical Physics, 25(4):1019–1027, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1992:SSD

[KM92] E. G. Kalnins and W. Miller, Jr. Series solutions for the Dirac equation in Kerr–Newman space–time. Journal of Mathematical Physics, 33(1):286–296, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kimberly:2004:VAE

[KM04] Dagny Kimberly and João Magueijo. Varying alpha and the electroweak model. Physics Letters B, 584(1–2):8–15, 2004. CODEN PYLBAJ. ISSN 0370-2693 (print), 1873-2445

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0370269304001832.

Kentosh:2012:GPS

[KM12] J. Kentosh and M. Mohageg. Global positioning system test of the local position invariance of Planck’s constant. Physical Review Letters, 108(11):110801:1–110801:4, March 16, 2012. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic),

1092-0145. URL http://journals.aps.org/prl/abstract/
10.1103/PhysRevLett.108.110801. See comment [BF12] and

news story [Ano12].

Kursunoglu:1985:HEP

[KMPD85] Behram Kurşunoğlu, Stephan L. Mintz, Arnold Perlmutter, and P. A. M. (Paul Adrien Maurice) Dirac, editors. High-energy physics: in honor of P. A. M. Dirac in his eightieth year, volume 20 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1985. ISBN 0-306-42070-8. LCCN QC793 .O7 1983.

Kalnins:1986:MOS

[KMW86] E. G. Kalnins, W. Miller, Jr., and G. C. Williams. Matrix operator symmetries of the Dirac equation and separation of variables. Journal of Mathematical Physics, 27(7):1893–1900, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Knox:2003:NHH

[KN03] Kevin C. Knox and Richard Noakes, editors. From Newton to Hawking: a History of Cambridge University’s Lucasian Professors of Mathematics. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-66310-5. xxiv + 486 pp. LCCN QA28

.F76 2003. URL http://www.loc.gov/catdir/description/
cam032/2003043584.html;   http://www.loc.gov/catdir/
enhancements/fy0731/2003043584-b.html;   http://www.
loc.gov/catdir/toc/cam032/2003043584.html.

Kolm:1967:SMM

[Kol67] Henry H. Kolm. Search for magnetic monopole in deep-sea sediment. Physics Today, 20(20):69–70, October 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Koplinger:2006:DEH

[Köp06] Jens Köplinger. Dirac equation on hyperbolic octonions. Applied Mathematics and Computation, 182(1):443–446, November 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kostin:1986:PDB

[Kos86] M. D. Kostin. Probability density for Bose–Einstein and Fermi– Dirac particles: Slater–Kahn functions. Journal of Mathematical Physics, 27(12):3014–3015, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kox:1997:DEI

[Kox97] A. J. Kox. The discovery of the electron: II. The Zeeman effect. European Journal of Physics, 18(3):139–144, May 1997. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL

http://iopscience.iop.org/0143-0807/18/3/003.

Kozhevnikov:1990:PDF

[Koz90] Aleksej Borisovich Kozhevnikov. Pol′′Dirak i fizika XX veka: sbornik nauchnych trudov. (Russian) [Paul Dirac and 20th Century Physics: collection of works]. Nauka, Moscow, Russia, 1990. ISBN 5-02-000702-1. 221 + 2 pp. LCCN QC16.D57 P65 1990.

Kursunoglu:1983:FTE

[KP83] Behram Kuršunoğlu and Arnold Perlmutter, editors. Field theory in elementary particles: Proceedings of Orbis Scientiae 1982, held January 18–21, 1982 by the Center for Theoretical Studies, University of Miami, Coral Gables, Florida, volume 19 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1983. ISBN 0-306-41345-0. LCCN QC793.3.F5 O7 1982.

Kursunoglu:1978:NFH

[KPS^{+}78] Kurşunoğlu, Arnold Perlmutter, Linda F. Scott, O. Kadiroğlu, J. Nowakowski, and F. Krausz, editors. New frontiers in highenergy physics: proceedings of Orbis Scientiae 1978 on New Frontiers in High-Energy Physics held by the Center for Theoretical Studies, University of Miami, Coral Gables, Florida, Jan. 16–19, 1978, volume 14 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1978. ISBN 0-306-40037-5. LCCN QC793 .O7 1978.

Kursunoglu:1975:TEH

[KPW75] Behram Kurşunoğlu, Arnold Perlmutter, and Susan M. Widmayer, editors. Theories and experiments in high-energy physics, volume 9 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1975. ISBN 0-306-36909-5. LCCN QC793 .O7 1975.

Kolb:1986:TVF

[KPW86] Edward W. Kolb, Malcolm J. Perry, and T. P. Walker. Time variation of fundamental constants, primordial nucleosynthesis, and the size of extra dimensions. Physical Review D (Particles and Fields), 33(4):869–871, February 15, 1986. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/

PhysRevD.33.869. See comment [Kha86].

Keller:1992:GSD

[KR92] Jaime Keller and Adán Rodríguez. Geometric superalgebra and the Dirac equation. Journal of Mathematical Physics, 33(1): 161–170, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kragh:1981:CMH

[Kra81a] Helge Kragh. The concept of the monopole. A historical and analytical case-study. Studies in History and Philosophy of Science Part A, 12(2):141–172, June 1981. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0039368181900170.

Kragh:1981:GDR

[Kra81b] Helge Kragh. The genesis of Dirac’s Relativistic Theory of Electrons. Archive for History of Exact Sciences, 24(1):31– 67, March 1981. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
24&issue=1&spage=31.

Kragh:1982:CPT

[Kra82] Helge Kragh. Cosmo-physics in the thirties: Towards a history of Dirac cosmology. Historical Studies in the Physical Sciences, 13(1):69–108, ???? 1982. CODEN HSPSAS. ISSN 0073-2672.

URL http://www.jstor.org/stable/27757506.

Kragh:1990:DSB

[Kra90] Helge Kragh. Dirac: a scientific biography. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-01756-4 (paperback), 0-521-38089-8 (hardcover). x + 389 pp. LCCN QC16.D57 K73 1990. URL http://www.loc.

gov/catdir/description/cam023/89017257.html;   http:
//www.loc.gov/catdir/toc/cam028/89017257.html.

Kragh:1991:CEDb

[Kra91a] Helge S. Kragh. Cosmonumerology and empiricism — the Dirac/Gamow dialogue. Astronomy Quarterly, 8:109–??, 1991. CODEN ASTQD2. ISSN 0364-9229 (print), 1878-0865 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1991AstQ...
.8..109K.

Kragh:1991:CEDa

[Kra91b] Helge S. Kragh. Cosmonumerology and empiricism: the Dirac– Gamow dialogue. Astronomy Quarterly, 8:107–??, 1991. CO-DEN ASTQD2. ISSN 0364-9229 (print), 1878-0865 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1991AstQ..
..8..107K.

Kragh:2006:DPA

[Kra06a] Helge Kragh. Dirac, Paul Adrien Maurice (1902–1984). In Oxford Dictionary of National Biography, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006.

Kragh:2006:CVS

[Kra06b] Helge S. Kragh. Cosmologies with varying speed of light: a historical perspective. Studies in History and Philosophy of Modern Physics, 37(4):726–737, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1355219806000530.

Kragh:2013:PDB

[Kra13] Helge Kragh. Paul Dirac and The principles of quantum mechanics. In Research and pedagogy: a history of quantum physics through its textbooks [BN13], page ?? ISBN 3-8442-5871-X.

LCCN QC173.98. URL http://www.edition-open-access.
de/studies/2/.

Kragh:2015:EED

[Kra15a] Helge Kragh. Expanding Earth and declining gravity: a chapter in the recent history of geophysics. History of Geo- and Space Sciences, 6(1):45–55, ???? 2015. CODEN ???? ISSN 2190-5010 (print), 2190-5029 (electronic). URL https://www.hist-geo-

space-sci.net/6/45/2015/.

Kragh:2015:PJV

[Kra15b] Helge Kragh. Pascual Jordan, varying gravity, and the expanding Earth. Physics in Perspective (PIP), 17(2):107–134, June 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-015-0157-9.

Kragh:2016:VGD

[Kra16] Helge Kragh, editor. Varying gravity: Dirac’s legacy in cosmology and geophysics, volume 54 of Science networks. Historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2016. ISBN 3-319-24377-2 (hardcover), 3-319-24379-9 (e-book). xi + 185 pp. LCCN QB981.

Kragh:2017:LSS

[Kra17] Helge Kragh. ‘Let the stars shine in peace!’ Niels Bohr and stellar energy, 1929–1934. Annals of Science, 74(2):126–148, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2019:VCN

[Kra19] Helge Kragh. Varying constants of nature: Fragments of a history. Physics in Perspective (PIP), 21(4):257–273, December 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Krisch:1987:EVP

[Kri87] A. D. Krisch. An experimenter’s view of P. A. M. Dirac. In Kurşunoğlu and Wigner [KW87], pages 46–52. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

Keller:1990:MDE

[KRR90] Jaime Keller and Suemi Rodríguez-Romo. A multivectorial Dirac equation. Journal of Mathematical Physics, 31(10):2501– 2510, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krupkova:1994:GSH

[Kru94] Olga Krupková. A geometric setting for higher-order Dirac– Bergmann theory of constraints. Journal of Mathematical Physics, 35(12):6557–6576, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kursunoglu:1973:FIP

[Kur73] Behram Kurşunoğlu, editor. Fundamental Interactions in Physics: proceedings of the Coral Gables Conference on Fundamental Interactions, January 22–26, 1973, volume 2 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1973. ISBN 0-306-36902-8. LCCN QC793.9 .C67 1973.

Kursunoglu:1987:DIU

[Kur87a] Behram Kurşunoğlu. Dirac’s influence on unified field theory. In Kurşunoğlu and Wigner [KW87], pages 276–

291. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Kursunoglu:1987:DCG

[Kur87b] Sevda A. Kurşunoğlu. Dirac in Coral Gables. In Kurşunoğlu and Wigner [KW87], pages 9–28. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Kursunoglu:1987:RAG

[KW87] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Ktitarev:1993:FPI

[KY93] D. V. Ktitarev and R. S. Yegikian. Feynman path integral for Dirac system with analytic potential. Journal of Mathematical Physics, 34(7):2821–2826, July 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalf:2001:ESA

[KY01] Hubert Kalf and Osanobu Yamada. Essential self-adjointness of n-dimensional Dirac operators with a variable mass term. Journal of Mathematical Physics, 42(6):2667–2676, June 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

L:1988:BRS

[L.88] K. L. Book review: Some other books of interest: Tributes to Paul Dirac, by J. G. Taylor; Reminiscences about a Great Physicist, by Behram N. Kurşunoğlu and Eugene P. Wigner; Instruments and Experiences, by R. V. Jones; Kelvin’s Baltimore Lectures and Modern Theoretical Physics, by Robert Kargon and Peter Achinstein. Science, 241(4870):1239–1240, September 2, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1702737.

Lakhtakia:1996:MMH

[Lak96] A. (Akhlesh) Lakhtakia, editor. Models and Modelers of Hydrogen: Thales, Thomson, Rutherford, Bohr, Sommerfeld, Goudsmit, Heisenberg, Schrödinger, Dirac, Sallhofer. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2302-1. xv + 424 pp. LCCN

QC171.2 .M63 1996. URL http://www.worldscientific.com/
worldscibooks/10.1142/2788.

Lamborn:1969:EDD

[Lam69] B. N. A. Lamborn. An expression for the Dirac delta function. I. SIAM Review, 11(4):603, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lamborn:1970:EDD

[Lam70] B. N. A. Lamborn. An expression for the Dirac delta function. II. SIAM Review, 12(4):567–569, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lamb:1987:SC

[Lam87] Willis E. Lamb, Jr. Schrödinger’s cat. In Kurşunoğlu and Wigner [KW87], pages 249–261. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Lanczos:1930:DWT

[Lan30] Cornelius Lanczos. Dirac’s wellenmechanische Theorie des Elektrons und ihre feldtheoretische Ausgestaltung. (German) [Dirac’s wave-mechanical theory of the electron and its fieldtheoretical interpretation]. Physikalische Zeitschrift, 31(??): 120–130, ???? 1930. CODEN PHZTAO. ISSN 0369-982X.

Lanczos:1972:PB

[Lan72] C. Lanczos. The Poisson bracket. In Salam and Wigner [SW72], pages 169–178. ISBN 0-521-08600-0. LCCN

QC174.1 .A85 1972.  URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Lannutti:1987:EPM

[Lan87a] Joseph E. Lannutti. Eulogy for Paul A. M. Dirac, 19 November 1984: “Who was this Guy?”. In Taylor [Tay87b], pages 43–47. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Lannutti:1987:RPD

[Lan87b] Joseph E. Lannutti. Recollections of Paul Dirac at Florida State University. In Kurşunoğlu and Wigner [KW87], pages 29–33. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Langmann:2001:GYM

[Lan01] Edwin Langmann. Generalized Yang–Mills actions from Dirac operator determinants. Journal of Mathematical Physics, 42 (11):5238–5256, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lanczos:2005:DWM

[Lan05a] C. Lanczos. Dirac’s wave mechanical theory of the electron and its field theoretical interpretation. ArXiv Physics e-prints, August 2005. CODEN APRHCB. URL http://adsabs.harvard.

edu/abs/2005physics...8009L.

Lanczos:2005:TAR

[Lan05b] C. Lanczos. The tensor analytical relationships of Dirac’s equation. ArXiv Physics e-prints, August 2005. CO-

DEN APRHCB. URL http://adsabs.harvard.edu/abs/
2005physics...8002L.

Lynden-Bell:1998:CMN

[LBNZ98] D. Lynden-Bell and M. Nouri-Zonoz. Classical monopoles: Newton, NUT space, gravomagnetic lensing, and atomic spectra. Reviews of Modern Physics, 70(2):427–445, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.70.427;   http://rmp.aps.org/
abstract/RMP/v70/i2/p427_1.

Long:1998:TBD

[LC98] Peter Long and Horace W. Crater. Two-body Dirac equations for general covariant interactions and their coupled Schrödingerlike forms. Journal of Mathematical Physics, 39(1):124–160, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2010:NMP

[LCH10] Xiang-Gui Li, C. K. Chan, and Y. Hou. A numerical method with particle conservation for the Maxwell–Dirac system. Applied Mathematics and Computation, 216(4):1096–1108, April 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:1995:DGM

[Lee95] Dae-Gyu Lee. The Dirac gamma matrices as “relics” of a hidden symmetry?: As fundamental representations of the algebra sp(4, R). Journal of Mathematical Physics, 36(1):524–530, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2021:CTV

[Lee21] Seokcheon Lee. Constraints on the time variation of the speed of light using Pantheon dataset. arxiv.org, ??(??):1–14, January

25, 2021. URL https://arxiv.org/abs/2101.09862.

Lee:2023:VVS

[Lee23] Seokcheon Lee. A viable varying speed of light model in the RW metric. Foundations of Physics, 53(2):??, April 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516

(electronic). URL https://link.springer.com/article/10.
1007/s10701-023-00682-1.

Legare:1995:ISR

[Lég95] M. Légaré. Invariant spinors and reduced Dirac equations under subgroups of the Euclidean group in four-dimensional Euclidean space. Journal of Mathematical Physics, 36(6):2777–2791, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lehmann:1991:PAE

[Leh91] Detlef Lehmann. A probabilistic approach to Euclidean Dirac fields. Journal of Mathematical Physics, 32(8):2158–2166, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenhard:2014:DMC

[Len14] Johannes Lenhard. Disciplines, models, and computers: the path to computational quantum chemistry. Studies in History and Philosophy of Science Part A, 48(??):89–96, December 2014. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368114000533.

Leon:1980:DIS

[Leo80] Jérôme Jp. Leon. The Dirac inverse spectral transform: kinks and boomerons. Journal of Mathematical Physics, 21(10):2572– 2578, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levitt:1980:GCT

[Lev80] L. S. Levitt. The gravitational constant at time zero. Lettere al Nuovo Cimento Series 2, 29(1):23–24, 1980. CODEN LNUCAE. ISSN 0375-930x (print), 1827-613x (electronic).

Legare:1986:RYM

[LH86] M. Légaré and J. Harnad. SO(4) reduction of the SU(N) Yang– Mills–Dirac equations. Journal of Mathematical Physics, 27(2): 620–626, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lighthill:1987:FAG

[Lig87] Sir James Lighthill. Fourier analysis and generalised functions. In Taylor [Tay87b], pages 96–103. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://w3.impa.br/~ jair/Lighthill.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Linnaeus:2010:PIS

[Lin10] Staffan Linnæus. Phase-integral solution of the radial Dirac equation. Journal of Mathematical Physics, 51(3):032304, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i3/p032304_s1.

Lipkin:1987:RQT

[Lip87] Harry J. Lipkin. From relativistic quantum theory to the human brain. In Kurşunoğlu and Wigner [KW87], pages 234–243. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Lennard-Jones:1931:BRP

[LJ31] J. E. Lennard-Jones. Book review: The Principles of Quantum Mechanics, by P. A. M. Dirac. Mathematical Gazette, 15(216): 505–506, December 1931. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.jstor.org/

stable/3605591.

Landau:2008:EUC

[LMSV08] Susana J. Landau, Mercedes E. Mosquera, Claudia G. Scóccola, and Héctor Vucetich. Early universe constraints on time variation of fundamental constants. Physical Review D (Particles and Fields), 78(8):083527, October 15, 2008. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
78.083527.

Labelle:1991:SDC

[LMV91] Simon Labelle, Michel Mayrand, and Luc Vinet. Symmetries and degeneracies of a charged oscillator in the field of a magnetic monopole. Journal of Mathematical Physics, 32(6):1516– 1521, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i6/p1516_s1.

Loudon:1983:QTL

[Lou83] Rodney Loudon. The quantum theory of light. Oxford science publications. Clarendon Press, Oxford, UK, second edition, 1983. ISBN 0-19-851155-8 (paperback), 0-19-851152-3 (hardcover). xiv + 393 pp. LCCN QC446.2 .L68 1983.

Lowdin:1964:SPTe

[Löw64] P.-O. Löwdin. Studies in perturbation theory. VIII. Separation of Dirac equation and study of the spin-orbit coupling and Fermi contact terms. Journal of Molecular Spectroscopy, 14(1):131– 144, October 1964. CODEN JMOSA3. ISSN 0022-2852 (print), 1096-083x (electronic).

Lass:1959:DMA

[LR59] Harry Lass and Oldwig Von Roos. The Dirac measure as applied to the solution of difference equations by means of transforms. American Mathematical Monthly, 66(6):483–485, June/ July 1959. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Landgren:1979:AMP

[LR79] J. J. Landgren and P. A. Rejto. An application of the maximum principle to the study of essential self-adjointness of Dirac operators. I. Journal of Mathematical Physics, 20(11):2204–2211, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landgren:1980:AMP

[LRK80] J. J. Landgren, P. A. Rejto, and Martin Klaus. An application of the maximum principle to the study of essential selfadjointness of Dirac operators. II. Journal of Mathematical Physics, 21(5): 1210–1217, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lichnerowicz:1962:TRG

[LT62] André Lichnerowicz and Marie A. Tonnelat, editors. Les théories relativistes de la gravitation (Royaumont, 21–27 juin, 1959), volume 91 of Colloques internationaux du Centre national de la recherche scientifique. Éditions du Centre National de la Recherche Scientifique (CNRS), Paris, France, 1962. LCCN ????

Lamoreaux:2004:NMO

[LT04] S. K. Lamoreaux and J. R. Torgerson. Neutron moderation in the Oklo natural reactor and the time variation of α. Physical Review D (Particles and Fields), 69(12):121701, June 15, 2004. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.

org/doi/10.1103/PhysRevD.69.121701.

Laporte:1931:ASA

[LU31] Otto Laporte and George E. Uhlenbeck. Application of spinor analysis to the Maxwell and Dirac equations. Physical Review, 37(11):1380–1397, June 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Lubkin:1975:MRM

[Lub75] Gloria B. Lubkin. Mixed reception for magnetic monopole announcement. Physics Today, 28(10):17–20, October 1975. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v28/i10/p17.

Lubkin:1982:CCF

[Lub82] Gloria B. Lubkin. Cabrera counts flux quanta to find a Dirac monopole. Physics Today, 35(6):17–19, June 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Lutzer:2007:MDF

[Lut07] Carl V. Lutzer. De-mystifying the Dirac δ-function. PRIMUS (Problems Resources and Issues in Mathematics Undergraduate Studies), 17(4):373–377, October 2007. ISSN 1051-1970 (print), 1935-4053 (electronic).

Ladenburg:1934:ANP

[LW34] Rudolf Ladenburg and Eugene Wigner. Award of the Nobel Prizes in Physics to Professors Heisenberg, Schrödinger and Dirac. The Scientific Monthly, 38(1):86–91, January 1934. CO-DEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

URL http://www.jstor.org/stable/15534.

Lightstone:1975:DDF

[LW75] A. H. Lightstone and Kam Wong. Dirac delta functions via nonstandard analysis. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 18(??):759–762, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Lannutti:1978:CTT

[LW78] J. E. Lannutti and P. K. Williams, editors. Current trends in the theory of fields: (Tallahassee 1978): a symposium in honor of P. A. M. Dirac, volume 48 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1978. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

MacLeod:1998:AFD

[Mac98] Allan J. MacLeod. Algorithm 779: Fermi–Dirac functions of order −1/2, 1/2, 3/2, 5/2. ACM Transactions on Mathematical Software, 24(1):1–12, March 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:/

/www.acm.org/pubs/citations/journals/toms/1998-24-1/
p1-macleod/.

Magueijo:2015:VVA

[Mag15] Joäao Magueijo. Varying c: vodka without alcohol? In Einstein: a hundred years of relativity [Rob15], page ?? ISBN 0-691-16989-6 (paperback). LCCN QC16.E5 R63 2015.

Majorana:1937:TSD

[Maj37] Ettore Majorana. Teoria simmetrica dell elettrone e del positrone. (Italian) [Symmetrical theory of the electron and the positron]. Il Nuovo Cimento (8), 14(4):171–184, April 1937. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (elec-

tronic). URL http://en.wikipedia.org/wiki/Majorana_
mass;   http://nl.wikipedia.org/wiki/Majorana-deeltje.

In this paper, Majorana predicted the existence of a new type of particle, now called a Majorana fermion, which is its own antiparticle, and whose existence may have finally been confirmed by experiment seventy years later [FK08, MZF^{+}12, Ser12]. See also comments in [ACM10]. Esposito [Esp08] reports about this paper “With amazing farsightedness Majorana suggested that the neutrino, which had just been postulated by Wolfgang Pauli and Fermi to explain puzzling features of radioactive beta decay, could be such a particle. This would make the neutrino unique among the elementary particles and, moreover, enable it to have mass. Today many experiments are still devoted to detect these peculiar properties, which include the phenomenon of neutrino oscillations: we have not yet succeeded to find a definite answer to Majorana’s proposal.”.

Malin:1975:EGF

[Mal75a] S. Malin. Erratum: Gravitational field equations and the possibility of time variation of all masses. Physical Review D (Particles and Fields), 12(2):628, July 15, 1975. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.

12.628. See [Mal75b].

Malin:1975:GFE

[Mal75b] S. Malin. Gravitational field equations and the possibility of time variation of all masses. Physical Review D (Particles and Fields), 11(4):707–710, February 15, 1975. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
11.707.

Malin:1975:WDE

[Mal75c] S. Malin. The Weyl and Dirac equations in terms of functions over the group SU_{2}. Journal of Mathematical Physics, 16(3): 679–684, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marlow:1965:UDN

[Mar65] A. R. Marlow. Unified Dirac–von Neumann formulation of quantum mechanics. I. Mathematical theory. Journal of Mathematical Physics, 6(6):919–927, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i6/p919_s1.

Marciano:1984:TVF

[Mar84] William J. Marciano. Time variation of the fundamental ‘constants’ and Kaluza–Klein theories. Physical Review Letters, 52 (7):489–491, February 13, 1984. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

link.aps.org/doi/10.1103/PhysRevLett.52.489.

Marques:1988:DEN

[Mar88] S. Marques. The Dirac equation in a non-Riemannian manifold. I. An analysis using the complex algebra. Journal of Mathematical Physics, 29(9):2127–2131, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matthews:1987:DFQ

[Mat87] P. T. Matthews. Dirac and the foundation of quantum mechanics. In Kurşunoğlu and Wigner [KW87], pages 199– 224. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Maurette:1976:FNR

[Mau76] Michel Maurette. Fossil nuclear reactors. Annual Review of Nuclear Science, 26(??):319–350, December 1976. CODEN AR-

NUA8. ISSN 0066-4243. URL http://www.annualreviews.
org/doi/abs/10.1146/annurev.ns.26.120176.001535.

Marques-Bonham:1990:DEN

[MB90] S. Marques-Bonham. The Dirac equation in a non-Riemannian manifold. II. An analysis using an internal local n-dimensional space of the Yang–Mills type. Journal of Mathematical Physics, 31(6):1478–1482, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marques-Bonham:1991:DEN

[MB91] Sirley Marques-Bonham. The Dirac equation in a non-Riemannian manifold. III. An analysis using the algebra of quaternions and octonions. Journal of Mathematical Physics, 32(5):1383–1394, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Montwill:2011:PDT

[MB11a] Alex Montwill and Ann Breslin. Paul Dirac: Tying things together. In The quantum adventure: does God play dice? [MB11b], chapter 13, pages 171–190. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (ebook). LCCN QC174.12 .M66 2012.

Montwill:2011:QAD

[MB11b] Alex Montwill and Ann Breslin. The quantum adventure: does God play dice? World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (ebook). ix + 248 pp. LCCN QC174.12 .M66 2012.

Magueijo:2002:IIE

[MBS02] João Magueijo, John D. Barrow, and Håvard Bunes Sandvik. Is it e or is it c ? Experimental tests of varying alpha. Physics Letters B, 549(3?4):284–289, 2002. CO-DEN PYLBAJ. ISSN 0370-2693 (print), 1873-2445 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0370269302029283.

McIntosh:1970:DPM

[MC70] Harold V. McIntosh and Arturo Cisneros. Degeneracy in the presence of a magnetic monopole. Journal of Mathematical Physics, 11(3):896–916, March 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i3/p896_s1.

McAllister:1990:DAE

[McA90] James W. McAllister. Dirac and the aesthetic evaluation of theories. Methodology and Science, 23(??):87–102, ????

1990. URL http://openaccess.leidenuniv.nl/bitstream/
1887/10436/1/9_57_018.pdf.

McCubbin:2004:BPL

[McC04] Norman McCubbin. Beauty in physics: the legacy of Paul Dirac. Contemporary Physics, 45(4):319–333, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McWeeny:1973:NUA

[McW73] Roy McWeeny. Natural units in atomic and molecular physics. Nature, 243(5404):196–198, May 25, 1973. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Ma:1988:IES

[MD88] Zhong Qi Ma and An Ying Dai. Integral equations with symmetrical kernel applied to a system with a Dirac-type spectrum. Journal of Mathematical Physics, 29(6):1321–1324, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Medvedev:1989:DLO

[Med89] Boris V. Medvedev. Dirak i logicheskie osnovy kvantovoj teorii. (Russian) [Dirac and the logical foundations of quantum theory]. TsNIIatominform, ????, 1989. ISBN ???? 28 + 1 pp. LCCN ????

Medvedev:1990:PMD

[Med90] B. V. Medvedev, editor. P. A. M. Dirak i Fizika XX Veka. (Russian) [P. A. M. Dirac and 20th Century Physics]. Nauka, Moscow, Russia, 1990. ISBN ???? ???? pp. LCCN ????

Mehra:1972:GAT

[Meh72] Jagdish Mehra. The golden age of theoretical physics: P. A. M. Dirac’s scientific work from 1924 to 1933. In Salam and Wigner [SW72], pages 17–59. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.

springer.com/us/book/9789027703453.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mehra:1987:DCP

[Meh87a] Jagdish Mehra. Dirac’s contribution to the early development of quantum mechanics. In Taylor [Tay87b], pages 63–75. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Mekhra:1987:GAT

[Meh87b] Jagdish Mehra. The Golden Age of theoretical physics: the scientific activity of P. A. M. Dirac from 1924 to 1933. Uspekhi Fizicheskikh Nauk, 153(9):135–165, September 1987. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1987/9/d/.

Mekhra:1987:LWP

[Meh87c] Jagdish Mehra. List of the works of P. A. M. Dirac. Uspekhi Fizicheskikh Nauk, 153(9):165–172, September 1987. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1987/9/e/.

Mehra:1991:D

[Meh91] Jagdish Mehra. Dirac. In Ferris and Fadiman [FF91], page 603. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Marciano:1987:DMM

[MG87] William J. Marciano and Maurice Goldhaber. Dirac’s magnetic monopole and the fine structure constant. In Kurşunoğlu and Wigner [KW87], pages 163–173. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

March:1998:SCT

[MHN98] N. H. March, A. Holas, and Á. Nagy. Self-consistent Thomas–Fermi–Dirac theory, extended by Gell-Mann and Brueckner correlation, in terms of density n and its two reduced gradients ∇^{2}n/n and ∇n/n. International Journal of Quantum Chemistry, 69(2):145–149, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL

http://www3.interscience.wiley.com/cgi-bin/abstract?
ID=29976;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=29976&PLACEBO=IE.pdf.

Miller:1997:DEC

[Mil97] David Miller. The Dirac equation. Cosmic implications of a tidy electron. Physics Education, 32(4):238–243, July 1997. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL

http://iopscience.iop.org/0031-9120/32/4/018/.

Milhorat:1998:SDO

[Mil98] Jean-Louis Milhorat. Spectrum of the Dirac operator on Gr_{2}(C^{m+2}). Journal of Mathematical Physics, 39(1):594–609, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milotti:2007:EFV

[Mil07] Edoardo Milotti. Enrico Fermi’s view of identical particles. arxiv.org, ??(??):9, May 9, 2007. URL http://arxiv.org/abs/ 0705.1363. See [Fer24, Fer26] for the original Italian and German versions.

Mitra:2010:DP

[Mit10] Asoke N. Mitra. Dirac postscript. Physics Today, 63(12):12, December 2010. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Mohankumar:2005:EFD

[MKK05] N. Mohankumar, T. Kannan, and S. Kanmani. On the evaluation of Fermi–Dirac integral and its derivatives by IMT and DE quadrature methods. Computer Physics Communications, 168(2):71–77, June 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505000639.

Milde:2013:USL

[MKS^{+}13] P. Milde, D. Köhler, J. Seidel, L. M. Eng, A. Bauer, A. Chacon, J. Kindervater, S. Mühlbauer, C. Pfleiderer, S. Buhrandt, C. Schütte, and A. Rosch. Unwinding of a skyrmion lattice by magnetic monopoles. Science, 340(6136):1076–1080, May 31, 2013. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.portal.uni-koeln.de/3963+
M51bdc3177e1.html;   http://www.rdmag.com/news/2013/
05/researchers-discover-artificial-magnetic-monopoles;
 http://www.sciencemag.org/content/340/6136/1076.abstract.

Mignaco:1993:LAD

[ML93] Juan Alberto Mignaco and Cesar Augusto Linhares. Lie algebras for the Dirac–Clifford ring. Journal of Mathematical Physics, 34 (5):2066–2074, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Mohankumar:1997:AEP

[MN97] N. Mohankumar and A. Natarajan. The accurate evaluation of a particular Fermi–Dirac integral. Computer Physics Communications, 101(1–2):47–53, April 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S001046559600166X.

Mohankumar:2007:TNS

[Moh07] N. Mohankumar. Two new series for the Fermi–Dirac integral. Computer Physics Communications, 176(11–12):665–669, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002007.

Monti:1995:DHM

[Mon95] Dalida Monti. Dirac’s holes model: From proton to positron. Nuncius, 10(1):99–130, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00044.

Moreno:1985:CFP

[Mor85] Matías Moreno. Closed formula for the product of n Dirac matrices. Journal of Mathematical Physics, 26(4):576–584, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mott:1987:RPD

[Mot87] Nevill Mott. Reminiscences of Paul Dirac. In Kurşunoğlu and Wigner [KW87], pages 230–233. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Moyer:1981:EDE

[Moy81a] Donald Franklin Moyer. Evaluations of Dirac’s electron, 1928– 1932. American Journal of Physics, 49(11):1055–1062, November 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v49/i11/p1055_s1. Part 2 of 3; see also part 1

[Moy81b] and part 3 [Moy81c]. The paper reprints two long letters between Niels Bohr and Paul Dirac from November– December 1929 on the problem of negative energies in Dirac’s quantum theory.

Moyer:1981:ODE

[Moy81b] Donald Franklin Moyer. Origins of Dirac’s electron, 1925– 1928. American Journal of Physics, 49(10):944–949, October 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://ajp.aapt.org/resource/1/

ajpias/v49/i10/p944_s1. Part 1 of 3; see also part 2 [Moy81a] and part 3 [Moy81c].

Moyer:1981:VDE

[Moy81c] Donald Franklin Moyer. Vindications of Dirac’s electron, 1932– 1934. American Journal of Physics, 49(12):1120–1125, December 1981. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/
1/ajpias/v49/i12/p1120_s1. Part 3 of 3; see also part 1

[Moy81b] and part 2 [Moy81a].

Metropolis:1951:SFT

[MR51] N. Metropolis and J. R. Reitz. Solutions of the Fermi–Thomas– Dirac equation. Journal of Chemical Physics, 19(??):555–573, 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Mehra:1982:DQM

[MR82a] Jagdish Mehra and Helmut Rechenberg. The discovery of quantum mechanics, 1925, volume 2 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90674-6. vi + 355 pp. LCCN QC173.98 .M44 vol. 2. See other volumes in this series [MR82c, MR82b, MR82d, MR82e, MR87, MR00].

Mehra:1982:FMM

[MR82b] Jagdish Mehra and Helmut Rechenberg. The Formulation of Matrix Mechanics and Its Modifications, 1925–1926, volume 3 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90675-4, 3-540-90675-4. viii + 334 pp. LCCN QC173.98 .M44

vol. 3. URL http://www.loc.gov/catdir/enhancements/
fy1005/82003253-d.html;   http://www.loc.gov/catdir/
enhancements/fy1005/82003253-t.html. See other volumes

in this series [MR82a, MR82c, MR82d, MR82e, MR87, MR00].

Mehra:1982:FEQ

[MR82c] Jagdish Mehra and Helmut Rechenberg. The fundamental equations of quantum mechanics, 1925–1926. The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0. viii + 322 pp. LCCN QC173.98 .M44 vol. 4. See other volumes in this series [MR82a, MR82b, MR82d, MR82e, MR87, MR00].

Mehra:1982:QTPa

[MR82d] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82e, MR87, MR00].

Mehra:1982:QTPb

[MR82e] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375– 878 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR87, MR00].

Mehra:1987:ESR

[MR87] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol.

5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR00].

Mehra:2000:CQM

[MR00] Jagdish Mehra and Helmut Rechenberg. The completion of quantum mechanics, 1926–1941, volume 6 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-95086-9 (part 1), 0-387-95182-2 (part 2). xxxvi + 1612 pp. LCCN QC173.98 .M44 vol. 6. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR87].

Mawhin:2010:SDE

[MR10] Jean Mawhin and André Ronveaux. Schrödinger and Dirac equations for the hydrogen atom, and Laguerre polynomials. Archive for History of Exact Sciences, 64(4):429–460, 2010. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Mignani:1975:ADL

[MRB75] R. Mignani, E. Recami, and M. Baldo. About a Dirac-like equation for the photon, according to Ettore Majorana. Technical Report Print-75-0054 (Catania), INFN-ROME-578, ????, Catania, Sicily, Italy, 1975.

McDougall:1938:CFD

[MS38] J. McDougall and Edmund C. Stoner. The computation of Fermi–Dirac functions. Philosophical Transactions of the Royal Society of London. Series A, Mathematical and Physical Sciences, 237(773):67–104, February 7, 1938. ISSN 0080-4614. URL

http://www.jstor.org/stable/91333.

Mukunda:1968:SDB

[MS68] N. Mukunda and E. C. G. Sudarshan. Structure of the Dirac bracket in classical mechanics. Journal of Mathematical Physics, 9(3):411–417, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Medvedev:1987:PMD

[MS87] B. V. Medvedev and D. V. Shirkov. P. A. M. Dirac: the formation of the basic ideas of quantum field theory. Uspekhi Fizicheskikh Nauk, 153(9):59–104, September 1987. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1987/9/b/. English translation

in Sov. Phys. Usp. 30 791–815 (1987).

Martin:1995:BCT

[MS95] A. Martin and J. Stubbe. Bargmann- and Calogero-type bounds for the Dirac equation. Journal of Mathematical Physics, 36(9): 4680–4690, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

MartinezyRomero:1992:CID

[MSB92] R. P. Martínez y Romero and A. L. Salas-Brito. Conformal invariance in a Dirac oscillator. Journal of Mathematical Physics, 33(5):1831–1836, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muraskin:1971:MPD

[Mur71] M. Muraskin. A mathematical property of Dirac plane waves. Journal of Mathematical Physics, 12(1):28–31, January 1971.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Murphy:2007:CLT

[MWF07] M. T. Murphy, J. K. Webb, and V. V. Flambaum. Comment on “Limits on the Time Variation of the Electromagnetic Fine-Structure Constant in the Low Energy Limit from Absorption Lines in the Spectra of Distant Quasars”. Physical Review Letters, 99(23):239001, December 7, 2007. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.99.239001. See [SCPA04].

Mourik:2012:SMF

[MZF^{+}12] V. Mourik, K. Zuo, S. M. Frolov, S. R. Plissard, E. P. A. M. Bakkers, and L. P. Kouwenhoven. Signatures of Majorana fermions in hybrid superconductor–semiconductor nanowire devices. Science, 336(6084):1003–1007, May 25, 2012. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://arxiv.org/abs/1204.2792;   http:/
/www.sciencemag.org/content/336/6084/1003.

Nakamura:1997:PSM

[Nak97] Toru Nakamura. Path space measures for Dirac and Schrödinger equations: nonstandard analytical approach. Journal of Mathematical Physics, 38(8):4052–4072, August 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakamura:2000:PSM

[Nak00] Toru Nakamura. Path space measure for the 3 + 1-dimensional Dirac equation in momentum space. Journal of Mathematical Physics, 41(8):5209–5222, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nash:1984:ISG

[Nas84] Patrick L. Nash. Identities satisfied by the generators of the Dirac algebra. Journal of Mathematical Physics, 25(2):204–209, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nash:1986:EEC

[Nas86] Patrick L. Nash. On the exceptional equivalence of complex Dirac spinors and complex space–time vectors. Journal of Mathematical Physics, 27(5):1185–1190, May 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nash:1997:CDS

[Nas97] Patrick L. Nash. On the construction of a Dirac spinor that generates a specified tetrad. Journal of Mathematical Physics, 38(11):6008–6017, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterov:2008:IDR

[NdlC08] Alexander I. Nesterov and Fermín Aceves de la Cruz. Infinitedimensional representations of the rotation group and Dirac monopole problem. Journal of Mathematical Physics, 49(1): 013505, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i1/p013505_s1.

Neumann:1933:KPH

[Neu33] Neumann János. A koordináta-mérés pontosságának határai az elektron dirac-félé elméletében. (Hungarian) [Über die Grenzen der Koordinatenmessungs-Genauigkeit in der Diracschen Theorie des Elektrons]. (German) [On the limits of coordinate measurement accuracy in the Dirac theory of the electron]. Mat. es Termeszettut … Ertesito, 50(??):366–385, ???? 1933. Reprinted in [Tau61b, Paper 20].

Neuberger:2000:ODO

[Neu00] H. Neuberger. The overlap Dirac operator. In Frommer et al. [F^{+}00], pages 1–17. ISBN 3-540-67732-1. ISSN 1439-7358. LCCN QC793.3.G38 N86 2000.

Nilsson:1982:RSP

[Nil82] Bengt E. W. Nilsson. Reduction of the super phase space for a massless Dirac particle. Journal of Mathematical Physics, 23 (3):451–455, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nilson:2003:TVF

[Nil03a] Tobias Nilson. On time-variation of the fundamental constants. Report, Upsala University, Uppsala, Sweden, January 31, 2003.

URL www.teorfys.uu.se/files/Tobias_Nilsson_varconst.

pdf. ii + 72.

Nilsson:2003:FAD

[Nil03b] Henrik Nilsson. Functional automatic differentiation with Dirac impulses. ACM SIGPLAN Notices, 38(9):153–164, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nishina:1932:HDL

[Nis32] Yoshio Nishina. Heisenberg and Dirac’s lectures on the problems of quantum mechanics. Keimeika Gaiyo, 11(??):3–??, ???? 1932.

Nishina:1935:PDBa

[Nis35a] Yoshio Nishina. Professors Dirac and Beck’s lectures (I). Kagaku, 5(??):359–361, ???? 1935.

Nishina:1935:PDBb

[Nis35b] Yoshio Nishina. Professors Dirac and Beck’s lectures (II). Kagaku, 5(??):400–402, ???? 1935.

Natarajan:1993:NEG

[NK93] A. Natarajan and N. Mohan Kumar. On the numerical evaluation of the generalised Fermi–Dirac integrals. Computer Physics Communications, 76(1):48–50, June 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/001046559390118V.

Nam:1984:GTM

[NM84] Kiwon Nam and Michael J. Moravcsik. General transformation matrix for Dirac spinors and the calculation of spinorial amplitudes. Journal of Mathematical Physics, 25(4):820–827, April 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Natarajan:2001:AMG

[NM01] A. Natarajan and N. Mohankumar. An accurate method for the generalized Fermi–Dirac integral. Computer Physics Communications, 137(3):361–365, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501001539.

Noerdlinger:1973:PRE

[Noe73] Peter D. Noerdlinger. Primordial 2.7^{∘} radiation as evidence against secular variation of Planck’s constant. Physical Review Letters, 30(16):761–762, April 16, 1973. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://journals.aps.org/prl/abstract/10.
1103/PhysRevLett.30.761.

North:1965:MUH

[Nor65] John David North. The measure of the universe: a history of modern cosmology. Clarendon Press, Oxford, UK, 1965. xxviii + 436 pp. LCCN QB981 .N77.

Norman:1986:FCR

[Nor86] Eric B. Norman. Are fundamental constants really constant? American Journal of Physics, 54(4):317–321, 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/54/4/10.1119/1.14847.

North:1990:MUH

[Nor90] John David North. The measure of the universe: a history of modern cosmology. Dover, New York, NY, USA, 1990. ISBN 0-486-66517-8 (paperback). xxviii + 436 pp. LCCN QB981 .N77 1990. US$12.95.

Novomestky:1974:AEU

[Nov74] F. Novomestky. Asymptotic expressions for the unit-step and Dirac delta functions. SIAM Journal on Applied Mathematics, 27(4):521–525, December 1974. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nenciu:1996:ODP

[NP96] Gheorghe Nenciu and Radu Purice. One-dimensional periodic Dirac Hamiltonians: semiclassical and high-energy asymptotics for gaps. Journal of Mathematical Physics, 37(7):3153–3167, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oakes:2000:PPP

[Oak00] Billie Blankenship Oakes. The personal papers of Paul A. M. Dirac: their history and preservation at the Florida State University. Thesis (Ph.D.), Florida State University, Tallahassee, FL, USA, 2000. ix + 209 pp.

Oevel:1988:DCF

[Oev88] Walter Oevel. Dirac constraints in field theory: lifts of Hamiltonian systems to the cotangent bundle. Journal of Mathematical Physics, 29(1):210–219, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ogurisu:1996:ESD

[Ogu96] Osamu Ogurisu. Essential spectrum of the Dirac Hamiltonian for a spin 1/2 neutral particle with an anomalous magnetic moment in an asymptotically constant magnetic field. Journal of Mathematical Physics, 37(3):1234–1243, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Olive:1998:M

[Oli98] David Olive. The monopole. In Goddard [God98], pages 88– 107. ISBN 0-521-58382-9 (hardcover), 0-521-01953-2 (paperback). LCCN QC16.D57 P38 1998. URL http://www.loc.

gov/catdir/description/cam028/97022443.html;   http:
//www.loc.gov/catdir/toc/cam021/97022443.html.

Olver:1986:DTC

[Olv86] Peter J. Olver. Dirac’s theory of constraints in field theory and the canonical form of Hamiltonian differential operators. Journal of Mathematical Physics, 27(10):2495–2501, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ordonez:1995:DRQ

[OP95] C. R. Ordóñez and J. M. Pons. Dirac and reduced quantization: a Lagrangian approach and application to coset spaces. Journal of Mathematical Physics, 36(3):1146–1165, March 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ogielski:1981:PIQ

[OS81] A. T. Ogielski and J. Sobczyk. Path integral quantization of the supersymmetric model of the Dirac particle. Journal of Mathematical Physics, 22(9):2060–2065, September 1981. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pais:1972:EHT

[Pai72] A. Pais. The early history of the theory of the electron: 1897– 1947. In Salam and Wigner [SW72], pages 79–93. ISBN 0-521-08600-0. LCCN QC174.1 .A85 1972. URL http://hooke.lib.

cam.ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;
http://www.loc.gov/catdir/enhancements/fy1001/72075298-
d.html;   http://www.loc.gov/catdir/enhancements/fy1001/
72075298-t.html.

Pais:1987:PED

[Pai87] Abraham Pais. Playing with equations, the Dirac way. In Kurşunoğlu and Wigner [KW87], pages 93–116. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

Paierls:1990:STD

[Paĭ90] Rudol’f Paĭerls. Some thought Dirac odd. In Paul Dirac and 20th century physics (Russian), pages 112–114. “Nauka”, Moscow, Russia, 1990. Translated from the English by T. B. Romanovskaya.

Pais:1998:PDA

[Pai98] Abraham Pais. Paul Dirac: aspects of his life and work. In Goddard [God98], pages 1–45. ISBN 0-521-58382-9 (hardcover), 0-521-01953-2 (paperback). LCCN QC16.D57

P38 1998. URL http://www.loc.gov/catdir/description/
cam028/97022443.html;   http://www.loc.gov/catdir/toc/
cam021/97022443.html.

Pais:2000:GSP

[Pai00a] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp.

LCCN Q141 .P29 2000.  URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/fy02/99046603.html.

Pais:2000:PD

[Pai00b] Abraham Pais. Paul Dirac. In The Genius of Science: a Portrait Gallery [Pai00a], pages 48–77. ISBN 0-19-850614-7 (hardcover). LCCN Q141 .P29 2000. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/
catdir/toc/fy02/99046603.html.

Pashby:2012:DPP

[Pas12] Thomas Pashby. Dirac’s prediction of the positron: a case study for the current realism debate. Perspectives on Science, 20(4):440–475, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v020/
20.4.pashby.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.4.pashby.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00081.

Pauli:1925:ZAE

[Pau25] Wolfgang Pauli. Über den Zusammenhang des Abschlusses der Elektronengruppen im Atom mit der Komplexstruktur der Spektren. (German) [On the relation of the completion of groups of electrons in the atom with the complex structure of spectra]. Zeitschrift für Physik, 31(1):765–783, February 1925. CODEN

ZEPYAA. ISSN 0044-3328. URL http://www.springerlink.
com/content/20w1m0vr050j033r/fulltext.pdf. This is the

paper in which Pauli introduced the famous Exclusion Principle, for which he received the Nobel Prize in Physics 1945 [Pau45].

Pauli:1930:BRP

[Pau30] Wolfgang Pauli. Book review: P. A. M. Dirac, The principles of quantum mechanics. Die Naturwissenschaften, 19(??):188, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1943:DNM

[Pau43] W. Pauli. On Dirac’s new method of field quantization. Reviews of Modern Physics, 15(3):175–207, July 1943. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.15.175;   http://rmp.aps.org/
abstract/RMP/v15/i3/p175_1.

Pauli:1945:NPP

[Pau45] Wolfgang Pauli. The Nobel Prize in Physics 1945. Nobel-

prize.org, 1945. URL http://www-groups.dcs.st-and.ac.
uk/~history/Extras/Pauli_Exclusion_Principle.html;
http://www.nobelprize.org/nobel_prizes/physics/laureates/

1945/pauli-bio.html. The Nobel Prize in Physics 1945 was awarded to Wolfgang Pauli “for the discovery of the Exclusion Principle, also called the Pauli Principle”.

Puppi:1968:ONP

[PB68] G. Puppi and Gilberto Bernardini, editors. Old and new problems in elementary particles. Academic Press, New York, USA, 1968. xii + 309 pp. LCCN QC721 .O4 1968. A volume dedicated to Gilberto Bernardini in his sixtieth birthday.

Peebles:1962:TMD

[PD62] P. J. E. Peebles and R. D. Dicke. The temperature of meteorites and Dirac’s cosmology and Mach’s principle. Journal of Geophysical Research, 67(10):4063–4070, September 1962. CODEN JGREA2. ISSN 0148-0227 (print), 2156-2202 (elec-

tronic). URL https://agupubs.onlinelibrary.wiley.com/
doi/10.1029/JZ067i010p04063.

Pegg:1977:FVP

[Peg77] D. T. Pegg. Future variation of Planck’s constant. Nature, 267 (5610):408–409, June 2, 1977. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Peierls:1934:VDT

[Pei34] R. Peierls. The vacuum in Dirac’s theory of the positive electron. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 146:420–441, September 1934. ISSN 0950-1207 (print), 2053-

9150 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1934RSPSA.146..420P.

Peierls:1935:DLL

[Pei35] R. Peierls. Die diracsche Löchertheorie und die Lichtgeschwindigkeit im Vakuum. (German) [Dirac’s hole theory and the speed of light in a vacuum]. Physica, 2:399–402, January 1935. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767

(electronic). URL https://ui.adsabs.harvard.edu/#abs/
1935Phy.....2..399P.

Peierls:1936:ISE

[Pei36] Rudolf Peierls. Interpretation of Shankland’s experiment. Nature, 137(3474):904, May 1, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Peierls:1942:NAP

[Pei42] Rudolf Peierls. Notes on the attached paper (AB4/42) by Dirac. Report BR42a (PRO AB4/42a), Directorate of Tube Alloys, ????, UK, June 1942.

Peierls:1972:FDS

[Pei72] R. Peierls. Fermi–Dirac statistics. In Salam and Wigner [SW72], pages 117–127. ISBN 0-521-08600-0. LCCN

QC174.1 .A85 1972.  URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Peierls:1985:OPA

[Pei85] Rudolf Peierls. Obituary: Paul Adrien Maurice Dirac. Physics Today, 38(1):111, January 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Peierls:1987:ADM

[Pei87a] Rudolf Peierls. Address to Dirac Memorial Meeting, Cambridge. In Taylor [Tay87b], pages 35–37. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://www.loc.gov/

catdir/enhancements/fy0745/87153334-d.html. Based on

the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Peierls:1987:DW

[Pei87b] Rudolf Peierls. Dirac’s way. In Kurşunoğlu and Wigner [KW87], pages 43–45. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Peierls:1992:BSD

[Pei92] R. Peierls. Broken symmetries: (Dirac memorial lecture, Cambridge, 15 June 1992). Contemporary Physics, 33(4):221–226, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:1997:AH

[Pei97] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Penrose:1997:MES

[Pen97] Roger Penrose. The mathematics of the electron’s spin. European Journal of Physics, 18(3):164–168, May 1997. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL

http://stacks.iop.org/0143-0807/18/i=3/a=006.

Pengpan:1999:KCD

[Pen99] Teparksorn Pengpan. Kostant’s cubic Dirac operator of Lie superalgebras. Journal of Mathematical Physics, 40(12):6577– 6588, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peres:1967:CFE

[Per67] Asher Peres. Constancy of the fundamental electric charge. Physical Review Letters, 19(22):1293–1294, November 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.19.1293.

Perlmutter:1980:RDH

[Per80] Arnold Perlmutter, editor. Recent developments in high-energy physics: [proceedings of Orbis Scientiae 1980, Coral Gables, Florida, Jan. 14–17, 1980], volume 17 of Studies in the natural science. Plenum Press, New York, NY, USA; London, UK, 1980. ISBN 0-306-40565-2. LCCN QC793.28 .O72 1980.

Petley:1985:FPC

[Pet85] B. W. (Brian William) Petley. The fundamental physical constants and the frontier of measurement. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-388-2. x + 347 pp. LCCN QC100.A2 P48 1985. This book has considerable discussion of Dirac’s proposal [Dir37b] that physical constants may vary slowly over cosmological time, and gives experimental evidence that severely limits such variations. See [Har28] for the original definition of atomic units, and [SH59, McW73] for updates.

Petley:1988:FPC

[Pet88] B. W. (Brian William) Petley. The fundamental physical constants and the frontier of measurement. Adam Hilger Ltd., Bristol, UK, 1988. ISBN 0-85274-388-2. x + 347 pp. LCCN QC100.A2 P48 1988.

Petsche:2011:PFG

[Pet11] Hans-Joachim Petsche, editor. From past to future: Graßmann’s work in context: Graßmann Bicentennial Conference, September 2009. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2011. ISBN 3-0346-0404-1 (print), 3-0346-0405-X (e-book). LCCN QA29.G73 F76 2011. URL http://

www.springerlink.com/content/978-3-0346-0405-5.

Peierls:1943:ARN

[PFD43] Rudolf Peierls, Klaus Fuchs, and P. A. M. Dirac. Approximate rate of neutron multiplication for a solid of arbitrary shape and uniform density: Part II, Application to the oblate spheroid, hemisphere and oblate hemispheroid. Report MS-D#.5 (PRO AB4/963), Directorate of Tube Alloys, ????, UK, 1943.

Phillips:1987:PRS

[Phi87] R. J. N. Phillips. Some words from a former student. In Taylor [Tay87b], pages 31–32. ISBN 0-85274-480-3. LCCN QC16.D57

T75 1987. US$10.00. URL http://www.loc.gov/catdir/
enhancements/fy0745/87153334-d.html. Based on the papers

presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Pickering:1981:CCC

[Pic81] Andrew Pickering. Constraints on controversy: The case of the magnetic monopole. Social Studies of Science, 11(1):63–93, February 1, 1981. CODEN SSSCDH. ISSN 0306-3127 (print),

1460-3659 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/030631278101100104.

Pichon:1989:NCG

[Pic89] Bernard Pichon. Numerical calculation of the generalized Fermi–Dirac integrals. Computer Physics Communications, 55(2):127–136, September 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465589900714.

Pickover:2008:AHL

[Pic08] Clifford A. Pickover. Archimedes to Hawking: laws of science and the great minds behind them. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-533611-9 (hardcover). ix + 514 pp. LCCN Q175.32.R45 P53

2008. URL http://catdir.loc.gov/catdir/toc/ecip087/
2007051167.html;   http://www.gbv.de/dms/goettingen/
555633241.pdf;   http://www.oup.com/us/catalog/general/
subject/Physics/?view=usa&ci=9780195336115.

Petsche:2009:HGR

[PKKL09] Hans-Joachim Petsche, Lloyd Kannenberg, Gottfried Keßler, and Jolanta Liskowacka, editors. Hermann Graßmann: roots and traces: autographs and unknown documents. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2009. ISBN 3-0346-0154-9 (hardcover), 3-0346-0155-7 (e-book). xi + 256 pp. LCCN QA28 .H47 2009.

Plotnitsky:2015:MPP

[Plo15] Arkady Plotnitsky. A matter of principle: The principles of quantum theory, Dirac’s equation, and quantum information. Foundations of Physics, 45(10):1222–1268, October 2015. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/s10701-
015-9928-z.

Podles:1997:DOG

[Pod97] P. Podleś. The Dirac operator and gamma matrices for quantum Minkowski spaces. Journal of Mathematical Physics, 38(9):

4474–4491, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Polkinghorne:1958:BRP

[Pol58] J. C. Polkinghorne. Book review: The Principles of Quantum Mechanics, by P. A. M. Dirac. Physics Today, 11(6):32–33, June 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Polkinghorne:1987:FD

[Pol87a] J. C. Polkinghorne. At the feet of Dirac. In Kurşunoğlu and Wigner [KW87], pages 227–229. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Polkinghorne:1987:BRD

[Pol87b] J. C. Polkinghorne. A brief reminiscence of Dirac. In Taylor [Tay87b], pages 33–34. ISBN 0-85274-480-3. LCCN QC16.D57

T75 1987. US$10.00. URL http://www.loc.gov/catdir/
enhancements/fy0745/87153334-d.html. Based on the papers

presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Polkinghorne:1987:DIQ

[Pol87c] J. C. Polkinghorne. Dirac and the interpretation of quantum mechanics. In Taylor [Tay87b], pages 76–83. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://www.loc.

gov/catdir/enhancements/fy0745/87153334-d.html. Based

on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Pons:2005:DIA

[Pon05] Josep M. Pons. On Dirac’s incomplete analysis of gauge transformations. Studies in History and Philosophy of Modern Physics, 36(3):491–518, September 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1355219805000456.

Portugal:1995:ESD

[Por95] Renato Portugal. Exact solutions of the Dirac equations in an anisotropic cosmological background. Journal of Mathematical Physics, 36(8):4296–4300, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Post:1984:DMD

[Pos84] E. J. Post. Is Dirac’s monopole a dipole? Physics Today, 37 (12):90–91, December 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Percus:1971:CND

[PP71] J. K. Percus and N. L. Petrakopoulos. On Clifford numbers, Dirac and relativistic Hamilton–Jacobi equations. Journal of Mathematical Physics, 12(12):2516–2520, December 1971. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patra:1986:SES

[PR86] Amar Chandra Patra and Dipankar Ray. Some exact solutions of Einstein–Dirac–Maxwell fields and a massive neutrino. Journal of Mathematical Physics, 27(2):568–572, February 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prugovecki:1973:BKF

[Pru73] Eduard Prugovecki. The bra and ket formalism in extended Hilbert space. Journal of Mathematical Physics, 14(10):1410– 1422, October 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v14/i10/p1410_s1.

Pauli:1932:TUEa

[PS32a] W. Pauli and J. Solomon. La théorie unitaire d’Einstein et Mayer et les équations de Dirac. (French) [Dirac’s equations and the unitary theory of Einstein and Mayer]. Journal de Physique et le Radium, 3(10):452–463, ???? 1932. CODEN JPRAAJ. ISSN 0368-3842.

Pauli:1932:TUEb

[PS32b] W. Pauli and J. Solomon. La théorie unitaire d’Einstein et Mayer et les équations de Dirac II. (French) [Dirac’s equations and the unitary theory of Einstein and Mayer. II]. Journal de Physique et le Radium, 3(12):582–589, ???? 1932. CODEN JPRAAJ. ISSN 0368-3842.

Perlmutter:1979:PAE

[PS79] Arnold Perlmutter and Linda F. Scott, editors. On the Path of Albert Einstein, volume 15 of Studies in the Natural Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-306-40296-3, 1-4684-3598-1 (print), 1-4684-3596-5 (e-book). LCCN QC19.2-20.85; QC178

.O72 1979. URL http://www.springerlink.com/content/
978-1-4684-3596-2.

Prykarpatskyj:1994:ASG

[PSA94] A. K. Prykarpatskyj, V. Hr. Samoilenko, and R. I. Andrushkiw. Algebraic structure of the gradient-holonomic algorithm for Lax integrable nonlinear dynamical systems. II. The reduction via Dirac and canonical quantization procedure. Journal of Mathematical Physics, 35(8):4088–4116, August 1994. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Panahi:1986:SDK

[PT86] M. Panahi and R. W. Tucker. Separation of Dirac and Kähler equations in spherically symmetric space–times. Journal of Mathematical Physics, 27(5):1398–1403, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quiney:2001:DEA

[QGW01] H. M. Quiney, V. N. Glushkov, and S. Wilson. The Dirac equation in the algebraic approximation. VII. A comparison of molecular finite difference and finite basis set calculations using distributed Gaussian basis sets. Advances in quantum chemistry, 39(??):241–259, 2001. CODEN AQCHA9. ISSN 0065-3276.

Rauscher:2011:OMP

[RA11] Elizabeth A. Rauscher and Richard L. Amoroso. Orbiting the moons of Pluto: complex solutions to the Einstein, Maxwell, Schrödinger, and Dirac equations, volume 45 of Series on knots and everything. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4324-24-8 (hardcover), 981-4324-25-6 (e-book). xvii + 391 pp. LCCN QC173.55 .R34 2011.

Radford:1996:LSD

[Rad96] C. J. Radford. Localized solutions of the Dirac–Maxwell equations. Journal of Mathematical Physics, 37(9):4418–4433, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramond:1989:BRB

[Ram89] Pierre Ramond. Book review: Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac, by Behram N. Kurşunoğlu and Eugene P. Wigner. American Scientist, 77 (1):71–72, January 1989. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/27855557.

Ranganathan:1987:GID

[Ran87] D. Ranganathan. A geometric interpretation for the Dirac field in curved space. Journal of Mathematical Physics, 28(10): 2437–2439, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reader:2013:WYN

[RC13] Joseph Reader and Charles W. Clark. 1932, a watershed year in nuclear physics. Physics Today, 66(3):44–49, March 2013. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See also [CR12], and corrections and comments [AG13].

Rapisarda:2006:ICE

[RCC^{+}06] A. Rapisarda, P. Castorina, F. Catara, E. Migneco, S. Lo Nigro, F. Porto, and E. Rimini, editors. International Conference on Ettore Majorana’s Legacy and the Physics the 21st Century in Commemoration of the Centennial of Majorana’s Birth: EMC2006, October 5–6, 2006, Dipartimento di Fisica e Astronomia, Università di Catania, Italy. ????, ????, 2006. LCCN

????  URL http://pos.sissa.it/cgi-bin/reader/conf.
cgi?confid=37;   http://www.ct.infn.it/Majorana2006/.

Rodrigues:2007:MFM

[RdO07] W. A. Rodrigues and Edmundo Capelas de Oliveira. The many faces of Maxwell, Dirac and Einstein equations: a Clifford bundle approach, volume 722 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-71292-5 (hardcover), 3-540-71293-3 (e-book). ISSN 0075-8450. xiv + 445 pp. LCCN

QC173.59.S65 R63 2007. URL http://www.loc.gov/catdir/
enhancements/fy0826/2007923174-d.html;   http://www.
loc.gov/catdir/enhancements/fy0826/2007923174-t.html.

Rodrigues:2016:MFM

[RdO16] W. A. Rodrigues and Edmundo Capelas de Oliveira. The many faces of Maxwell, Dirac and Einstein equations: a Clifford bundle approach, volume 922 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2016. ISBN 3-319-27636-0 (print), 3-319-27637-9 (e-book). ISSN 0075-8450 (print), 1616-6361 (electronic). xvi + 587 pp. LCCN QC173.59.S65. URL http://

link.springer.com/10.1007/978-3-319-27637-3.

Rechenberg:1997:EPS

[Rec97] H. Rechenberg. The electron in physics — selection from a chronology of the last 100 years. European Journal of Physics, 18(3):145–149, May 1997. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://stacks.iop.org/

0143-0807/18/i=3/a=004.

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Reines:1972:CFOa

[Rei72a] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reines:1972:CFOb

[Rei72b] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Adam Hilger Ltd., Bristol, UK, 1972. ISBN 0-85274-223-1. xiv + 320 pp. LCCN QC780.

URL http://adsabs.harvard.edu/abs/1972cht..conf.....
R;   http://adsabs.harvard.edu/abs/1972cofu.book....
.R.

Reuter:1999:SDK

[Reu99] M. Reuter. Symplectic Dirac–Kähler fields. Journal of Mathematical Physics, 40(11):5593–5640, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rhodes:1950:FDF

[Rho50] P. Rhodes. Fermi–Dirac functions of integral order. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 204(1078):396–405, December 22, 1950. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.jstor.

org/stable/98693.

Richard:2003:DDC

[Ric03] Christoph Richard. Dense Dirac combs in Euclidean space with pure point diffraction. Journal of Mathematical Physics, 44(10): 4436–4449, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rigden:2002:HEE

[Rig02] John S. Rigden. Hydrogen: the essential element. Harvard University Press, Cambridge, MA, USA, 2002. ISBN 0-674-00738-7. 280 pp. LCCN QD181.H1 R54 2002. URL http://catdir.loc.

gov/catdir/toc/fy022/2001051708.html.

Rijnierse:1966:HDE

[Rij66] P. J. Rijnierse. High density expansion for the Thomas–Fermi– Dirac function. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 292(1429):288– 297, May 17, 1966. CODEN PRLAAZ. ISSN 0080-4630. URL

http://www.jstor.org/stable/2415718.

Rivas:1994:QGS

[Riv94] Martín Rivas. Quantization of generalized spinning particles: new derivation of Dirac’s equation. Journal of Mathematical Physics, 35(7):3380–3399, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reed:1996:RSP

[RK96] Irving S. Reed and Wolfgang F. Kraske. Reflections, spinors, and projections on a Minkowski space underlie Dirac’s equation. Linear Algebra and its Applications, 239 (1–3):227–262, May ??, 1996. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.elsevier.com/cgi-bin/cas/tree/store/laa/cas_sub/
browse/browse.cgi?year=1996&volume=239&issue=1-3&aid=
9500197.

Rosenbaum:1968:IST

[RL68] M. Rosenbaum and C. P. Luehr. Intrinsic spinor techniques with applications to the Lorentz group and the Dirac equation. Journal of Mathematical Physics, 9(12):2225–2236, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reifler:1995:UDE

[RM95] Frank Reifler and Randall Morris. Unification of the Dirac and Einstein Lagrangians in a tetrad model. Journal of Mathematical Physics, 36(4):1741–1752, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reifler:1996:IGB

[RM96] Frank Reifler and Randall Morris. Inclusion of gauge bosons in the tensor formulation of the Dirac theory. Journal of Mathematical Physics, 37(7):3630–3640, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reynolds:1964:SER

[ROB64] J. T. Reynolds, D. S. Onley, and L. C. Biedenharn. Some exact radial integrals for Dirac–Coulomb functions. Journal of Mathematical Physics, 5(3):411–419, March 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i3/p411_s1.

Roberts:1966:DBK

[Rob66] J. E. Roberts. The Dirac bra and ket formalism. Journal of Mathematical Physics, 7(6):1097–1104, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i6/
p1097_s1.

Robinson:1983:NTV

[Rob83] Arthur L. Robinson. New test of variable gravitational constant. Science, 222(4630):1316–1317, December 23, 1983. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See [Dir74b, Dir79g, Jor59a, Rob85].

Robinson:1985:APT

[Rob85] Arthur L. Robinson. Atomic physics tests Lorentz invariance: Accurate measurements of the frequencies of atomic transitions lead to tighter limits on the velocity dependence of the laws of physics. Science, 229(4715):745–747, August 23, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/cgi/

reprint/229/4715/745. See [Dir74b, Dir79g, Jor59a, Rob83].

Robotti:1997:DE

[Rob97] Nadia Robotti. The discovery of the electron: I. European Journal of Physics, 18(3):133–138, May 1997. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://

stacks.iop.org/0143-0807/18/i=3/a=002.

Robinson:2015:EHY

[Rob15] Andrew Robinson. Einstein: a hundred years of relativity. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16989-6 (paperback). 256 pp. LCCN QC16.E5 R63 2015.

Rodrigues:1980:CPK

[Rod80] A. J. Rodrigues. Convergence of Padé kernel approximants to the delayed Dirac function. Journal of the Institute of Mathematics and its Applications, 25(1):17–27, 1980. CODEN JM-TAA8. ISSN 0020-2932.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (ebook). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rohrlich:1960:SES

[Roh60] F. Rohrlich. Self-energy and stability of the classical electron. American Journal of Physics, 28(7):639–643, October 1960. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v28/i7/p639_s1.

Ross:1983:DGM

[Ros83] D. K. Ross. Dirac gravitational magnetic monopoles do not exist. Journal of Mathematical Physics, 24(7):1814–1816, July 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roux:1984:EDH

[Rou84] Jean-François Roux. Exact diagonalization of the Hamiltonian of a Dirac particle with anomalous moment, interacting with an external magnetic field. Journal of Mathematical Physics, 25 (6):2087–2089, June 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rozental:1967:NBH

[Roz67] S. (Stefan) Rozental, editor. Niels Bohr: his life and work as seen by his friends and colleagues. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. 355 pp. LCCN QC16.B63 N53.

Roza:2020:SDM

[Roz20] Engel Roza. On the second dipole moment of Dirac’s particle. Foundations of Physics, 50(8):828–849, August 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Rodriguez-Romo:1993:ASD

[RR93] Suemi Rodríguez-Romo. Algebraic spinors in Dirac propagators as a directed random walk. Journal of Mathematical Physics, 34 (10):4590–4600, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosenfeld:1964:NBH

[RRK^{+}64] Léon Rosenfeld, Erik Rüdinger, Oskar Klein, Werner Heisenberg Hendrik G. B. Casimir, Otto Robert Frisch, Stefan Rozental, Aage Bohr, Abraham Pais, Jørgen Kalckar, Christian Møller, Mogens Pihl, Viktor F. Weisskopf, Johannes Pedersen, Viggo Kampmann, Richard Courant, Paul A. M. Dirac, Hans Henrik Koch, William Scharff, Mogens Andersen, Hans Bohr, and Niels Bohr. Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers]. J. H. Schultz Forlag, Copenhagen, DK, 1964. 341 pp. LCCN ????

Rodl:2006:DTT

[RRS06] Vojtech Rödl, Andrzej Rucinski, and Endre Szemerédi. A Dirac-type theorem for 3-uniform hypergraphs. Combinatorics, Probability and Computing, 15(1–2):229–251, January 2006. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163

(electronic). URL http://journals.cambridge.org/action/
displayIssue?jid=CPC&volumeId=15&issueId=01.

Rebbi:1984:SP

[RS84] Claudio Rebbi and Giulio Soliani, editors. Solitons and particles. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1984. ISBN 9971-966-42-5, 9971-966-43-3 (paperback). xiii + 819 pp. LCCN QC174.26.W28 S635 1984.

Ruotsalainen:1987:SBE

[RS87] K. Ruotsalainen and J. Saranen. Some boundary element methods using Dirac’s distributions as trial functions. SIAM Journal on Numerical Analysis, 24(4):816–827, August 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rigden:2010:BNG

[RS10] John S. Rigden and Roger H. Stuewer. Book notes [Graham Farmelo, The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom (Basic Books, 2009, 539 pages, $29.95); Vlatko Vedral, Decoding Reality: The Universe as Quantum Information (Oxford University Press, 2010, x + 229 pages, $29.95)]. Physics in Perspective (PIP), 12(3):350–352, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-010-0031-8.

Rathe:1999:CSS

[RSK99] U. W. Rathe, P. Sanders, and P. L. Knight. A case study in scalability: An ADI method for the two-dimensional timedependent Dirac equation. Parallel Computing, 25(5):525–533, May 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-

7336 (electronic). URL http://www.elsevier.com/cas/tree/
store/parco/sub/1999/25/5/1403.pdf.

Rudolph:1999:ESE

[RTV99] Gerd Rudolph, Torsten Tok, and Igor Volobuev. Exact solutions in Einstein–Yang–Mills–Dirac systems. Journal of Mathematical Physics, 40(11):5890–5904, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:1981:BSC

[RU81] Antonio F. Rañada and Juan M. Usón. Bound states of a classical charged nonlinear Dirac field in a Coulomb potential. Journal of Mathematical Physics, 22(11):2533–2538, November 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudiger:1984:SSD

[Rüd84] Roland Rüdiger. Separable systems for the Dirac equation in curved space–times. Journal of Mathematical Physics, 25(3): 649–654, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruffini:1982:PSM

[Ruf82] Remo Ruffini, editor. Proceedings of the Second Marcel Grossmann Meeting on General Relativity: organized and held at the International Centre for Theoretical Physics, Trieste 5–11 July, 1979. North-Holland Publishing Co., Amsterdam, The Netherlands, 1982. ISBN 0-444-86357-5 (set). LCCN QC173.6 .M37 1979.

Ruggeri:1988:DBR

[Rug88] G. J. Ruggeri. Dirac bracket revisited. Journal of Mathematical Physics, 29(2):362–364, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rumpf:1986:SGP

[Rum86] Helmut Rumpf. A simple Grassmannian path integral representation of the Dirac propagator. Journal of Mathematical Physics, 27(6):1649–1654, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ryckman:2017:BRV

[Ryc17] Thomas Ryckman. Book review: Void: The Strange Physics of Nothing, James Owen Weatherall, Yale U. Press, 2016, 196 p, $26.00, ISBN 978-0-300-20998-3. Physics Today, 70(9):59–60, September 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Ruzicka:1980:FYD

[RZ80] J. Ružička and V. P. Zrelov. Fifty years of Dirac Monopole: Complete Bibliography. ????, ????, 1980. ISBN ???? 232

pp. LCCN ???? URL http://lss.fnal.gov/archive/other/
jinr-51-2-80-850-pt1.pdf.

Sagar:1991:GQC

[Sag91] Robin P. Sagar. A Gaussian quadrature for the calculation of generalized Fermi–Dirac integrals. Computer Physics Communications, 66(2–3):271–275, September/October 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/001046559190076W.

Salam:1969:CPT

[Sal69] Abdus Salam, editor. Contemporary physics: Trieste Symposium 1968: proceedings of the International Symposium on Contemporary Physics organized by and held at the International Centre for Theoretical Physics, Trieste, from 7 to 28 June 1968, volume 214 of Publication / Division of Scientific and Technical Information, International Atomic Energy Agency. International Atomic Energy Agency, Vienna, Austra, 1969. LCCN ????

Salam:1987:DFFa

[Sal87a] Abdus Salam. Dirac and finite field theories. In Kurşunoğlu and Wigner [KW87], pages 262–275. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Salam:1987:DFFb

[Sal87b] Abdus Salam. Dirac and finite field theories. In Taylor [Tay87b], pages 84–95. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Sanz:1979:LDE

[San79] J. L. Sanz. On the Lorentz–Dirac equation. Journal of Mathematical Physics, 20(11):2334–2338, November 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Synge:1962:SAC

[SBB^{+}62] J. L. Synge, H. Bondi, W. B. Bonnor, C. B. Rayner, W. Wilson, P. A. M. Dirac, and A. Lichnerowicz. Systematic approximations in the calculation of gravitational fields [and discussion]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 270(1342):315–326, November 27, 1962. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/2414530.

Salumbides:2008:ILV

[SBK^{+}08] E. J. Salumbides, D. Bailly, A. Khramov, A. L. Wolf, K. S. E. Eikema, M. Vervloet, and W. Ubachs. Improved laboratory values of the h_{2} Lyman and Werner lines for constraining time variation of the proton-to-electron mass ratio. Physical Review Letters, 101(22):223001, November 28, 2008. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.101.223001.

Shaarawi:1990:NAS

[SBZ90] Amr M. Shaarawi, Ioannis M. Besieris, and Richard W. Ziolkowski. A novel approach to the synthesis of nondispersive wave packet solutions to the Klein–Gordon and Dirac equations. Journal of Mathematical Physics, 31(10):2511–2519, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:1978:CLD

[sC78] Kuo shung Cheng. Constraints of the Lorentz–Dirac equation. Journal of Mathematical Physics, 19(8):1656–1657, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shishkin:1991:DEE

[SC91] German V. Shishkin and William D. Cabos. The Dirac equation in external fields: variable separation in Cartesian coordinates.

Journal of Mathematical Physics, 32(11):3184–3188, November 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shishkin:1992:DEEb

[SC92a] German V. Shishkin and William D. Cabos. Dirac equation in external fields: separation of variables in curvilinear coordinates. Journal of Mathematical Physics, 33(3):914–925, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shishkin:1992:DEEa

[SC92b] German V. Shishkin and William D. Cabos. Dirac equation in external fields: separation of variables in nondiagonal metrics. Journal of Mathematical Physics, 33(1):297–303, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schouten:1931:DEGa

[Sch31a] J. A. Schouten. Dirac equations in general relativity: 1. Four dimensional theory. Journal of Mathematics and Physics (MIT), 10(1–4):239–271, April 1931. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/
10.1002/sapm1931101239.

Schouten:1931:DEGb

[Sch31b] J. A. Schouten. Dirac equations in general relativity: 2. Five dimensional theory. Journal of Mathematics and Physics (MIT), 10(1–4):272–283, April 1931. CODEN JMPHA9. ISSN 0097-

1421. URL https://onlinelibrary.wiley.com/doi/epdf/
10.1002/sapm1931101272.

Scherzer:1938:IED

[Sch38] O. Scherzer. Die imaginäre Einheit in der Diracgleichung. (German) [The imaginary unit in the Dirac equation]. Annalen der Physik (1900), 425(7):591–593, 1938. ISSN 1521-3889.

Schrodinger:1940:MDE

[Sch41] Erwin Schrödinger. Maxwell’s and Dirac’s equations in the expanding Universe. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46(??):25–47, ???? 1940–1941. CODEN PRIAAK. ISSN 0035-8975. URL http://

www.jstor.org/stable/20490746.

Schrodinger:1952:DNE

[Sch52] Erwin Schrödinger. Dirac’s new electrodynamics. Nature, 169 (4300):538, March 29, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v169/n4300/pdf/169538a0.pdf.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schechter:1969:BRL

[Sch69] Joseph Schechter. Book review: Lectures on Quantum Field Theory, by P. A. M. Dirac. American Journal of Physics, 37(2): 233, February 1969. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/
1/ajpias/v37/i2/p233_s1.

Schwartz:1972:FNF

[Sch72] L. Schwartz. La fonction δ et les noyaux. (French) [The δ function and nuclei]. In Salam and Wigner [SW72], pages 179–182. ISBN 0-521-08600-0. LCCN QC174.1 .A85 1972. URL http://

hooke.lib.cam.ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=
1733506;   http://www.loc.gov/catdir/enhancements/fy1001/
72075298-d.html;   http://www.loc.gov/catdir/enhancements/
fy1001/72075298-t.html.

Schopf:1985:BRB

[Sch85] H.-G. Schöpf. Book review: Barut, A. O. /van der Merwe, A. /Vigier, J.-P. (eds.), Quantum, Space and Time — The Quest Continues. Studies and Essays in Honour of Louis de Broglie, Paul Dirac and Eugene Wigner. Cambridge et al., Cambridge University Press 1984. Vii, 662 S., £25.00 A P/ b. US $49.50. ISBN 0-521-31911-0. (Cambridge Monographs on Physics). Zeitschrift für Angewandte Mathematik und Mechanik, 65(10), 1985. CODEN ZAMMAX. ISSN 0044-2267 (print),

1521-4001 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/zamm.19850651014/abstract.

Schweber:1988:BRB

[Sch88] S. S. Schweber. Book review: Behram N. Kurşunoğlu and Eugene P. Wigner: Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac. Isis, 79(2):356–357, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233680.

Schmutzer:1991:PUF

[Sch91a] E. Schmutzer. Projective unified field theory in context with the cosmological term and the variability of the gravitational constant. In Zichichi et al. [ZdSS91], page ?? ISBN 0-306-44054-7. LCCN QC178 .G63 1991. URL http://www.loc.gov/

catdir/enhancements/fy0820/91040369-t.html.

Schweber:1991:BRD

[Sch91b] Sylvan S. Schweber. Book review: Dirac: A Scientific Biography by Helge S. Kragh. Physics Today, 44(4):99–100, April 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Schweber:1994:QMW

[Sch94a] S. S. (Silvan S.) Schweber. QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). xxviii + 732 pp. LCCN QC680 .S34 1994. US$72.50.

URL http://www.loc.gov/catdir/description/prin021/
93033550.html;   http://www.loc.gov/catdir/toc/prin031/
93033550.html.

Schweber:1994:PMD

[Sch94b] Sylvan S. Schweber. P. A. M. Dirac and the birth of quantum electrodynamics. In QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga [Sch94a], pages 11–32. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). LCCN QC680

.S34 1994. US$72.50. URL http://www.loc.gov/catdir/
description/prin021/93033550.html;   http://www.loc.
gov/catdir/toc/prin031/93033550.html.

Schweber:1994:PDS

[Sch94c] Sylvan S. Schweber. Postscript: Dirac and scientific creativity. In QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga [Sch94a], pages 70–71. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). LCCN QC680

.S34 1994. US$72.50. URL http://www.loc.gov/catdir/
description/prin021/93033550.html;   http://www.loc.
gov/catdir/toc/prin031/93033550.html.

Schucking:1999:JPP

[Sch99] Engelbert L. Schucking. Jordan, Pauli, politics, Brecht, and a variable gravitational constant. Physics Today, 52 (10):26–31, October 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/adsabs.harvard.edu/abs/1999PhT....52j..26S;   http:
//link.aip.org/link/phtoad/v52/i10/p26/s1;   http://
www.physicstoday.org/resource/1/phtoad/v52/i10/p26_s1.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schwarzschild:2006:SMM

[Sch06] Bertram Schwarzschild. Search for magnetic monopoles at the Tevatron sets new upper limit on their production. Physics Today, 59(7):16–18, July 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Srianand:2004:LTV

[SCPA04] R. Srianand, H. Chand, P. Petitjean, and B. Aracil. Limits on the time variation of the electromagnetic fine-structure constant in the low energy limit from absorption lines in the spectra of distant quasars. Physical Review Letters, 92(12):121302, March 26, 2004. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.92.121302. See comment [MWF07].

Semay:1993:TBD

[SCSB93] C. Semay, R. Ceuleneer, and B. Silvestre-Brac. Two-body Dirac equation with diagonal central potentials. Journal of Mathematical Physics, 34(6):2215–2225, June 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salam:1990:UFF

[SDEW90] Abdus Salam, P. A. M. (Paul Adrien Maurice) Dirac, Jonathan (Jonathan M.) Evans, and Gerard Watts. Unification of fundamental forces: the first of the 1988 Dirac Memorial Lectures. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37140-6 (hard covers). ix + 143 pp. LCCN QC794.6.G7 S25 1989. URL http://www.loc.

gov/catdir/description/cam023/89015861.html;   http:
//www.loc.gov/catdir/toc/cam028/89015861.html.

Sebens:2019:HES

[Seb19] Charles T. Sebens. How electrons spin. Studies in History and Philosophy of Modern Physics, 68(??):40–50, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981830087X.

Seiler:1936:ADA

[Sei36] Karl Seiler. Zur atomaren Dispersion und Absorption von Röntgenstrahlen nach der relativistischen Wellenmechanik von Dirac. (German) [On atomic dispersion and X-ray absorption in the relativistic wave mechanics of Dirac]. Annalen der Physik (1900), 5(27)(4):329–372, 1936. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Seiler:1999:NIC

[Sei99] Werner M. Seiler. Numerical integration of constrained Hamiltonian systems using Dirac brackets. Mathematics of Computation, 68(226):661–681, April 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/jourcgi/jour-pbprocess?fn=110&arg1=S0025-5718-
99-01010-8&u=/mcom/1999-68-226/.

Selesnick:1994:DEQ

[Sel94] S. A. Selesnick. Dirac’s equation on the quantum net. Journal of Mathematical Physics, 35(8):3936–3958, August 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Semay:1993:VTT

[Sem93] Claude Semay. Virial theorem for two-body Dirac equation. Journal of Mathematical Physics, 34(5):1791–1793, May 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Service:2012:PDN

[Ser12] Robert F. Service. Physicists discover new type of particle—sort of. Science Now, ??(??):??, April 12, 2012. URL http://news.

sciencemag.org/sciencenow/2012/04/majorana-fermions-
found.html.

Steeb:1981:SDE

[SES81] W.-H. Steeb, W. Erig, and W. Strampp. Symmetries and the Dirac equation. Journal of Mathematical Physics, 22(5):970– 973, May 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steeb:1982:SSN

[SES82] W.-H. Steeb, W. Erig, and W. Strampp. Similarity solutions of nonlinear Dirac equations and conserved currents. Journal of Mathematical Physics, 23(1):145–153, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salvat:1995:ANS

[SFVW95] F. Salvat, J. M. Fernández-Varea, and W. Williamson, Jr. Accurate numerical solution of the radial Schrödinger and Dirac wave equations. Computer Physics Communications, 90(1):151–168, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559500039I.

Silbar:2011:SRD

[SG11] Richard R. Silbar and T. Goldman. Solving the radial Dirac equations: a numerical odyssey. European Journal of Physics, 32(1):217, 2011. CODEN EJPHD4. ISSN 0143-0807 (print),

1361-6404 (electronic). URL http://stacks.iop.org/0143-
0807/32/i=1/a=021.

Shull:1959:AU

[SH59] H. Shull and G. G. Hall. Atomic units. Nature, 184(4698):1559– 1560, November 14, 1959. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Schucking:1999:EPE

[SH99] E. L. (Engelbert L.) Schucking and Alex Harvey, editors. On Einstein’s path: essays in honor of Engelbert Schucking. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98564-6 (hardcover). LCCN QC173.5 .O6 1999. URL http://www.loc.gov/

catdir/enhancements/fy0816/98020292-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0816/98020292-t.html.

Shanmugadhasan:1987:DRS

[Sha87] S. Shanmugadhasan. Dirac as research supervisor and other remembrances. In Taylor [Tay87b], pages 48–57. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http:

//www.loc.gov/catdir/enhancements/fy0745/87153334-d.

html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Salam:1991:UFF

[SHD91] Abdus Salam, W. Heisenberg, and P. Dirac. La unificaciön de las fuerzas fundamentales. (Spanish) [Unification of fundamental forces]. Gedisa, Barcelona, Spain, 1991. ISBN 84-7432-413-0. ???? pp. LCCN ???? URL http://

148.201.96.14/dc/ver.aspx?ns=000033655;   http://en.
scientificcommons.org/8713570. Notes by J. Evans and G.

Watts. Preface by John C. Taylor. Translated to Spanish by Alberto L. Bixio.

Simoes:2002:DCC

[Sim02] Ana Simões. Dirac’s claim and the chemists. Physics in Perspective (PIP), 4(3):253–266, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-002-8369-1.

Skyrme:1971:KDE

[Sky71] T. H. R. Skyrme. Kinks and the Dirac equation. Journal of Mathematical Physics, 12(8):1735–1742, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Slater:1929:TCS

[Sla29] J. C. Slater. The theory of complex spectra. Physical Review (2), 34(10):1293–1322, November 15, 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v34/i10/p1293_1.

Sommerfeld:1935:VNA

[SM35]  Arnold Sommerfeld and A. W. Maue.  Verfahren zur

näherungsweisen Anpassung einer Lösung der Schrödinger- an die Dirac-Gleichung. (German) [Method for the approximate adaptation of a solution of the Schrödinger equation to the Dirac equation]. Annalen der Physik (1900), 414(7):629–642, ???? 1935. ISSN 0003-3804 (print), 1521-3889 (electronic).

Salvat:1991:ANS

[SM91] Francesc Salvat and Ricardo Mayol. Accurate numerical solution of the Schrödinger and Dirac wave equations for central fields. Computer Physics Communications, 62(1):65–79, January 1991. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465591901222.

Sparber:2003:SAM

[SM03] C. Sparber and P. Markowich. Semiclassical asymptotics for the Maxwell–Dirac system. Journal of Mathematical Physics, 44 (10):4555–4572, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sharma:1983:SDE

[SMI83] L. K. Sharma, Sheilly Mehta, and V. P. Iyer. Solution of the Dirac equation for the general even power potential with application to particle spectroscopy. Journal of Mathematical Physics, 24(10):2521–2527, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smorodinskiy:1987:PMD

[Smo84] Ya. A. Smorodinskiy. P. A. M. Dirac (8.VIII.1902–20.X.1984). Uspekhi Fizicheskikh Nauk, 148(3):527–534, March 1984. CO-DEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

URL http://ufn.ru/ru/articles/1986/3/e/. English trans-

lation in Sov. Phys. Usp. 29 281–285 (1986).

Smorodinskiy:1987:SEA

[Smo87] Ya. A. Smorodinskiy. Several episodes [about Dirac]. Uspekhi Fizicheskikh Nauk, 153(9):187–190, September 1987. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1987/9/g/. English translation

in Sov. Phys. Usp. 30 823–825 (1987).

Smorodinski:1993:HDB

[Smo93] Yakov A. Smorodinski. Heisenberg und Dirac: Die Bedeutung des Schönen in der Naturwissenschaft. (German) [Heisenberg und Dirac: The meaning of beauty in science].

Physikalische Blätter, 49(5):436–438, May 1993. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19930490518/abstract.

Snow:1973:BSH

[Sno73] C. P. Snow. The banquet of the symposium — in honour of Paul Dirac, including an address on: The classical mind. In Mehra [Meh73], chapter 44, pages 805–819. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://

www.springer.com/us/book/9789027703453.

Sommerfeld:1934:LEM

[Som34] Arnold Sommerfeld. Literaturberichte: L’électron magnétique. (German) [Literature report: the magnetic electron]. Monatshefte für Mathematik und Physik, 41(1):A24, December 1934. CODEN MMPHA8. ISSN 1812-8076. URL http://link.

springer.com/article/10.1007/BF01697933. L. de Broglie,

(Théorie de Dirac). 315 S. Hermann et Cie., Paris 1934. Preis Frs. 100.

Sommerfeld:1936:KPI

[Som36] Arnold Sommerfeld. Über die Klein’schen Parameter α, β, γ, δ und ihre Bedeutung für die Dirac-Theorie. (German) [On the Klein parameters α, β, γ, δ and their meaning for the Dirac theory]. Sitzungsberichte der Akademie der Wissenschaften in Wien, Mathemat.-naturwissenschaftl. Kl. IIa, Wien, 145(9–10): 639–650, 1936.

Sonego:1999:CBL

[Son99] Sebastiano Sonego. Conformal behavior of the Lorentz–Dirac equation and Machian particle dynamics. Journal of Mathematical Physics, 40(8):3918–3924, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steeb:1984:IND

[SOS84] W.-H. Steeb, W. Oevel, and W. Strampp. On the integrability of nonlinear Dirac equations. Journal of Mathematical Physics, 25 (7):2331–2335, July 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Srivastava:1983:FUS

[Sri83] S. K. Srivastava. Friedmann Universe and superluminal Dirac particles. Journal of Mathematical Physics, 24(4):996–1000, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Srivastava:1989:SDE

[Sri89] Sushil K. Srivastava. Solution of the Dirac equation in Kasner’s space–time. Journal of Mathematical Physics, 30(12): 2838–2844, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salam:1972:PIQ

[SS72] A. Salam and J. Strathdee. The path-integral quantization of gravity. In Salam and Wigner [SW72], pages 249–268. ISBN 0-521-08600-0. LCCN QC174.1 .A85 1972. URL http://

hooke.lib.cam.ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=
1733506;   http://www.loc.gov/catdir/enhancements/fy1001/
72075298-d.html;   http://www.loc.gov/catdir/enhancements/
fy1001/72075298-t.html.

Shishkin:1998:DPP

[SS98] G. V. Shishkin and W. H. Abdel Salam. The Dirac particle in the presence of the tensor wave on the background of gravitation. Journal of Mathematical Physics, 39(1):454–466, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sanyuk:2003:DCP

[SS03] V. I. Sanyuk and A. D. Sukhanov. Dirac in 20th century physics: a centenary assessment. Physics-Uspekhi, 46(9):937– 956, September 2003. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://iopscience.iop.

org/1063-7869/46/9/A03/pdf/1063-7869_46_9_A03.pdf.

Sidje:2011:RAF

[SS11] Roger B. Sidje and Yousef Saad. Rational approximation to the Fermi–Dirac function with applications in density functional theory. Numerical Algorithms, 56(3):455–479, March 2011. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=3&spage=455.

Stanciu:1967:FES

[Sta67] George N. Stanciu. Further exact solutions of the Dirac equation. Journal of Mathematical Physics, 8(10):2043–2047, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i10/p2043_s1.

Stanford:1987:DUM

[Sta87] Henry King Stanford. Dirac at the University of Miami. In Kurşunoğlu and Wigner [KW87], pages 53–56. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://

adsabs.harvard.edu/abs/1987ragp.book.....D;   http:/
/www.loc.gov/catdir/description/cam023/86033409.html;
 http://www.loc.gov/catdir/toc/cam028/86033409.html.

Staley:2012:DFT

[Sta12] Kent W. Staley. Dirac’s ‘Fine-Tuning Problem’: a constructive use of anachronism? Perspectives on Science, 20(4):476–503, Winter 2012. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http:/

/muse.jhu.edu/journals/perspectives_on_science/v020/
20.4.staley.html;   http://muse.jhu.edu/journals/perspectives_
on_science/v020/20.4.staley.pdf;   http://www.mitpressjournals.
org/doi/abs/10.1162/POSC_a_00082.

Stinner:2013:BRG

[Sti13] Arthur Stinner. Book review: Graham Farmelo: The Strangest Man: The Hidden Life of Paul Dirac. Science & Education (Springer), 22(4):887–891, April 2013. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Stoney:1894:XEA

[Sto94] G. Johnstone Stoney. XLIX. of the ‘electron,’ or atom of electricity. Philosophical Magazine, 38(233):418–420, October 1894. CODEN PHMAA4. ISSN 0031-8086. URL http://www.

tandfonline.com/doi/abs/10.1080/14786449408620653. In

this article, the author introduces the term ‘electron’ into the English language, from Greek ‘elektron’ (‘amber’, a substance that could be rubbed with a cloth to produce an electric charge).

Stricker:1971:SDB

[Str71] F. Stricker. [summary of Dirac’s Fundamental problems in physics]. Chemiker Zeitung, 95(??):880–881, ???? 1971.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Stuewer:1988:BRB

[Stu88] Roger H. Stuewer. Book review: Behram N. Kurşunoğlu, Eugene P. Wigner, and John G. Taylor, Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac and Tributes to Paul Dirac. Physics Today, 41(2):84, February 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v41/i2/p84/s2.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Sucher:1967:BRF

[Suc67] J. Sucher. Book review: a formalism for quantum physics [Lectures on Quantum Field Theory, by P. A. M. Dirac, Belfer Graduate School of Science, Yeshiva University, New York; Academic Press, New York, 1966. 159 pp. $7.50]. Science, 158(3801):623, November 3, 1967. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1722626.

Shishkin:1989:DEEb

[SV89a] German V. Shishkin and Víctor M. Villalba. Dirac equation in external vector fields: new exact solutions. Journal of Mathematical Physics, 30(10):2373–2381, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shishkin:1989:DEEa

[SV89b] German V. Shishkin and Víctor M. Villalba. Dirac equation in external vector fields: separation of variables. Journal of Mathematical Physics, 30(9):2132–2142, September 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sisterna:1990:TVF

[SV90] P. Sisterna and H. Vucetich. Time variation of fundamental constants: Bounds from geophysical and astronomical data. Physical Review D (Particles and Fields), 41(4):1034–1046, February 15, 1990. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.41.1034.

Shishkin:1993:END

[SV93] German V. Shishkin and Víctor M. Villalba. Electrically neutral Dirac particles in the presence of external fields: exact solutions. Journal of Mathematical Physics, 34(11):5037–5049, November 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sisterna:1995:TVF

[SV95] P. D. Sisterna and H. Vucetich. Time variation of fundamental constants: Bounds from local data. In Barbour and Pfister [BP95], pages 403–421. ISBN 0-8176-3823-7, 3-7643-3823-7 (paperback). LCCN QC137 .M33 1995. URL http://

www.gbv.de/dms/bowker/toc/9780817638238.pdf;   http:
//www.zentralblattmath.org/zmath/en/search/?an=0846.
01008.

Salam:1972:AQT

[SW72] Abdus Salam and Eugene Paul Wigner, editors. Aspects of quantum theory. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08600-0. xvi + 268 pp.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Sud:1976:RIF

[SWO76] K. Sud, L. E. Wright, and D. S. Onley. Radial integrals with finite energy loss for Dirac–Coulomb functions. Journal of Mathematical Physics, 17(12):2175–2181, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Szmytkowski:1998:OFW

[Szm98] Radosław Szmytkowski. Operator formulation of Wigner’s Rmatrix theories for the Schrödinger and Dirac equations. Journal of Mathematical Physics, 39(10):5231–5252, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [Szm99].

Szmytkowski:1999:EOF

[Szm99] Radosław Szmytkowski. Erratum: “Operator formulation of Wigner’s R-matrix theories for the Schrödinger and Dirac equations” [J. Math. Phys. 39, 5231 (1998)]. Journal of Mathematical Physics, 40(8):4181, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [Szm98].

Szmytkowski:2001:DDE

[Szm01] Radosław Szmytkowski. Discontinuities in Dirac eigenfunction expansions. Journal of Mathematical Physics, 42(9):4606–4617, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takeuchi:1931:ALG

[Tak31] Tokio Takéuchi. Über die Abnahme der Lichtgeschwindigkeit. (German) [On the decrease in the speed of light]. Zeitschrift für Physik, 69(11–12):857–859, November 1931. CODEN ZEPYAA.

ISSN 0044-3328. URL https://link.springer.com/article/
10.1007/BF01339470.

Takahashi:1979:SSN

[Tak79] K. Takahashi. Soliton solutions of nonlinear Dirac equations. Journal of Mathematical Physics, 20(6):1232–1238, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Talman:2004:OID

[Tal04] James D. Talman. Overlap integrals for Dirac–Slater orbitals. International Journal of Quantum Chemistry, 100(2):109–113, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tamura:2003:RCN

[Tam03] Hideo Tamura. Resolvent convergence in norm for Dirac operator with Aharonov–Bohm field. Journal of Mathematical Physics, 44(7):2967–2993, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taub:1949:SMS

[Tau49] A. H. Taub. A special method for solving the Dirac equations. Reviews of Modern Physics, 21(3):388–392, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.21.388;   http://rmp.aps.org/
abstract/RMP/v21/i3/p388_1.

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon, New York, NY, USA, 1961–1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Taylor:1987:CD

[Tay87a] J. G. Taylor. Constrained dynamics. In Taylor [Tay87b], pages 114–123. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987.

US$10.00. URL http://www.loc.gov/catdir/enhancements/

fy0745/87153334-d.html. Based on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Taylor:1987:TPD

[Tay87b] John Gerald Taylor, editor. Tributes to Paul Dirac. Adam Hilger Ltd., Bristol, UK, 1987. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://www.loc.gov/

catdir/enhancements/fy0745/87153334-d.html. Based on

the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

TorresdelCastillo:1997:SDE

[TCC97] G. F. Torres del Castillo and L. C. Cortés-Cuautli. Solution of the Dirac equation in the field of a magnetic monopole. Journal of Mathematical Physics, 38(6):2996–3006, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teller:1948:CPC

[Tel48] Edward Teller. On the change of physical constants. Physical Review, 73(7):801–802, April 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v73/i7/p801_1.

Teller:1972:CC

[Tel72] E. Teller. Are the constants constant? In Reines [Rei72a], pages 60–66. ISBN 0-87081-025-1. LCCN QC780 .C65.

Teller:1973:DH

[Tel73] E. Teller. The Dirac hypothesis. In Kurşunoğlu [Kur73], pages 351–352. ISBN 0-306-36902-8. LCCN QC793.9 .C67 1973.

Temple:1935:BRP

[Tem35] G. Temple. Book review: The Principles of Quantum Mechanics, by P. A. M. Dirac. Mathematical Gazette, 19(235):301–302, October 1935. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL http://www.jstor.org/stable/
3606137.

TenKate:1968:DAS

[Ten68] A. Ten Kate. Dirac algebra and the six-dimensional Lorentz group. Journal of Mathematical Physics, 9(2):181–185, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thaller:1988:NFA

[Tha88] Bernd Thaller. Normal forms of an abstract Dirac operator and applications to scattering theory. Journal of Mathematical Physics, 29(1):249–257, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thaller:1992:DE

[Tha92] Bernd Thaller. The Dirac equation. Texts and monographs in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-54883-1 (Berlin), 0-387-54883-1 (New York). xvii + 357 pp. LCCN QC174.26.W28

T43 1992. URL http://www.loc.gov/catdir/enhancements/
fy0815/92012288-d.html.

Thaller:2006:DE

[Tha06] Bernd Thaller. Dirac equation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-642-08134-7. xvii + 357 pp. LCCN QC174.26.W28.

Tiwari:2012:SDE

[Tiw12] Suresh Chandra Tiwari. The Supersymmetric Dirac Equation: The Application to Hydrogenic Atoms, by Allen Hirshfeld, Scope: textbook. Level: undergraduates. Contemporary Physics, 53(6):543–544, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tjostheim:1975:NUD

[Tjo75] D. Tjostheim. A note on the unified Dirac–von Neumann formulation of quantum mechanics. Journal of Mathematical Physics, 16(4):766–767, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsoulos:2019:DTS

[TKS19] Ioannis G. Tsoulos, O. T. Kosmas, and V. N. Stavrou. Dirac-Solver: a tool for solving the Dirac equation. Computer Physics Communications, 236(??):237–243, March 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465518303588.

Tombal:1985:MCD

[TM85] Ph. Tombal and A. Moussiaux. MACSYMA computation of the Dirac–Bergmann algorithm for Hamiltonian systems with constraints. Journal of Symbolic Computation, 1(4):419–421, December 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Tolksdorf:1998:EHY

[Tol98] Júrgen Tolksdorf. The Einstein–Hilbert–Yang–Mills–Higgs action and the Dirac–Yukawa operator. Journal of Mathematical Physics, 39(4):2213–2241, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torretti:1987:BRA

[Tor87] Roberto Torretti. Book review: Asim O. Barut, Alwyn van der Merwe, and Jean-Pierre Vigier, editors, Quantum, Space, and Time — The Quest Continues, Studies and Essays in Honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Noûs, 21 (3):442–444, September 1987. CODEN ???? ISSN ???? URL

http://www.jstor.org/stable/2215196.

Trautman:1992:SDO

[Tra92] Andrzej Trautman. Spinors and the Dirac operator on hypersurfaces. I. General theory. Journal of Mathematical Physics, 33(12):4011–4019, December 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tu:1991:AST

[Tu91] Khiet Tu. Analytic solution to the Thomas–Fermi and Thomas– Fermi–Dirac–Weizsäcker equations. Journal of Mathematical Physics, 32(8):2250–2253, August 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Taub:1934:DEP

[TVvN34] A. H. Taub, O. Veblen, and J. von Neumann. The Dirac equation in projective relativity. Proceedings of the National Academy of Sciences of the United States of America, 20(??):383–388, ???? 1934. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). Reprinted in [Tau61b, Paper 24].

Urani:1983:DGC

[UB83] John R. Urani and George P. Barker. Dirac general covariance and tetrads. I. Clifford and Lie bundles and torsion. Journal of Mathematical Physics, 24(10):2407–2410, October 1983. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Urani:1982:GDE

[UK82] John R. Urani and Marilyn H. Kemp. A generalization of the Dirac equation to accelerating reference frames. Journal of Mathematical Physics, 23(3):423–424, March 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Uhlenbeck:1931:NCR

[UL31] G. E. Uhlenbeck and Otto Laporte. New covariant relations following from the Dirac equations. Physical Review, 37(11): 1552–1554, June 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ullmann-Margalit:1986:PSI

[UM86] Edna Ullmann-Margalit, editor. The Prism of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 2, volume 95(2) of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. ix + 250 pp. LCCN Q174 .B67

vol. 95 Q175. URL http://www.springerlink.com/content/
978-94-009-4566-1;   https://link.springer.com/book/
10.1007/978-94-009-4566-1.

Uzan:2003:FCT

[Uza03] Jean-Philippe Uzan. The fundamental constants and their variation: observational and theoretical status. Reviews of Modern Physics, 75(2):403–455, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.75.403;
http://rmp.aps.org/abstract/RMP/v75/i2/p403_1.

VanVleck:1972:TDR

[Van72] John H. Van Vleck. Travels with Dirac in the Rockies. In Salam and Wigner [SW72], pages 7–16. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Villalba:2002:SVE

[VC02] Víctor M. Villalba and Esteban Isasi Catalá. Separation of variables and exact solution of the Klein–Gordon and Dirac equations in an open universe. Journal of Mathematical Physics, 43 (10):4909–4920, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

vanderWaerden:1968:SQM

[vdW68] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics, volume 5 of Classics of science. Dover, New York, NY, USA, 1968. ISBN 0-486-61881-1. xi + 430 pp. LCCN QC174.1 .W3 1968.

vanderWaerden:2007:SQM

[vdW07] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics. Dover, New York, NY, USA, 2007.

ISBN 0-486-45892-X (paperback). xi + 430 pp. LCCN

QC174.12 .W34 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006050791-d.html.

vanEijndhoven:1985:MID

[vEdG85] S. J. L. van Eijndhoven and J. de Graaf. A mathematical interpretation of Dirac’s formalism for quantum mechanics. Lecture Notes in Mathematics, 1162:209–251, 1985. CODEN LN-MAA2. ISBN 3-540-16065-5 (print), 3-540-39747-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081155/.

Vilkas:1998:SOM

[VIK98] Marius Jonas Vilkas, Yasuyuki Ishikawa, and Konrad Koc. Second-order multiconfigurational Dirac–Fock calculations on boronlike ions. International Journal of Quantum Chemistry, 70(4-5):813–823, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=75060;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=75060&
PLACEBO=IE.pdf.

Villalba:1990:ESD

[Vil90a] Víctor M. Villalba. Exact solution of the Dirac equation in a reducible Einstein space. Journal of Mathematical Physics, 31 (6):1483–1486, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Villalba:1990:SVEb

[Vil90b] Víctor M. Villalba. Separation of variables and exact solution to the Dirac equation in curvilinear orthogonal coordinates with cylindrical symmetry. Journal of Mathematical Physics, 31(11): 2702–2707, November 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Villalba:1995:ESD

[Vil95] Víctor M. Villalba. Exact solution of the Dirac equation for a Coulomb and scalar potential in the presence of an Aharonov– Bohm and a magnetic monopole fields. Journal of Mathematical Physics, 36(7):3332–3344, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vijaykumar:2021:CTV

[VKA21] Aditya Vijaykumar, Shasvath J. Kapadia, and Parameswaran Ajith. Constraints on the time variation of the gravitational constant using gravitational wave observations of binary neutron stars. Physical Review Letters, 126(14), April 2021. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

vonNeumann:1928:EBD

[vN28] John von Neumann. Einige Bemerkungen zur Diracschen Theorie des Drehelektrons. (German) [Some remarks on the Dirac theory of electron spin]. Zeitschrift für Physik, 48(??):868–881, ???? 1928. CODEN ZEPYAA. ISSN 0044-3328. Reprinted in [Tau61a, Paper 17].

vonNeumann:1940:DSS

[vN40] John von Neumann. Discussion of De Sitter’s space and of Dirac’s equation in it. Unpublished manuscript. Reviewed by A. H. Taub. Reprinted in [Tau63c, Paper 18, p. 177]., 1940.

vonNeumann:1996:PJN

[vN96] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974–1975. Gift, Nicholas A. Vonneuman, 1993.

Villalba:1990:SVEa

[VP90] V. M. Villalba and U. Percoco. Separation of variables and exact solution to Dirac and Weyl equations in Robertson–Walker space–times. Journal of Mathematical Physics, 31(3):715–720, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vrbik:1994:DEC

[Vrb94] Jan Vrbik. Dirac equation and Clifford algebra. Journal of Mathematical Physics, 35(5):2309–2314, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanEijndhoven:1986:MID

[vSJLEdJG86] S. J. L. van (Stephanus J. L.) Eijndhoven and J. de (Johannes) Graaf. A mathematical introduction to Dirac’s formalism, volume 36 of North-Holland mathematical library. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-70127-3. xv + 430 pp. LCCN QC174.12 .E33 1986.

Vulcanov:2003:CDE

[Vul03] Dumitru N. Vulcanov. Calculation of the Dirac equation in curved spacetimes with possible torsion using MAPLE and RE-DUCE. Computer Physics Communications, 154(3):205–218, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

vonWeizsacker:2002:GPA

[vW02] Carl Friedrich von Weizsäcker. Grosse Physiker: von Aristoteles bis Werner Heisenberg. (German) [Great Physicists: from Aristole to Werner Heisenberg], volume 33078 of dtv. Deutscher Taschenbuch-Verlag, München, Germany, unabridged edition, 2002. ISBN 3-423-33078-3. 376 pp. LCCN QC6. EUR-D 12.50.

Walton:1976:NCA

[Wal76] J. R. Walton. A note on certain asymptotic expressions for the unit-step and Dirac delta functions. SIAM Journal on Applied Mathematics, 31(2):304–306, September 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Waldegrave:2009:PDO

[Wal09] William Waldegrave. Paul Dirac: One of the greatest British minds of the 20th century. The Telegraph, ??(??):??, May

16, 2009. URL http://www.telegraph.co.uk/technology/
4436537/Paul-Dirac-One-of-the-greatest-British-minds-
of-the-20th-century.html.

Warwick:1996:BRR

[War96] Andrew Warwick. Book review: R. H. Dalitz (ed.), The Collected Works of P. A. M. Dirac, 1924–48. Cambridge: Cambridge University Press, 1995. Pp. xxiv + 1310. ISBN 0-521-36231-8. £175.00, $250.00. British Journal for the History of Science, 29(4):488–489, December 1996. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4027707.

Watt:1991:IPI

[Wat91] Stephen M. Watt, editor. ISSAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15–17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Wasserman:1978:TIP

[WB78] Ira Wasserman and Kenneth Brecher. Time invariance of Planck’s constant. Physical Review Letters, 41(14):920–923, October 1978. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://journals.aps.org/

prl/abstract/10.1103/PhysRevLett.41.920.

Wasserman:1970:ESF

[WBD70] A. Wasserman, T. J. Buckholtz, and H. E. DeWitt. Evaluation of some Fermi–Dirac integrals. Journal of Mathematical Physics, 11(2):477–482, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wolfe:1976:LVF

[WBR76] A. M. Wolfe, Robert L. Brown, and Morton S. Roberts. Limits on the variation of fundamental atomic quantities over cosmic time scales. Physical Review Letters, 37(4):179–181, July 26, 1976. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.37.179.

Wehling:2014:DM

[WBSB14] T. O. Wehling, A. M. Black-Schaffer, and A. V. Balatsky. Dirac materials. Advances in Physics, 63(1):1–76, January 2014. CO-DEN ADPHAH. ISSN 0001-8732 (print), 1460-6976 (electronic).

Wilczek:2006:FRM

[WD06] Frank Wilczek and Betsy Devine. Fantastic realities: 49 mind journeys and a trip to Stockholm. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-649-X (hardcover), 981-256-655-4 (paperback), 981-277-430-0 (ebook). ix + 522 pp. LCCN QC75 .W55 2006.

Weatherall:2016:VSP

[Wea16] James Owen Weatherall. Void: the strange physics of nothing. Foundational Questions in Science. Yale University Press, New Haven, CT, USA, 2016. ISBN 0-300-20998-3 (hardcover). 196 pp. LCCN QC28 .W35 2016.

Weber:1987:DWG

[Web87] Joseph Weber. Dirac in 1962, weak and gravitational radiation interactions. In Kurşunoğlu and Wigner [KW87], pages 244–248. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987.

URL http://adsabs.harvard.edu/abs/1987ragp.book...
..D;   http://www.loc.gov/catdir/description/cam023/
86033409.html;   http://www.loc.gov/catdir/toc/cam028/
86033409.html.

Weiner:1977:HTC

[Wei77] Charles Weiner, editor. History of twentieth century physics: Storia della fisica del XX secolo, Proceedings of the International School of Physics “Enrico Fermi” = Rendiconti della Scuola internazionale di fisica “Enrico Fermi”, course 57, July 31–August 12, 1972. Academic Press, New York, USA, 1977. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Wesson:1980:DGC

[Wes80] Paul S. Wesson. Does gravity change with time? Physics Today, 33(7):32–37, July 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See comments [AWW^{+}81].

Weyl:1931:GQ

[Wey31a] Hermann Weyl. Gruppentheorie und Quantenmechanik. S. Hirzel, Leipzig, Germany, second edition, 1931. xi + 366 pp. LCCN QA171 .W5 1931.

Weyl:1931:TGQ

[Wey31b] Hermann Weyl. The Theory of Groups and Quantum Mechanics. Methuen and Co. Ltd., London, UK, 1931. xxii + 1 + 422 + 1 pp. LCCN QA171 .W54. Translated from the second (revised) German edition to English by H. P. (Howard Percy) Robertson.

Weyl:1950:TGQ

[Wey50] Hermann Weyl. The Theory of Groups and Quantum Mechanics. Dover, New York, NY, USA, 1950. xxii + 422 pp. LCCN QA171 .W54. Translated from the second (revised) German edition to English by H. P. (Howard Percy) Robertson.

Webb:1999:STV

[WFC^{+}99] John K. Webb, Victor V. Flambaum, Christopher W. Churchill, Michael J. Drinkwater, and John D. Barrow. Search for time variation of the fine structure constant. Physical Review Letters, 82(5):884–887, February 1, 1999. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

link.aps.org/doi/10.1103/PhysRevLett.82.884.

Wiener:1929:DEE

[Wie29] Norbert Wiener. Dirac equations and Einstein theory. Nature, 123(3112):944–945, June 22, 1929. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v123/n3112/pdf/
123944c0.pdf.

Wigner:1939:URI

[Wig39] E. P. Wigner. On unitary representations of the inhomogeneous Lorentz group. Annals of Mathematics (2), 40(1):149–204, January 1939. CODEN ANMAAH. ISSN 0003-486X (print),

1939-8980 (electronic). URL http://www.jstor.org/stable/

1968551. Reprinted in [Wig89].

Wightman:1972:DE

[Wig72a] A. S. Wightman. The Dirac equation. In Salam and Wigner [SW72], pages 95–115. ISBN 0-521-08600-0. LCCN

QC174.1 .A85 1972.  URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Wigner:1972:TEU

[Wig72b] E. P. Wigner. On the time-energy uncertainty relation. In Salam and Wigner [SW72], pages 237–247. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http:/
/www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Wigner:1987:ADM

[Wig87a] Eugene Wigner. Address delivered at Memorial Meeting in Tallahassee. In Taylor [Tay87b], pages 40–42. ISBN 0-85274-480-3. LCCN QC16.D57 T75 1987. US$10.00. URL http://www.loc.

gov/catdir/enhancements/fy0745/87153334-d.html. Based

on the papers presented at the Memorial Meeting for Paul Adrien Maurice Dirac which was held in Cambridge on 19 April 1985 and on the speeches made at the dinner in St John’s College on the same evening.

Wigner:1987:RPD

[Wig87b] Eugene P. Wigner. Remembering Paul Dirac. In Kurşunoğlu and Wigner [KW87], pages 57–65. ISBN 0-521-34013-6. LCCN

QC16.D57 R46 1987. URL http://adsabs.harvard.edu/abs/
1987ragp.book.....D;   http://www.loc.gov/catdir/description/
cam023/86033409.html;   http://www.loc.gov/catdir/toc/
cam028/86033409.html.

Wigner:1985:RPA

[Wig88] Eugene P. Wigner. Remembering Paul Adrien Maurice Dirac. In Zichichi [Zic88], pages 269–274. ISBN 0-306-43235-8. LCCN QC793.3.B5 I57 1985.

Wigner:1989:URI

[Wig89] E. Wigner. On unitary representations of the inhomogeneous Lorentz group. Nuclear Physics B, Proceedings Supplements, 6(1):9–64, March 1989. CODEN NPBSE7. ISSN 0920-5632 (print), 1873-3832 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0920563289904027.

Reprint of [Wig39].

Wigner:1996:ADM

[Wig96a] E. P. Wigner. Address delivered to the Memorial Meeting [for Paul Dirac]. In Wigner [Wig96d], pages 214–215. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CRA

[Wig96b] E. P. Wigner. Concluding remarks (address at the Dirac Symposium). In Wigner [Wig96d], pages 195–196. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:RPD

[Wig96c] E. P. Wigner. Remembering Paul Dirac. In Wigner [Wig96d], pages 219–230. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96d] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wightman:2001:TFD

[Wig01] A. S. Wightman. Three formulas for the off-diagonal density matrix of a Dirac spinor, with an application. Journal of Mathematical Physics, 42(2):674–685, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Willers:1936:BWHb

[Wil36]  R. A. Willers.  Buchbesprechung: W. Heisenberg, E.

Schrödinger, P. A. M. Dirac, Die moderne Atomtheorie. Die bei der Entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. 45 S. u. 6 Fig. Leipzig 1934, Verlag von S. Hirzel. Preis kart. 2,50 M. Zeitschrift für Angewandte Mathematik und Mechanik, 16(2):126, 1936. CODEN ZAMMAX. ISSN 1521-4001.

Wilkinson:1958:DCN

[Wil58] D. H. Wilkinson. Do the ‘constants of nature’ change with time? Philosophical Magazine, 3(30):582–585, 1958. CODEN PHMAA4. ISSN 0031-8086. URL http://www.tandfonline.

com/doi/abs/10.1080/14786435808565799.

Wils:1991:CES

[Wil91] Patrick Wils. A class of exact solutions of the Einstein–Dirac equations. Journal of Mathematical Physics, 32(1):231–233, January 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Will:1993:GCC

[Wil93a] Clifford M. Will. Is the gravitational constant constant? In Was Einstein Right?: Putting General Relativity to the Test [Wil93b], chapter 9, pages 160–180. ISBN 0-465-09086-9 (paperback). LCCN QC173.6 .W55 1993. URL http://www.loc.gov/

catdir/enhancements/fy0832/92053250-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy0832/92053250-d.html.

Will:1993:WER

[Wil93b] Clifford M. Will. Was Einstein Right?: Putting General Relativity to the Test. Basic Books, New York, NY, USA, second edition, 1993. ISBN 0-465-09086-9 (paperback). xii + 290 pp. LCCN QC173.6 .W55 1993. URL http://www.loc.gov/

catdir/enhancements/fy0832/92053250-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy0832/92053250-d.html.

Willis:1998:SID

[Wil98] Barton L. Willis. Similarity and intertwining of Dirac operators. Journal of Mathematical Physics, 39(3):1451–1457, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wilczek:2003:PMD

[Wil03] Frank Wilczek. A piece of magic: the Dirac equation. In Farmelo [Far03], pages 132–160. ISBN 1-86207-555-7. LCCN Q125 .I88 2003.

Will:2005:RC

[Wil05] Clifford M. Will. Relativity at the centenary. Physics World, 18(1):27–32, January 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/18/1/phwv18i1a27.
pdf;   http://www.nobelprize.org/nobel_prizes/physics/
laureates/1993/.

Wong:1990:ESD

[Won90] M. K. F. Wong. Exact solution of the n-dimensional Dirac– Coulomb equation. Journal of Mathematical Physics, 31(7): 1677–1680, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wotzasek:1991:DWZ

[Wot91] C. Wotzasek. Derivation of the Wess–Zumino term for the chiral Schwinger model using Dirac bracket formalism. Journal of Mathematical Physics, 32(2):540–543, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1996:ERD

[WP96] Y. Wang and Fu-Cho Pu. Exact results on a Dirac-like Lee model. Journal of Mathematical Physics, 37(11):5424–5428, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wright:2016:BSP

[Wri16] Aaron Sidney Wright. A beautiful sea: P. A. M. Dirac’s epistemology and ontology of the vacuum. Annals of Science, 73 (3):225–256, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wuthrich:2011:DEF

[Wüt11a] Adrian Wüthrich. The Dirac equation: Feynman’s great struggle. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 65–111. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064

(electronic). URL https://link.springer.com/chapter/10.
1007/978-90-481-9228-1_4.

Wuthrich:2011:GFD

[Wüt11b] Adrian Wüthrich. The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2011. ISBN 90-481-9227-7 (hardcover), 90-481-9228-5 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 208 pp. LCCN QC794.6.F4 W88 2010.

Wu:1986:TVN

[WW86] Yong-Shi Wu and Zi Wang. Time variation of Newton’s gravitational constant in superstring theories. Physical Review Letters, 57(16):1978–1981, October 20, 1986. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/10.1103/PhysRevLett.57.1978.

Wong:1984:GTT

[WY84] M. K. F. Wong and Hsin Yang Yeh. Group theoretic treatment of the Dirac–Coulomb equation and matrix elements of its tensor operators. Journal of Mathematical Physics, 25(9):2602–2613, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wong:1985:DCG

[WY85] M. K. F. Wong and E. H. Y. Yeh. The Dirac Coulomb Green’s function and its application to relativistic Rayleigh scattering. Journal of Mathematical Physics, 26(7):1701–1710, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:1978:MH

[Yan78a] Chen Ning Yang. SU_{2} monopole harmonics. Journal of Mathematical Physics, 19(12):2622–2627, December 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v19/
i12/p2622_s1.

Yang:1978:GDM

[Yan78b] Chen Ning Yang. Generalization of Dirac’s monopole to SU_{2} gauge fields. Journal of Mathematical Physics, 19(1):320–328, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yonei:1977:ATF

[YG77] K. Yonei and J. Goodisman. Accurate Thomas–Fermi–Dirac calculations for diatomic systems. International Journal of Quantum Chemistry, 11(1):163–178, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yoshioka:1990:QCP

[YI90] Akira Yoshioka and Kiyotaka Ii. The quantization condition in the presence of a magnetic field and quasiclassical eigenvalues of the Kepler problem with a centrifugal potential and Dirac’s monopole field. Journal of Mathematical Physics, 31(6):1388– 1394, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:2009:ICD

[YL09] Fajun Yu and Li Li. An integrable couplings of Dirac soliton hierarchy with self-consistent sources. Applied Mathematics and Computation, 207(1):171–177, January 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martinez-y-Romero:1999:NRA

[yRSBSV99] R. P. Martínez y Romero, A. L. Salas-Brito, and Jaime Saldaña-Vega. Nonunitary representations of the SU (2) algebra in the Dirac equation with a Coulomb potential. Journal of Mathematical Physics, 40(5):2324–2336, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoo:2003:BBN

[YS03] Jerry Jaiyul Yoo and Robert J. Scherrer. Big bang nucleosynthesis and cosmic microwave background constraints on the time variation of the Higgs vacuum expectation value. Physical Review D (Particles and Fields), 67(4):043517, February 15, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.67.043517.

Yang:2001:CDE

[YYH^{+}01] Lihua Yang, Xinge You, Robert M. Haralick, Ihsin T. Phillips, and Yuan Y. Tang. Characterization of Dirac edge with new wavelet transform. Lecture Notes in Computer Science, 2251 (??):129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2251/22510129.htm;
 http://link.springer-ny.com/link/service/series/0558/
papers/2251/22510129.pdf.

Yu:2008:HSI

[YZ08] Fajun Yu and Hongqing Zhang. Hamiltonian structure of the integrable couplings for the multicomponent Dirac hierarchy. Applied Mathematics and Computation, 197(2):828–835, April 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zastawniak:1989:NPS

[Zas89] Tomasz Zastawniak. The nonexistence of the path-space measure for the Dirac equation in four space–time dimensions. Journal of Mathematical Physics, 30(6):1354–1358, June 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zichichi:1984:SSS

[ZDF^{+}84] A. Zichichi, Paul A. M. Dirac, S. Ferrara, H. Kleinert, Andre Martin, Eugene P. Wigner, and Chen-Ning Yang. Special session on symmetries and gauge invariance. In Zichichi [Zic84], pages 725–745. ISBN 0-306-41738-3. LCCN QC793.3.F5;

19845. URL http://usparc.ihep.su/spires/find/hep/www?
irn=1397141.

Zichichi:1991:GMC

[ZdSS91] Antonino Zichichi, Venzo de Sabbata, and Norma Sánchez, editors. Gravitation and modern cosmology: the cosmological constant problem, volume 56 of Ettore Majorana international science series. Physical sciences. Plenum Press, New York, NY, USA; London, UK, 1991. ISBN 0-306-44054-7. xiii + 228 pp.

LCCN QC178 .G63 1991. URL http://www.loc.gov/catdir/
enhancements/fy0820/91040369-t.html.

Zecca:1996:DER

[Zec96] Antonio Zecca. The Dirac equation in the Robertson–Walker space–time. Journal of Mathematical Physics, 37(2):874–879, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zichichi:1984:GIT

[Zic84] Antonino Zichichi, editor. Gauge interactions: theory and experiment, volume 20 of Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1984. ISBN 0-306-41738-3. LCCN QC793.3.F5; 19845.

Zichichi:1988:ONF

[Zic88] Antonino Zichichi, editor. Old and new forces of nature: Proceedings of the twenty-third course of the International School of Subnuclear Physics on Old and new forces of nature, held August 4–14, 1985, in Erice, Sicily, Italy, volume 23 of The Subnuclear series. Plenum Press, New York, NY, USA; London, UK, 1988. ISBN 0-306-43235-8. LCCN QC793.3.B5 I57 1985.

Zichichi:2000:DEP

[Zic00] Antonino Zichichi. Dirac, Einstein and physics. Physics World, 13(3):17–18, March 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

physicsworldarchive.iop.org/full/pwa-pdf/13/3/phwv13i3a17.
pdf.

Zorbas:1980:PSA

[Zor80] J. Zorbas. Perturbation of self-adjoint operators by Dirac distributions. Journal of Mathematical Physics, 21(4):840–847, April 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1990:WFB

[ZQ90] Jian Zu Zhang and Yong Chang Qi. Wave functions of bound states of a fermion and a Dirac dyon and matrix elements in an external electromagnetic field. Journal of Mathematical Physics, 31(7):1796–1799, July 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Xing:2006:WDM

[zXZ06] Zhi zhong Xing and He Zhang. Why do Majorana neutrinos run faster than Dirac neutrinos? Electronic Journal of Theoretical Physics, 3(10):191–209, ???? 2006. CODEN EJTPBC. ISSN 1729-5254. URL http://www.ejtp.com/

articles/ejtpv3i10p191.pdf. Special Majorana Centenary

Issue: Majorana Legacy in Contemporary Physics.